A Complete Bibliography of Publications in

Marine Resource Economics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

01 April 2025 Version 1.12

Title word cross-reference

$49.85 [All98]. 0 [All98]. 0-8133-8942-9 [All98]. 1 [Tow03]. 11 [Kna95]. 13 [Gat86]. 1955 [Sch95]. 1990 [Poo97]. 1994 [Ano95g]. 1995 [Ano95g, Ano96b, Mun00]. 1996 [Ano96b, Ano97b]. 1997 [Ano97b, Ano98b]. 1998 [Ano98b, Ano99b]. 1999 [Ano99b]. 2000 [Cha09]. 2009 [And10]. 2030 [CER21]. 21st [SH04]. 26 [Com89]. 42-Year [ASS18]. 86-4 [And87c]. 9 [All98]. 90 [Tow03]. 90-5166-882-1 [Tow03]. Abalone [HGJH15]. Ability [KC95]. Acceptance [EK95, LAM21]. Access [And87c, And02b, Chr96, Gom92, KH06, KBS01, KBS02, Mai19, Mer23, Sto86b, And04, Kro04]. Accident [BP97]. Accidental [AKKM86]. Accommodating [HBB07a]. Accomplish [AR01, Han98]. Account [Arn93, ZAS^{+}24]. Accounting [DTS20, Edw05, EAÖH25, PH20, WB09, ZFG20]. Achieved [FJ03]. Achievement [And04]. ACIAR [Com89]. Acidification [MF22]. Acres [HD25]. across [KBK^{+}21]. Act [Cro16, Fel02, MSI01, CSW23, WR97]. Activities [JL99]. Actual [Cav23]. Adam [And15c]. Adaptive [Wil87]. ADB [Jin96]. ADB/INFOFISH [Jin96]. Adding [Tro01]. Address [OLM^{+}21]. Addressing [LLLV16]. Adjustment [Asc97a, NTS11]. Adjustments [DG94]. Administrative [CS13]. Administrator [SP85]. Advances [Arn07a]. Advancing [CKMT15]. Advent [Ale87, Chr96]. Advertising [BSL92, KNX95, KP01, KM02]. Advisories [Ali11]. AFA [And02c]. Afar [ARA^{+}25]. Affect [Álv22, LL13]. Affected [RNFH22]. Affecting [AMW^{+}99, HNE^{+}89, NGBM95]. Affordable [Cro16, CSW23]. African [HLH^{+}14]. after [SRS^{+}13]. against [MUWB23]. Age [Gil88, QT19, Tah09]. Age-Structured [Gil88, QT19, Tah09]. Agency [And87b]. Agenda [CER21]. Agent [JV01]. Agglomeration [Tve02a]. Aggregate [TP85]. Aggregation [XM11]. Agreement [BDCDeS00, Li98a, Mun00]. Agricultural [DS88]. Agriculture [AMW^{+}99, Ken06, Swa94]. AIDS [HCC23]. Aimed [PCAIGO^{+}14]. Air [MFF^{+}23]. Alan [Bos13, Poo97]. Alaska [Ano93b, BHBG93, Boy93, CHW09a, GHM95, Her93, Kna96, Kna97, LS14, Mat09, Mer23, PA09, SCK^{+}15, WS10, And87c]. Alaskan [Hol00]. Aleutian [AGYW10a]. Algae [MJ20, WMB95]. Algal [MFF^{+}23]. Alie [Com93]. Alien [FPA13]. Aliger [MVRRP21]. Allen [Gat86]. Alligators [HF01]. Allocate [Row04]. Allocating [CN95]. Allocation [Abb15, ALP^{+}21, AH02, BC84, Eas92, EA89, Gar88, HJKP03, KG98, KM02]. Allowances [AS05]. Allsopp [Pol86]. Almon [HMMhL92]. Almon-Type [HMMhL92]. Almost [Tho12]. along [Bea25]. Alternative [AS89, Kin03, LLR10, MS01, MSI01, SYN^{+}19, Sto87, SC95, THM98, YH96]. Alternatives [LKK03a, LLA00, McT99]. Altruistic [SS01]. Amenity [BS06a]. America [PIH14, TB01]. American [And10, Fel02, MSI01, RG86, WK88, dFdOdA24, SGC^{+}16]. americanus [GWB94a]. Amoco [GAB^{+}86a]. Among [BHG^{+}16, BM96, PPV96, Str04]. ANA [ZFG20]. Anadromous [Ano89a, Hup89]. Analysis [AMW^{+}99, And94b, And99, And02b, And02c, And07, And15c, Ano95b, AKT^{+}22, BaUS04, BM96, BHBG93, BA20, CM85, CV93, CYP92, Cox07, DKK11, DAP11, Dia96, EU99, Egg06, EK95, EOA^{+}17, FW10, GDF04, GD05, HLH^{+}14, Hee89, Her88, HBCP09, mHS06, JV01, JHT06, JKPHS12, JLT19, JRBH06, Kat07, KNH93, Koe84, Kum02, KGW22, LLLV16, LS99, LA11, LLA00, hLRT87a, LF11, MCFH09, Mat09, MF22, NP97, NQ06, NDW22, PC94, Pas07, PGB00, PA20, Poo97, Qua06, RLK22, Roh07, RGA07, RNFH22, RL14, SLW85, STNY23, Sal89, San04, SW18, SCK^{+}15, Seu24, SDS12, Smi23, SE94, TKMA17, TYH15, Ter07, TH12, TBA95, THM98, TFB18, VD05, Wal06, WA23, WTW^{+}19, WS10, Wau94, WA92, Wil07, WG03]. Analyzing [PSL03]. Anchoaveta [Sai88]. Anchovies [TPPT11]. Anchovy [ALP^{+}21, ESC18, KB05]. Anderson [Wau87]. Angler [AW88, CHLH03, HHSW12, LL13, SB85, WCH01]. Anglerfish [GIP02]. Anglers [Ano89a, DNA^{+}12, Hup89, Mil88, Ola09, TB97]. Angling [Bel89]. Animal [AYJO^{+}19a]. Ann [Cha95]. Announcements [SHL^{+}20b]. Annual [And15c, Ken92]. Antarctic [KT94]. Anthonie [Gat86]. Anthony [MDC^{+}15]. Antidumping [KP08]. Any [WM16]. Applicable [Swi85]. Application [AH93, And15a, Ano89a, BT14, BLS16, BBI05a, DAP11, DBP00, Fou02, GA95, GDF04, GD05, HLM20, Hup89, Hut14, LRCL25, MB10, PS03, SB85, SRTS86, SL98, Smi96, SE94, WB09, Yoh84]. Applications [And85b, GOFR88, Jen02, MP99]. Applied [AKV^{+}17, SRA^{+}23, Swa94]. Applying [HMMhL92, LSW17]. Appraisal [GN88]. Approach [AB93, Bos13, DAP11, Fel09, GBDT08, GMS97a, HF07, HH09, HCC23, Ken87, Kli87, LHT96, LG07, LP01, NP97, NSV^{+}22, PCAIGO^{+}14, PWL05a, STY20, Smi96, SF19, TKMA17, TS02, VH11, ZAS^{+}24, dFdOdA24]. Approaches [NNNJ22, OLM^{+}21, SC95, WT10, ZS16]. approaches. [GPC^{+}13]. Appropriate [ST87]. Approximation [Hof04]. Aquaculture [AR10, AGR19, ACDL21, And85a, And02a, Arn92, AT02, ABS03, ART07, AGT08, AG14, AGC^{+}22, ARA^{+}25, CDNS17, CAAT10, FJ23, Gat86, Gut02, Gut08, HO11, HJKP03, Joh85, KC95, Kin93, LA11, MW02, Mik07, MOOT23, NAB14, PDS^{+}21, PWC22, RNKA21, RS96, SJQ21, Sal93, Sha11, TY05, Tve02b, Tve02a, YB96, YL06a, YL06b, ZWS21, All98, All95]. Aquaculture-Fisheries [Mik07]. Aquatic [GGKW17, Smi23]. ARCH [Ali11]. Arctic [KHB^{+}25]. Arcto [Arm94]. ARDL [Gor20]. ARDL-Bounds [Gor20]. Area [AAO^{+}09, CN00, DR04]. Areas [DAP11, GM07, Hal04, HNAY^{+}21, NF10]. Arid [MS01]. ARMAX [Ali11]. Arnason [Rus00]. Art [HH99]. Arthur [And15b]. article [HBB^{+}07b, Ano09g]. Artificial [Mil88, MMH09a]. Artisanal [ATR^{+}23, AAC^{+}21, Cav23, DCB15, Esp15, PCAIGO^{+}14, Sei93, TH12]. Asche [Kri09]. ASEAN [WTW^{+}19]. Asia [AK97, BP94, VM86, ZS09, Mar86a]. Asian [Bro86, Tis86]. Aspatial [AAC^{+}21]. aspects [Poo94]. Assessing [LSH^{+}18]. Assessment [AS05, GST22, GW92, Opa87, PCAIGO^{+}14, Ter07, VD05, Wil88]. Assessments [Poo97]. Asset [RCLL17, SK21]. Assigning [Lib07]. Assistant [ARS14]. Associated [HDR96, Kah87]. Association [And10]. Asymmetric [FPdR20, Jaf04]. At-Home [TYH15]. Atlantic [AL18, ADG^{+}11, BB06, BER^{+}14, BT14, BDCDeS00, BM00, DBP00, Edw01, Edw05, ES16, Gar88, Gar93, GH86, Ken03, LLA00, Lee14, Ola09, PTE04, PD02, Str04, THM98, WGM86, ZT22, ZSG23]. Attendance [Lew19, AK20, HLM20, LW20, LWL^{+}20, PH20, ZFG20]. Attitudes [BP06, Kna96, RAC^{+}17]. Attribute [AK20, HLM20, Lew19, LW20, LWL^{+}20, PH20, WCFZS03, ZFG20]. Attributes [NSV^{+}22, RGA07, SVS99, Wes02]. Auction [Álv22, Cav23, HVTY05, WAUR17, ZSG23]. Auctions [EVH05, FW10, PTMQ22]. Australia [Bos04, SPC18, SPZ19]. Australian [Sta87, Tis86]. Author [Wil18a]. Average [DF16, Ham15]. Avoidance [BWR^{+}23, SW15]. Award [Ano09g, SHL^{+}20b, Ano06c, Ano07c, Ano08c, HBB^{+}07b]. Away [ZFG20]. Away-from-Home [ZFG20].

B [BP94, Com93, Geo87, Geo96, Kin93, SYN^{+}19]. Back [AKP^{+}98b, ASP^{+}84, ASB^{+}85, ASR^{+}86, ABS^{+}93, ALL^{+}94, ASH^{+}92, AKN^{+}93, AKG^{+}97, AGK^{+}01, AAR^{+}08, AHT^{+}09, BCH^{+}00, BLE^{+}05, BSS^{+}86, BCA^{+}87, BYD^{+}88, BBH^{+}06, BMS^{+}89d, BMS^{+}89c, BHB^{+}93, BMD^{+}00, BBI^{+}05b, CUW^{+}07b, CHW^{+}09b, CDT^{+}95, CGP^{+}03b, CMS^{+}85, CMC^{+}99b, tCTH^{+}93, CMGB^{+}07b, CPG^{+}98, CPB^{+}92, EWR^{+}86, ELP^{+}07, ESO^{+}03, FAH^{+}02, FRK^{+}04, FFH^{+}95, GWB^{+}94b, GNC^{+}88, GSEM^{+}97, GMS^{+}97b, GAB^{+}86b, GWR^{+}00, HDS^{+}85, HMS^{+}86, HCL^{+}98, HBB^{+}96, HBB^{+}07b, HHN^{+}89, IOE^{+}99, JMF^{+}02, JRB^{+}06, KHD^{+}96, KAC^{+}97, LKK^{+}03b, LPV^{+}01, hLRT^{+}87b, LVV^{+}05, LAG^{+}04, LGO^{+}88, MPH^{+}99, MCC^{+}92, MMP^{+}04, MDS^{+}88, MMH^{+}09b, MBD^{+}00, NGB^{+}95, OBB^{+}84, PBA^{+}94, PMT^{+}99b, PWL^{+}05b, RCN^{+}09, RHJ^{+}92b, SSH^{+}89, SBK^{+}87, SLM^{+}98, SAS^{+}94, TBH^{+}01, TBC^{+}86, TAW^{+}02, TTS^{+}02, VGH^{+}96, WMS^{+}01, WDS^{+}85, WKGK87, WSW^{+}04, WGW^{+}03, YCN^{+}95, YBG^{+}96, YLB^{+}06, ZBP^{+}06b]. Bad [Ham15]. Bag [CLL22, WDLH11, WG03]. Bailey [All98, Com88, Com93]. Bait [RHH10]. Bait-Constrained [RHH10]. Baitfish [ALP^{+}21]. Baltic [BHW15, GSJ13, Ham15, Hut14, Kro04, NLH12, OGL^{+}16, Gri00]. Ban [CL98]. Bandwagon [Ola09]. Bangladesh [AGR19, CL98, DAP11, GBDT08, HNAY^{+}21]. Bankruptcy [GIP02]. Banks [SS16]. Barange [Sch07]. Barbara [ARS14]. Barbed [Hol00]. Barents [SK21, Sum97]. Bargaining [BHW15, TPC03]. Barrier [PgK09]. Barriers [HD25]. Barter [Hol13]. Based [Arn07a, CKMT15, CP92, DF16, ESC18, LRCL25, LAM21, MJ11, Neh93, OHS14, Sha11, SF19, VW13, WH99, Yam95, CNAB09, HRSW18, Tow03, Tur96]. Baseline [LLR10, Sai97]. Basic [AS84, And15a]. Basin [AMW^{+}99, Boo89, CC86, Rud88, WT05]. Bass [BK89, RS96]. Bay [FGMW03, GMS97a, ATR^{+}23, BMS89a, Cum87a, Cum87b, GKK^{+}88, HTTA92, Hua15, KW09, Kli87, Lip04, MB10, MPE10]. Bayesian [HT06, SZC07]. Be [SDQ21, WB86, Cum87b, FJ03, Mat96]. Beach [AED11, cHPZ07, LKK03a, PMT99a, PKLB09, PCH^{+}13, PHCK16, PPQ^{+}19]. Beaches [BLEV05, LSH^{+}18]. Beal [Loo01]. Beam [VD05]. Beam-trawler [VD05]. Behavior [And99, And04, DNA^{+}12, Jen98, MEA^{+}23, MWH^{+}18, NQ06, Roy98, SS01, Ves96]. Behavioral [CMC99a, MS11, OB84]. Behaviour [NTS11, OS08]. Benefit [BHBG93, JL99, Kat07, ZS09].

Benefits [AED11, BHK97, BMS89a, ES16, Fre95, GM07, HTTA92, Her06, HS86, LSH^{+}18, MPTB22, OGL^{+}16, PWC22, RG86, Sch98]. Bengal [ATR^{+}23]. Bergstrom [Bos13]. Bering [AGYW10a, CH25]. Better [Han10, MU16]. Between [And85a, AH02, BC84, CN95, Tve02b, YB96, BAYN16, CHN12, FKAvR10, GLN^{+}23, HVTY05, HGJH15, HNAY^{+}21, JHT06, NL09]. Beyond [ALS^{+}18]. Bibliography [WBC^{+}13]. Bicriteria [PC94]. Bidder [ZSG23]. Bids [KTR00]. Big [Ham15, Mat96, MN18]. Bigeye [OLM^{+}21]. Bighead [EK95]. Bight [Con85]. Bimatrix [Sum95]. Bio [PRS92]. Bio-Economic [PRS92]. Biodiversity [GS98, Sai97, SS16]. Bioeconomic [AKV^{+}17, BSA^{+}22, BaUS04, CHS93, CV93, CYP92, CPW03, EU99, Gil88, GST22, GDF04, JH05, JKPHS12, JLMR96, LSHQ06, LLA00, LSW17, Loo88, MB10, NLH12, PC94, Res95, RHH10, Sam92, San04, SE94, TP85, THM98, War94, Wau94, YB96]. Bioeconomics [AS84, CR15, Ken06, NC07, SW04, Smi23, Gat86]. Biography [Gud98]. Biological [KB05]. Biology [ADK^{+}11, Swi85]. Biomass [CLSA22]. Biscay [MB10]. Bixby [Tow06]. Bjørn [Tow03]. Bjørndal [Arm13]. Black [KBS01, KBS02, KB05]. Blooms [MFF^{+}23]. Blows [HBW22]. Blue [BER^{+}14, Hua15, HD25]. Bluefin [BB06, BDCDeS00, DBP00, GN88, HH86, KW86, PD02, Sha11, Tok18]. Boat [Álv22, CMLC16, Ham98, HT06, WDLH11]. Boaters [Lip04]. Boating [TS02]. Boats [Kva18]. Bonnie [Geo06]. Bonny [AAO^{+}09]. Book [ALS^{+}18, Ale87, All95, All98, And87a, ADvH^{+}18, And96, ACH^{+}12, Arm13, ALG^{+}19, BP94, Bos13, BHH^{+}12, Cha95, Cha00, Cha09, Com88, Com89, Com93, Cur85, DPP^{+}22, Fel94, Gat86, Gat96, Geo87, Geo96, Geo06, GPC^{+}13, GHH^{+}15, Gre99, Gri93, Gri00, Her88, Her95, KFS^{+}15, Kin93, Kna95, Kri09, Kri19, Kva18, Lin13, Loo01, Mar86a, McC02, Meu95, Neh93, Pan87, Pol86, Poo97, Ret88, Rob93, Roh03a, Rus00, Sai88, Sai97, Sch95, Sch07, Sla07, SP85, SHK^{+}13, SGC^{+}16, SMRK85, Tow03, Tow06, Tow09, Tyr01, Van96, Vis01, Vis02, Wau87, Wau88, WBC^{+}13, Yoh03, HRA^{+}23]. Books [GTP^{+}11, HLA^{+}11, HHN^{+}12, Pas12, SAB^{+}12]. Bootstrapped [Wal06]. Borg [Lin13]. Borgese [Kna95]. Both [SK21, YM02]. Botsford [Gat86]. Bottom [ESOF03, PWC22, SW18]. Boulder [All98]. Boundary [SS16]. Bounds [Gor20]. Bowhead [Dow93]. Brands [RGA07]. Branzino [AGC^{+}22]. Brazil [Cav23]. Breadth [Smi17]. Britain [Bur92, FR04]. British [CDTW95, Hol13]. Brown [HNCS12, SABO14]. Bruce [Tow09, Roh03a]. Budgeting [DAP11]. Building [Abb17]. Bumpy [BT14]. Business [MUWB23, Tow03, ZT22]. Buyback [FW10, FGMW03, WJ12]. Buyout [KTR00]. Bycatch [AS94, HJ97, LHT96, SW15, War94].

C [Bos13, Fel94, DBP00]. C-Games [DBP00]. C. [Com88]. Cadiz [GAB^{+}86a]. Cake [FPA13]. California [AW88, Ano89a, DR04, Hup89, LSH^{+}18, LZ87, MJ20, RLK22, SW04]. Campbell [Com89, Her95]. Can [Asc07, MUWB23, PTSB99, WB86]. CAN$72.00 [All98]. Canada [Cha09, Tow06, CC87, Coo88, CP92, EQJ94, Gar88, HG89]. Canadian [DS85a, DS85a]. Cancelled [WHL^{+}18]. Canned [EK95]. Cap [SW15]. Cap-and-trade [SW15]. Capacity [AMM98, And07, Asc07, CGS05, CCBM85, Cox07, Fel02, GD05, Han93, HF07, HG89, Hol99, KSV05, KPS04, Lin05a, Lin05b, NTS11, Pas07, Roh07, RJY16, Ter07, TP05, VD05, Ves05, WT10, WA23, WMT04, Wil07]. Cape [DvHM18, LLLV16]. Capital [BC87, CM17, HKPW13, HdSK15, LN16, Mat10, RJY16, Sto86a, Sum95, Wil09, WCA^{+}10]. Captain [Kna96]. Capture [AAO^{+}09, Com88, Sha11, YB96]. Capture-Based [Sha11]. Captured [HNAY^{+}21]. Care [And02a, Cro16, CSW23]. Career [SRA^{+}23]. Carl [Wil87]. Carmel [Kva18]. Carolina [BLEV05, Cro16, HNCS12, PS03, WDLH11]. Carp [EK95]. Carson [WBC^{+}13]. Case [ASS18, ATR^{+}23, And05, AYNN19, AYAB^{+}20, Ano19a, Ano20e, Ano20f, Ano20g, AMO19, AGC^{+}22, BBA^{+}10, BSL92, DAAY^{+}20, Dow93, ESC18, FHS95, GR11, GD05, Han85b, HNCS12, Jaf04, Kat07, Ken92, KP01, KTR00, LLR10, LCB92, Loo88, MEA^{+}23, MJ20, Mat10, MG16, NK18, NLH12, OGL^{+}16, PT97, PRS92, San05, SPC18, SYN^{+}19, SABO14, Sum97, Tow06, Ves96, WJN^{+}16, WTW^{+}19, WMB95, WJP^{+}02, Wil09, Wil18a, Wil18b, ZSG23, WCA^{+}10]. Cases [Tis86, Van96]. Cash [MUK23]. Catch [Arn94, BP97, BKS^{+}23, Cla85, CC87, CP98, EG13, GAW^{+}18, Han88, mHS06, MKN23, NV12, PPV22, SAU12, Sch98, SGC^{+}16, WH99]. Catch-Effort [CP98]. Catcher [CH25]. Catcher/Processor [CH25]. Catches [KT94]. catching [MSBh95]. Catfish [KP01, Kou95, NQ06, Qua06]. Caught [PDS^{+}21, RLA10]. Cause [AMO19]. Causes [BS89]. Cave [HM11]. Centenary [Egg12]. Center [dFdOdA24]. Central [BCHH00, CGP03a, OLM^{+}21, Ano89a, Hup89, QXH11]. Cents [Hol13]. Century [SH04]. Cephalopod [DK96]. CERCLA [GOFR88]. Certainty [HLM20]. Certification [CDNS17, MW02, Wak14]. Certified [Wak14]. CES [Hof04]. CGE [WS10]. Chaffee [Roh03a]. Chain [ART07, FPdR20, HLH^{+}14, Jaf04]. Chains [Kri09]. Challenge [CER21]. Challenges [AGT08, GS98, Wil13]. Change [Cro16, Edw92, Gre99, GMT11, Kva16, MKN23, Sch07, SPC18, Sai88]. Changed [Gud98]. Changes [AA00, HM95, Hut14, Ken92, XZ14, vBDSD10]. Changing [Bea25, Joh96b, MF02, NSG12]. Channel [BBA^{+}10]. Chaos [CK94, GSE97]. Characteristic [Li98a]. Characteristics [BWR^{+}23, BLS16, HDR96, vBDSD10]. Charges [And88]. Charr [KHB^{+}25]. Charter [WDLH11]. Chaston [Joh85]. Chennat [Bro85]. Chesapeake [KW09, BMS89a, BK89, Cum87a, Cum87b, GKK^{+}88, Hua15, Kli87, Lip04]. Chet [Roh03a]. Chicken [Ola07]. Chile [ACDL21, AHTT09, Esp15, ESC18, PT97, PTSB99]. Chilean [Gon96, KSPT^{+}17, LG98, PTDV17]. Chin [GPC^{+}13]. China [Chu86, Jin96, YAA20]. Chinese [ZFG20, ZWS21]. Chinook [Boo89]. Choice [AK20, AYJO^{+}19a, BHK97, Ber06, BCSA22, BM00, DBR^{+}19, EW86, GLL^{+}10, HH99, HLM20, HBCP09, Lew19, LW20, LWL^{+}20, Mil88, NGBM95, NSV^{+}22, PH20, PH99, RAW17, RCLL17, WH99, ZWS21]. Choices [BHG^{+}16, Dup93, Esp15]. Christopher [Poo97, Vis01]. Christy [Pon89]. Citation [Egg06]. Clam [Wan95]. clarkii [JLMR96]. Classes [Mai19]. Classifying [Ham98]. Clem [Ken06, Tyr01]. Climate [Gre99, ILRS18, JBH24, MKN23, MM04, SPC18, Sch07]. Climate-Friendly [ILRS18]. Clonts [All95]. Closure [CN00]. Closures [AP17, tCT93, DTS20, PKLB09]. cloth [All98]. Club [GLM^{+}25]. Clupea [KJE03, San05]. CO [All98, AKP98a, Arm94]. Co-Management [AKP98a, Arm94]. Coalition [KL07, LK00]. Coasean [OLM^{+}21]. Coast [RCLL17, SYN^{+}19, FKAvR10, HS86, LSJN^{+}24, LSW09, Mat10, Sha11, WJ12, Wil09, WCA^{+}10, Tow06]. Coastal [ALP^{+}21, AA00, BHJ03, BC13, BCP23, EAÖH25, FP25, HD25, HLM20, KNYA23, NDW22, Par87, PIH14, PH20, SB87, Smi17, Swa94, UUL^{+}10, Van96, Whi93, ZS09, Kin93, Yoh03]. Coasts [Bea25]. Cod [Arm94, BHW15, DSBT97, ESOF03, GHB93, GH96, Gud98, Ham15, Kro04, Lee14, PA20, Poo97, SRS^{+}13, ZSG23]. Coefficients [RLK23]. Cognition [PPV96]. Coho [BHJ03]. Cohorts [Sch01]. Cointegration [BM96, RNFH22]. Cold [AYSN17, FKAvR10]. Collapse [BB06, LS21, MMSS12, PSK15]. Collective [CUW07a, ESC18, HRSW18, PTDQ^{+}19, ZS16]. Collusion [PTMQ22]. Colombia [MVRRP21]. Color [SAR05]. Columbia [AMW^{+}99, Boo89, CDTW95, Hol13]. Columbian [Ale87]. Combined [PCAIGO^{+}14, WDLH11]. Combining [GPC^{+}13, PCH^{+}13, Smi23]. Comment [And94a, AR01, Boy00, Gra96, Li09, NC07, RC89, WCA^{+}10, Wil18a]. Comments [Gul89, Kna97, Pon89, ST89, And07]. Commercial [AY12, Abb15, And85a, BHK97, Boy93, Cro16, CSW23, HLH^{+}14, KNYA23, KHB^{+}25, LSJN^{+}24, LLS25, Loo88, NF10, RCLL17, Sto87, aMUSS02, Wat01, WJ12, WDJK20]. Commercially [SRTS86]. Commission [DS85a]. Common [ASS18, And85a, BSL92, ESC18, GC16, Lib07, Maz04, Meu95, SS01, Smi97, Wil18a, Wil18b]. Common-Pool [GC16]. Common-Property [And85a]. Commons [Hol00]. Communities [ALP^{+}21, BHJ03, BCP23]. Community [Bea25, Neh93, Sei93, WB09, Yam95, Fel94]. Community-Based [Neh93, Yam95]. Companies [MN18]. Comparing [Kli87, KGW22]. Comparison [And02b, AS89, CAAT10, hS98, SC95, WT10, WCH01]. Compensating [TS02]. Compensation [DS87, EC89]. Compensatory [MPTB22]. Competing [NC07]. Competition [NL09, SK21, XZ14]. Competitive [Ken87]. Competitiveness [ART07]. Complex [DBR^{+}19, HBB07a]. Compliance [DCB15]. Comprehensive [OAA04, WBC^{+}13]. Computable [Ken87, WS10]. Concept [SK99]. Conceptual [Par87]. Conch [MVRRP21]. Condition [PPQ^{+}19]. Conditional [AY12, MUK23]. Conditions [DKK04]. Conflating [RAH17]. Conflict [NF10]. Conflicts [Tis86]. Congestion [LRCL25]. Congressional [Abr85]. Conjoint [AB93]. Connecticut [GMTW19]. Connections [AKV^{+}17]. Conner [Com93]. Connor [All98].

Consequences [GST22, Mat97]. Consequentiality [ZWS21]. Conservation [AMM98, ACH^{+}12, DR04, HW12, MMCM13, MP97, NSV^{+}22, OLM^{+}21, Rus99, WR97, War94, Lin13]. Considerations [HH99, KT94]. Consistent [XM11]. Consolidation [AY12]. Constrained [RHH10]. Constraint [Gom92]. Constraints [Fel94, Yah88]. Construction [Res95]. Consumer [AYAB^{+}20, AW18, BLEV05, Boc84, EK95, LAM21, MWH^{+}18, NGBM95, OHS14, RLK23, RAC^{+}17, SAR05, ZFG20, ZWS21]. Consumers [AYJO^{+}19a]. Consumption [Boc84, BA20, LGL19, MUK23, OHS14, TYH15, ZFG20]. Contamination [AED11, LZ87, MWH^{+}18, WMB95]. Contemporary [And85b]. Contents [Ano95a, Ano96a, Ano97a, Ano98a, Ano99a, Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a]. Context [DSBT97, EOA^{+}17, RHJ92a]. Continential [RRS88]. Contingent [LL94, WBC^{+}13]. Continuous [And86, And88]. Continuous-Time [And86]. Contract [Esp15, HRK86]. Contracts [LA11]. Contradictory [HH09]. Contributions [BS13, ZBP06a, Com93]. Control [And87c, AS94, dLS92, DS88, cHPZ07, Seg87, SHL^{+}20a, WT05]. Controlled [GHM06]. Controls [And88]. Convenience [CDNS17, EOA^{+}17]. Convex [Ber23]. Cook [ALG^{+}19]. Cooperation [CGP03a, GLM^{+}25, LF11, MM04, SK21, YGB97]. Cooperative [Arm94, FLS14, KNX95, Li98a, PD00, mHS06]. Cooperatives [HKPW13, HdSK15, MSI01, PPF88, Sto87]. Coops [And02c]. Coordination [HRK86]. Cope [OG03]. Coral [FKAvR10, MMCM13, WCFZS03]. Corporation [Cha95]. Correct [MU16]. Correction [BLS16, CL09]. Cost [AAW^{+}17, And15c, BHBG93, EAÖH25, GSJ13, GAB^{+}86a, Kat07, KG97, LSH^{+}18, LHCL15, Lew19, Loo01, MB10, PAC^{+}22, Sal93, SB85, San04, Smi89, TB97]. Cost-Benefit [Kat07]. Cost-Effective [EAÖH25, GSJ13, San04, TB97]. Cost-Recovery [KG97]. Cost/Benefit [BHBG93]. Costly [SW15]. Costs [AKP98a, AR01, Boy93, CS13, CBI92, DKK04, DvHM18, FGM^{+}98, GH86, GAW^{+}18, Han88, MMSS12, Par87, PWC22, Tve02a, WDJK20]. Could [FJ03]. Council [Roh03b]. Councils [Pow04]. Count [BA20]. Counters [HDR96]. Countries [ALP^{+}21, Arn23, GST22, HNAY^{+}21, Kin93, NFW12, YGB97]. Country [VOJ^{+}01]. County [NDW22]. Crab [AGYW10a, FPA13, GHM95, Hua15, MJ20, Mat09, SK21]. Crayfish [JLMR96, Kat07]. Crisis [AHTT09, Han15, Rob93]. Criteria [dLS92, MP99]. Critical [PSK15]. Critique [GHH^{+}15]. Crop [And94a, PS89, PS94]. Crossroads [BCP23]. Crutchfield [And15b]. Culling [Hea95]. Culture [BS89, HRA^{+}23]. Cultured [HNAY^{+}21]. Cunningham [And87a]. Currency [Str14]. Current [HH99, LWL^{+}20]. Curse [Asc97b]. Curtis [All95]. Curves [AY12]. Customary [RHJ92a]. Cutter [Smi96]. Cycling [MS01].

D [ALS^{+}18, Her95, Sch95, Vis02, Yoh03]. D. [Sai88, Sai97, SP85]. Dale [ACH^{+}12]. Dalton [MDC^{+}15]. Damage [Opa87, WCFZS03]. Damages [GOFR88, Kah87, SB87]. Dams [Boo89]. Danger [PPV96]. Daniel [Cha00]. Danish [And05, AYAB^{+}20, HF07, NAB14]. Data [Boy93, BA20, CL09, CS86, FP25, FJ03, GPC^{+}13, GD05, Hol13, Kli87, LHCL15, PCH^{+}13, PGB00, SDS12, TYH15, Wal06, WA23, WDJK20, WDLH11, WHL^{+}18, VD05]. Database [Wan92]. Dataset [CLSA22]. David [And87a, Cha00, Geo87, Geo96, Her88, Vis02, Wau88]. Davis [CV93]. Day [LSBH21]. Day-to-Day [LSBH21]. days [Cha09]. DEA [VD05]. Dealing [MS01]. Death [Chr96]. Debris [LSH^{+}18]. Decades [Sai88]. Decision [CUW07a, MP99, Tom10]. Decision-Making [MP99]. Decisions [CCBM85, CHLH03, CM04, LGOS21, NGBM95, PTDV17]. Decline [Boo89, BS89]. Decommissioning [GDF04]. Deepwater [WHL^{+}18]. Defend [MUWB23]. Deffeyes [Sla07]. Defined [MU16]. Definition [PH99]. Deforestation [PB94]. Degradation [LLS25]. Delaware [MPE10]. Delay [KW09]. Delft [Tow03]. Delineating [CD07]. Deliver [Smi23]. Delivering [SAA^{+}20]. Delta [NSV^{+}22, vBDSD10]. Demand [AP17, Asc97a, ADG^{+}11, AKT^{+}22, BWR^{+}23, Bir86, BT14, BM00, BLS16, Bur92, CHW09a, CMLC16, DKK11, DAP11, FR04, GHM95, GMS97a, HH99, HHSW12, Ham15, HBW22, HMMhL92, HCC23, Hua15, KBK^{+}21, Kli87, LLLV16, hLRT87a, MPTB22, MUWB23, Mil88, MMH09a, PA09, PH99, PS03, RLK23, RCLL17, SD97, SPZ19, SDS12, Smi89, TKMA17, Tho12, Tok18, TFB18, WA92, WMB95, XM11, ZKD^{+}14]. Demands [PTE04]. Demographic [SD97]. Demon [SHK^{+}13]. Denitrification [DLL17]. Denmark [AR10, AAF10, AYNN19]. Dennis [Van96]. Density [Han25, Hea95]. Dependency [WB09]. Dependent [Han88, Hea95, KW86, MS98, MMSS12]. Depends [San05]. Depensation [PSK15]. Derby [HKBH05]. Deriving [Sch98]. Description [GWB94a]. Descriptive [Arn93]. Deseasonalized [GA95]. Design [LSW09, MC03]. Designations [HKS04]. Designing [San04]. Detailed [And99]. Determinants [SAA^{+}20, YAA20]. Determination [PTDQ^{+}19]. Determines [JRBH06]. Determining [WB09]. Deterministic [KPS04]. Deterrence [PTMQ22]. Developed [Neh93]. Developing [ALP^{+}21, Arn23, CCBM85, GST22, HNAY^{+}21, Kin93, VOJ^{+}01]. Development [All98, ABS03, Chu86, CLSA22, CMM88, Com93, Cum87b, Her88, HBAY^{+}16, Pol86, Sco88, SB87, TgKtC99, VM86, Yam95, Pan87]. Developments [And85b, Gat86]. Device [MS11, War94]. Dewar [SMRK85]. Diana [Loo01]. Dicentrarchus [RS96]. Diets [ILRS18]. Differences [FP25, KBK^{+}21]. Different [AG01, LSHQ06, NL09, hS98]. Differentiated [Hua15, NK18]. Differing [FM96, HH86]. Dimensions [All98]. Direct [HVTY05, Hol99]. Directions [LWL^{+}20, Yoh03]. Disaggregated [AKT^{+}22]. Disaster [WM16]. Discard [Ves96]. Discarding [Arn94]. Discards [And94c]. Discount [Coo88, Han86, PB84, Row04]. Discounting [BS06a, SS97a]. Discrete [AK20, And86, BHK97, BCSA22, BM00, Smi85]. Discrete-Choice [BM00]. Discrete-Time [And86]. Discriminated [VSSH05]. Discussion [Dol88, Hol04]. Disease [AHTT09, CR15, PDS^{+}21]. Diseases [HO11]. Dispersion [RL14]. Displacement [Par87]. Disposal [Con85, Jin94]. Dispute [EQJ94]. Disputes [Abb15, Asc97b]. Dissipated [KH06]. Dissipation [Hal04, KNX95, VW13, WW93]. Distance [Fou02, WH99]. Distribution [And02b, GMM^{+}15, GSKA20, WJ12]. Distributional [GST22, GC16]. Diversity [BWR^{+}23]. Diving [HM11, MMH09a]. Do [AL18, BKS^{+}23, DCB15, LL13, LGOS21, NV12, Vis02, WM16]. Doeringer [Geo87, Geo96]. Does [Álv22, And07, AYNN19, LA06, PSKS13, Roh07, Seu24, Sjö15, UUL^{+}10]. Dog [Cox07]. Dolphin [HBCP09]. Domestic [SVS99, ZBP06a]. Dominated [KJE03]. Dominica [TH12]. Douglas [Yoh03]. Doulman [Her88, Wau88]. Downstream [GAW^{+}18]. Dr [Ano09g]. Dr. [Ano06c, Ano07c, Ano08c, HBB^{+}07b]. Drain [Han10]. Drawback [LG98]. Driven [LP01]. Drivers [CMGB07a, LSJN^{+}24, Mer23]. Drives [CDNS17]. Driving [ADK^{+}11]. Dual [HF07, Jen02]. Due [Hut14, Par87, WHL^{+}18]. Dugong [Tis86]. Dungeness [MJ20]. Dunn [And87a]. Duration [WTW^{+}19]. During [LCB92, And15c]. Dutch [Smi96, VD05]. Duties [KP08]. Dwiponggo [Com88]. Dynamic [AS94, Arn13, Asc97a, Dow93, FM96, GSJ13, HCC23, Koe84, NAB14, SF19, Sum97]. Dynamics [ASS18, And92, BC87, Han88, Kro04, Sam92, SPC18, SYN^{+}19, aMUSS02, Wil18a, Wil18b, YH96, SABO14].

E. [Kna95]. Early [Roh03b, Cha09]. Earnings [GAW^{+}18, LHCL15]. Easements [WT05]. East [BER^{+}14, Sha11, BB06, Bos04, Joh96b, Ken03]. Eastern [MKN23]. Eat [FPA13]. Eburon [Tow03]. Eco [LGOS21, Roh03a]. Eco-label [LGOS21]. Eco-labelling [Roh03a]. Ecofundamentalism [GHH^{+}15]. Ecolabeled [AYAB^{+}20, WAUR17]. Ecolabeling [GW00, Roh03b]. Ecological [Gri00, KW09]. Ecology [Geo06, Ken06, McC02]. Econometric [Ano93b, Bir86, Gom92, Her93, HC06, SRTS86, WK88]. Economic [AAO^{+}09, ATR^{+}23, AMK06, And94b, And15a, Ano95b, AKKM86, BG00, Bos13, CL98, Chu86, CMGB07a, CHLH03, Cum87b, Dia96, Dol88, DS85b, ES16, Esp15, Gar93, GN88, Geo96, GBL^{+}25, GOFR88, HGV^{+}99, HdSK15, HBCP09, cHPZ07, HS86, Kah87, KG98, LKK03a, LLLV16, LSH^{+}18, LP01, LS14, LL94, MPTB22, MFF^{+}23, MF22, NDW22, NFW12, PFN13, PRS92, PSL03, RG86, RRS88, ST87, SAU12, SPC18, SW06, SCK^{+}15, Sha11, STY20, Smi96, SV16, TH12, TBA95, Whi93, WJP^{+}02, WL94, vBDSD10]. Economically [HHP10]. Economics [Abb15, And87a, And13, And87b, And94c, Ano93a, Ano19a, Ano20e, Ano20f, Ano20g, Arm13, AT02, BS13, Bos13, Cha95, CMC99a, CER21, CM17, Cum87a, DPP^{+}22, Egg06, Gri00, HCT96, HDK^{+}11, Her95, HSCH04, KC95, Ken06, Kna12, KT94, PWC22, PB84, Sch07, Sco11, SB87, Smi17, Smi23, SV16, SCS06, SRA^{+}23, Swi85, Tyr01, Vis02, WGM86, Wan92, Wau87, Wes02, Poo94, All95, Bjø93, Com89, Pan87, ADvH^{+}18, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano00b, Ano00a, Ano01b, Ano01a, Ano02b, Ano02a, Ano03b, Ano03a, Ano04b, Ano04a, Ano05b, Ano05a, Ano06b, Ano06a, Ano07b, Ano07a, Ano08b, Ano08a, Ano09g, Ano09b, Ano09a, Egg06, HBB^{+}07b, KFS^{+}15, Smi12, Sut11]. Economies [Jin12, Sal89]. Economists [And02a, And10]. Economy [Bro85, Geo87, McC02, MOOT23, PT97, SRS^{+}13, ZAS^{+}24, Gre99]. Ecosystem [BS06a, CKMT15, DLL17, LW20, PWC22, RHH14, Sai88]. Ecosystem-Based [CKMT15]. Ecosystems [Kah87, MD17]. Ecuador [PB94, PPF88]. Edited [ACH^{+}12, GPC^{+}13]. Editor [ARS14, Mar86b, Smi12]. Editorial [Abb17, And12, ARS14, Ano93a, Cum87a, Jar24, Smi17, Sut84, Sut92, Sut95, Wes98, SS97b]. EEC [Fro84]. Eel [Sta15]. Effect [Cla85, DKK04, EVH05, ESC18, Han86, Han07a, JL99, SAR05, ZWS21]. Effective [EAÖH25, GSJ13, San04, TB97]. Effects [AGYW10a, And99, CDTW95, CN00, CP16, Chu86, Coo88, CSW23, DCB15, DF16, Fel02, FLS14, GC16, GMM^{+}15, HH86, Han85a, HO11, HR99, KP01, Kna97, Lee14, Loo88, NSG12, SD97, SS97a, SAU12, Sch95, Sch97, Sta87, WJP^{+}02, WCFZS03, XKM08]. Efficiencies [HT06]. Efficiency [AGR19, And88, AC97, Boy00, CD07, Dan00a, Dan00b, ESC18, Fel02, GHM06, GMTW19, GR11, Hel22, KBBN23, KNYA23, Mai19, OA10, SL98, Wal06]. Efficient [Ken03, LHT96]. Efficiently [Hal16]. Effort [CL89, CP98, EG13, MB10, ST87, ST89, SYN^{+}19, Smi96, TBA95, YH96]. Egoistic [SS01]. Eilperin [SHK^{+}13]. Elasticity [EG13, Tok18]. Eliciting [CKMT15, DBR^{+}19]. Eliminating [TP05]. Elinor [GLM^{+}25]. Elmer [Ret88]. Emerging [AR10, All98, Bro85, DS87]. Emission [NAB14]. Emphasis [HGV^{+}99, Neh93]. Empirical [AYJO^{+}19a, Bos04, Esp15, Fre95, Hut14, JLT19, Kou95, LSW09, PTE04, RAH17, Sal89, SW04, WG03]. Employment [AGYW10a, AK20, SRTS86]. Endangered [DvHM18, LLR10, Lew19, Tom10]. Endogenous [BSA^{+}22, SW04]. Energy [FR92, GLN^{+}23, MOJG04]. Enforcement [And89a, Arn13, LN16]. England [AB93, CG85, DG94, EA89, Geo87, Hol04, HKPW13, HdSK15, HRSW18, JHT06, KTR00, PPV96, SMRK85, WR97]. English [ZSG23]. Enhance [UUL^{+}10]. Enhancement [BHJ03, BHBG93]. Enhancements [LLR10]. Enhancing [GRA^{+}25]. Enrichment [KBS01, KBS02, KB05]. Enterococcus [KGW22]. Enterprise [Gar88]. Entry [FHS95, GDF04, LSBH21, MGC89, PTMQ22, Wil88]. Envelopment [GD05, VD05, Wal06, WA23]. Environment [BP94, Ret94, Sis84, ZWS21, vBDSD10, Ale87]. Environmental [AK20, AP17, AYJO^{+}19a, Bos13, CP16, Cha00, Chu86, DKK04, FJ23, GPC^{+}13, GRA^{+}25, JKP95, JLMR96, LA06, LZ87, Opa87, Roh09, Tve02b, Vis01, Wan92]. Environmentalism [GHH^{+}15]. Equations [Asc97a]. Equilibria [HRK86, LK00]. Equilibrium [And94a, GST22, Mai10, PS89, PS94, RHH10, Seu24, WS10]. Equity [AC97]. Equivalency [MPTB22]. Erosion [cHPZ07, LKK03a]. Errata [AGYW^{+}10b, Ano89a, Ano95b, KBS02, YL06a]. Erratum [AYJO^{+}19b, Ano93b]. Error [CL09, SLW85]. Essential [HKS04]. Estimate [Boy93, Gil88]. Estimates [BLEV05, FR92, JL99, PgK09]. Estimating [BHK97, CS86, DTS20, EC89, FKAvR10, GAB^{+}86a, HTTA92, SB85, Tok18, Wal06, WT10, WHL^{+}18]. Estimation [Bir86, KPS04, SZC07, WDJK20]. Estonia [EVH05]. EU [ADG^{+}11, JV01, Lin05b]. Europe [AYSN17, Tho12, Ves05]. European [CL98, GIP02, Gar93, GH96, Nie05, NSG12, VSSH05]. Eutrophicated [EAÖH25]. Eutrophied [GSJ13]. Evaluating [STY20]. Evaluation [GWB94a, KSPT^{+}17, LKK03a, Meu95, Roh09, Sha11, Wan95, YC15]. Event [MJ20, MWH^{+}18]. Evergreen [TY05]. Evidence [AW88, BAYN16, Cav23, CLSA22, DvHM18, Edw92, Esp15, Fre95, HD25, HRSW18, JA89, LPV21, Li98b, LSW09, McT99, MUK23, RCLL17, SS16, TPC03, ZS09, BKS^{+}23, LGOS21, Ola09]. Evoked [KNH93]. Evolution [HdSK15, Rus00, WL94]. Evolutionary [Gat96]. Ex [Fel09, GCG98, PTDQ^{+}19, PA20]. Ex-Vessel [GCG98, PTDQ^{+}19, PA20, Fel09]. Examination [EW86, UUL^{+}10]. Examine [AY12]. Examining [HGJH15]. Example [TL12]. Examples [Sch07]. Excess [RJY16, TP05, WMT04]. Exchange [XKM08, ZKD^{+}14]. Excluder [MS11]. Exclusion [WJP^{+}02]. Excursions [HBCP09]. Exit [IO99]. Exogenous [BSA^{+}22]. Expansion [Abr85, SS16]. Expectations [CM04, Dup93, RJY16]. Expected [Cla85, Sch98]. Expenditure [TYH15]. Expenditures [KM02]. Experience [Tow03]. Experiences [Pol86]. Experiment [DBR^{+}19, Lew19, LAM21, NSV^{+}22, RCLL17, WAUR17, ZWS21]. Experimental [FW10, HO11]. Experiments [AK20, GLL^{+}10, HLM20, LPV21, LW20, LWL^{+}20]. Explicit [CPW03]. Exploitation [Han86, Ken03, Maz04]. Exploiting [SS01]. Exploration [FR92, GAW^{+}18]. Exploratory [OG03]. Exploring [Cha00]. Export [Jin96, Str14, XKM08]. Exports [LA11, SAA^{+}20]. Extended [And02b, CG85, Mun86, VM86]. Extension [MCFH09]. Extensive [FJ03]. Externalities [Boy00, CPW03, Dan00a, Dan00b, FR92, JLMR96, KL07, RHH14, Seg87, WJN^{+}16]. Externality [BS06b]. Extreme [GHH^{+}15]. Exvessel [hLRT87a].

F [Com88, Her95, Sch07]. F. [Gri93]. Facing [Kri09]. Factor [Gut02, Sal93, Seu24]. Factors [GBDT08, HNE^{+}89, NGBM95, PPF88, QXH11]. Failure [Hal16, ZT22]. Falkland [Mai19]. Familiar [PMT99a]. Far [Joh96b]. Farm [NQ06, GBDT08]. Farm-raised [NQ06]. Farmed [AAW^{+}17, Asc97b, AG01, ADK^{+}11, Bjø88, BLS16, BAYN16, Hea93, Hea95, NSL^{+}07, NL09, RAC^{+}17, RNFH22, SAR05, XKM08]. Farmer [AR10].

Farmers [LPG13, MUWB23, O’H20, SJQ21]. Farming [Asc08, GRA^{+}25, MS01, MN18, Øst99, Sal89, VH11, vBDSD10]. Farms [GBDT08, OA10]. Faroese [DAAY^{+}20]. Faustmann [Gut08]. Favorite [PMT99a]. Feasible [O’H20]. Federal [SMRK85, Ter07, DS87, WJ12]. Fee [LRCL25]. Fee-Based [LRCL25]. Feeding [Arn92, Hea93, MS98]. Fees [Mai19]. Fencing [Hol00]. Field [LPV21, LAM21]. Fight [SGC^{+}16]. Fighting [Abb15]. Fillet [Qua06]. Fin [CMGB07a]. Financial [ZT22]. Financing [SB87]. Finfish [YC15]. Finite [GMTW19]. Finlayson [Poo97]. Finley [Kva18]. Finnish [VSSH05]. Firm [LS99, LGOS21]. Firm-Level [LS99]. Firms [OG03]. First [ASS^{+}14, Lib07, Nie05]. First-hand [Nie05]. First-Possession [Lib07]. Fish [Álv22, And96, AYJO^{+}19a, Ano89a, AG01, AH02, AKT^{+}22, BC84, Bjø88, BDCDeS00, BAYN16, BHH^{+}12, BA20, Bur92, CHN12, DO14, DS85a, DAP11, EOA^{+}17, FPdR20, FR04, GH86, GSA93, GHB93, GH96, Gud98, GLN^{+}23, HKBH05, HRA^{+}23, Han85a, Han13, HRA^{+}17, Hea93, Hea95, HVTY05, HKS04, HG89, Hup89, ILRS18, JHT06, JA89, JRBH06, KJE03, Ken92, KNH93, Kin03, LSHQ06, MM04, MS98, MN18, MUK23, Mun00, Mun07, MF02, NP97, NGBM95, NSG12, OA10, OZAo96, OAA04, PTE04, PTSB99, PY89, Res95, RAC^{+}17, RGA07, Sai97, Sal89, SD97, Sjö15, SHK^{+}13, Tah09, TYH15, Tok18, Tow06, WMT04, WJ12, XM11, YM02, MSBh95, Tow09]. Fisher [BCSA22, CLSA22, Dup93, GBL^{+}25, MEA^{+}23]. Fisherie [BP97, PR99]. Fisheries [AGYW10a, Abd95, AK97, AKP98a, AAO^{+}09, ATR^{+}23, AAC^{+}21, And05, And13, And84, And85b, And86, And87b, And87c, And88, And89a, And94b, And02a, And10, ACH^{+}12, And15a, ABW21, AS89, AS94, Ano95b, AC97, AKV^{+}17, Arn94, Arn07a, Arn07b, Arn13, Arn23, Asc07, ARA^{+}25, BC84, Ber23, Ber06, BCHH00, BKS^{+}23, Bjø93, BB06, BS13, BER^{+}14, Boy93, BDCDeS00, BBI05a, CGP03a, CM85, Cha09, Chr96, CM17, CPW03, Cox07, Cro16, CN06, CP92, Cun94, CNAB09, DAAY^{+}20, DS85a, DK96, Dow93, DCB15, Egg06, Egg12, ESOF03, ES16, EAÖH25, EA89, Esp15, Fel02, FLS14, FHS95, FHM17, FR04, Fro84, FGM^{+}98, Gar93, Gat86, Geo96, GST22, Gon96, GW92, GHM06, GM07, GLM^{+}25, GW00, GMM^{+}15, GSKA20, GR11, HKBH05, HDK^{+}11, Han84]. Fisheries [Han15, Hat05, HGV^{+}99, HKS04, HJKP03, HB96, Hol00, HSCH04, HH09, HRSW18, IO99, JV01, Jen02, JKPHS12, Jin12, Joh96b, Joh85, KM93, KG97, KBBN23, Kna12, KBS01, KBS02, Koe84, Kva16, LHT96, LG07, LSW17, LP01, Lib07, LLS25, LF11, Mai19, MP85, MP99, MMSS12, MSI01, Mat10, Meu95, Mik07, Mun07, Neh93, NF10, NFW12, NLH12, NTS11, OGL^{+}16, OB84, PC94, Pan87, PTE04, Pas07, PCAIGO^{+}14, PWL05a, PPV22, PFN13, PFNV21, PD00, PB84, PSKS13, RAH17, RAW17, Roh03a, RCLL17, Rus99, Rus00, RHH14, Sch95, Sch07, SW06, SCK^{+}15, Smi97, SABO14, Ste98, Sto87, Sum97, hS98, SRA^{+}23, Swi85, SC95, TL12, TB01, TFB18, Tow03, Tro01, UUL^{+}10, VW13, VOJ^{+}01, WJN^{+}16, WS10, Wau87, Wau94, WJP^{+}02, Wil04, Wil09, WCA^{+}10, Wil13, WL94]. Fisheries [WG03, YH96, Yoh84, ZBP06a, ZS16, And87a, ALS^{+}18, Arm13, Com88, Her95, Rob93, SMRK85].

Fishermen [AMK06, AA00, Cla85, Cun94, CM04, EL07, Ham98, Han85a, IO99, Kna97, PTDQ^{+}19, PPF88, SJQ21, Str04]. Fishers [MEA^{+}23, PPV96, TH12]. Fishery [Abr85, And85a, And87c, And94a, And04, And07, AS89, AL18, Arm94, AMA00, AA00, BS06b, BC87, BG00, BBA^{+}10, BHW15, Bos04, CL89, CDTW95, CN00, CP16, CS13, CH25, tCT93, CV93, CCBM85, CYP92, Com89, Com93, Coo88, CD07, CG85, DTS20, Dia96, dLS92, DS85b, DF16, EW86, Eas92, Edw01, Edw05, EU99, ESC18, Fou02, FGMW03, GIP02, GWB94a, GN88, GOG00, Gom92, Gra96, GHM06, Gri93, Gul89, GMTW19, GR11, GDF04, HH86, HRK86, Han85b, HC06, Hol04, HKPW13, HHP10, HCC23, HRSW18, mHS06, HW12, Jin96, JA89, Joh95, Joh96a, KW09, KW86, KH06, KPS04, KB05, KNYA23, KHB^{+}25, Kro04, LLA00, LHCL15, LS21, LSW09, Lin05a, LCB92, LK00, LSBH21, MB10, MSI01, Mat09, MGC89, MC92, Mer23, Mun86, NP97, PTDV17]. Fishery [PCAIGO^{+}14, PRS92, PS89, PS94, Pon88, Pon89, Pow04, PSL03, Res95, Ret88, RW00, RG86, RC89, Roh07, Rud88, RJY16, RHH10, SLW85, SH04, ST87, SRTS86, SRS^{+}13, Sco88, Sco93, Sco11, Sei93, Seu24, SL98, SYN^{+}19, Sis84, Smi97, SP85, SW04, Spu85, Sta15, Sta87, Sto86b, Sto86a, Sum95, SCS06, Swa94, SE94, Ter07, TBA95, THM98, TPC03, Ves96, VOJ^{+}01, Wak14, WR97, Wat01, WJ12, WH99, WDLH11, Wil85, Yah88, Yam95, YC15, YB96, Pol86]. Fishing [AY12, AH93, AS05, AC97, BWR^{+}23, BHK97, BHW15, Boy93, BM00, CHW09a, Cav23, CMC99a, CSW23, Dol88, EW86, EG13, Fel94, FHJ87, GIP02, Gat84, Geo87, GMS97a, HHSW12, Han93, Hol99, HT06, HM95, HR99, HBB07a, Hut14, JBH24, JLT19, JRBH06, Ken87, KTR00, LS14, LSJN^{+}24, LSW09, Loo88, LGOS21, MC99, Mat96, Ola09, PA09, PT97, PTSB99, PTDV17, PPV22, PAC^{+}22, RL14, Roy98, Rus00, Sam92, SW18, Sch95, Smi96, SDQ21, hS98, TP85, TP05, aMUSS02, Ves05, VOJ^{+}01, WT10, WA23, WDJK20, WCH01, ZWS21, GBL^{+}25, MSBh95, Poo94, Poo97, DPP^{+}22]. Fishnet [Cav23]. Five [ASS^{+}14]. Fixed [And88, Han88]. Flatfish [Lin05a]. Fleet [CGS05, CSW23, CG85, DTS20, EVH05, GD05, HF07, HS86, LSJN^{+}24, LSW09, Lin05b, Sam92, Smi96, hS98, VD05, WB86]. Fleets [CN95, FHJ87, Mat96]. Flexibility [Maz04]. Flexible [Qua06]. Flood [FP25]. Florida [BC13, BMS89b, GMS97a, RLA10, TS02, WHL^{+}18]. Flower [SS16]. Fluctuating [BC84]. Fluke [SAU12]. Flying [Res95]. Following [MWH^{+}18]. Fond [BA20]. Food [ALP^{+}21, OHS14, TFB18, BHH^{+}12, Kri09]. Foodborne [MUWB23]. Foods [Smi23]. Forecast [Ano93b, Her93]. Forecasting [Ali11, Gor20, GA95, MOJG04, dFdOdA24]. Forecasts [MFF^{+}23]. Foreign [BC84]. Foreman [Vis01]. Forest [Loo88, Swa94]. Formation [And02c, Dup93, Edw01, FLS14, Nie05]. Formats [CDNS17]. Forms [AH02]. Forum [And10, SLA14]. Forward [AK97, SF19]. Four [BHH^{+}12]. Fowl [YM02]. Framed [LAM21]. Framework [CN06, DS87, Par87, Sch98]. Framing [GLL^{+}10]. France [AGH04, BA20, KM02, LGL19, MW02]. Francis [Pon89]. Frank [Kri09, Loo01]. Fredrik [Gri00]. French [ADG^{+}11, GDF04, GD05, Jaf04, PGB00, RAC^{+}17, XM11]. Frequency [OHS14, KJE03]. Fresh [AB93, BHG^{+}16, DG94, FPdR20, GH86, GH96, HVTY05, NGBM95, OZAo96, OAA04]. Freshness [Lee14, STNY23]. Freshwater [MC99, NAB14]. Friedman [Vis02]. Friendly [ILRS18]. Front [Ano84a, Ano84b, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89b, Ano89c, Ano89d, Ano89e, Ano92a, Ano92b, Ano92c, Ano92d, Ano93c, Ano93d, Ano93e, Ano93f, Ano94a, Ano94b, Ano94c, Ano94d, Ano95c, Ano95d, Ano95e, Ano95f, Ano96c, Ano96d, Ano96e, Ano96f, Ano97c, Ano97d, Ano97e, Ano97f, Ano98c, Ano98d, Ano98e, Ano98f, Ano99c, Ano99d, Ano99e, Ano99f, Ano00c, Ano00d, Ano00e, Ano00f, Ano01c, Ano01d, Ano01e, Ano01f, Ano02c, Ano02d, Ano02e, Ano02f, Ano03c, Ano03d, Ano03e, Ano03f, Ano04c, Ano04d, Ano04e, Ano04f, Ano05c, Ano05d, Ano05e, Ano05f, Ano06d, Ano06e, Ano06f, Ano06g, Ano07d, Ano07e, Ano07f, Ano07g]. Front [Ano08d, Ano09c, Ano09d, Ano09e, Ano09f, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano24, Ano25a, Ano25b, Ano25c, Ano25d]. Frontier [SL98, WT10]. Frontiers [Mai19]. Frozen [AB93, GH96, HVTY05, RGA07, SDS12]. Fuel [CMLC16, HW12]. Fukushima [AW18, WM16]. Fully [CD07]. Function [Hof04, Li98a, Sal93]. Functions [ESOF03, Fou02, NB03, RLK22, TP85]. Fundy [DG00]. Futility [Kin03]. Future [And02a, BHH^{+}12, CAAT10, CN06, Cur85, LWL^{+}20, Yoh03, Meu95]. Futures [LLR10].

G [ALS^{+}18, Com89, Geo87, Geo96, Wau87]. G. [Bjø93, Tow09]. Gains [AAF10, Asc97b, CD07, Han10, YGB97]. Game [AMA00, FM96, Gat96, mHS06, Ken87, Li98a, LK00, NP97, PY89, Sum95]. Games [DBP00, KL07]. GARCH [Ali11]. Garden [SS16]. Gas [Cur85, MOJG04, RRS88]. Gates [Roh11, Sut11]. Gatornomics [HF01]. Gear [DG00, Lee14, PRS92]. General [GST22, Seu24, WS10]. Generating [SC95]. Generation [LSW09, LCB92]. Generic [BSL92, KP01, KM02]. Geoffrey [And85b, Gat86]. Geographic [WJ12]. George [Mar86a]. Georgia [HHP10, STY20]. Gerard [Cur85]. Gerhardsen [Bjø93]. German [AYJO^{+}19a]. Germany [BLS16, BAYN16, NSL^{+}07]. Gibraltar [CGS05]. gigas [MVRRP21]. Ginsburg [Kna95]. Giulio [Ale87, Gul89]. Giving [Wil07]. Global [AAC^{+}14, ALS^{+}18, mHS06, KBK^{+}21, Kri09, Kva18, CER21]. Globalisation [TB01]. Goal [And04, DS85b]. Good [GLN^{+}23, Han10]. Goodlad [HRA^{+}23]. Goods [GLM^{+}25, LW20, SAA^{+}20]. Gopalakrishnan [Bro85]. Gordon [Arm13]. Gough [Cha09]. Governance [CGP03a, CN06]. Government [And96, Chu86, Han85b, Kva18, Mat97, Sco93, SMRK85]. Grafton [ACH^{+}12]. Grant [And87c]. Grasses [GGKW17]. Gray [LL94].

Great [Bur92, FR04]. Greece [Fou02]. Greenberg [BHH^{+}12, Kri19, SGC^{+}16]. Greenland [CV93, Ves96]. Gren [Gri00]. Grey [mHS06]. Grocery [HDR96]. Gross [ZBP06a]. Groundfish [DR04, DG94, Gar88, GR11, HG89, Hol00, Hol13, HKPW13, HBAY^{+}16, HRSW18, KTR00, LSW17, LSW09, Mat10, WR97, WJ12, Wil09, WCA^{+}10]. Groundfish/Whiting [Mat10, Wil09, WCA^{+}10]. Groundfishery [DG00, SRTS86]. Grounds [Sei93]. Grouper [RLA10]. Groups [Roh09, WCH01]. Grow [CT14]. Growing [ZWS21]. Growth [And88, AYNN19, ART07, ADG^{+}11, BT14, Hea95, KBK^{+}21, Tve02b]. Guangdong [Chu86]. Guest [Cum87a, Mar86b, SS97b]. Guide [SP85]. Guides [Roh09]. Guinea [Her95]. Gulf [ASS^{+}14, GCG98, GOG00, Gri93, PTE04, SABO14, SS16, Str04, WT05]. Gulland [Pon89].

H [Pol86, Vis01]. H. [Com89, Her95]. Haab [GPC^{+}13]. Habitat [AKV^{+}17, HKS04, HLM20, Mil88, Tom10]. Haji [And96]. Hake [Álv22, Jaf04, THM98]. Halibut [AS89, CDTW95, CK94, CC87, Coo88, HC06, Kna96, Kna97, hLRT87a, MC03, Sto86a]. Hallwood [KFS^{+}15]. hand [Nie05]. Handbook [ACH^{+}12]. Hannesson [Sch07, Wau94, And13, BS13, DPP^{+}22, GHH^{+}15]. Happened [YC15]. Harbor [DR04]. Harbors [ABW21]. harengus [KJE03, San05]. Harmful [MFF^{+}23]. Harmonizing [Lin13]. Harold [Gri93]. Harrison [ARS14]. Harvest [CL10, CMLC16, CLL22, ESOF03, Han93, HJKP03, HKPW13, HdSK15, LL13, LS14, MS98, MS11, NC07, San05, Sch01, SP85]. Harvested [PWC22]. Harvesting [And89b, AP17, Arn92, Bjø88, BG00, DS85a, JHT06, KJE03, KW09, Li98b, Li09, Loo88, PY89, QT19, RW00, Ste98, Tah09, WMT04, YL06a, YL06b, Sar09]. Harvests [Eas92, Ken92]. Hassenzahl [Cha00]. Hatchery [AL13]. Hatteras [DvHM18, LLLV16]. Hawaii [CN00, CP16, HCT96, Ham98, HBCP09, PHCK16, Poo94, PSL03, SL98]. Hawaiian [CYP92, TPC03]. Haystack [TB97]. Hazard [BP97, Yoh03]. Hazardous [Cum87b]. Hazards [GRA^{+}25]. Health [Cha00, Cum87b, LSJN^{+}24]. Healthier [Kri19]. Healthy [ILRS18]. Hedonic [CL10, Ham15, KGW22, LS99, Lee14, PA20, RLK22, RGA07, STNY23, Álv22]. Help [And07, PFNV21, Roh07]. Herbicide [GKK^{+}88]. Herrick [Sch07]. Herring [Abb15, AMA00, BC87, BG00, BaUS04, BER^{+}14, EQJ94, KJE03, LK00, NB03, San05]. Hersoug [Tow03]. Heterogeneity [AYAB^{+}20, BM00, DTS20, HHSW12, OHS14]. Heterogeneous [CH25, FHJ87, GR11, WAUR17]. Heteroskedastic [MC99]. Hidden [SHK^{+}13]. High [CN06, KM93, Li98a, Lin13, Mun07, PTSB99, PD00]. High-Seas [Li98a]. Highgrading [And94b, Ano95b, Ves96]. Highly [NP97]. Hilborn [ACH^{+}12]. Hiring [MG16]. History [HH99, Mun00, Geo06, WBC^{+}13]. Hit [HO11]. Hoar [Gri93]. Holden [Meu95]. Home [TYH15, ZFG20]. Hong [Ano06c, Ano07c, Ano08c, Ano09g, HBB^{+}07b]. Honor [And13]. Hooked [Tow09]. Hoonaard [Fel94]. Horizon [WHL^{+}18]. Household [SD97, TYH15]. Households [LL94]. Housing [FP25, LA06, NDW22]. Howard [All95]. Huang [GPC^{+}13]. Hubbert [Sla07]. Hurt [PFNV21]. Hybrid [And86]. Hydroelectric [Boo89]. Hypothesis [Bel89]. Hypoxia [HNCS12, SABO14, WT05].

Ian [Joh85]. Iceland [GSKA20]. Icelandic [Arn93, DSBT97, Fel94, Mat97, Rus00]. Ideal [Tho12]. Identification [aMUSS02]. Identifying [Cav23]. IFQ [ASS^{+}14, HBAY^{+}16, Kna97, MC03]. IFQs [AGYW10a, AL18, Fel09, Kna96]. II [Tyr01]. Ill [MU16]. Ill-Defined [MU16]. Illegal [CMC99a, HW12]. Illness [MUWB23]. Illness-Related [MUWB23]. Impact [AK20, AYJO^{+}19a, AL18, BC13, CL98, tCT93, CHLH03, CG85, Dia96, Gre99, GDF04, Hee89, HRSW18, mHS06, KBBN23, LLLV16, LG98, MC92, MWH^{+}18, PGB00, PAC^{+}22, SPC18, Wak14]. Impacts [AAW^{+}17, Bea25, BHG^{+}16, DTS20, FJ23, LP01, LS14, LLS25, MJ20, MC03, MUK23, MF22, Roh03b, SCK^{+}15, WS10]. Impending [Sla07]. Imperfect [XZ14]. Implementation [CMM88, Mun00, Smi89]. Implications [AW88, CAAT10, CM17, FR04, GKK^{+}88, Gra96, GW00, HH09, Joh95, Joh96a, LS99, Lib07, MPTB22, MSI01, PDS^{+}21, RLK23, Roh09, Ste98, TP05, TFB18, War94]. Implicit [EG13, GGKW17, SVS99]. Import [HCC23, XZ14, ZKD^{+}14]. Importance [Han88]. Imported [HGJH15]. Imports [HG89, MJ11]. Improved [Lip04, MFF^{+}23]. Improvements [BMS89a, Fre95, HTTA92, Sch98]. Improving [Arn23, MVRRP21]. Imputation [LHCL15]. Imputing [FGM^{+}98]. In-Season [BS06b]. Incentives [CH25, MD17, RAH17]. Incident [GAB^{+}86a]. Incidents [GOFR88]. Including [AK20]. Inclusion [And05]. Income [Geo87, KBK^{+}21]. Incomes [Cun94, GBL^{+}25]. Incorporating [CM04, DS85b, HLM20]. Increase [BKS^{+}23]. Increases [NV12]. Increasing [AED11, Hut14, LL94, MMSS12]. Index [EG13, GBDT08, Ano87e, VH11]. Indications [Roh03b]. Indicators [Ali11, TL12]. Indirect [Hol99]. Individual [And89b, Arn93, BP06, CUW07a, CDTW95, DG00, GWB94a, GN88, Gra96, HKBH05, Hut14, JLT19, Joh95, Joh96a, LSW09, Mat09, Mer23, Sei93, TPPT11, ZS16]. Indonesia [MEA^{+}23, Com93, Sai97]. Indonesian [Com88, Res95]. Induce [GLL^{+}10]. Induced [PSK15]. Industrial [Tve02a]. Industries [Kri09, LN16]. Industry [AR10, And94a, And96, Bro86, HKBH05, Her88, HM95, Jin96, KP08, KSV05, Kou95, LA11, LG98, Lin05a, LGOS21, MOJG04, MG16, Øst99, OG03, PS89, PS94, Roh09, Sal89, Sal93, ST89, SPC18, TP05, Tve02b, ZT22, All98, SMRK85, Rus00]. Inefficiency [GD05, Han85b]. Influence [DKK11, WM16, ZT22, vBDSD10]. Influencing [PPF88, QXH11]. INFOFISH [Jin96]. Inform [Smi23]. Information [Abd95, AYJO^{+}19a, BHG^{+}16, CM04, DOY21, FR92, FJ03, LWL^{+}20, MP85, MWH^{+}18, PPQ^{+}19, SP85, SC95, Wes02, Wil07]. Infrastructure [NDW22]. Ing [Gri00]. Ing-Marie [Gri00]. Initial [Kna97].

Injury [DS87, EC89]. Inland [Bea25, DAP11, HNAY^{+}21, STY20]. Innovation [GRA^{+}25]. Innovations [GMT11, WA92]. Input [GHM06, Gut02, Hof04, Øst99]. Input-Controlled [GHM06]. Inshore [Fou02, Rud88]. Insights [CPW03, PCAIGO^{+}14, WJ12]. Institutional [LG07]. Institutions [RAW17]. Instrument [And88]. Instrumental [HBW22]. Insurance [CSW23, FP25, JKP95, LSJN^{+}24, NNNJ22]. Integrated [Abd95, CL09, DBR^{+}19, GOFR88, NSG12]. Integration [AYSN17, AGH04, ABOS12, BAYN16, HNAY^{+}21, Nie05, NSL^{+}07, SPZ19]. Interaction [Pow04, Sum97]. Interactions [And85a, Mik07, YB96]. Interdependencies [TBA95]. Interests [Abr85]. Interfaces [Chu86]. International [Ano93b, GHM95, Hal16, Her93, Lin13, OGL^{+}16, ZS09, Neh93]. Internationally [Mun07]. Interrelationships [PSL03]. Interseasonal [And89b]. Intra [And89b, GBDT08]. Intra- [And89b]. Intra-farm [GBDT08]. Intracultural [PPV96]. Introducing [Ano19a, Ano20e, Ano20f, Ano20g, Kat07]. Introduction [And10, Arn07a, AT02, AGT08, AG14, Bjø00, GMT11, HGH99, HGV^{+}99, HSCH04, LH12, LW20, Mar86b, NA88, Opa87, SLA14, Smi12, SCS06, Ves05, And87a]. Introductory [Wil18a]. Intuitive [Yoh84]. Invader [KB05]. Invariance [RAW17]. Invasive [FPA13, SHL^{+}20a]. Inventory [WW93]. Inverse [Ham15, Hua15, PTE04, Tho12]. Investigating [BM00]. Investigation [AYJO^{+}19a, Bos04]. Investment [CM85, Jen98, LN16, NTS11, PSKS13, Sum95]. Investments [Han07b, SSH07]. Investors [LGOS21]. Invisible [FHS95]. Invoicing [Str14]. Involvement [Mat97]. Involving [HJ97, Lew19]. Iran [MMCM13]. Irreversible [CM85, PSKS13, SSH07, Sum95]. ISBN [All98, Tow03]. Ishak [And96]. Island [GMTW19, KGW22, MMCM13, PKLB09, OZAo96, OAA04, SAU12]. Islands [PgK09, AGYW10a, CYP92, Com89, Her88, Mai19, SHL^{+}20a, TPC03, Wau88]. Issue [And10, AT02, AGT08, AG14, Bjø00, Cum87a, GMT11, HGH99, LH12, LW20, SLA14, Ves05]. Issues [And89a, Ano93a, Eas92, KP08, PH99, Wau88, Whi93, Yah88]. Italian [PRS92]. ITQ [And94b, Ano95b, BS06b, CS13, CMM88, CD07, Hol13, KSPT^{+}17, LCB92, Mat97, Tur96, Wan95, YC15]. ITQs [AAF10, Arn07b, Boy00, CUW07a, Dan00a, Dan00b, Han07b, Hel22, Lib07, MSI01, PTMQ22]. Ivorian [Dia96].

J [Cur85, Geo06, Kna95, Meu95, Wau88]. Jack [KSPT^{+}17, PTDV17]. James [Gre99, And15b, BP94, Rob93]. Japan [HGJH15, HCC23, STNY23, WAUR17, Yam95]. Japanese [BM96, HMMhL92, Wak14, WM16, WW93]. Japen [Neh93]. Jens [Egg12]. Jentoft [All98]. Jersey [Geo06, HD25, Yoh03]. Jobs [Geo87]. Johansen [KSV05]. John [Bos13, GPC^{+}13, HRA^{+}23, Pon89, Roh11, Sut11]. Johnston [Gat86]. Joint [Kum02]. Jolly [All95]. Jones [Tow06]. Jorgen [Gat96].

Joseph [Cha09]. Jr [Mar86a, Pon89, Sch07, Vis01]. Ju [GPC^{+}13]. Ju-Chin [GPC^{+}13]. Juliet [SHK^{+}13]. July [Ano95g, Ano96b, Ano97b, Ano98b, Ano99b]. June [Ano95g, Ano96b, Ano97b, Ano98b]. Jurisdiction [Abr85, CG85, Mun86, VM86]. Just [AC97]. Justice [Vis01].

Kammen [Cha00]. Karyn [ADvH^{+}18]. Kedah [VOJ^{+}01]. Keen [Ret88]. Kenneth [Sla07]. Kent [Mar86a]. Kerry [Gri00]. Kevin [Neh93]. King [FPA13, Mat09, SK21]. Kinship [Geo87]. Kish [MMCM13]. Knecht [Tow09]. Know [PPQ^{+}19]. Korea [TgKtC99, UUL^{+}10]. Korean [PR99]. Kuhn [PH99, PS03]. Kurlansky [Gud98]. Kuwait [Cha95].

L [Pol86]. L. [San05]. Lab [LPV21]. Lab-in-the-Field [LPV21]. Label [LGL19, LGOS21]. Labeling [DKK11, RLA10]. labelling [Roh03a]. Labels [CDNS17, FS15]. Labor [MGC89, MG16]. Laboratory [AS05]. labrax [RS96]. Lake [Bea25, BBI05a, CBI92]. Land [Chu86, Par87, SJQ21]. Landed [GCG98]. Landing [BCSA22, DR04, Egg12]. Landings [HNE^{+}89, WJ12]. Langdon [ALG^{+}19]. Large [DTS20, MMH09a]. Large-Scale [DTS20]. Largest [MN18]. Last [BHH^{+}12]. Latent [Mai19]. Latin [TB01]. Laura [Tow06]. Law [AGH04, Geo06, Hal16, Van96, Vis02, Vis02]. Lawson [Pan87]. Laying [KNYA23]. Laying-Up [KNYA23]. Lead [AYNN19, NV12]. Learn [Asc07]. Learning [GLL^{+}10, NNNJ22, dFdOdA24]. Lease [RL14]. Leasing [TY05]. Led [Kva18]. Lee [Wau87]. Legacy [ASS18]. Legal [AKKM86]. Legitimacy [DCB15]. Length [PAC^{+}22, Roy98]. Lessons [EQJ94, GAB^{+}86a, HNAY^{+}21, LSHQ06, Yoh03]. Level [CHN12, LS99, WB09, SDS12]. Levels [And88, CC87, KBK^{+}21]. Liability [JKP95, Seg87]. Lice [AAW^{+}17]. License [FW10, FGMW03, WCH01]. Licensing [Wil88]. Life [Kri19]. Lifestyle [OHS14]. Limited [FHS95, GDF04, LSBH21, MGC89, Wil88]. Limited-Entry [GDF04, LSBH21]. Limits [CLL22, HR99, LS14, WDLH11, WG03]. Linear [dLS92, Hof04, Mai10]. Linefish [HLH^{+}14]. Linkages [And96, FKAvR10, GSA93, Swa94]. Links [HVTY05]. Lion [LLR10]. Lionfish [SHL^{+}20a]. Literature [NTS11]. Litigation [KG98]. Litter [KBBN23]. Live [KGW22, QXH11]. Living [Hal16, Lin13, PWC22, Tis86]. Lobster [tCT93, CYP92, EU99, Gor20, GMTW19, HBW22, RG86, Sta87, TPC03, WK88]. Local [ALP^{+}21, ABW21, BHJ03, GST22, HNE^{+}89, Mat97, SS01, SGC^{+}16]. Locally [RLA10]. Location [Dup93, EW86, PTDV17]. Locations [ACDL21, BCSA22]. Logan [SP85]. Logistic [Qua06]. Logit [CHW09a, McT99, PA09]. Long [ART08, Gil88, GMTW19, HGJH15, Kri19, Sch97, KGW22]. Long-Run [Gil88, HGJH15, Sch97]. Longline [And87c, CN95, CN00, CP16, HCT96, PSL03, SL98, Str04, Poo94]. Looking [CT14]. Lorenz [AY12]. Loss [Hel22]. Losses [HKS04, HNCS12]. Lost [WHL^{+}18]. Lot [LA06]. Louis [Gat86]. Louisiana [DKK04, PIH14]. Low [KJE03]. Low-frequency [KJE03].

M [All95, Bjø93, Cha00, Gat86, Kna95, Roh11, Sut11]. M. [Meu95]. Machine [NNNJ22, dFdOdA24]. Mackerel [AA00, BER^{+}14, Ken92, Ken03, KSPT^{+}17, PTDV17]. Made [HRA^{+}23]. Magnuson [WR97]. Main [Bel89]. Makers [Wil07]. Making [CUW07a, Hol13, MP99]. Malaysia [Abd95, SYN^{+}19, TYH15, VOJ^{+}01, Yah88, YH96, And96]. Maldivian [CL89, ST87, ST89]. Malleability [RJY16]. Malmquist [VH11]. Mammals [Sch97]. Manage [GSE97, Hal16]. Managed [Ale87, FHS95, HD25, LCB92, Sei93]. Management [AK97, AKP98a, AR10, And13, And84, And85b, And87b, And87c, And89a, And07, And15a, AW88, Arm94, AC97, AR01, Arm13, Arn07a, Arn23, Bea25, BP97, Ber23, BBA^{+}10, BER^{+}14, CHS93, CGP03a, CM85, CKMT15, CMM88, Com89, Con85, Cox07, CP92, Cum87a, Cun94, CNAB09, DBR^{+}19, dLS92, DK96, DS85b, DBP00, ES16, EAÖH25, ESC18, FM96, FHJ87, FR04, Fro84, GW92, GSJ13, GLM^{+}25, GW00, GMM^{+}15, HH86, Han84, Han10, Hat05, Hee89, HB96, Hol99, Hol04, HH09, HHP10, HRSW18, JV01, Jin12, JLMR96, KM93, KG97, KC95, KSV05, LKK03a, LHT96, LPG13, LLA00, LG07, LSW17, LRCL25, Lin05b, Maz04, MM04, MP97, Mun86, Neh93, NF10, NFW12, NLH12, OGL^{+}16, OB84, PC94, PPV22, PD00, PD02, PSKS13, Pow04, RNKA21]. Management [RAW17, RG86, RS96, Roh07, Rud88, RHJ92a, Rus99, Rus00, SH04, SAU12, SW06, SYN^{+}19, SW15, Smi97, Smi85, Smi86, SF19, SC95, Ter07, Tom10, Tro01, UUL^{+}10, WGM86, Wan95, WR97, War94, Wat01, Wil87, Wil04, WL94, Yah88, Yam95, YH96, Yoh03, ZS16, Arn07b, Tow03, ACH^{+}12, Wau87]. Managers [LS99, PCAIGO^{+}14, Sis84]. Managing [BDCDeS00, FPA13, GBL^{+}25, KB05, Poo94, SK21, TFB18, Cha09, Gri00, Tow06]. Mangone [Cur85]. Mangrove [NSV^{+}22, PB94]. Manual [Loo01]. Manuel [Sch07]. Many [KHB^{+}25]. Maori [MC92]. Marahudin [Com88]. Maree [ACH^{+}12]. Margaret [SMRK85]. Marginal [JA89]. Mariculture [PB94]. Marie [Gri00]. Marine [AAO^{+}09, Ali11, ADvH^{+}18, And02b, ACH^{+}12, Ano93a, Ano19a, Ano20e, Ano20f, Ano20g, AR01, BWR^{+}23, BS06a, BP94, Bro85, Bro86, CER21, CKMT15, CC86, Com88, CR15, DBR^{+}19, EC89, Fre95, GOFR88, GS98, HHSW12, Han98, Hee89, HGV^{+}99, HB96, Hol04, JKP95, Kah87, Ken06, KBBN23, LSH^{+}18, Lew19, LW20, LZ87, MD17, Mil88, Opa87, Ret88, Ret94, Row04, RHJ92a, San04, Sch97, SW04, SZC07, SS16, TB97, Tis86, TgKtC99, VM86, WB09, Whi93, WH99, ZAS^{+}24, Mar86a, Roh03b, Van96, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano00b, Ano00a, Ano01b, Ano01a, Ano02b, Ano02a, Ano03b, Ano03a, Ano04b, Ano04a, Ano05b, Ano05a, Ano06b, Ano06a, Ano07b, Ano07a, Ano08b, Ano08a, Ano09g, Ano09b]. Marine [Ano09a, Egg06, HBB^{+}07b, Smi12, Sut11]. Maritime [VM86]. Mark [Gud98, Mar86a]. Market [AMM98, And85a, And92, AB93, AYSN17, AGH04, ABOS12, AGC^{+}22, BM96, BAYN16, CDNS17, DOY21, EOA^{+}17, FS15, GSA93, GA95, Han85a, Hel22, HC06, Hol13, HBAY^{+}16, HNAY^{+}21, Hua15, JLT19, KOZ25, LAM21, MJ20, Mat97, MD17, MS11, NQ06, NDW22, NK18, Nie05, NSL^{+}07, NSG12, NL09, PGB00, PWC22, Qua06, RLK22, RW00, Roh03b, Rus99, STNY23, SVS99, SPC18, SPZ19, Sjö15, Smi23, TKMA17, Tok18, Tro01, VSSH05, WM16, WA92, WW93, XZ14, vBDSD10, And96]. Market-oriented [Tro01]. Marketable [NAB14]. Marketing [HDR96, Joh85, MCFH09, OZAo96, OAA04, Pro84]. Markets [Álv22, AH02, ADG^{+}11, AG14, FP25, GH96, GLN^{+}23, GMT11, Kri09, LA06, O’H20, PA20, RL14, STY20, SHL^{+}20a, Wes02]. Markov [Tom10, VW13]. Markov-Perfect [VW13]. Marlin [DNA^{+}12]. Marsh [BP94]. Martosubroto [Sai97]. Mary [Cha95]. Masirah [OZAo96, OAA04]. Masthead [Ano16a, Ano16b, Ano17a, Ano17b, Ano17c, Ano17d, Ano18a, Ano18b, Ano18c, Ano18d, Ano19b, Ano19c, Ano19d, Ano19e]. Materials [Van96]. Matrices [WB09]. Matter [AKP^{+}98b, ASP^{+}84, ASB^{+}85, ASR^{+}86, ABS^{+}93, ALL^{+}94, Ano84a, Ano84b, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89b, Ano89c, Ano89d, Ano89e, Ano92a, Ano92b, Ano92c, Ano92d, Ano93c, Ano93d, Ano93e, Ano93f, Ano94a, Ano94b, Ano94c, Ano94d, Ano95c, Ano95d, Ano95e, Ano95f, Ano96c, Ano96d, Ano96e, Ano96f, Ano97c, Ano97d, Ano97e, Ano97f, Ano98c, Ano98d, Ano98e, Ano98f, Ano99c, Ano99d, Ano99e, Ano99f, Ano00c, Ano00d, Ano00e, Ano00f, Ano01c, Ano01d, Ano01e, Ano01f, Ano02c, Ano02d, Ano02e, Ano02f, Ano03c, Ano03d, Ano03e, Ano03f, Ano04c, Ano04d, Ano04e, Ano04f, Ano05c, Ano05d, Ano05e, Ano05f, Ano06d, Ano06e]. Matter [Ano06f, Ano06g, Ano07d, Ano07e, Ano07f, Ano07g, Ano08d, Ano09c, Ano09d, Ano09e, Ano09f, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano24, Ano25a, Ano25b, Ano25c, Ano25d, ASH^{+}92, AKN^{+}93, AKG^{+}97, AGK^{+}01, AAR^{+}08, AHT^{+}09, BCH^{+}00, BLE^{+}05, BSS^{+}86, BCA^{+}87, BYD^{+}88, BBH^{+}06, BMS^{+}89d, BMS^{+}89c, BHB^{+}93, BMD^{+}00, BBI^{+}05b, CUW^{+}07b, CHW^{+}09b, CDT^{+}95, CGP^{+}03b, CMS^{+}85, CMC^{+}99b, tCTH^{+}93, CMGB^{+}07b, CPG^{+}98, CPB^{+}92, DCB15, EWR^{+}86, ELP^{+}07, ESO^{+}03, FAH^{+}02, FRK^{+}04, FFH^{+}95, GWB^{+}94b, GNC^{+}88, GSEM^{+}97, GMS^{+}97b, GAB^{+}86b, GWR^{+}00, HDS^{+}85, HMS^{+}86, HCL^{+}98, HBB^{+}96, HBB^{+}07b, HHN^{+}89, IOE^{+}99, JMF^{+}02, JRB^{+}06, KHD^{+}96, KAC^{+}97, LKK^{+}03b, LPV^{+}01, LA06, hLRT^{+}87b, LVV^{+}05, LAG^{+}04, LGO^{+}88, MPH^{+}99, MCC^{+}92, MMP^{+}04, MDS^{+}88, MMH^{+}09b, MBD^{+}00, NGB^{+}95]. Matter [OBB^{+}84, PBA^{+}94, PMT^{+}99b, PWL^{+}05b, PSKS13, RCN^{+}09, RHJ^{+}92b, SSH^{+}89, Seu24, SBK^{+}87, SLM^{+}98, Sjö15, SAS^{+}94, TBH^{+}01, TBC^{+}86, TAW^{+}02, TTS^{+}02, VGH^{+}96, WMS^{+}01, WDS^{+}85, WKGK87, WSW^{+}04, WGW^{+}03, YCN^{+}95, YBG^{+}96, YLB^{+}06, ZBP^{+}06b]. Matters [Vis02]. Mauritanian [DK96]. Maximised [KH06]. Maximizing [BCHH00]. Maximum [Mai10, SV16]. May [And10]. McCay [Geo06]. McGoodwin [Rob93]. Measure [Han10, War94]. Measured [SDQ21]. Measurement [Asc07, PTE04]. Measures [AH93, AMW^{+}99, BC13, LRCL25, Sch98]. Measuring [Ano89a, BMS89a, FJ08, GOFR88, GMTW19, HNCS12, Hup89, HM11, Kah87, MWH^{+}18, MPE10, Sch95, Smi96, WA23, ZAS^{+}24]. Meat [SD97]. Mechanisms [LG98, MP85]. Meet [And07, Roh07]. Meeting [CER21]. Mekong [vBDSD10]. Melissa [Gri93]. Member [PD00]. Members [Kna12]. Memorial [Bjø93]. Mendelsohn [Gre99]. Menz [Com89]. Meryl [ACH^{+}12]. Meta [JRBH06, MF22]. Meta-Analysis [JRBH06, MF22]. Metallic [CC86]. Metapopulation [San04]. Method [FGM^{+}98, SB85]. Methodological [KSV05]. Methodology [WL94]. Methods [AS84, AS89, Smi89, Swi85]. Mexico [ASS^{+}14, GCG98, GOG00, Gri93, LPV21, MCFH09, MS01, SABO14, SS16, Str04, WT05]. Michael [And87a, Pon89]. Michigan [Bea25, CBI92]. Microeconomic [And02c]. Microeconomics [And99, Hat05]. Migrating [Han13]. Migration [And02b]. Migratory [MPE10, NP97]. Mineral [CC86]. Minerals [Bro86]. Minimum [HR99]. Minke [KT94]. Miriam [Tow06]. Misallocation [Sal93]. Mislabeling [RLA10]. Missing [LHCL15]. Mississippi [WT05]. Mitigating [SB87]. Mixed [LSBH21, TKMA17]. Mixture [GMTW19]. Mobile [DG00]. Mobility [Seu24]. Model [AH93, And86, AB93, AS94, Ano93b, AKV^{+}17, AMA00, BMS89b, BM00, BBI05a, CHS93, CHW09a, CN00, CS86, CMC99a, CPW03, CP98, dLS92, EA89, EK95, FGMW03, Gil88, Gor20, GHM95, GOFR88, GMTW19, Her93, HC06, HJ97, HBB07a, JH05, KBS01, KBS02, Kou95, Kva16, LSW17, LSH^{+}18, LRCL25, MB10, Mil88, PC94, PMT99a, PA09, PH99, PS03, PRS92, Qua06, RLK23, Res95, RS96, Sam92, ST87, SRTS86, Smi85, Sto86a, Sum95, Wal06, WK88, WS10, KSV05]. Modeling [Ber06, BCSA22, BLS16, CL09, CN00, DNA^{+}12, Eas92, HSCH04, HCC23, JL99, OA10, OB84, PTE04, SZC07]. Modelling [ADvH^{+}18, HMMhL92, Ken87]. Models [Ali11, BSA^{+}22, BHK97, Dol88, Gom92, GSE97, HH99, Kli87, Loo01, MPTB22, RAH17, SLW85, SW06, TP85, MSBh95]. Modest [Gul89, Pon88, Pon89, RC89]. Mollett [And87c]. Molluscan [FJ23]. Monopolies [MU16]. Monopoly [Han85a]. Monopsony [Sto86b]. Montreal [WMB95]. Moral [BP97]. Morality [DCB15]. Moratorium [SRS^{+}13]. Morgen [Kna95]. Morrissey [ADvH^{+}18]. Mortality [BSA^{+}22]. Moss [Geo87, DR04]. Motivated [RLK22]. Motives [MP97]. MOVA [Tro01]. Moving [SF19]. MPAs [AAC^{+}21]. MRE [Ano95g, ARS14]. MSC [Wak14]. Mullet [mHS06]. Multi [And05, dLS92, DG00, ES16, Lin05a, PRS92, TBA95]. Multi-Criteria [dLS92]. Multi-Gear [PRS92]. Multi-national [Lin05a]. Multi-Species [DG00, ES16, PRS92, TBA95, And05]. Multicohort [Ken92, Spu85].

Multilateral [CGP03a]. Multimedia [Jin94]. Multinational [Ken03]. Multiobjective [SE94]. Multiple [FS15, Kli87, MP99, SB85, Sch01, WJN^{+}16]. Multiple-Criteria [MP99]. Multiple-Site [SB85]. Multipurpose [FHJ87]. Multispecies [CN95, DSBT97, Hee89, Her06, LHT96, NLH12, PWL05a, SW18]. Multistage [DAP11]. Multiunit [PTMQ22]. Munro [Arm13]. Musings [SRA^{+}23]. Mussel [AR10]. Mussels [Tho12, WMB95]. Myanmar [AKT^{+}22]. Myopic [HRK86, RJY16].

N [Kna95, Hof04]. NAAFE [And10, Kna12, SLA14]. Narragansett [HTTA92]. Nash [LK00]. National [BC84, CHN12, DvHM18, LLLV16, Loo88, PKLB09, PIH14, Smi97, SS16, WB86, Lin05a]. Nations [LSHQ06, Mun00]. native [Kat07]. Natural [ADvH^{+}18, BSA^{+}22, BCP01, Bos13, DS87, Gri93, GOFR88, LN16, SK21, Wan92]. Nature [GSE97, Loo01, MF02]. Navigating [KHB^{+}25, SH04]. Near [ARA^{+}25, AW18]. Need [Cum87b, Wil07]. Needles [TB97]. Negative [HO11]. Nephrops [EU99, MB10]. Nested [CHW09a, MC99, Mil88, PA09]. Net [BHK97]. Netherlands [Tow03]. Network [RL14, San04]. Neumann [Gre99]. New-Member [PD00]. Newfoundland [SRTS86, SRS^{+}13, McC02]. Newport [And10]. News [KOZ25]. Niche [FS15]. Nigeria [AAO^{+}09, CLSA22, OA10]. Nina [And87c]. Niño [PTDV17]. Nitrogen [NAB14, WT05]. Nixon [Van96]. NMFS [HGV^{+}99]. No [And87c, Com89]. Non [AK20, AKV^{+}17, BC13, dLS92, GBL^{+}25, HLM20, mHS06, Kat07, Lew19, LW20, LWL^{+}20, Mai10, PH20, ZFG20]. Non-Attendance [Lew19, AK20, HLM20, LW20, LWL^{+}20, PH20, ZFG20]. Non-cooperative [mHS06]. Non-fishing [GBL^{+}25]. Non-Linear [dLS92, Mai10]. Non-native [Kat07]. Non-technical [BC13]. Non-Use [AKV^{+}17]. Noncooperative [FM96, PY89]. None [MU16]. Nonfee [LRCL25]. Nonharvested [PWC22]. Nonimmigrant [MG16]. Nonlinear [CP98]. Nonmarket [AKKM86, PWC22]. Nonparametric [LHT96, WT10]. Nonpoint [DS88]. Nonrenewable [Swa94]. Nonsustainable [PB94]. Norbert [Yoh03]. Nordic [NFW12, WJN^{+}16]. North [LLA00, ASS18, And87c, And10, BLEV05, BC87, BER^{+}14, Cro16, HVTY05, HNCS12, Ken03, Lin05a, MC03, NB03, PS03, QXH11, WDLH11, Wil18a, Wil18b]. North-East [Ken03]. Northeast [ES16, LSW17, SW18, ZS09]. Northern [Álv22, BDCDeS00, DBP00, GIP02, KW09, PD02, Poo97, SRS^{+}13]. Northwest [WHL^{+}18, YH96]. Northwestern [CYP92, TPC03]. norvegicus [EU99]. Norway [CAAT10, EU99, GHB93, Han85b, HRA^{+}17, HVTY05, MOOT23, PA20, ZSG23]. Norwegian [Arm94, AMA00, AA00, BG00, BaUS04, BER^{+}14, ESOF03, FHS95, FKAvR10, GR11, HBW22, HM95, KJE03, Kri09, LA11, LK00, Sal89, Sal93, San05, Str14, SAA^{+}20, Tve02b, Tve02a, VH11]. Note [And13, And94c, CL89, Han07a, Øst99, ST89]. Novel [GMTW19]. NSSH [BaUS04]. NSW [GHM06]. Nuclear [AW18]. Nuisance [SK21]. Number [hS98]. Numbers [SDQ21]. Nunatsiavut [KHB^{+}25]. Nutrient [DLL17, KBS01, KBS02, KB05, NDW22]. Nutrition [BHG^{+}16]. NY [NDW22].

Oahu [PHCK16]. Objectives [And07, DS85b, Roh07]. Observable [Tom10]. Observed [FP25]. Obstacle [Gom92]. Obstacles [Sco93]. Obtaining [LGOS21]. Ocean [Ber06, Edw94, GHM06, GOFR88, Hal16, HKS04, HJKP03, Jin94, Kva18, MF22, Smi17, Kna95, MKN23, OLM^{+}21]. Oceans [AAC^{+}14, Ale87, KFS^{+}15]. Off [PWC22]. Off-Bottom [PWC22]. Offshore [FR92, Gar88, MOJG04, hS98]. Ofiara [Yoh03]. Oil [AKKM86, Cur85, GAB^{+}86a, GLN^{+}23, JKP95, MOJG04, RRS88, Sla07, WHL^{+}18]. Oldest [YC15]. Oligopsony [FJ08]. Oman [AMK06, OZAo96, OAA04]. Omar [And96]. Omega [Kri19]. Ommer [McC02]. One [AGH04, FPA13]. Onsite [BWR^{+}23, LLLV16]. Open [And02b, And04, Chr96, Gom92, Jin12, KH06, KBS01, KBS02, Kro04, Sto86b]. Open-Access [Gom92, KH06, KBS01, KBS02, Sto86b, And04, Kro04]. Operations [Gat84, YL06a, YL06b]. Opportunities [AGT08, AGC^{+}22, Cro16, Fel94]. Opportunity [Par87]. Optimal [ALP^{+}21, And89b, Arn92, Arn13, Bjø88, CL89, CM85, CCBM85, CC87, Egg12, Gil88, Gut08, Hal04, Han86, Hea93, HJKP03, HF07, HHP10, JKP95, KW09, Ken92, KM02, LSHQ06, MB10, Maz04, MS98, NLH12, PWL05a, PD02, RS96, SS97a, ST87, ST89, Sch01, hS98, Tah09, WB86, YL06a, YL06b]. Optimality [FJ03]. Optimization [dLS92, Jin94, PRS92]. Optimum [HS86]. Option [CLL22, Jin94, Li98b, Li09, Sar09]. Options [And87c, PC94, Her88]. Orange [CHS93]. Orchestrate [CNAB09]. Order [Ale87, Álv22, Vis02, Cha95]. Organic [AYNN19, AYJO^{+}19a]. Organization [HM95]. Organizations [PTDQ^{+}19]. oriented [Tro01]. Origin [Meu95]. Origins [LWL^{+}20]. Ostrom [GLM^{+}25]. Other [RGA07]. Otter [CG85]. Our [SGC^{+}16]. Outdoor [LRCL25]. Outer [RRS88]. Outport [McC02]. Output [GHB93]. Outreach [MCFH09]. Outstanding [Ano09g, HBB^{+}07b]. Overcapacity [WMT04]. Overexploitation [ASS18, Wil18a, Wil18b]. Overfishing [Cav23, LPV21, OLM^{+}21, Kva18]. Owen [Her95]. Owned [EC89]. Ownership [Edw94, Ret88]. Oyster [BS89, DKK11, DKK04, KW09, LPG13, MUWB23, MWH^{+}18, PWC22, PECY25, SPC18, Geo06]. Oysters [DLL17, MCFH09].

P [Kin93, Sai97, SP85, Yoh03]. P. [Gat86]. Pacific [DS85a, MKN23, OLM^{+}21, ASS18, And87c, AS89, BCHH00, Bro85, Bro86, CGP03a, CC86, Com89, CC87, Coo88, DNA^{+}12, Her88, HC06, HBAY^{+}16, HS86, LSW09, hLRT87a, MC03, Mat10, Rud88, Sto86a, SE94, Tok18, Tow06, Wau88, Wil09, WCA^{+}10, Wil18a, Wil18b, ZBP06a]. Padre [PKLB09]. Panel [PGB00]. Pangasius [vBDSD10]. Papers [And10]. Papua [Her95]. Paradigms [PR99]. Parameter [AH93, SLW85, aMUSS02]. Parameters [CS86, McT99]. Parasitic [AAW^{+}17]. Parents [BHG^{+}16]. Paris [GSA93].

Partial [DS85a, YL06a, YL06b]. Partially [Tom10]. Participate [KTR00]. Participation [Boy93, CHLH03, ZSG23]. Partner [GBL^{+}25]. Partnership [AAC^{+}14]. Passive [HJ97]. Patagonian [HHP10]. Patchy [And02b]. Path [SCK^{+}15]. Pathogen [Ali11]. Paths [Mai10]. Patterns [AG01, HBB07a, LL13, RLK23, SJQ21, THM98, YAA20]. Paul [BHH^{+}12, KFS^{+}15, Kri19, SGC^{+}16]. Pauly [Sai88, Sai97]. Pavel [Gar93]. Pay [BHJ03, CS13, FS15, HM11, JRBH06, NNNJ22, PgK09, PIH14, Ret94, WCH01]. Payoffs [FM96]. Peak [Sla07]. Pedigree [Sco11]. Peer [DCB15, FLS14]. Pelagic [BER^{+}14, ES16, Sch07, Str04, YH96, dFdOdA24]. Peninsular [YH96, And96, SYN^{+}19, Yah88]. Perceived [AR10, RNKA21]. Perception [LPG13]. Perceptions [AW18, CAAT10, OHS14]. Perfect [Tow09, VW13]. Performance [AMK06, And04, Esp15, Gar93, HdSK15, PFN13]. Peril [Vis01]. Periodically [And88]. Permit [WJ12]. Permits [NAB14]. Perspective [And87b, KG98, MM04, Wau88, WL94]. Perspectives [Neh93, Kin93]. Peru [TPPT11]. Peruvian [Sai88]. Peter [All98, Gri93, Geo87, Geo96]. Petroleum [Cha95]. Philip [Geo87]. Philippine [SVS99]. Phillips [Roh03a]. Pillars [DAAY^{+}20]. Pilot [SAU12]. Pink [EW86]. Pioneers [Fel94]. Pirates [Tow09]. Place [Lew19]. Planet [Kri19]. Planning [NDW22]. Plant [AW18]. Play [KGW22]. Playing [Ola07]. Poaching [Tow09]. Poernomo [Com93]. Poisson [MSBh95]. Policies [AAC^{+}21, BS89, CN00, Chu86, MC92, SYN^{+}19, Smi86, hS98, YH96]. Policing [Hal04]. Policy [And89a, And96, And07, CKMT15, CM17, Com93, CNAB09, Gar93, HDK^{+}11, Han85b, Jen98, Kna12, Mar86a, MC03, Meu95, PR99, Pas07, RAW17, Roh07, SB87, Smi97, SC95, TS02, WJ12, Wil07, DS85a, Bos13]. Political [MOOT23, PT97]. Pollnac [Com93, Kin93, SP85]. Pollock [CH25]. Pollution [CPW03, DS88, GOFR88, Kah87, KGW22, MOJG04, NDW22, Opa87]. Polynomials [HMMhL92]. Polyprion [GWB94a]. Pond [RNKA21]. Pontecorvo [Ale87, Gul89, RC89]. Pool [EOA^{+}17, GC16, Lib07]. Poor [Wil13]. Population [CS86, Han88, LG07, Ste98]. Populations [LL94, QT19, Tah09]. Portfolio [JBH24, PWL05a]. Ports [SW04]. Possession [Lib07]. Possibilities [GHB93, Sal89]. Possible [Chu86]. Pot [Sta87]. Potential [And02c, tCT93, EK95, HH86, HS86, LAM21, LF11, MMSS12, YGB97]. Power [And96, BHW15, FJ08, Hel22, Rus99, TP85]. pp [All98, Tow03]. practical [Poo94]. Practice [CNAB09, DS88, HDK^{+}11, Pas07]. Practices [LWL^{+}20, PR99]. Prawn [GHM06, SVS99]. Prawns [SPZ19]. Predator [BBI05a, EAÖH25, GSE97]. Predicting [NNNJ22, PWL05a]. Predictions [Sam92]. Predictors [GLN^{+}23]. Preface [PA09]. Preference [AYAB^{+}20, BM00, GLL^{+}10, Lew19, LWL^{+}20, PCH^{+}13, WDLH11, ZWS21, GPC^{+}13]. Preferences [AB93, BLS16, CKMT15, Edw92, EL07, FJ23, KNH93, Kum02, LGL19, MCFH09, MG16, MJ11, RLK23, RAC^{+}17, SJQ21, Str04, WAUR17].

Preliminary [BMS89b, MCFH09, Mat09, McT99]. Premiums [AYNN19, FP25]. Presence [Boy00, Dan00a, Dan00b, JLMR96]. Preserve [Ret94]. Press [Ale87, All98]. Pressure [Hut14]. Prevalence [Lib07]. Previously [PH20]. Prey [BBI05a, EAÖH25, GSE97]. Price [Álv22, ART08, And89b, AYNN19, AL18, AG01, AGH04, AMO19, Bos04, CLL22, DO14, DOY21, Dup93, FPdR20, Gor20, GGKW17, GLN^{+}23, HVTY05, Jaf04, NK18, Nie05, NSG12, OS08, Øst99, PTDQ^{+}19, PA20, PAC^{+}22, RLK22, RL14, STNY23, San05, ZT22, dFdOdA24]. Price-quantity [Bos04]. Prices [ADK^{+}11, BKS^{+}23, DOY21, GCG98, GH96, GLN^{+}23, Ham15, HRA^{+}17, HGJH15, HNE^{+}89, MS98, PGB00, SVS99, XKM08, ZSG23]. Pricing [Álv22, Fel09, Kou95, Lee14, Sjö15, WAUR17]. Primary [OG03, Kri09]. Principal [Gat84, JV01, Pro84]. Principal-Agent [JV01]. Principle [Kri19]. Principles [And15a, AKKM86, Rud88]. Private [BS89, FP25]. Prize [Bjø93]. Pro [Wil13]. Pro-Poor [Wil13]. Problem [Cha00, PD00]. Problems [And87c, BC84, Kin93, Mun00, SF19, SCS06, Swi85]. Procambarus [JLMR96]. Proceedings [And87c, Com89]. Process [And84, MGC89, Tom10, MSBh95]. Processing [CL98, DG94, GH86, GHB93, HKBH05, HG89, JHT06, LWL^{+}20, MP85, Mat09, MG16, OG03, Pro84, RW00, Sto86b]. Processor [CH25]. Processors [GAW^{+}18, LS99, MC03]. Produced [AW18, LS99]. Producer [TPC03]. Producers [AGC^{+}22]. Product [AB93, AH02, KM02, OHS14, ZBP06a]. Production [AGR19, Asc97b, AG14, Boy00, CP98, Dan00a, Dan00b, FJ23, GH86, GBDT08, GW00, GMT11, KC95, Ken06, KBBN23, LP01, OA10, RAH17, RAW17, Sal93, SL98, Tve02a, vBDSD10]. Productive [GMTW19]. Productivity [Asc97b, ABS03, ART07, GBDT08, Ret88]. Products [CDNS17, CHN12, HCC23, LAM21, NK18, OZAo96, OAA04, RNFH22, SD97, TYH15, WAUR17]. Professor [BS13, RC89]. Profile [Gri93, Jin96]. Profit [FHM17]. Profitability [DF16, GRA^{+}25, Ste98, UUL^{+}10]. Profitable [HF01]. Program [ASS^{+}14, BHBG93, KTR00, Kna97, RLA10, SAU12, YC15, HD25]. Programming [AS94, PC94]. Programs [And94b, Ano95b, DLL17, DS85b, GDF04, HH86, LSW09]. Progress [Smi89, VH11, WA92]. Promise [Vis01]. Pronged [STY20]. Propagation [SLW85]. Propensity [DF16]. Property [ASS18, And85a, Arn07a, BC13, BSL92, Chr96, Edw01, Geo06, GC16, Lib07, MU16, Maz04, Mun07, NV12, Sco88, Wil18a, Wil18b]. Proposal [Pon88, Pon89, RC89, Gul89]. Proposed [SS16]. Prospects [ACDL21, SRTS86]. Protected [GM07, Hal04, NF10]. Protecting [MD17]. Protection [CP16, DvHM18, LLR10]. Providing [PCAIGO^{+}14]. Province [Chu86]. Proxies [Øst99]. Psuty [Yoh03]. Public [CKMT15, DBR^{+}19, FJ23, FR92, FP25, Geo06, Hal16, Sal93]. Publicity [HO11]. Publicly [EC89]. Published [PH20]. Pulsed [BSL92]. Purchase [NGBM95, QXH11]. Purse [CN95, FHS95, PFN13]. Pursuit [DAAY^{+}20].

Putting [SV16].

Quahog [HKS04]. QUAIDS [BLS16]. Quality [BC13, BMS89a, Fre95, HTTA92, LA06, Lip04, MW02, PHCK16, Tve02b]. Quantifying [AAW^{+}17, MC03, SPC18]. Quantile [WA23]. Quantities [NSG12]. Quantity [And88, San05, Bos04]. Quarry [Bel89]. Queen [MVRRP21]. Quentin [ACH^{+}12]. Quest [Kri19, RAW17]. Quota [Arn93, Bos04, EVH05, FHS95, Gra96, HRA^{+}17, Hat05, JLT19, Joh95, Joh96a, KG98, LN16, LSW09, Mat09, RL14, Wat01]. Quotas [AMM98, And86, BP06, CDTW95, Cla85, DG00, GWB94a, GN88, HKBH05, Han88, HF07, Hut14, KW86, PTDQ^{+}19, PAC^{+}22, SDQ21, TPPT11, WB86]. Quotes [GW92].

R [And87a, ACH^{+}12, Arm13, Gri00, Kna95, Rob93, SP85]. R. [Kin93]. Race [TPPT11]. Radioactive [WM16]. Ragnar [Rus00]. raised [NQ06]. Randall [Bos13]. Random [BMS89b, GMS97a, HBB07a, LSH^{+}18, LRCL25, MSBh95, McT99, PMT99a, RLK23, Sch98, TS02]. Ranking [HT06]. Rate [Han86, PB84, Row04, ZKD^{+}14]. Rates [Cla85, GLN^{+}23, LL13, SB85, XKM08]. Rationalization [Seu24]. Ratios [ZT22]. Rattle [Chr96]. Ray [ACH^{+}12, Fou02]. Real [And15a]. Real-World [And15a]. Reallocation [BHK97, NAB14]. Reassessment [OAA04]. Rebuilding [LSHQ06]. Recirculating [KC95]. Recovery [AMW^{+}99, BB06, KG97]. Recreation [AED11, BM00, DvHM18, Fre95, HH99, Kli87, LLLV16, LRCL25, MPTB22, PMT99a, PH99, Smi89]. Recreational [Abb15, AH93, Ali11, AL13, AP17, And15c, BWR^{+}23, CHW09a, CLL22, GOG00, GMS97a, HHSW12, HGV^{+}99, HR99, HBB07a, JRBH06, JL99, LSW17, LS14, Loo88, MPE10, NF10, OGL^{+}16, Ola09, PA09, PCH^{+}13, PHCK16, PCAIGO^{+}14, RCLL17, TS02, WH99, WCH01, WHL^{+}18, WCFZS03, WG03, MSBh95]. Recreational-Commercial [Abb15]. Red [Abb15, ASS^{+}14, FPA13, GOG00, JLMR96, Mat09, SK21, Wat01, JKPHS12, NSV^{+}22]. Redfish [FKAvR10]. Reduce [AED11, Cav23, LPV21]. Reduced [AYNN19]. Reducing [GRA^{+}25, WT05]. Reduction [KPS04, Sta87, hS98, War94]. Reductions [LSH^{+}18, LP01]. Reef [LLS25, PECY25, WCFZS03]. Reefs [MMCM13, MMH09a, PWC22]. Reference [Hol04]. Reflections [KSV05, Smi12]. Reflective [And13]. Reform [Geo96, GST22, Wil13]. Regimes [AC97, BBA^{+}10, Chr96, Lin13, Rus99]. Region [GKK^{+}88, Her88, QXH11, Com89, Wau88]. Regional [CHLH03, LP01, LS14, SW06]. Regions [KBK^{+}21]. Regress [VH11]. Regrouping [LPV21]. Regulated [Wil85]. Regulating [And86, LRCL25, WJN^{+}16]. Regulation [And87b, And94b, And99, And15a, AS89, Ano95b, BP97, BCP01, CP16, Dow93, Her06, Koe84, LZ87, Maz04, MD17, MS11, PT97, Sal93, Sto87, Ves96, Yah88, Yoh84]. Regulations [Han85b, RG86, Smi23]. Regulators [LSHQ06]. Regulatory [CMGB07a, DKK04, Geo96, Hol00, MP85, Poo94]. Related [KP08, MUWB23, ZS09]. Relation [KSV05]. Relations [Bos04]. Relationship [CP98, HGJH15, JHT06, Tve02b, WMT04]. Relative [And88, AG01, ABS03]. Relaxing [Gom92]. Relaying [DKK04]. Reluctant [Fel94]. Remedial [BS89]. Remembering [Roh11]. Removing [Gom92]. Remuneration [AGYW10a, GMM^{+}15, PFN13]. Renewable [CS86, Edw94, FM96, Han86, MP97, PB94, Swa94, Wil87]. Rent [AAF10, BCP01, BCHH00, Edw01, FHS95, FHM17, GMM^{+}15, GSKA20, Hal04, Han10, KNX95, LSW09, LCB92, VW13]. Rent-Seeking [Edw01]. Rents [BBA^{+}10, DS85a, Edw05, KFS^{+}15, KH06, MU16]. Reorganization [And02c]. Repeated [BM00, LRCL25, PTMQ22]. Replenishable [PY89, Smi85, Smi86]. Reply [Dan00b, Joh96a, Li09, PS94, Pon89, WCA^{+}10]. Report [And87c, DS85a]. Requirements [DKK11]. Research [JH05, SRA^{+}23]. Reserve [San04, ZSG23]. Reserves [And02b, AR01, Han98, HB96, SW04, SZC07]. Residence [Lew19]. Residential [Par87]. Residuals [Con85]. Resilient [McC02]. Resistance [IO99]. Resolution [NF10]. Resource [AMM98, AAF10, Ano93a, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano00b, Ano00a, Ano01b, Ano01a, Ano02b, Ano02a, Ano03b, Ano03a, Ano04b, Ano04a, Ano05b, Ano05a, Ano06b, Ano06a, Ano07b, Ano07a, Ano08b, Ano08a, Ano09g, Ano09b, Ano09a, Ano19a, Ano20e, Ano20f, Ano20g, BC84, BCP01, BCHH00, BBA^{+}10, Bos13, CS86, DS87, Egg06, FM96, FHS95, FHM17, GC16, GSKA20, Han10, Hee89, Ken06, LS99, LN16, MVRRP21, Mer23, MP97, Row04, Smi85, Smi86, Smi12, SF19, Sut11, Swa94, TgKtC99, Wan92, WB09, Gri93, Bos13]. Resources [ASS18, CER21, CC86, EC89, Edw94, GIP02, GOFR88, Hal16, Han86, Hol04, Mun86, PB94, PH20, RRS88, RHJ92a, Sai97, Tis86, VM86, Wil87, Wil18a, Wil18b, BP94, ADvH^{+}18, KFS^{+}15, Lin13, Ret88]. Resoure [HBB^{+}07b]. Response [AW88, AA00, JBH24]. Responses [LSBH21, MS11]. Responsible [PR99]. Responsiveness [ART08]. Restoration [BS06a, MPTB22, PgK09, PIH14, PECY25]. Restricted [GKK^{+}88, Sal93]. Restrictions [EQJ94, Par87]. Restructuring [LSW09]. Results [AS84, BMS89b, LL94]. Retail [AYAB^{+}20, CHN12, FR04, RGA07, STNY23]. Retailers [FJ08]. Retailing [MF02]. Retreat [HD25]. Retrospective [Wil88]. Returns [FGM^{+}98, KOZ25, MMSS12, NFW12, PWC22]. Revealed [PCH^{+}13, WDLH11, GPC^{+}13]. Revealing [DOY21]. Revenue [AA00, Han07b]. Revenues [Her93, Ano93b]. Reversal [EQJ94]. Review [ALS^{+}18, Ale87, All95, All98, AS84, And87a, ADvH^{+}18, And96, ACH^{+}12, Arm13, ALG^{+}19, BP94, Bos13, BHH^{+}12, Cha95, Cha00, Cha09, Com88, Com89, Com93, Cur85, DS85a, Dol88, DPP^{+}22, Fel94, Fre95, Gat86, Gat96, Geo87, Geo96, Geo06, GPC^{+}13, GHH^{+}15, Gre99, Gri93, Gri00, HRA^{+}23, Her88, Her95, KFS^{+}15, Kin93, Kna95, Kri09, Kri19, Kva18, Lin13, Loo01, MP99, Mar86a, McC02, Meu95, Neh93, Pan87, Pol86, Poo97, Ret88, Rob93, Roh03a, Rus00, Sai88, Sai97, Sch95, Sch07, SW06, Sla07, SP85, SHK^{+}13, SGC^{+}16, SMRK85, Tow03, Tow06, Tow09, Tyr01, Van96, Vis01, Vis02, Wau87, Wau88, WBC^{+}13, Yoh03]. Reviewed [GTP^{+}11, HLA^{+}11, HHN^{+}12, Pas12, SAB^{+}12]. Reviewers [Ano95g, Ano96b, Ano97b, Ano98b, Ano99b, Ano00b, Ano01b, Ano02b, Ano03b, Ano04b, Ano05b, Ano06b, Ano07b, Ano08b, Ano09b]. Revisited [Li09, LS21, SSH07, Sar09]. Rhode [SAU12]. RI [And10]. Richard [WBC^{+}13, Com93]. Right [Mer23]. Rights [ACDL21, Arn07a, Chr96, CD07, CP92, Edw01, ESC18, Gon96, GC16, Hol04, HRSW18, KFS^{+}15, Lib07, MU16, Mun07, NV12, Tow03, VW13]. Rights-Based [ESC18, VW13, HRSW18, Tow03]. Risk [AR10, AGR19, BP97, BP06, Cha00, EL07, GBL^{+}25, JKP95, KBBN23, Kum02, LPG13, MS01, NQ06, OA10, PSK15, RNKA21, SJQ21, Seg87, Str04]. Risk-Sharing [Seg87]. Risks [AR10]. Risky [MUWB23]. River [AMW^{+}99, ALG^{+}19, Boo89, NSV^{+}22, WT05, And15c]. Rivers [AAO^{+}09]. Road [BT14]. Roads [HRA^{+}23]. Robert [Gre99]. Robinson [Pon89]. Rock [SPC18, Sta87]. Rockfish [DR04]. Rognvaldur [Wau94, BS13, DPP^{+}22, GHH^{+}15, And13, Sch07]. Role [Abb15, FJ23, GBL^{+}25, KW09, Lin05b, Mer23, MS11, OHS14, PTDQ^{+}19, PPV22, PB84, SJQ21]. Rosemary [McC02]. Rotation [Gut08]. Roughy [CHS93]. Rowena [Pan87]. Rules [Han93, KL07, Lib07, San05]. RUM [MC99]. Run [ART08, And99, And15c, Gil88, Gor20, HGJH15, Sch97, Sto86a, HKS04, KSV05]. Runoff [WT05]. Rural [McC02]. Russian [Arm94, Joh96b].

S [Ano06c, Ano07c, Ano08c, Ano09g, HBB^{+}07b, Sla07]. S.-Y [Ano06c, Ano07c, Ano08c, Ano09g, HBB^{+}07b]. Sablefish [MC03]. Safer [PPV22]. Safety [CL98, PPV22]. Sales [AYAB^{+}20, HVTY05, Mer23]. Salmon [AAW^{+}17, AMW^{+}99, ART08, And92, AB93, AL13, AW88, AYAB^{+}20, Ano93b, Asc97b, ABS03, AHTT09, ADG^{+}11, ALG^{+}19, AMO19, ADK^{+}11, Bel89, BHJ03, Bir86, Boo89, BHBG93, Boy93, BT14, BM00, DOY21, FJ08, GRA^{+}25, GA95, Gut02, HMMhL92, Her93, KM02, KOZ25, LA11, LG98, Loo88, Mer23, MN18, MOOT23, MJ11, OS08, OGL^{+}16, Ola07, Ola09, OHS14, Øst99, RLK22, RNFH22, SAR05, Str14, Tve02b, Tve02a, VH11, VSSH05, WW93, XKM08, XM11, XZ14, ZKD^{+}14, ZFG20, ZT22, ZWS21]. Salmonid [CBI92]. Salmonids [AYNN19]. Salt [HRA^{+}23]. Saltwater [WCH01]. Salz [Gar93]. Sample [BLS16]. Sampling [LLLV16, McT99, TB97, WDJK20]. Samuel [Sch07]. Sanctions [Sch97]. Sanctuary [SS16]. Sardine [ESC18]. Sardinella [Dia96]. Satellite [ZAS^{+}24]. Scale [ACDL21, DTS20, GST22, GD05, SJQ21, Sal89, SP85, AAO^{+}09, EL07, PFN13, ZBP06a]. Scaling [Sch95]. Scallop [AL18, Edw01, Edw05, GDF04, WGM86]. Scanner [SDS12]. Schedule [YL06a, YL06b]. Schedules [Arn92]. Schuur [Gat86]. Science [Sch95, Gat86]. Scientists [Sis84]. Scope [Ken03]. Scores [DF16]. Scotia [DG00]. Scott [MDC^{+}15]. Scottish [TP05]. Sea [AGYW10a, AAW^{+}17, And87c, AL18, Asc08, CP16, Edw01, Edw05, Gri00, Gut08, LLR10, MKN23, RW00, RS96, SW04, WGM86, WDJK20, BC87, CH25, GSJ13, Hut14, JKPHS12, KBS01, KB05, Kro04, Lin05a, NLH12, NB03, SK21, Sum97, Cur85, KBS02]. Sea/Aleutian [AGYW10a]. Seabird [Tom10]. Seafood [AW18, AG14, ARA^{+}25, BHG^{+}16, BLS16, BA20, CL98, CHN12, Edw92, FS15, GW00, GMT11, HDR96, KBK^{+}21, KNH93, Kri19, MG16, O’H20, OG03, PGB00, Pro84, QXH11, RLK23, Roh03b, Roh09, RLA10, STNY23, STY20, SDS12, SAA^{+}20, WM16, WAUR17, WTW^{+}19, WA92, Wes02, YAA20]. Seafood-Safety [CL98]. Seafood. [SGC^{+}16]. Seal [ASS18, Wil18a, Wil18b]. Searching [ALG^{+}19]. Seas [CN06, KM93, Li98a, Lin13, Mun07, PTSB99, PD00]. Seashore [DvHM18, LLLV16, PKLB09]. Season [BS06b, KH06, PAC^{+}22, Roy98]. Seasonal [tCT93, KM02]. Seasonality [EOA^{+}17, HMMhL92]. Seaweed [GD05, LAM21, LGL19, SJQ21]. Seaweed-Based [LAM21]. Section [Ano19a, Ano20e, Ano20f, Ano20g]. Sector [ABW21, BS89, Esp15, Gar88, GAW^{+}18, MW02, SAU12, BP94, SP85]. Sectoral [GW92]. Sectors [JHT06]. Security [TFB18]. Seeking [BCP01, Edw01]. Segment [HF07]. Segmentation [AYAB^{+}20]. Segments [OHS14]. Seine [CN95, FHS95, PFN13]. Seiners [And05]. Selected [And10, Eas92, NFW12, Swi85]. Selecting [And89a]. Selection [BLS16, HT06, Sch01, THM98]. Selectivity [MB10, Mai19, SW18]. Self [Arn07b, Sco93, UUL^{+}10]. Self-Government [Sco93]. Self-management [Arn07b]. Senegalese [TL12]. Separability [SD97]. September [Ano99b]. Sequential [FW10]. Series [GA95, LPV21]. Service [HDK^{+}11]. Services [DLL17, KW09, LW20]. Set [HH99, KNH93, PH99]. Sets [WH99]. Sewage [Con85]. Share [AMM98, Esp15, GMM^{+}15, HRK86, PTDQ^{+}19, Qua06, SAU12]. Shared [MM04, Mun86, Mun07]. Shares [BKS^{+}23, GAW^{+}18, Mil88]. Sharing [And94a, Han07b, Han13, JKP95, KL07, MOOT23, PS89, PS94, Seg87]. Shark [CMGB07a]. Sharks [SHK^{+}13]. Shelf [RRS88]. Shellfish [AP17, FJ23, LZ87, WMB95]. Ship [MMH09a]. Shock [JBH24]. Shocks [MUWB23, Mer23, WS10]. Shorebirds [MPE10]. Shoreline [Bea25]. Shorelines [PWC22]. Shoreside [GAW^{+}18]. Short [ART08, And99, ADK^{+}11, Gor20, HKS04, Neh93, Sto86a, KSV05]. Short-Run [And99, Gor20, Sto86a, HKS04, KSV05]. Short-Term [ADK^{+}11]. Shortage [Sla07]. Should [And02a, CS13, Ret94, SDQ21, Cha00]. Shrimp [AYSN17, ABOS12, CL98, CV93, EW86, FGMW03, GCG98, GBDT08, Gri93, HNE^{+}89, HNCS12, KP08, MS01, MS11, NK18, NNNJ22, PDS^{+}21, SVS99, SABO14, TKMA17, Ves96]. Sicilian [WJP^{+}02]. Signal [Kat07]. Signals [MW02]. Significance [LS14]. Silver [THM98]. Simone [Lin13]. Simple [Gil88, MB10, SLW85]. Simulated [PWC22]. Simulating [FGM^{+}98, Gor20]. Simulation [FGMW03, Kli87, SF19]. Simulation-Based [SF19]. Simultaneous [KHB^{+}25]. Sinaloa [MCFH09]. Sinclair [All98]. Site [AH93, Kli87, SB85]. Sites [Ali11, PMT99a, TY05]. Siuslaw [Loo88]. Size [And89b, AL18, CH25, Han25, HR99, HS86, Lee14, LA06, Sjö15, hS98, WB86, WG03]. Size-Targeting [CH25]. Sizes [AG01]. Skill [VOJ^{+}01]. Sludge [Con85]. Small [AAO^{+}09, ACDL21, EL07, GST22, Ham98, PFN13, SJQ21, Sch07, SP85, WB09, YH96, ZBP06a, dFdOdA24]. Small-Scale [ACDL21, GST22, SJQ21, SP85, AAO^{+}09, EL07, PFN13, ZBP06a]. Smith [Sch95]. Snapper [ASS^{+}14, GOG00, MEA^{+}23, Wat01]. Snob [Ola09]. Snoek [HLH^{+}14]. Snow [GHM95]. Social [All98, CMGB07a, HRK86, HKPW13, HdSK15, RL14, Rud88, WB09]. Societal [MOOT23]. Society [McC02]. Socio [AMK06]. Socio-Economic [AMK06]. Sociological [Poo97]. Sockeye [And15c]. Sole [BBA^{+}10]. Solution [BDCDeS00, PY89]. Solutions [Arm94]. Solving [SF19, Cha00]. Some [BLEV05, BMS89b, JA89, McT99]. Sound [CNAB09, GMTW19, KGW22]. Source [ART07, DS88, MJ11]. Source-Based [MJ11]. South [BC13, Bos04, CER21, HLH^{+}14, HHP10, PTE04, Sta87, dFdOdA24]. Southeast [KP08, Mar86a, Tis86, VM86, WH99]. Southeastern [HF01]. Southern [Sta87, GN88, HH86, KW86, LSH^{+}18, LS14, PPV96]. Space [GA95, HJKP03]. Spain [Álv22]. Spanish [FPdR20]. Spatial [Ber06, CN00, CM04, EA89, GSJ13, Her06, HSCH04, Hol04, RAW17, SABO14, Str04, Wil04]. Spatial-dynamics [SABO14]. Spatial-Temporal [CN00]. Spatially [CPW03]. Spawning [AMA00, BG00, BaUS04, BER^{+}14, San05, KJE03, LK00]. Special [And10, Ano93a, AT02, AGT08, AG14, Bjø00, Cum87a, GMT11, HGH99, LH12, LW20, Neh93, SLA14, Ves05]. Species [Ano89a, BM96, Bos04, DvHM18, DG00, ES16, HHSW12, HNAY^{+}21, Hup89, Hut14, Kat07, LLR10, Lew19, NC07, NL09, PRS92, Sch98, SHL^{+}20a, TBA95, dFdOdA24, And05]. Species-Specific [HHSW12, Sch98]. Specific [HHSW12, Sch98]. Specification [Whi93]. Spill [WHL^{+}18]. Spillover [CP16, HO11]. Spills [AKKM86, GAB^{+}86a, WM16]. Spirit [GLM^{+}25]. Sport [BHK97, Mil88, TB97]. Sportfishery [AW88, CL09]. Sportfishing [BMS89b, CL10, CMLC16, CHLH03]. Spring [AMA00, BG00, BaUS04, BER^{+}14, KJE03, LK00, San05]. Spring-Spawning [AMA00, BG00, KJE03, LK00]. Squires [ACH^{+}12]. Stability [BDCDeS00, KL07]. Stackelberg [HRK86, LF11]. Stage [Sum95]. Stakeholder [CAAT10]. State [AAO^{+}09, GA95, Gul89, HH99, Pon88, Pon89, RC89, Smi86, Mai10]. State-Space [GA95]. Stated [HBCP09, Lew19, LWL^{+}20, PCH^{+}13, WDLH11, ZWS21, GPC^{+}13]. States [SS97a, CT14, Gre99, HF01, LSW17, Pow04, QXH11, SDS12, THM98, YC15]. Statistical [LS14, Swi85, WH99]. Statistics [Gul89, HGV^{+}99, Pon88, Pon89, RC89]. Status [LSJN^{+}24]. Steady [Mai10, SS97a, Smi86]. Steady-State [Smi86, Mai10]. Steelhead [JA89, Loo88]. Steering [AAC^{+}14]. Steller [LLR10]. Stephen [And87a].

Stevens [WR97]. Stevenson [SP85]. Stewardship [Roh03b]. Stochastic [AS84, AS89, AS94, CM85, EOA^{+}17, JH05, KPS04, Kva16, Seg87, SL98, Smi86]. Stochastically [PSK15]. Stochasticity [PSKS13]. Stock [And88, Arm94, BS06b, BaUS04, BB06, BDCDeS00, DOY21, DSBT97, EG13, Han88, Han07a, Han13, Han25, HJ97, KJE03, KOZ25, Li98a, LF11, MMSS12, PY89, Poo97, PSK15, SW15, Smi89, War94]. Stock-dependent [MMSS12]. Stocking [CBI92]. Stocks [ALP^{+}21, And05, And02b, CHS93, Ken92, LSHQ06, MM04, Mun00, Mun07, NP97, PTSB99, Sch97, Sch07]. Stopping [TPPT11]. Store [SDS12]. Store-level [SDS12]. Stores [HDR96]. Straddling [Li98a, LF11, PTSB99, PTDV17]. Strait [CGS05, CV93]. Stranded [Mat10, Wil09, WCA^{+}10]. Strategic [Fel09, Pow04, QT19, Rus99, Sum97]. Strategies [AR10, And89b, HHP10, KJE03, LPG13, MS01, PWL05a, RNKA21, Ste98, GSE97]. Strategy [JKP95, LWL^{+}20, TH12]. Strength [Abb17]. Striped [BK89]. Structural [CL09, Edw92, EK95, GMT11, SCK^{+}15, SZC07, XZ14]. Structure [AMK06, BG00, Bro86, BLS16, Gar93, Geo96, RLK22, Sal89]. Structured [Gil88, QT19, Tah09]. Studies [Ano19a, Ano20e, Ano20f, Ano20g, Øst99, PH20, Sai97, Tow06]. Study [DAAY^{+}20, GBDT08, HO11, Jaf04, KTR00, Loo88, NLH12, OA10, OG03, PFNV21, WMB95, WJP^{+}02, XM11]. Stuffing [Sto86a]. Stylized [AMA00]. Subjectivist [WL94]. Submarkets [RLK22]. Subsidies [CSW23, DF16, HW12, Jin12, KNYA23, Kva18, Lin05b, PFNV21, SK99]. Subsistence [KHB^{+}25]. Substitutability [BM96, Gut02]. Substitute [AH93, JL99]. Substitution [GHB93, HBB07a, KM02, RLK23, RLA10, Sal89]. Substitutions [CHN12]. Success [AW88, HKPW13, PPF88, SB85]. Suffolk [NDW22]. Sultanate [AMK06, OZAo96, OAA04]. Supplied [NSG12]. Supply [ART08, AYNN19, ART07, ADK^{+}11, FPdR20, GHM95, Kri09, LP01, LSBH21, MOJG04, NB03, RLK22, dFdOdA24]. Supply-Driven [LP01]. Supply-Motivated [RLK22]. Support [MOOT23]. Surf [Wan95]. Surface [MKN23]. Surfclam [HKS04]. Surimi [LS99]. Surplus [BLEV05, CP98]. Surpluses [Gil88]. Survey [AKKM86, CLSA22, FP25, HGV^{+}99, Jen02, Kli87, Kna97, Kna12, LHCL15, LL94, NTS11, PIH14, TYH15, WH99, Yoh84]. Sustainability [ATR^{+}23, DAAY^{+}20, KT94, TL12, TB01, Tve02b]. Sustainable [GW00, HF01, Lin13, Mai10, NC07, NV12, Roh09, ZFG20]. Svein [All98]. Swamp [JLMR96]. Sweden [Kat07]. Swedish [BHW15, EU99, Sta15]. Swordfish [CP16, LLA00]. Sydney [SPC18]. System [Abd95, CMM88, Gar88, GW92, GMM^{+}15, Ham98, Her06, Hua15, KSPT^{+}17, LCB92, OZAo96, OAA04, SDS12, TL12, Tho12, Yam95, Arn93]. Systems [DBR^{+}19, GOFR88, KC95, Neh93, PFN13, Rud88]. Systems/Economic [GOFR88].

T [Pon89, WBC^{+}13]. Table [ALG^{+}19]. Tadashi [Neh93]. Tail [Cox07]. Tait [ACH^{+}12]. Taiwan [mHS06, hS98]. Taking [Smi89]. Tampa [GMS97a].

Tanzania [MUK23]. Targeted [Kin03]. Targeting [CH25, PWL05a, TH12]. Targets [GMM^{+}15]. Tariffs [Kin03]. Tax [Ber23, CN00, Egg12, Jen98]. Taxes [And86, Han07b, WT05]. Taxing [Gra96, Joh95, Joh96a]. Technical [AGR19, ARS14, BC13, ESC18, Fel02, GHM06, GR11, KBBN23, Kva16, OA10, Øst99, PSL03, SL98, TBA95, VH11]. Techniques [MP99]. Technological [Cha00]. Technology [DK96, FJ23, Kum02, RAH17, Sam92]. Temperature [MKN23]. Temporal [CN00]. Temporary [KNYA23]. Ten [Tow06]. Tend [Mat96]. Tenure [RHJ92a]. Terkla [Geo87, Geo96]. Term [ADK^{+}11]. Terms [SDQ21]. Terrestrial [Kah87]. Territorial [Gon96]. Testing [GHB93, GSA93]. Tests [AGH04]. Tetreault [Cha95]. Texas [FGMW03]. Thalassorama [Pas93]. their [EVH05, MC92, WB09]. Themselves [MUWB23]. Theoretic [Ken87, NP97]. Theoretical [And85b, WG03]. Theory [CM17, DS88, Jen02, Li98b, Pas07, PFN13, Smi89, Swa94, Wil85, Gat96]. There [ABW21, LAM21]. Thoughts [Wil07]. Three [BG00, DAAY^{+}20, PH20, Sai88]. Thriving [ABW21]. Thy [PPQ^{+}19]. Tilapia [AGR19, NL09, YM02]. Tim [GPC^{+}13, Sch95]. Timber [Loo88]. Time [And86, Arn92, BHW15, GA95, KW86, KBK^{+}21, MS98, SJQ21]. Time-Dependent [KW86]. Tisdell [Ken06, Tyr01]. Too [FPA13, Mat96]. Tool [KG97]. Toothfish [HHP10]. Topical [SCS06]. Total [LL94, Whi93]. Tourism [ABW21, TgKtC99, Tyr01]. Toxic [MJ20, MFF^{+}23, WMB95]. Traceability [MWH^{+}18]. Tracing [HBW22]. Tradable [AMM98, AS05]. Trade [Asc97b, CMGB07a, EQJ94, KP08, PGB00, Sch97, WTW^{+}19, YAA20, SW15, Geo96]. Traded [WB86]. Tradeoffs [PTSB99]. Trading [DLL17, HW12, JLT19]. Traditional [AMK06, Dol88, HLH^{+}14, JKPHS12, Rud88]. Tragedy [ALS^{+}18]. Trail [PS03]. Training [GLL^{+}10]. Transaction [AKP98a]. Transboundary [Arm94, FM96, MP97, PY89]. Transfer [MPTB22, ZS09]. Transferable [Arn93, BP06, DG00, GWB94a, GN88, Gra96, Joh95, Joh96a, PAC^{+}22, Sei93]. Transfers [MUK23]. Transition [GC16, HKBH05, LCB92, PR99, SJQ21]. Transmission [FPdR20, GLN^{+}23, Jaf04, NK18]. Transport [JKP95]. Travel [And15c, LSH^{+}18, Loo01, SB85, Smi89]. Travels [SHK^{+}13]. Trawl [CG85, ESOF03, GHM06, HBAY^{+}16, HS86, KBBN23, SYN^{+}19, VOJ^{+}01]. Trawler [EVH05, HF07, Hut14, VD05]. Trawlers [DR04, SW18]. Treatment [DF16]. Trend [Kva16]. Trevor [Roh03a]. Triazine [GKK^{+}88]. Tribute [And15b, MDC^{+}15]. Trip [LL13, MPTB22, WDJK20, WHL^{+}18]. Trips [PS03]. Trond [Arm13]. Tropical [dLS92, PC94]. Trouble [SMRK85]. Troubled [Geo96]. Trout [AYJO^{+}19a, NSL^{+}07, VSSH05]. Trust [Geo06]. Truth [Poo97]. Tsukayama [Sai88]. Tsukiji [Tok18]. Tucker [PH99, PS03]. Tuna [ALP^{+}21, BCHH00, BB06, BM96, BDCDeS00, CL89, CN95, CGP03a, DBP00, GN88, GLN^{+}23, HH86, Her88, Her95, KW86, MKN23, OLM^{+}21, PD02, ST87, ST89, Sha11, Tok18, Wau88]. TURFs [CUW07a]. Turner [Gri00]. Turtle [CP16, MS11]. Turtles [Tis86]. Twenty [SRS^{+}13]. Two [FP25, MU16, STY20, Sum95]. Two-Pronged [STY20]. Two-Stage [Sum95]. Type [HMMhL92]. Types [BG00, Gat84, Pro84].

U.S. [ABOS12, tCT93, Edw01, Edw05, EQJ94, GH96, HG89, KNH93, LS99, LLA00, MJ11, NQ06, Qua06, SH04, SW06, Ter07, WR97, WJ12]. U.S.A. [CAAT10, KP08]. UK [FJ08, GW92, MF02, RGA07]. Unbreaded [SDS12]. Uncertain [MP85, Sis84]. Uncertainty [And84, And86, Ber23, Boc84, Dup93, FHJ87, Fro84, Gat84, Han84, HH09, Jin94, Koe84, LZ87, MS01, OG03, PB84, Pro84, SW15, Smi85, Spu85, Yoh84]. Underestimating [ZBP06a]. Underlying [Rud88]. Understanding [FJ23, MEA^{+}23]. Unfinished [Tow03]. Union [CL98]. United [HF01, LSW17, CT14, Gre99, Mun00, Pow04, QXH11, SDS12, THM98, YC15]. Units [LG07]. University [Ale87]. Unknowns [KHB^{+}25]. Unprocessed [Han85a]. Unregulated [Hut14]. Unsustainability [ATR^{+}23]. Untangling [AL13, GM07, SCK^{+}15]. Upper [HTTA92, WT05]. Upstream [mHS06, KP01, ALG^{+}19]. Uptake [FP25]. Upton [Gri93]. Upwelling [Sai88]. Urchin [RW00, SW04]. Use [And04, AKV^{+}17, GKK^{+}88, Gil88, Gon96, HF01, HJ97, Ken06, Lin13, MPE10, PB94, Par87, PCH^{+}13, Ter07, WHL^{+}18]. Used [Øst99]. User [ACDL21]. Using [AY12, AED11, Ber23, Boy93, CNAB09, FP25, GMTW19, HBW22, Her93, Kli87, PC94, PCAIGO^{+}14, Qua06, Row04, SHL^{+}20a, SDS12, Wal06, WA23, WDJK20, WHL^{+}18, MSBh95, VD05, Ano93b]. Utility [BMS89b, GMS97a, HBB07a, LSH^{+}18, LRCL25, PMT99a, Sch98, TS02, MSBh95]. Utilization [AMM98, And02b, CGS05, DSBT97, Fel02, MVRRP21].

Valencia [Mar86a]. Validity [Whi93]. Valuable [Gri93]. Valuation [CL10, GPC^{+}13, HGV^{+}99, HLM20, cHPZ07, LL94, MPTB22, Maz04, MN18, NDW22, Whi93, WBC^{+}13]. Valuations [AKKM86]. Value [AL13, ABW21, Ano89a, BK89, CLL22, GOG00, Gra96, HLH^{+}14, Hup89, Jaf04, JA89, Joh95, Joh96a, LLLV16, Li98b, Li09, Lip04, LGOS21, MMCM13, Mat09, MOOT23, MFF^{+}23, MPE10, PECY25, RGA07, RRS88, STNY23, SB85, Sar09, SK21, Smi23, Sta15, TBA95, Tro01, Tur96]. Value-based [Tur96]. Values [AKV^{+}17, BS06a, BC13, DBR^{+}19, HBCP09, JRBH06, LL94, PHCK16, PPQ^{+}19, Whi93, WHL^{+}18]. Valuing [CMLC16, DLL17, LLR10, NSV^{+}22, PKLB09, PCH^{+}13, SS16, WDLH11, Loo01]. Variability [KJE03, PPV96, ZT22]. Variable [Han88]. Variables [HBW22, Hof04]. Variation [Str04, TS02]. Variations [SB85]. Varies [And89b]. Varying [AH93]. Vector [CL09]. Versus [CDNS17, MD17, BSA^{+}22, FP25, LRCL25, Pas07, SK21]. Vertical [And96]. Vessel [And99, BG00, CDTW95, GCG98, HCT96, HF07, Hut14, KTR00, PTDQ^{+}19, PA20, TPPT11, Wal06, WT10, Fel09]. Vessels [And04, Dol88, hS98, YGB97]. Viability [AAO^{+}09, STY20]. Viable [SRTS86]. Victoria [BBI05a]. Vietnam [vBDSD10, KBBN23, NNNJ22, NSV^{+}22]. Vietnamese [PFN13, PFNV21].

View [TB01]. Villages [Cav23]. Virgin [SHL^{+}20a]. Virginia [BS89]. Virtual [LG07]. Visitation [LLLV16]. Visitors [LL94, WHL^{+}18]. Visits [And15c]. Vol [Gat86]. Volatility [AMO19, DO14, FPdR20, Mer23, OS08, ZKD^{+}14]. Volume [Ano87e, Ano95a, Ano96a, Ano97a, Ano98a, Ano99a, Ano00a, Ano01a, Ano02a, Ano03a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a]. Volumes [Tyr01]. Voracera [CGS05]. Voting [Abr85]. vs [And86, CUW07a, Fou02, SS01, ZS16]. Vulnerability [CLSA22].

W [Gat86, Gat96, Van96]. W. [Pol86]. Wagging [Cox07]. Walters [Wil87]. Ward [Loo01, Roh03a]. Warming [mHS06, Egg12]. Warmwater [AYSN17]. Warren [Gat86]. Wars [Geo06]. Waste [Jin94]. Water [BC13, BMS89a, Chu86, FKAvR10, Fre95, HTTA92, Lip04, NDW22, PS03]. Waterfront [BC13]. Waters [EAÖH25, Geo96, Tis86]. Waugh [And85b, Com89]. Way [AK97, HBW22]. Wealth [CNAB09]. Wealth-based [CNAB09]. Webster [ALS^{+}18]. Weeks [Kin93]. Weibull [Gat96]. Weight [BSA^{+}22, MS98]. Weight-Dependent [MS98]. Weighting [LHCL15]. Welfare [AYJO^{+}19a, CHLH03, FJ03, GKK^{+}88, HKS04, HNCS12, KT94, PTE04, STNY23, WCFZS03]. West [LSJN^{+}24, WJ12, SYN^{+}19]. Western [BCHH00, BBA^{+}10, CGP03a, DNA^{+}12, Ken92, OLM^{+}21, Sai97]. Westview [All98]. Wet [Bur92]. Wetland [PIH14, Swa94]. Wetlands [AED11, PIH14, SB87]. Whale [Dow93, KT94, LL94]. Whatever [YC15]. Where [KGW22, O’H20, PTSB99]. Whitefish [AGH04, BAYN16, Nie05]. Whitehead [GPC^{+}13]. Whiting [BER^{+}14, Mat10, SE94, Wil09, WCA^{+}10]. Whitmarsh [And87a]. Who [CS13, PTSB99, Ret94]. Whole [HG89]. Wholesale [WW93]. Width [PCH^{+}13]. Wild [AL13, ALG^{+}19, BLS16, BAYN16, BHH^{+}12, FPdR20, HJKP03, NL09, PDS^{+}21, RAC^{+}17, RNFH22]. Wild-Caught [PDS^{+}21]. Wild-Harvest [HJKP03]. Wildlife [Whi93]. Will [AR01]. Will. [Fel94]. Williams [ACH^{+}12]. Willingness [BHJ03, Cav23, FS15, HM11, JRBH06, KTR00, NNNJ22, PgK09, PIH14, WCH01]. Wind [HBW22]. Wire [Hol00]. Wisconsin [CBI92]. Within [Fro84, KH06]. Within-Season [KH06]. Without [CS86]. Workshop [And87c]. World [And15a, Arm13, Bir86, Cha95, Gud98, Han15, MN18, NK18, Rob93, Sch07, Sla07, SHK^{+}13, TB01]. Worldwide [And87c, Gul89, Pon88, Pon89, RC89]. Worth [PgK09]. Would [Han98]. Wreckfish [GWB94a, YC15]. Wrong [Row04]. WTP [RLA10, SAR05]. Wulff [Gri00].

Yamamoto [Neh93]. Year [ASS18, Cha09]. Yearbook [Kna95]. Years [ASS^{+}14, Egg06, SRS^{+}13]. Yellowfin [CN95]. Yield [EOA^{+}17, SV16]. Yields [Mai10]. York [Ale87, Con85, BK89].

Zealand [AY12, BS06b, CMM88, LCB92, MC92, Tow03]. Zone [SYN^{+}19, Sta87]. Zones [MS01, WJP^{+}02, ZS09]. Zoning [Hol04].

References

Asche:2000:FRR

[AA00] Frank Asche and Kristin Aarland. Fishermen’s response to revenue changes: The Norwegian coastal mackerel fishery. Marine Resource Economics, 15(1):67–71, Spring 2000. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abbott:2014:SGP

[AAC^{+}14] Joshua Abbott, James L. Anderson, Liam Campling, Rögnvaldur Hannesson, Elizabeth Havice, M. Susan Lozier, Martin D. Smith, and Michael J. Wilberg. Steering the global partnership for oceans. Marine Resource Economics, 29(1):1–16, March 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Albers:2021:MAP

[AAC^{+}21] Heidi J. Albers, Madison Ashworth, Tabaré Capitán, Róger Madrigal-Ballestero, and Louis Preonas. MPAs and aspatial policies in artisanal fisheries. Marine Resource Economics, 36 (4):341–367, October 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:2010:IDR

[AAF10] Peder Andersen, Jesper Levring Andersen, and Hans Frost. ITQs in Denmark and resource rent gains. Marine Resource Economics, 25(1):11–22, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Adeogun:2009:EVS

[AAO^{+}09] O. A. Adeogun, P. O. Abohweyere, H. K. Ogunbadejo, Alhaji Tanko, and L. Jim-saiki. Economic viability of small-scale marine capture fisheries in the Bonny Area, Rivers State, Nigeria. Marine Resource Economics, 24(2):195–203, ???? 2009. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2008:BM

[AAR^{+}08] Frank Asche, Trude B. Andersen, Kristin H. Roll, Sigbjørn Tveterås, Atle Oglend, Marius Sikveland, Frank Asche, Atle Oglend, Marius Sikveland, Walter R. Keithly, Jr., and Pawan Poudel. Back matter. Marine Resource Economics, 23(4):??, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abolofia:2017:CLQ

[AAW^{+}17] Jay Abolofia, Frank Asche, James E. Wilen, Jay Abolofia, Frank Asche, James E. Wilen, Joshua K. Abbott, Ola Flaaten, Knut Heen, Thórólfur Matthíasson, Jay Abolofia, Frank Asche, James E. Wilen, Ola Flaaten, Knut Heen, Thórólfur Matthíasson, Domenico Carlucci, Biagia De Devitiis, Gianluca Nardone, and Fabio Gaetano Santeramo. The cost of lice: Quantifying the impacts of parasitic sea lice on farmed salmon. Marine Resource Economics, 32(3):329–349, July 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1993:CAM

[AB93] James L. Anderson and Sofia U. Bettencourt. A conjoint approach to model product preferences: The New England market for fresh and frozen salmon. Marine Resource Economics, 8(1): 31–49, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abbott:2015:FRH

[Abb15] Joshua K. Abbott. Fighting over a red herring: The role of economics in recreational-commercial allocation disputes. Marine Resource Economics, 30(1):1–20, January 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abbott:2017:EBS

[Abb17] Joshua K. Abbott. Editorial: Building on strength. Marine Resource Economics, 32(3):243–248, July 2017. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abdullah:1995:TIF

[Abd95] Nik Mustapha R. Abdullah. Towards an integrated fisheries information system in Malaysia. Marine Resource Economics, 10(3):312–320, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2012:USM

[ABOS12] Frank Asche, Lori S. Bennear, Atle Oglend, and Martin D. Smith. U.S. shrimp market integration. Marine Resource Economics, 27(2):181–192, June 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abrams:1985:EFJ

[Abr85] Burton A. Abrams. The expansion of fishery jurisdiction: Fishery interests and Congressional voting. Marine Resource Economics, 2(2):143–152, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1993:BM

[ABS^{+}93] James L. Anderson, Sofia U. Bettencourt, Juan Carlos Seijo, Cathy Roheim Wessells, James E. Wilen, James L. Anderson, Sofia U. Bettencourt, Kjell G. Salvanes, Cathy Roheim Wessells, and James E. Wilen. Back matter. Marine Resource Economics, 8(1):??, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2003:RPD

[ABS03] Frank Asche, Trond Bjørndal, and Elin H. Sissener. Relative productivity development in salmon aquaculture. Marine Resource Economics, 18(2):205–210, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersson:2021:LFT

[ABW21] Anna Andersson, Johan Blomquist, and Staffan Waldo. Local fisheries and thriving harbors: Is there a value for the tourism sector? Marine Resource Economics, 36(2):111–131, April 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Armstrong:1997:JFE

[AC97] Claire W. Armstrong and Derek J. Clark. Just fishing? Equity and efficiency in fisheries management regimes. Marine Resource Economics, 12(3):203–220, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Albers:2021:PSS

[ACDL21] Heidi J. Albers, Carlos Chávez, Jorge Dresdner, and Mauricio Leiva. Prospects for small-scale aquaculture in Chile: User rights and locations. Marine Resource Economics, 36(4):389– 410, October 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2012:BRH

[ACH^{+}12] Lee G. Anderson, Benaissa Chidmi, Terry Hanson, Giap Nguyen, Kehar Singh, Madan M. Dey, Prasanna Surathkal, Kobi Abayomi, Tracy Yandle, Benaissa Chidmi, Terry Hanson, Giap Nguyen, Kehar Singh, Madan M. Dey, Prasanna Surathkal, Kobi Abayomi, and Tracy Yandle. Book review: Handbook of Marine Fisheries Conservation and Management. Edited by R. Quentin Grafton, Ray Hilborn, Dale Squires, Maree Tait, and Meryl Williams. Marine Resource Economics, 27(4):397–399, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2011:DGA

[ADG^{+}11] Frank Asche, Roy E. Dahl, Daniel V. Gordon, Trine Trollvik, and Paul Aandahl. Demand growth for Atlantic salmon: The EU and French markets. Marine Resource Economics, 26(4): 255–265, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asheim:2011:PBD

[ADK^{+}11] Leif Jarle Asheim, Roy Endre Dahl, Subal C. Kumbhakar, Atle Oglend, Ragnar Tveteras, Frank Asche, Roy E. Dahl, Daniel V. Gordon, Trine Trollvik, Paul Aandahl, Terje Vassdal, Helen Marita Sørensen Holst, Jinghua Xie, Øystein Myrland, Terje Vassdal, Helen Marita Sørensen Holst, Atle G. Guttormsen, Kristin H. Roll, Frank Asche, Roy E. Dahl, Daniel V. Gordon, Trine Trollvik, and Paul Aandahl. Are prices or biology driving the short-term supply of farmed salmon? Marine Resource Economics, 26(4):343–357, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:2018:BRE

[ADvH^{+}18] Peder Andersen, Steven J. Dundas, Roger H. von Haefen, Carol Mansfield, Kanae Tokunaga, Kazi Ali Toufique, Sami Naim Farook, Ben Belton, Kanae Tokunaga, Kazi Ali Toufique, Sami Naim Farook, Ben Belton, Steven J. Dundas, Roger H. von Haefen, and Carol Mansfield. Book review: Economics of the Marine. Modelling Natural Resources. By Karyn Morrissey. Marine Resource Economics, 33(1):117–118, January 2018. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Awondo:2011:IBR

[AED11] Sebastain N. Awondo, Kevin J. Egan, and Daryl F. Dwyer. Increasing beach recreation benefits by using wetlands to reduce contamination. Marine Resource Economics, 26(1):1–15, March 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2001:PRP

[AG01] Frank Asche and Atle G. Guttormsen. Patterns in the relative price for different sizes of farmed fish. Marine Resource Economics, 16(3):235–247, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2014:SMA

[AG14] Frank Asche and Atle G. Guttormsen. Seafood markets and aquaculture production: Special issue introduction. Marine Resource Economics, 29(4):301–304, December 2014. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2022:MOU

[AGC^{+}22] Frank Asche, Taryn Garlock, Edward Camp, Jordi Guillen, Ganesh Kumar, Ignacio Llorente, Gina Shamshak, Chris Moore, Jasmine Fuller, Ted E. Gilliland, James N. Sanchirico, J. Edward Taylor, Frank Asche, Taryn Garlock, Edward Camp, Jordi Guillen, Ganesh Kumar, Ignacio Llorente, and Gina Shamshak. Market opportunities for US aquaculture producers: The case of Branzino. Marine Resource Economics, 37(2):221–233, April 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2004:TMI

[AGH04] Frank Asche, Daniel V. Gordon, and Rögnvaldur Hannesson. Tests for market integration and the law of one price: The market for whitefish in France. Marine Resource Economics, 19 (2):195–210, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2001:BM

[AGK^{+}01] Frank Asche, Atle G. Guttormsen, Duncan Knowler, Edward B. Barbier, Ivar Strand, John C. Whitehead, William B. Clifford, Thomas J. Hoban, Frank Asche, Atle G. Guttormsen, Harald Bergland, Derek J. Clark, Pål Andreas Pedersen, John C. Whitehead, William B. Clifford, and Thomas J. Hoban. Back matter. Marine Resource Economics, 16(3):??, ???? 2001. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Alam:2019:PRT

[AGR19] Md. Akhtarul Alam, Atle G. Guttormsen, and Kristin H. Roll. Production risk and technical efficiency of tilapia aquaculture in Bangladesh. Marine Resource Economics, 34(2):123–141, April 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2008:AOC

[AGT08] Frank Asche, Atle G. Guttormsen, and Ragnar Tveteras. Aquaculture — opportunities and challenges special issue introduction. Marine Resource Economics, 23(4):395–400, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abbott:2010:ERE

[AGYW10a] Joshua K. Abbott, Brian Garber-Yonts, and James E. Wilen. Employment and remuneration effects of IFQs in the Bering Sea/Aleutian Islands crab fisheries. Marine Resource Economics, 25(4):333–354, December 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abbott:2010:E

[AGYW^{+}10b] Joshua K. Abbott, Brian Garber-Yonts, James E. Wilen, David W. Carter, Christopher Liese, Tu van Binh, Marijke D’Haese, Stijn Speelman, Luc D’Haese, Joshua K. Abbott, Brian Garber-Yonts, James E. Wilen, Matthew A. Freeman, Richard T. Woodward, Quach Thi Khanh Ngoc, and Ola Flaaten. Errata. Marine Resource Economics, 25(4):??, December 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Agnello:1993:SSM

[AH93] Richard J. Agnello and Yunqi Han. Substitute site measures in a varying parameter model with application to recreational fishing. Marine Resource Economics, 8(1):65–77, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2002:AFB

[AH02] Frank Asche and Rögnvaldur Hannesson. Allocation of fish between markets and product forms. Marine Resource Economics, 17(3):225–238, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2009:BM

[AHT^{+}09] Frank Asche, Håvard Hansen, Ragnar Tveteras, Sigbjørn Tveterås, Stephen Kasperski, Robert Wieland, Carl Lian, Rajesh Singh, Quinn Weninger, Frank Asche, Håvard Hansen, Ragnar Tveteras, Sigbjørn Tveterås, Stephen Kasperski, Robert Wieland, Carl Lian, Rajesh Singh, and Quinn Weninger. Back matter. Marine Resource Economics, 24(4):??, ???? 2009. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2009:SDC

[AHTT09] Frank Asche, Håvard Hansen, Ragnar Tveteras, and Sigbjørn Tveterås. The salmon disease crisis in Chile. Marine Resource Economics, 24(4):405–411, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abdullah:1997:FMA

[AK97] Nik Mustapha R. Abdullah and K. Kuperan. Fisheries management in Asia: The way forward. Marine Resource Economics, 12 (4):345–353, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ahi:2020:ANA

[AK20] Julide Ceren Ahi and Gorm Kipperberg. Attribute nonattendance in environmental discrete choice experiments: The impact of including an employment attribute. Marine Resource Economics, 35(3):201–218, July 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:1997:BM

[AKG^{+}97] Frank Asche, Barry Kaufmann, Gerry Geen, Kjell G. Salvanes, Don J. Devoretz, Frank Asche, Porter Hoagland, Di Jin, Kjell G. Salvanes, and Don J. Devoretz. Back matter. Marine Resource Economics, 12(1):??, Spring 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Assaf:1986:NVA

[AKKM86] George B. Assaf, Brent G. Kroetch, Brent C. Kroetch, and Subodh C. Mathur. Nonmarket valuations of accidental oil spills: a survey of economic and legal principles. Marine Resource Economics, 2(3):211–237, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:1993:BM

[AKN^{+}93] Ragnar Arnason, Henry W. Kinnucan, Robert G. Nelson, Johanis Hiariey, Anthony Scott, Ragnar Arnason, Diane P. Dupont, Henry W. Kinnucan, Robert G. Nelson, and Johanis Hiariey. Back matter. Marine Resource Economics, 8(3):??, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abdullah:1998:TCF

[AKP98a] Nik Mustapha Raja Abdullah, K. Kuperan, and Robert S. Pomeroy. Transaction costs and fisheries co-management. Marine Resource Economics, 13(2):103–114, Summer 1998. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abdullah:1998:BM

[AKP^{+}98b] Nik Mustapha Raja Abdullah, K. Kuperan, Robert S. Pomeroy, Barry Kaufmann, Gerry Geen, Adesoji Adelaja, Bonnie Mc-Cay, Julia Menzo, Adesoji Adelaja, Bonnie McCay, Julia Menzo, Dhazn Gillig, Oral Capps, Jr., Wade L. Griffin, Nik Mustapha Raja Abdullah, K. Kuperan, and Robert S. Pomeroy. Back matter. Marine Resource Economics, 13(2):??, Summer 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Aung:2022:DAF

[AKT^{+}22] Yee Mon Aung, Ling Yee Khor, Nhuong Tran, Michael Akester, and Manfred Zeller. A disaggregated analysis of fish demand in Myanmar. Marine Resource Economics, 37(4):467–490, October 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Armstrong:2017:UNU

[AKV^{+}17] Claire W. Armstrong, Viktoria Kahui, Godwin K. Vondolia, Margrethe Aanesen, and Mikołaj Czajkowski. Use and non-use values in an applied bioeconomic model of fisheries and habitat connections. Marine Resource Economics, 32(4):351–369, October 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2013:URV

[AL13] Leif E. Anderson and S. Todd Lee. Untangling the recreational value of wild and hatchery salmon. Marine Resource Economics, 28(2):175–197, June 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ardini:2018:DIU

[AL18] Greg Ardini and Min-Yang Lee. Do IFQs in the US Atlantic sea scallop fishery impact price and size? Marine Resource Economics, 33(3):263–288, July 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Alexander:1987:GPN

[Ale87] Lewis M. Alexander. Book review: Giulio Pontecorvo, The New Order of the Oceans: The Advent of a Managed Environment (New York: Columbian University Press, 1986). Marine Resource Economics, 4(2):145–149, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2019:BRU

[ALG^{+}19] Frank Asche, Sterenn Lucas, Stéphane Gouin, Marie Lesueur, Md. Akhtarul Alam, Atle G. Guttormsen, Kristin H. Roll, Sterenn Lucas, Stéphane Gouin, Marie Lesueur, Md. Akhtarul Alam, Atle G. Guttormsen, Kristin H. Roll, and Stefano Mainardi. Book review: Upstream: Searching for Wild Salmon, from River to Table. By Langdon Cook. Marine Resource Economics, 34(2):197–198, April 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ali:2011:AGA

[Ali11] Ghulam Ali. ARCH, GARCH, and ARMAX models for forecasting pathogen indicators and advisories at marine recreational sites. Marine Resource Economics, 26(3):211–224, September 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1994:BM

[ALL^{+}94] Lee G. Anderson, John B. Loomis, Douglas M. Larson, Ragnar Arnason, John B. Loomis, Douglas M. Larson, Lee G. Anderson, and Ragnar Arnason. Back matter. Marine Resource Economics, 9(3):??, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Allen:1995:BRC

[All95] P. Geoffrey Allen. Book review: Curtis M. Jolly and Howard A. Clonts, Economics of Aquaculture. Marine Resource Economics, 10(4):407–408, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Allen:1998:BRC

[All98] P. Geoffrey Allen. Book review: Connor Bailey, Svein Jentoft, and Peter Sinclair, Aquaculture Development: Social Dimensions of an Emerging Industry, Boulder CO: Westview Press, 1996, pp. 287, $49.85, £36.95, CAN$72.00, cloth. ISBN 0-8133-8942-9. Marine Resource Economics, 13(1):87–88, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Akpalu:2021:OAA

[ALP^{+}21] Wisdom Akpalu, Andreas Leibbrandt, Sergio Puerto, Maria Alejandra Vélez, Thuy Thi Thanh Pham, Ola Flaaten, Luong Trong Nguyen, Nghiep Ke Vu, Carlos Chávez, Håkan Eggert, Matthew Reimer, Andreas Leibbrandt, Sergio Puerto, Maria Alejandra Vélez, Jorge Marco, Diego Valderrama, Mario Rueda, Maykol Rodríguez-Prieto, Heidi J. Albers, Madison Ashworth, Tabaré Capitán, Róger Madrigal-Ballestero, and Louis Preonas. Optimal allocation of anchovy stocks as baitfish for tuna and as food for local communities in developing coastal countries. Marine Resource Economics, 36(4):439–461, October 2021. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abbott:2018:BRB

[ALS^{+}18] Joshua K. Abbott, Christopher G. Leggett, Nora Scherer, Timothy C. Haab, Ryan Bailey, Jason P. Landrum, Adam Domanski, John C. Whitehead, Tim Haab, Sherry L. Larkin, John B. Loomis, Sergio Alvarez, Andrew Ropicki, Peggy Schrobback, Sean Pascoe, Louisa Coglan, Christopher G. Leggett, Nora Scherer, Timothy C. Haab, Ryan Bailey, Jason P. Landrum, Adam Domanski, Ling Huang, Subhash Ray, Kathleen Segerson, and John Walden. Book review: Beyond the Tragedy in Global Fisheries. By D. G. Webster. Marine Resource Economics, 33 (2):203–208, April 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Alvarez:2022:DBO

[Álv22] Antonio Álvarez. Does boat order in the auction affect the price in fish markets? Hedonic pricing for hake in Northern Spain. Marine Resource Economics, 37(3):305–323, July 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:2000:NSS

[AMA00] Ragnar Arnason, Gylfi Magnusson, and Sveinn Agnarsson. The Norwegian spring-spawning herring fishery: a stylized game model. Marine Resource Economics, 15(4):293–319, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Al-Marshudi:2006:SES

[AMK06] Ahmed Salim Al-Marshudi and Hemesiri Kotagama. Socioeconomic structure and performance of traditional fishermen in the Sultanate of Oman. Marine Resource Economics, 21(2):221– 230, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Adelaja:1998:MSC

[AMM98] Adesoji Adelaja, Bonnie McCay, and Julia Menzo. Market share, capacity utilization, resource conservation, and tradable quotas. Marine Resource Economics, 13(2):115–134, Summer 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2019:CCS

[AMO19] Frank Asche, Bård Misund, and Atle Oglend. The case and cause of salmon price volatility. Marine Resource Economics, 34 (1):23–38, January 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ussif:2002:DCF

[aMUSS02] Al amin M. Ussif, Leif K. Sandal, and Stein I. Steinshamn. On the dynamics of commercial fishing and parameter identification.

Marine Resource Economics, 17(1):35–46, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Aillery:1999:SRC

[AMW^{+}99] Marcel Aillery, Michael R. Moore, Marca Weinberg, Glenn Schaible, and Noel Gollehon. Salmon recovery in the Columbia River Basin: Analysis of measures affecting agriculture. Marine Resource Economics, 14(1):15–40, Spring 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1984:UFM

[And84] Lee G. Anderson. Uncertainty in the fisheries management process. Marine Resource Economics, 1(1):77–87, ???? 1984. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1985:MIB

[And85a] James L. Anderson. Market interactions between aquaculture and the common-property commercial fishery. Marine Resource Economics, 2(1):1–24, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1985:GWF

[And85b] Lee G. Anderson. Geoffrey waugh, fisheries management: Theoretical developments and contemporary applications. Marine Resource Economics, 2(1):109–113, ???? 1985. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1986:TVQ

[And86] Eric E. Anderson. Taxes vs. quotas for regulating fisheries under uncertainty: a hybrid discrete-time continuous-time model. Marine Resource Economics, 3(3):183–207, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:1987:BRS

[And87a] Peder Andersen. Book review: Stephen Cunningham, Michael R. Dunn, and David Whitmarsh, Fisheries Economics: an Introduction. Marine Resource Economics, 4(1):75–78, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1987:MAP

[And87b] Lee G. Anderson. A management agency perspective of the economics of fisheries regulation. Marine Resource Economics, 4(2):123–131, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1987:NMF

[And87c] Lee G. Anderson. Nina mollett, fishery access control problems worldwide: Proceedings of the workshop on management options for the North Pacific longline fisheries, Alaska Sea Grant report no. 86-4. Marine Resource Economics, 4(4):315–318, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1988:REC

[And88] Eric E. Anderson. Relative efficiency of charges and quantity controls in fisheries with continuous stock growth and periodically fixed instrument levels. Marine Resource Economics, 5 (3):215–230, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1989:EIS

[And89a] Lee G. Anderson. Enforcement issues in selecting fisheries management policy. Marine Resource Economics, 6(3):261–277, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1989:OII

[And89b] Lee G. Anderson. Optimal intra- and interseasonal harvesting strategies when price varies with individual size. Marine Resource Economics, 6(2):145–162, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1992:SMD

[And92] James L. Anderson. Salmon market dynamics. Marine Resource Economics, 7(1):87–88, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1994:CSF

[And94a] Eric E. Anderson. Crop sharing in the fishery and industry equilibrium: Comment. Marine Resource Economics, 9(1):67– 72, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1994:EAH

[And94b] Lee G. Anderson. An economic analysis of highgrading in ITQ fisheries regulation programs. Marine Resource Economics, 9 (3):209–226, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1994:NED

[And94c] Lee G. Anderson. A note on the economics of discards. Marine Resource Economics, 9(2):183–186, Summer 1994. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1996:BRI

[And96] Lee G. Anderson. Book review: Ishak Haji Omar, Market Power, Vertical Linkages, and Government Policy: The Fish Industry in Peninsular Malaysia. Marine Resource Economics, 11(3):220–222, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1999:MVB

[And99] Lee G. Anderson. The microeconomics of vessel behavior: a detailed short-run analysis of the effects of regulation. Marine Resource Economics, 14(2):129–150, Summer 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2002:AFW

[And02a] James L. Anderson. Aquaculture and the future: Why fisheries economists should care. Marine Resource Economics, 17(2):133– 151, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2002:CUS

[And02b] Lee G. Anderson. A comparison of the utilization of stocks with patchy distribution and migration under open access and marine reserves: an extended analysis. Marine Resource Economics, 17 (4):269–289, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2002:MAF

[And02c] Lee G. Anderson. A microeconomic analysis of the formation and potential reorganization of AFA coops. Marine Resource Economics, 17(3):207–224, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2004:OAF

[And04] Lee G. Anderson. Open-access fishery performance when vessels use goal achievement behavior. Marine Resource Economics, 19 (4):439–458, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:2005:ISM

[And05] Jesper Levring Andersen. The inclusion of stocks in multispecies fisheries: The case of Danish seiners. Marine Resource Economics, 20(2):163–184, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2007:DCA

[And07] Lee G. Anderson. Does capacity analysis help us meet fishery policy and management objectives? Comments. Marine Resource Economics, 22(1):89–93, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2010:SII

[And10] James L. Anderson. Special issue introduction: Selected papers from the North American Association of Fisheries Economists (NAAFE) Forum, Newport, RI, May 17–20, 2009. Marine Resource Economics, 25(1):??, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2012:E

[And12] James L. Anderson. Editorial. Marine Resource Economics, 27(1):1–2, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:2013:FEF

[And13] Peder Andersen. Fisheries economics and fisheries management: a reflective note in honor of Rögnvaldur Hannesson. Marine Resource Economics, 28(4):351–359, December 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2015:ABE

[And15a] Lee G. Anderson. The application of basic economic principles to real-world fisheries management and regulation. Marine Resource Economics, 30(3):235–249, July 2015. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2015:TJA

[And15b] Lee G. Anderson. Tribute to James Arthur Crutchfield. Marine Resource Economics, 30(4):435–437, October 2015. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Androkovich:2015:RVA

[And15c] Robert A. Androkovich. Recreational visits to the Adam’s River during the annual sockeye run: a travel cost analysis. Marine Resource Economics, 30(1):35–49, January 2015. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1984:FMa

[Ano84a] Anonymous. Front matter. Marine Resource Economics, 1(1): ??, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1984:FMb

[Ano84b] Anonymous. Front matter. Marine Resource Economics, 1(2): ??, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1985:FMa

[Ano85a] Anonymous. Front matter. Marine Resource Economics, 1(3): ??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1985:FMb

[Ano85b] Anonymous. Front matter. Marine Resource Economics, 1(4): ??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1985:FMc

[Ano85c] Anonymous. Front matter. Marine Resource Economics, 2(1): ??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1985:FMd

[Ano85d] Anonymous. Front matter. Marine Resource Economics, 2(2): ??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1986:FMa

[Ano86a] Anonymous. Front matter. Marine Resource Economics, 2(3): ??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1986:FMb

[Ano86b] Anonymous. Front matter. Marine Resource Economics, 2(4): ??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1986:FMc

[Ano86c] Anonymous. Front matter. Marine Resource Economics, 3(1): ??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1986:FMd

[Ano86d] Anonymous. Front matter. Marine Resource Economics, 3(2): ??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1986:FMe

[Ano86e] Anonymous. Front matter. Marine Resource Economics, 3(3): ??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1986:FMf

[Ano86f] Anonymous. Front matter. Marine Resource Economics, 3(4): ??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1987:FMa

[Ano87a] Anonymous. Front matter. Marine Resource Economics, 4(1): ??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1987:FMb

[Ano87b] Anonymous. Front matter. Marine Resource Economics, 4(2): ??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1987:FMc

[Ano87c] Anonymous. Front matter. Marine Resource Economics, 4(3): ??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1987:FMd

[Ano87d] Anonymous. Front matter. Marine Resource Economics, 4(4): ??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1987:IV

[Ano87e] Anonymous. [Index to volume 4, 1987]. Marine Resource Economics, 4(4):329–334, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1988:FMa

[Ano88a] Anonymous. Front matter. Marine Resource Economics, 5(1): ??, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1988:FMb

[Ano88b] Anonymous. Front matter. Marine Resource Economics, 5(2): ??, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1988:FMc

[Ano88c] Anonymous. Front matter. Marine Resource Economics, 5(3): ??, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1988:FMd

[Ano88d] Anonymous. Front matter. Marine Resource Economics, 5(4): ??, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1989:EMV

[Ano89a] Anonymous. Errata: Measuring the value of fish to anglers: Application to Central California anadromous species. Marine Resource Economics, 6(4):??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1989:FMa

[Ano89b] Anonymous. Front matter. Marine Resource Economics, 6(1): ??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1989:FMb

[Ano89c] Anonymous. Front matter. Marine Resource Economics, 6(2): ??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1989:FMc

[Ano89d] Anonymous. Front matter. Marine Resource Economics, 6(3): ??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1989:FMd

[Ano89e] Anonymous. Front matter. Marine Resource Economics, 6(4): ??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1992:FMa

[Ano92a] Anonymous. Front matter. Marine Resource Economics, 7(1): ??, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1992:FMb

[Ano92b] Anonymous. Front matter. Marine Resource Economics, 7(2): ??, Summer 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1992:FMc

[Ano92c] Anonymous. Front matter. Marine Resource Economics, 7(3): ??, Fall 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1992:FMd

[Ano92d] Anonymous. Front matter. Marine Resource Economics, 7(4): ??, Winter 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1993:ESI

[Ano93a] Anonymous. Editorial: Special issues of marine resource economics. Marine Resource Economics, 8(3):??, Fall 1993. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1993:EUI

[Ano93b] Anonymous. Erratum: Using an International Econometric Model to Forecast Alaska Salmon Revenues. Marine Resource Economics, 8(4):??, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [Her93].

Anonymous:1993:FMa

[Ano93c] Anonymous. Front matter. Marine Resource Economics, 8(1): ??, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1993:FMb

[Ano93d] Anonymous. Front matter. Marine Resource Economics, 8(2): ??, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1993:FMc

[Ano93e] Anonymous. Front matter. Marine Resource Economics, 8(3): ??, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1993:FMd

[Ano93f] Anonymous. Front matter. Marine Resource Economics, 8(4): ??, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1994:FMa

[Ano94a] Anonymous. Front matter. Marine Resource Economics, 9(1): ??, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1994:FMb

[Ano94b] Anonymous. Front matter. Marine Resource Economics, 9(2): ??, Summer 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1994:FMc

[Ano94c] Anonymous. Front matter. Marine Resource Economics, 9(3): ??, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1994:FMd

[Ano94d] Anonymous. Front matter. Marine Resource Economics, 9(4): ??, Winter 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1995:MRE

[Ano95a] Anonymous. Marine Resource Economics: Contents of volume 10. Marine Resource Economics, 10(4):416–418, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1995:EEA

[Ano95b] Anonymous. Errata: an economic analysis of highgrading in ITQ fisheries regulation programs. Marine Resource Economics, 10 (1):99–100, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1995:FMa

[Ano95c] Anonymous. Front matter. Marine Resource Economics, 10(1): ??, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1995:FMb

[Ano95d] Anonymous. Front matter. Marine Resource Economics, 10(2): ??, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1995:FMc

[Ano95e] Anonymous. Front matter. Marine Resource Economics, 10 (3):??, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1995:FMd

[Ano95f] Anonymous. Front matter. Marine Resource Economics, 10(4): ??, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1995:MRJ

[Ano95g] Anonymous. MRE reviewers: July 1994 through June 1995. Marine Resource Economics, 10(4):??, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1996:MREb

[Ano96a] Anonymous. Marine Resource Economics: Contents of volume 11. Marine Resource Economics, 11(4):319–320, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1996:MREa

[Ano96b] Anonymous. Marine Resource Economics: Reviewers July 1995 through June 1996. Marine Resource Economics, 11(4):??, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1996:FMa

[Ano96c] Anonymous. Front matter. Marine Resource Economics, 11(1): ??, Spring 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1996:FMb

[Ano96d] Anonymous. Front matter. Marine Resource Economics, 11(2): ??, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1996:FMc

[Ano96e] Anonymous. Front matter. Marine Resource Economics, 11 (3):??, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1996:FMd

[Ano96f] Anonymous. Front matter. Marine Resource Economics, 11(4): ??, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1997:MREb

[Ano97a] Anonymous. Marine Resource Economics: Contents of volume 12. Marine Resource Economics, 12(4):368–369, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1997:MREa

[Ano97b] Anonymous. Marine Resource Economics: Reviewers July 1996 through June 1997. Marine Resource Economics, 12(4):??, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1997:FMa

[Ano97c] Anonymous. Front matter. Marine Resource Economics, 12(1): ??, Spring 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1997:FMb

[Ano97d] Anonymous. Front matter. Marine Resource Economics, 12(2): ??, Summer 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1997:FMc

[Ano97e] Anonymous. Front matter. Marine Resource Economics, 12 (3):??, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1997:FMd

[Ano97f] Anonymous. Front matter. Marine Resource Economics, 12(4): ??, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1998:MREb

[Ano98a] Anonymous. Marine Resource Economics: Contents of volume 13. Marine Resource Economics, 13(4):312–313, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1998:MREa

[Ano98b] Anonymous. Marine Resource Economics: Reviewers July 1997 through June 1998. Marine Resource Economics, 13(4):??, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1998:FMa

[Ano98c] Anonymous. Front matter. Marine Resource Economics, 13(1): ??, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1998:FMb

[Ano98d] Anonymous. Front matter. Marine Resource Economics, 13(2): ??, Summer 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1998:FMc

[Ano98e] Anonymous. Front matter. Marine Resource Economics, 13 (3):??, Fall 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1998:FMd

[Ano98f] Anonymous. Front matter. Marine Resource Economics, 13(4): ??, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1999:MREb

[Ano99a] Anonymous. Marine Resource Economics: Contents of volume 14. Marine Resource Economics, 14(4):388–389, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1999:MREa

[Ano99b] Anonymous. Marine Resource Economics: Reviewers July 1998 through September 1999. Marine Resource Economics, 14(4): ??, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1999:FMa

[Ano99c] Anonymous. Front matter. Marine Resource Economics, 14(1): ??, Spring 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1999:FMb

[Ano99d] Anonymous. Front matter. Marine Resource Economics, 14(2): ??, Summer 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1999:FMc

[Ano99e] Anonymous. Front matter. Marine Resource Economics, 14 (3):??, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:1999:FMd

[Ano99f] Anonymous. Front matter. Marine Resource Economics, 14(4): ??, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2000:MREb

[Ano00a] Anonymous. Marine Resource Economics: Contents of volume 15. Marine Resource Economics, 15(4):380–381, ???? 2000. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2000:MREa

[Ano00b] Anonymous. Marine Resource Economics: Reviewers 2000. Marine Resource Economics, 15(4):??, ???? 2000. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2000:FMa

[Ano00c] Anonymous. Front matter. Marine Resource Economics, 15(1): ??, Spring 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2000:FMb

[Ano00d] Anonymous. Front matter. Marine Resource Economics, 15(2): ??, Summer 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2000:FMc

[Ano00e] Anonymous. Front matter. Marine Resource Economics, 15(3): ??, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2000:FMd

[Ano00f] Anonymous. Front matter. Marine Resource Economics, 15(4): ??, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2001:MREb

[Ano01a] Anonymous. Marine Resource Economics: Contents of volume 16. Marine Resource Economics, 16(4):319–320, ???? 2001. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2001:MREa

[Ano01b] Anonymous. Marine Resource Economics: Reviewers 2001. Marine Resource Economics, 16(4):??, ???? 2001. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2001:FMa

[Ano01c] Anonymous. Front matter. Marine Resource Economics, 16(1): ??, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2001:FMb

[Ano01d] Anonymous. Front matter. Marine Resource Economics, 16(2): ??, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2001:FMc

[Ano01e] Anonymous. Front matter. Marine Resource Economics, 16(3): ??, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2001:FMd

[Ano01f] Anonymous. Front matter. Marine Resource Economics, 16(4): ??, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2002:MREb

[Ano02a] Anonymous. Marine Resource Economics: Contents of volume 17. Marine Resource Economics, 17(4):345–346, ???? 2002. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2002:MREa

[Ano02b] Anonymous. Marine Resource Economics: Reviewers 2002. Marine Resource Economics, 17(4):??, ???? 2002. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2002:FMa

[Ano02c] Anonymous. Front matter. Marine Resource Economics, 17(1): ??, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2002:FMb

[Ano02d] Anonymous. Front matter. Marine Resource Economics, 17(2): ??, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2002:FMc

[Ano02e] Anonymous. Front matter. Marine Resource Economics, 17(3): ??, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2002:FMd

[Ano02f] Anonymous. Front matter. Marine Resource Economics, 17(4): ??, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2003:MREb

[Ano03a] Anonymous. Marine Resource Economics: Contents of volume 18. Marine Resource Economics, 18(4):380–381, ???? 2003. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2003:MREa

[Ano03b] Anonymous. Marine Resource Economics: Reviewers 2003. Marine Resource Economics, 18(4):??, ???? 2003. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2003:FMa

[Ano03c] Anonymous. Front matter. Marine Resource Economics, 18(1): ??, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2003:FMb

[Ano03d] Anonymous. Front matter. Marine Resource Economics, 18(2): ??, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2003:FMc

[Ano03e] Anonymous. Front matter. Marine Resource Economics, 18(3): ??, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2003:FMd

[Ano03f] Anonymous. Front matter. Marine Resource Economics, 18(4): ??, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2004:MREb

[Ano04a] Anonymous. Marine Resource Economics: Contents of volume 19. Marine Resource Economics, 19(4):532–533, ???? 2004. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2004:MREa

[Ano04b] Anonymous. Marine Resource Economics: Reviewers 2004. Marine Resource Economics, 19(4):??, ???? 2004. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2004:FMa

[Ano04c] Anonymous. Front matter. Marine Resource Economics, 19(1): ??, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2004:FMb

[Ano04d] Anonymous. Front matter. Marine Resource Economics, 19(2): ??, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2004:FMc

[Ano04e] Anonymous. Front matter. Marine Resource Economics, 19(3): ??, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2004:FMd

[Ano04f] Anonymous. Front matter. Marine Resource Economics, 19(4): ??, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2005:MREb

[Ano05a] Anonymous. Marine Resource Economics: Contents of volume 20. Marine Resource Economics, 20(4):468–469, ???? 2005. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2005:MREa

[Ano05b] Anonymous. Marine Resource Economics: Reviewers 2005. Marine Resource Economics, 20(4):??, ???? 2005. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2005:FMa

[Ano05c] Anonymous. Front matter. Marine Resource Economics, 20(1): ??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2005:FMb

[Ano05d] Anonymous. Front matter. Marine Resource Economics, 20(2): ??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2005:FMc

[Ano05e] Anonymous. Front matter. Marine Resource Economics, 20(3): ??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2005:FMd

[Ano05f] Anonymous. Front matter. Marine Resource Economics, 20(4): ??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2006:MREb

[Ano06a] Anonymous. Marine Resource Economics: Contents of volume 21. Marine Resource Economics, 21(4):448–449, ???? 2006. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2006:MREa

[Ano06b] Anonymous. Marine Resource Economics: Reviewers 2006. Marine Resource Economics, 21(4):??, ???? 2006. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2006:DHA

[Ano06c] Anonymous. The Dr. S.-Y. Hong Award. Marine Resource Economics, 21(4):450–451, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2006:FMa

[Ano06d] Anonymous. Front matter. Marine Resource Economics, 21(1): ??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2006:FMb

[Ano06e] Anonymous. Front matter. Marine Resource Economics, 21(2): ??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2006:FMc

[Ano06f] Anonymous. Front matter. Marine Resource Economics, 21(3): ??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2006:FMd

[Ano06g] Anonymous. Front matter. Marine Resource Economics, 21(4): ??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2007:MREb

[Ano07a] Anonymous. Marine Resource Economics: Contents of volume 22. Marine Resource Economics, 22(4):446–448, ???? 2007. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2007:MREa

[Ano07b] Anonymous. Marine Resource Economics: Reviewers 2007. Marine Resource Economics, 22(4):??, ???? 2007. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2007:DHA

[Ano07c] Anonymous. The Dr. S.-Y. Hong Award. Marine Resource Economics, 22(4):450–451, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2007:FMa

[Ano07d] Anonymous. Front matter. Marine Resource Economics, 22(1): ??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2007:FMb

[Ano07e] Anonymous. Front matter. Marine Resource Economics, 22(2): ??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2007:FMc

[Ano07f] Anonymous. Front matter. Marine Resource Economics, 22(3): ??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2007:FMd

[Ano07g] Anonymous. Front matter. Marine Resource Economics, 22(4): ??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2008:MREb

[Ano08a] Anonymous. Marine Resource Economics: Contents of volume 23. Marine Resource Economics, 23(4):550–552, ???? 2008. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2008:MREa

[Ano08b] Anonymous. Marine Resource Economics: Reviewers 2008. Marine Resource Economics, 23(4):??, ???? 2008. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2008:DHA

[Ano08c] Anonymous. Dr. S.-Y. Hong Award. Marine Resource Economics, 23(4):554–555, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2008:FM

[Ano08d] Anonymous. Front matter. Marine Resource Economics, 23(4): ??, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2009:MREb

[Ano09a] Anonymous. Marine Resource Economics: Contents of volume 24. Marine Resource Economics, 24(4):414–416, ???? 2009. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2009:MREa

[Ano09b] Anonymous. Marine Resource Economics: Reviewers 2009. Marine Resource Economics, 24(4):??, ???? 2009. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2009:FMa

[Ano09c] Anonymous. Front matter. Marine Resource Economics, 24(1): ??, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2009:FMb

[Ano09d] Anonymous. Front matter. Marine Resource Economics, 24(2): ??, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2009:FMc

[Ano09e] Anonymous. Front matter. Marine Resource Economics, 24(3): ??, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2009:FMd

[Ano09f] Anonymous. Front matter. Marine Resource Economics, 24(4): ??, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2009:DHA

[Ano09g] Anonymous. The Dr. S.-Y. Hong Award for Outstanding Article in: Marine Resource Economics. Marine Resource Economics, 24(4):417–418, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2016:Ma

[Ano16a] Anonymous. Masthead. Marine Resource Economics, 31(3):??, July 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2016:Mb

[Ano16b] Anonymous. Masthead. Marine Resource Economics, 31(4): ??, October 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2017:Ma

[Ano17a] Anonymous. Masthead. Marine Resource Economics, 32(1): ??, January 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2017:Mb

[Ano17b] Anonymous. Masthead. Marine Resource Economics, 32(2):??, April 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2017:Mc

[Ano17c] Anonymous. Masthead. Marine Resource Economics, 32(3):??, July 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2017:Md

[Ano17d] Anonymous. Masthead. Marine Resource Economics, 32(4): ??, October 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2018:Ma

[Ano18a] Anonymous. Masthead. Marine Resource Economics, 33(1): ??, January 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2018:Mb

[Ano18b] Anonymous. Masthead. Marine Resource Economics, 33(2):??, April 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2018:Mc

[Ano18c] Anonymous. Masthead. Marine Resource Economics, 33(3):??, July 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2018:Md

[Ano18d] Anonymous. Masthead. Marine Resource Economics, 33(4): ??, October 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2019:INS

[Ano19a] Anonymous. Introducing a new section in marine resource economics: Case studies. Marine Resource Economics, 34(4):??, October 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2019:Ma

[Ano19b] Anonymous. Masthead. Marine Resource Economics, 34(1): ??, January 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2019:Mb

[Ano19c] Anonymous. Masthead. Marine Resource Economics, 34(2):??, April 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2019:Mc

[Ano19d] Anonymous. Masthead. Marine Resource Economics, 34(3):??, July 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2019:Md

[Ano19e] Anonymous. Masthead. Marine Resource Economics, 34(4): ??, October 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2020:FMa

[Ano20a] Anonymous. Front matter. Marine Resource Economics, 35(1): ??, January 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2020:FMb

[Ano20b] Anonymous. Front matter. Marine Resource Economics, 35(2): ??, April 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2020:FMc

[Ano20c] Anonymous. Front matter. Marine Resource Economics, 35 (3):??, July 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2020:FMd

[Ano20d] Anonymous. Front matter. Marine Resource Economics, 35(4): ??, October 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2020:INSa

[Ano20e] Anonymous. Introducing a new section in marine resource economics: Case studies. Marine Resource Economics, 35(1):??, January 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2020:INSb

[Ano20f] Anonymous. Introducing a new section in marine resource economics: Case studies. Marine Resource Economics, 35(2):??, April 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2020:INSc

[Ano20g] Anonymous. Introducing a new section in marine resource economics: Case studies. Marine Resource Economics, 35(3):??, July 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2021:FMa

[Ano21a] Anonymous. Front matter. Marine Resource Economics, 36(1): ??, January 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2021:FMb

[Ano21b] Anonymous. Front matter. Marine Resource Economics, 36(2): ??, April 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2021:FMc

[Ano21c] Anonymous. Front matter. Marine Resource Economics, 36 (3):??, July 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2021:FMd

[Ano21d] Anonymous. Front matter. Marine Resource Economics, 36(4): ??, October 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2022:FMa

[Ano22a] Anonymous. Front matter. Marine Resource Economics, 37(1): ??, January 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2022:FMb

[Ano22b] Anonymous. Front matter. Marine Resource Economics, 37(2): ??, April 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2022:FMc

[Ano22c] Anonymous. Front matter. Marine Resource Economics, 37 (3):??, July 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2022:FMd

[Ano22d] Anonymous. Front matter. Marine Resource Economics, 37(4): ??, October 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2023:FMa

[Ano23a] Anonymous. Front matter. Marine Resource Economics, 38(1): ??, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2023:FMb

[Ano23b] Anonymous. Front matter. Marine Resource Economics, 38(2): ??, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2023:FMc

[Ano23c] Anonymous. Front matter. Marine Resource Economics, 38(3): ??, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2023:FMd

[Ano23d] Anonymous. Front matter. Marine Resource Economics, 38(4): ??, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2023:FMe

[Ano23e] Anonymous. Front matter. Marine Resource Economics, 39(1): ??, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2024:FM

[Ano24] Anonymous. Front matter. Marine Resource Economics, 39(2): ??, ???? 2024. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2025:FMa

[Ano25a] Anonymous. Front matter. Marine Resource Economics, 39(3): ??, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2025:FMb

[Ano25b] Anonymous. Front matter. Marine Resource Economics, 39(4): ??, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2025:FMc

[Ano25c] Anonymous. Front matter. Marine Resource Economics, 40(1): ??, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anonymous:2025:FMd

[Ano25d] Anonymous. Front matter. Marine Resource Economics, 40(2): ??, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2017:RDS

[AP17] Leif E. Anderson and Mark L. Plummer. Recreational demand for shellfish harvesting under environmental closures. Marine Resource Economics, 32(1):43–57, January 2017. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Armstrong:2001:CMR

[AR01] Claire W. Armstrong and Siv Reithe. Comment: Marine reserves: Will they accomplish more with management costs? Marine Resource Economics, 16(2):165–175, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ahsan:2010:FPR

[AR10] Dewan A. Ahsan and Eva Roth. Farmers’ perceived risks and risk management strategies in an emerging mussel aquaculture industry in Denmark. Marine Resource Economics, 25(3):309– 323, September 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2025:SAF

[ARA^{+}25] Frank Asche, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, FRANK ASCHE, Jay Abolofia, Frank Asche, James E. Wilen, Robert Fonner, and Gil Sylvia. Seafood from aquaculture and fisheries, from near and afar. Marine Resource Economics, 40(2):141–170, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Armstrong:1994:CST

[Arm94] C. W. Armstrong. Cooperative solutions in a transboundary fishery: The Russian–Norwegian co-management of the Arcto– Norwegian cod stock. Marine Resource Economics, 9(4):329– 351, Winter 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Armstrong:2013:BRE

[Arm13] Claire W. Armstrong. Book review: The Economics and Management of World Fisheries. By Trond Bjørndal and Gordon R. Munro. Marine Resource Economics, 28(3):305–306, September 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:1992:OFS

[Arn92] Ragnar Arnason. Optimal feeding schedules and harvesting time in aquaculture. Marine Resource Economics, 7(1):15–35, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:1993:IIT

[Arn93] Ragnar Arnason. The Icelandic Individual Transferable Quota System: a descriptive account. Marine Resource Economics, 8 (3):201–218, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:1994:CDF

[Arn94] Ragnar Arnason. On catch discarding in fisheries. Marine Resource Economics, 9(3):189–207, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:2007:APR

[Arn07a] Ragnar Arnason. Advances in property rights based fisheries management: an introduction. Marine Resource Economics, 22 (4):335–346, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:2007:FSM

[Arn07b] Ragnar Arnason. Fisheries self-management under ITQs. Marine Resource Economics, 22(4):373–390, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:2013:ODF

[Arn13] Ragnar Arnason. On optimal dynamic fisheries enforcement. Marine Resource Economics, 28(4):361–377, December 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:2023:IFM

[Arn23] Ragnar Arnason. Improving fisheries management in developing countries. Marine Resource Economics, 38(3):263–283, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2014:BHM

[ARS14] James Anderson, Cathy Roheim, and Marty Smith. Barbara Harrison MRE editorial assistant and technical editor 1998– 2014. Marine Resource Economics, 29(3):??, September 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2007:PGS

[ART07] Frank Asche, Kristin H. Roll, and Ragnar Tveteras. Productivity growth in the supply chain — another source of competitiveness for aquaculture. Marine Resource Economics, 22 (3):329–334, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:2008:PRS

[ART08] Trude B. Andersen, Kristin H. Roll, and Sigbjørn Tveterås. The price responsiveness of salmon supply in the short and long run. Marine Resource Economics, 23(4):425–437, ???? 2008. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:1984:SBR

[AS84] Peder Andersen and Jon G. Sutinen. Stochastic bioeconomics: a review of basic methods and results. Marine Resource Economics, 1(2):117–136, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Androkovich:1989:RSF

[AS89] R. A. Androkovich and K. R. Stollery. Regulation of stochastic fisheries: a comparison of alternative methods in the Pacific halibut fishery. Marine Resource Economics, 6(2):109–122, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Androkovich:1994:SDP

[AS94] Robert A. Androkovich and Kenneth R. Stollery. A stochastic dynamic programming model of bycatch control in fisheries. Marine Resource Economics, 9(1):19–30, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:2005:LAT

[AS05] Christopher M. Anderson and Jon G. Sutinen. A laboratory assessment of tradable fishing allowances. Marine Resource Economics, 20(1):1–23, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1985:BM

[ASB^{+}85] James L. Anderson, Karl C. Samples, Richard C. Bishop, Rögnvaldur Hannesson, James L. Anderson, Karl C. Samples, Richard C. Bishop, Colin W. Clark, Anthony T. Charles, John R. Beddington, and Marc Mangel. Back matter. Marine Resource Economics, 2(1):??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:1997:DAD

[Asc97a] Frank Asche. Dynamic adjustment in demand equations. Marine Resource Economics, 12(3):221–237, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:1997:TDP

[Asc97b] Frank Asche. Trade disputes and productivity gains: The curse of farmed salmon production? Marine Resource Economics, 12 (1):67–73, Spring 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2007:CMF

[Asc07] Frank Asche. Capacity measurement in fisheries: What can we learn? Marine Resource Economics, 22(1):105–108, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Asche:2008:FS

[Asc08] Frank Asche. Farming the sea. Marine Resource Economics, 23 (4):527–547, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Arnason:1992:BM

[ASH^{+}92] Ragnar Arnason, David B. Sampson, Karen M. Hayes, Timothy J. Tyrrell, Timothy J. Tyrell, Glen Anderson, Ragnar Arnason, David B. Sampson, and James L. Anderson. Back matter. Marine Resource Economics, 7(1):??, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andersen:1984:BM

[ASP^{+}84] Peder Andersen, Jon G. Sutinen, Charles Plourde, Richard Bodell, Evan F. Koenig, Evan F. Koenig, Peder Andersen, and Jon G. Sutinen. Back matter. Marine Resource Economics, 1 (2):??, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Anderson:1986:BM

[ASR^{+}86] Eric E. Anderson, William E. Schrank, Noel Roy, Eugene Tsoa, Wiliam E. Shrank, Raimo P. Hämäläinen, Jukka Ruusunen, Veijo Kaitala, Raimo P. Hämäläinen, Jukka Ruusunen, Veijo Kaitala, and Eric E. Anderson. Back matter. Marine Resource Economics, 3(3):??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Agar:2014:GMR

[ASS^{+}14] Juan J. Agar, Jessica A. Stephen, Andy Strelcheck, Assane Diagne, Juan J. Agar, Jessica A. Stephen, Andy Strelcheck, Assane Diagne, Martin D. Smith, Frank Asche, Lori S. Bennear, Atle Oglend, Andrew J. Ropicki, Sherry L. Larkin, Martin D. Smith, Frank Asche, Lori S. Bennear, Atle Oglend, Juan J. Agar, Jessica A. Stephen, Andy Strelcheck, Assane Diagne, Andrew J. Ropicki, and Sherry L. Larkin. The Gulf of Mexico red snapper IFQ program: The first five years. Marine Resource Economics, 29(2):177–198, June 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abbott:2018:CPR

[ASS18] Joshua K. Abbott, James N. Sanchirico, and Martin D. Smith. Common property resources and the dynamics of overexploitation: The case of the North Pacific fur seal — a 42-year legacy. Marine Resource Economics, 33(3):209–212, July 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See comment [Wil18a].

Asche:2002:EAS

[AT02] Frank Asche and Ragnar Tveteras. Economics of aquaculture: Special issue introduction. Marine Resource Economics, 17(2): 73–75, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

AkhtaruzzamanKhan:2023:TES

[ATR^{+}23] Md. Akhtaruzzaman Khan, Md. Takibur Rahman, Kristin Helen Roll, Atle Guttormsen, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Min-Yang Lee, Scott Steinback, and Kristy Wallmo. Towards economic sustainability or unsustainability: The case of artisanal fisheries in the Bay of Bengal. Marine Resource Economics, 38 (2):181–201, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Andrews:1988:ARS

[AW88] Elizabeth J. Andrews and James E. Wilen. Angler response to success in the California salmon sportfishery: Evidence and management implications. Marine Resource Economics, 5(2): 125–138, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Aruga:2018:CPT

[AW18] Kentaka Aruga and Hiroki Wakamatsu. Consumer perceptions toward seafood produced near the Fukushima nuclear plant. Marine Resource Economics, 33(4):373–386, October 2018. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Abayomi:2012:UCL

[AY12] Kobi Abayomi and Tracy Yandle. Using conditional Lorenz curves to examine consolidation in New Zealand commercial fishing. Marine Resource Economics, 27(4):303–321, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ankamah-Yeboah:2020:CPH

[AYAB^{+}20] Isaac Ankamah-Yeboah, Frank Asche, Julia Bronnmann, Max Nielsen, and Rasmus Nielsen. Consumer preference heterogeneity and preference segmentation: The case of ecolabeled salmon in Danish retail sales. Marine Resource Economics, 35(2):159– 176, April 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ankamah-Yeboah:2019:IAW

[AYJO^{+}19a] Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, and Rasmus Nielsen. The impact of animal welfare and environmental information on the choice of organic fish: an empirical investigation of German trout consumers. Marine Resource Economics, 34(3):247–266, July 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ankamah-Yeboah:2019:E

[AYJO^{+}19b] Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, Daniel K. Lew, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, Daniel K. Lew, Michael R. Springborn, and Amanda Faig. Erratum. Marine Resource Economics, 34(3):??, July 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ankamah-Yeboah:2019:DOS

[AYNN19] Isaac Ankamah-Yeboah, Max Nielsen, and Rasmus Nielsen. Does organic supply growth lead to reduced price premiums? The case of salmonids in Denmark. Marine Resource Economics, 34(2):105–121, April 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ankamah-Yeboah:2017:MIC

[AYSN17] Isaac Ankamah-Yeboah, Lisa Ståhl, and Max Nielsen. Market integration of cold and warmwater shrimp in Europe. Marine Resource Economics, 32(4):371–390, October 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Buason:2020:FFC

[BA20] Arnar Buason and Sveinn Agnarsson. Fond of fish? Count data analysis of seafood consumption in France. Marine Resource Economics, 35(2):137–157, April 2020. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2004:BAN

[BaUS04] Trond Bjørndal, Al amin Ussif, and Ussif Rashid Sumaila. A bioeconomic analysis of the Norwegian spring spawning herring (NSSH) stock. Marine Resource Economics, 19(3):353–365, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bronnmann:2016:MIB

[BAYN16] Julia Bronnmann, Isaac Ankamah-Yeboah, and Max Nielsen. Market integration between farmed and wild fish: Evidence from the whitefish market in Germany. Marine Resource Economics, 31(4):421–432, October 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2006:EAB

[BB06] Trond Bjørndal and Ana Brasão. The East Atlantic bluefin tuna fisheries: Stock collapse or recovery? Marine Resource Economics, 21(2):193–210, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2010:RRM

[BBA^{+}10] Trond Bjorndal, Mintewab Bezabih, Dewan A. Ahsan, Eva Roth, Kolawole Ogundari, Olajide O. Akinbogun, Richard W.

Ryan, Daniel S. Holland, Guillermo E. Herrera, Dewan A. Ahsan, Eva Roth, Kolawole Ogundari, Olajide O. Akinbogun, Eriko Hoshino, Richard M. Hillary, and John Pearce. Resource rents and management regimes: The case of the Western Channel sole fishery. Marine Resource Economics, 25(3):325–331, September 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2006:BM

[BBH^{+}06] Trond Bjørndal, Ana Brasão, Mark Herrmann, Keith Criddle, John B. Walden, Trond Bjørndal, Ana Brasão, John B. Walden, Jared Greenville, Jessica Hartmann, and T. Gordon Macaulay. Back matter. Marine Resource Economics, 21(2):??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Brown:2005:PPM

[BBI05a] Gardner Brown, Brett Berger, and Moses Ikiara. A predator– prey model with an application to Lake Victoria fisheries. Marine Resource Economics, 20(3):221–248, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Brown:2005:BM

[BBI^{+}05b] Gardner Brown, Brett Berger, Moses Ikiara, Kjetil Helstad, Terje Vassdal, Torbjørn Trondsen, James A. Young, Di Jin, Guillermo E. Herrera, Kjetil Helstad, Terje Vassdal, Torbjørn Trondsen, James A. Young, Duncan Knowler, Edward B. Barbier, Di Jin, and Guillermo E. Herrera. Back matter. Marine Resource Economics, 20(3):??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Beddington:1984:APB

[BC84] John R. Beddington and Colin W. Clark. Allocation problems between national and foreign fisheries with a fluctuating fish resource. Marine Resource Economics, 1(2):137–154, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:1987:CDN

[BC87] Trond Bjørndal and Jon M. Conrad. Capital dynamics in the North Sea herring fishery. Marine Resource Economics, 4(1): 63–74, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bin:2013:ITN

[BC13] Okmyung Bin and Jeffrey Czajkowski. The impact of technical and non-technical measures of water quality on coastal waterfront property values in South Florida. Marine Resource Economics, 28(1):43–63, January 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:1987:BM

[BCA^{+}87] Trond Bjørndal, Jon M. Conrad, Peder Andersen, R. Sathiendrakumar, C. A. Tisdell, Trond Bjørndal, Jon M. Conrad, John O. S. Kennedy, B. A. Cook, and Parzival Copes. Back matter. Marine Resource Economics, 4(1):??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bertignac:2000:BM

[BCH^{+}00] Michel Bertignac, Harry F. Campbell, John Hampton, Anthony J. Hand, Diane P. Dupont, R. Quentin Grafton, Ujjayant Chakravorty, Keiichi Nemoto, Diane P. Dupont, R. Quentin Grafton, Ujjayant Chakravorty, Keiichi Nemoto, and John R. Boyce. Back matter. Marine Resource Economics, 15(3):??, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bertignac:2000:MRR

[BCHH00] Michel Bertignac, Harry F. Campbell, John Hampton, and Anthony J. Hand. Maximizing resource rent from the Western and Central Pacific tuna fisheries. Marine Resource Economics, 15 (3):151–177, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bergland:2001:RSR

[BCP01] Harald Bergland, Derek J. Clark, and Pål Andreas Pedersen. Rent seeking and the regulation of a natural resource. Marine Resource Economics, 16(3):219–233, ???? 2001. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Birkenbach:2023:CCC

[BCP23] Anna M. Birkenbach, Andreea L. Cojocaru, and Ruth B. M. Pincinato. Coastal communities at a crossroads. Marine Resource Economics, 38(4):353–364, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Birkenbach:2022:DCM

[BCSA22] Anna M. Birkenbach, Andreea L. Cojocaru, Martin D. Smith, and Frank Asche. Discrete choice modeling of fishers’ landing locations. Marine Resource Economics, 37(3):235–262, July 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Brasao:2000:MNA

[BDCDeS00] Ana Brasão, Clara Costa Duarte, Clara Costa-Duarte, and Maria A. Cunha e Sá. Managing the Northern Atlantic bluefin tuna fisheries: The stability of the UN fish stock agreement solution. Marine Resource Economics, 15(4):341–360, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Beasley:2025:CCA

[Bea25] W. Jason Beasley. Changing coasts along Lake Michigan: Shoreline management impacts on the inland community. Marine Resource Economics, 39(4):263–283, ???? 2025. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bell:1989:MQH

[Bel89] Frederick W. Bell. Main quarry hypothesis and salmon angling. Marine Resource Economics, 6(1):71–82, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Berman:2006:MSC

[Ber06] Matthew Berman. Modeling spatial choice in ocean fisheries. Marine Resource Economics, 21(4):375–394, ???? 2006. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2014:MPF

[BER^{+}14] Trond Bjørndal, Nils-Arne Ekerhovd, Richard W. Ryan, Daniel S. Holland, Guillermo E. Herrera, Daniel R. Petrolia, Matthew G. Interis, Joonghyun Hwang, Hiroki Wakamatsu, Daniel R. Petrolia, Matthew G. Interis, Joonghyun Hwang, Hiroki Wakamatsu, Trond Bjørndal, and Nils-Arne Ekerhovd. Management of pelagic fisheries in the North East Atlantic: Norwegian spring spawning herring, mackerel, and blue whiting. Marine Resource Economics, 29(1):69–83, March 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Berglann:2023:FMU

[Ber23] Helge Berglann. Fisheries management under uncertainty using a convex tax. Marine Resource Economics, 38(1):47–64, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2000:ESH

[BG00] Trond Bjørndal and Daniel V. Gordon. The economic structure of harvesting for three vessel types in the Norwegian springspawning herring fishery. Marine Resource Economics, 15(4): 281–292, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Boyce:1993:BM

[BHB^{+}93] John Boyce, Mark Herrmann, Diane Bischak, Joshua Greenberg, Veijo Kaitala, Gordon R. Munro, Daniel V. Gordon, Kjell G. Salvanes, Frank Atkins, Veijo Kaitala, Gordon R. Munro, Daniel V. Gordon, Kjell G. Salvanes, Frank Atkins, John Boyce, Mark Herrmann, Diane Bischak, and Joshua Greenberg. Back matter. Marine Resource Economics, 8(4):??, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Boyce:1993:ASE

[BHBG93] John Boyce, Mark Herrmann, Diane Bischak, and Joshua Greenberg. The Alaska salmon enhancement program: a cost/benefit analysis. Marine Resource Economics, 8(4):293–312, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bi:2016:INI

[BHG^{+}16] Xiang Bi, Lisa House, Zhifeng Gao, Hing Ling Chan, Minling Pan, Xiang Bi, Lisa House, Zhifeng Gao, Craig E. Landry, Alyson R. Lewis, Haiyong Liu, Hans Vogelsong, Xiang Bi, Lisa House, Zhifeng Gao, Hing Ling Chan, Minling Pan, Julia Bronnmann, Jens-Peter Loy, and Karen J. Schroeder. Impacts of nutrition information on choices of fresh seafood among parents. Marine Resource Economics, 31(3):355–372, July 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Brown:2012:BRF

[BHH^{+}12] Gardner Brown, Timothy Haab, Robert Hicks, Kurt Schnier, John C. Whitehead, Andrew M. Scheld, Christopher M. Anderson, Hirotsugu Uchida, Gardner Brown, Timothy Haab, Robert Hicks, Kurt Schnier, John C. Whitehead, Andrew M. Scheld, Christopher M. Anderson, Hirotsugu Uchida, Djiga Thiao, and Francis Laloë. Book review: Four Fish: The Future of the Last Wild Food By Paul Greenberg. Marine Resource Economics, 27 (3):283–285, September 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bell:2003:WPL

[BHJ03] Kathleen P. Bell, Daniel Huppert, and Rebecca L. Johnson. Willingness to pay for local Coho salmon enhancement in coastal communities. Marine Resource Economics, 18(1):15–31, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Berman:1997:ENB

[BHK97] Matthew Berman, Sharman Haley, and Hongjin Kim. Estimating net benefits of reallocation: Discrete choice models of sport and commercial fishing. Marine Resource Economics, 12(4):307– 327, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Blomquist:2015:TFB

[BHW15] Johan Blomquist, Cecilia Hammarlund, and Staffan Waldo. Time for fishing: Bargaining power in the Swedish Baltic cod fishery. Marine Resource Economics, 30(3):315–329, July 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bird:1986:EEW

[Bir86] Peter J. W. N. Bird. Econometric estimation of world salmon demand. Marine Resource Economics, 3(2):169–182, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:1988:OHF

[Bjø88] Trond Bjørndal. Optimal harvesting of farmed fish. Marine Resource Economics, 5(2):139–159, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:1993:GMG

[Bjø93] Trond Bjørndal. The G. M. Gerhardsen Memorial Prize in Fisheries Economics. Marine Resource Economics, 8(3):183–184, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2000:SII

[Bjø00] Trond Bjørndal. Special issue introduction. Marine Resource Economics, 15(4):261–263, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Buerger:1989:NYV

[BK89] Robert Buerger and James R. Kahn. New York value of Chesapeake striped bass. Marine Resource Economics, 6(1):19–25, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Birkenbach:2023:DCS

[BKS^{+}23] Anna M. Birkenbach, David J. Kaczan, Martin D. Smith, Greg Ardini, Daniel S. Holland, Min-Yang Lee, Doug Lipton, and Michael D. Travis. Do catch shares increase prices? Evidence from US fisheries. Marine Resource Economics, 38(3):203–228, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bin:2005:BM

[BLE^{+}05] Okmyung Bin, Craig E. Landry, Christopher L. Ellis, Hans Vogelsong, Gro Steine, Frode Alfnes, Mia Bencze Rørå, Junjie Wu, Katsuya Tanaka, Max Nielsen, Gro Steine, Frode Alfnes, Mia Bencze Rørå, Okmyung Bin, Craig E. Landry, Christopher L. Ellis, and Hans Vogelsong. Back matter. Marine Resource Economics, 20(2):??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bin:2005:SCS

[BLEV05] Okmyung Bin, Craig E. Landry, Christopher L. Ellis, and Hans Vogelsong. Some consumer surplus estimates for North Carolina beaches. Marine Resource Economics, 20(2):145–161, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bronnmann:2016:CDS

[BLS16] Julia Bronnmann, Jens-Peter Loy, and Karen J. Schroeder. Characteristics of demand structure and preferences for wild and farmed seafood in Germany: an application of QUAIDS modeling with correction for sample selection. Marine Resource Economics, 31(3):281–300, July 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bose:1996:SAS

[BM96] Shekar Bose and Alistair McIlgorm. Substitutability among species in the Japanese tuna market: a cointegration analysis. Marine Resource Economics, 11(3):143–155, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Breffle:2000:IPH

[BM00] William S. Breffle and Edward R. Morey. Investigating preference heterogeneity in a repeated discrete-choice recreation demand model of Atlantic salmon fishing. Marine Resource Economics, 15(1):1–20, Spring 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Breffle:2000:BM

[BMD^{+}00] William S. Breffle, Edward R. Morey, Clara Costa Duarte, Ana Brasão, Pedro Pintassilgo, Nicolas Péridy, Patrice Guillotreau, Pascal Bernard, William S. Breffle, Edward R. Morey, Clara Costa Duarte, Ana Brasão, Pedro Pintassilgo, and Asgeir Danielsson. Back matter. Marine Resource Economics, 15(1): ??, Spring 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bockstael:1989:MBI

[BMS89a] N. E. Bockstael, K. E. McConnell, and I. E. Strand. Measuring the benefits of improvements in water quality: The Chesapeake Bay. Marine Resource Economics, 6(1):1–18, ???? 1989. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bockstael:1989:RUM

[BMS89b] N. E. Bockstael, K. E. McConnell, and I. E. Strand. A random utility model for sportfishing: Some preliminary results for Florida. Marine Resource Economics, 6(3):245–260, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bockstael:1989:BMb

[BMS^{+}89c] N. E. Bockstael, K. E. McConnell, I. E. Strand, Lee G. Anderson, Douglas E. Booth, Lee G. Anderson, N. E. Bockstael, K. E. McConnell, I. E. Strand, Charles Plourde, and J. Barry Smith. Back matter. Marine Resource Economics, 6(3):??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bockstael:1989:BMa

[BMS^{+}89d] N. E. Bockstael, K. E. McConnell, I. E. Strand, Charles Plourde, David Yeung, Neal S. Johnson, Richard M. Adams, N. E. Bockstael, K. E. McConnell, I. E. Strand, Charles Plourde, and David Yeung. Back matter. Marine Resource Economics, 6(1):??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bockstael:1984:UAC

[Boc84] Nancy E. Bockstael. Uncertainty about consumption and consumer uncertainty. Marine Resource Economics, 1(1):67–76, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Booth:1989:HDD

[Boo89] Douglas E. Booth. Hydroelectric dams and the decline of Chinook salmon in the Columbia River Basin. Marine Resource Economics, 6(3):195–211, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bose:2004:EIP

[Bos04] Shekar Bose. An empirical investigation of price-quantity relations of the quota species of Australia’s South East fishery. Marine Resource Economics, 19(2):161–172, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bostedt:2013:BRR

[Bos13] Göran Bostedt. Book review: Resource Economics: an Economic Approach to Natural Resource and Environmental Policy. By John C. Bergstrom and Alan Randall. Marine Resource Economics, 28(1):105–106, January 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Boyce:1993:UPD

[Boy93] John R. Boyce. Using participation data to estimate fishing costs for commercial salmon fisheries in Alaska. Marine Resource Economics, 8(4):367–394, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Boyce:2000:CEI

[Boy00] John R. Boyce. Comment: Efficiency of ITQs in the presence of production externalities. Marine Resource Economics, 15 (3):233–243, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bluffstone:1994:BRJ

[BP94] Randall Bluffstone and Theodore Panayotou. Book review: James B. Marsh, Resources and Environment in Asia’s Marine Sector. Marine Resource Economics, 9(2):187–188, Summer 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bergland:1997:CRA

[BP97] Harald Bergland and Pål Andreas Pedersen. Catch regulation and accident risk: The moral hazard of fisheries’ management. Marine Resource Economics, 12(4):281–292, Winter 1997. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bergland:2006:RAI

[BP06] Harald Bergland and Pål Andreas Pedersen. Risk attitudes and individual transferable quotas. Marine Resource Economics, 21 (1):81–100, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Broadus:1985:CGE

[Bro85] J. M. Broadus. Chennat gopalakrishnan, the emerging marine economy of the Pacific. Marine Resource Economics, 2(2):203– 209, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Broadus:1986:APM

[Bro86] J. M. Broadus. Asian Pacific marine minerals and industry structure. Marine Resource Economics, 3(1):63–88, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bosch:1989:DPS

[BS89] Darrell J. Bosch and Leonard A. Shabman. The decline of private sector oyster culture in Virginia: Causes and remedial policies. Marine Resource Economics, 6(3):227–243, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Berman:2006:DAV

[BS06a] Matthew Berman and Ussif Rashid Sumaila. Discounting, amenity values, and marine ecosystem restoration. Marine Resource Economics, 21(2):211–219, ???? 2006. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bisack:2006:NZI

[BS06b] Kathryn D. Bisack and Jon G. Sutinen. A New Zealand ITQ fishery with an in-season stock externality. Marine Resource Economics, 21(3):231–249, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:2013:CPR

[BS13] Trond Bjørndal and Stein Ivar Steinshamn. On the contributions of Professor Rögnvaldur Hannesson to fisheries economics. Marine Resource Economics, 28(4):345–350, December 2013. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bang:2022:EVE

[BSA^{+}22] Rasmus Noss Bang, Stein Ivar Steinshamn, Yee Mon Aung, Ling Yee Khor, Nhuong Tran, Michael Akester, Manfred Zeller, Irmelin Slettemoen Helgesen, Rasmus Noss Bang, and Stein Ivar Steinshamn. Endogenous versus exogenous natural mortality and weight in bioeconomic models. Marine Resource Economics, 37(4):491–514, October 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bockstael:1992:PGA

[BSL92] Nancy E. Bockstael, Ivar E. Strand, and Douglas W. Lipton. Pulsed generic advertising: The case of common property. Marine Resource Economics, 7(4):189–208, Winter 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bird:1986:BM

[BSS^{+}86] Peter J. W. N. Bird, Kenneth Stollery, J. Barry Smith, J. Barry Smith, and Peter J. W. N. Bird. Back matter. Marine Resource Economics, 3(2):??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Braekkan:2014:BRD

[BT14] Eivind Hestvik Brækkan and Sverre Braathen Thyholdt. The bumpy road of demand growth — an application to Atlantic salmon. Marine Resource Economics, 29(4):339–350, December 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Burton:1992:DWF

[Bur92] M. P. Burton. The demand for wet fish in Great Britain. Marine Resource Economics, 7(2):57–66, Summer 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Backstrom:2023:OCD

[BWR^{+}23] Jesse D. Backstrom, Richard T. Woodward, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, Jesse D. Backstrom, Richard T. Woodward, Kyle Montanio, Emi Uchida, Katrina Kosec, and Jennifer Meredith. Onsite characteristics and diversity avoidance in marine recreational fishing demand. Marine Resource Economics, 38(4):435–463, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Bjorndal:1988:BM

[BYD^{+}88] Trond Bjørndal, Jahara Yahaya, John P. Doll, Trond Bjørndal, John P. Doll, Elizabeth J. Andrews, and James E. Wilen. Back matter. Marine Resource Economics, 5(2):??, ???? 1988. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chu:2010:SPA

[CAAT10] Jingjie Chu, James L. Anderson, Frank Asche, and Lacey Tudur. Stakeholders’ perceptions of aquaculture and implications for its future: a comparison of the U.S.A. and Norway. Marine Resource Economics, 25(1):61–76, March 2010. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cavalcanti:2023:IAW

[Cav23] Carina Cavalcanti. Identifying the actual willingness to reduce overfishing with a fishnet auction: Evidence from artisanal fishing villages in Brazil. Marine Resource Economics, 38(2):101– 111, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cochrane:1992:LMS

[CBI92] Jeffrey A. Cochrane, Richard C. Bishop, and David B. Ives. Lake Michigan salmonid stocking costs in Wisconsin. Marine Resource Economics, 7(3):169–185, Fall 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clark:1986:MMM

[CC86] Allen L. Clark and Jennifer Cook Clark. Marine metallic mineral resources of the Pacific Basin. Marine Resource Economics, 3(1): 45–62, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cook:1987:OLC

[CC87] B. A. Cook and Parzival Copes. Optimal levels for Canada’s Pacific halibut catch. Marine Resource Economics, 4(1):45–61, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clark:1985:OCD

[CCBM85] Colin W. Clark, Anthony T. Charles, John R. Beddington, and Marc Mangel. Optimal capacity decisions in a developing fishery. Marine Resource Economics, 2(1):25–53, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Costello:2007:EGF

[CD07] Christopher Costello and Robert Deacon. The efficiency gains from fully delineating rights in an ITQ fishery. Marine Resource Economics, 22(4):347–361, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Carlucci:2017:CLV

[CDNS17] Domenico Carlucci, Biagia De Devitiis, Gianluca Nardone, and Fabio Gaetano Santeramo. Certification labels versus convenience formats: What drives the market in aquaculture products? Marine Resource Economics, 32(3):295–310, July 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Casey:1995:BM

[CDT^{+}95] Keith E. Casey, Christopher M. Dewees, Bruce R. Turris, James E. Wilen, Kenneth E. McConnell, Ivar E. Strand, Lynne Blake-hedges, Terry Heaps, Keith E. Casey, Christopher M. Dewees, Bruce R. Turris, James E. Wilen, Kenneth E. McConnell, Ivar E. Strand, Lynne Blake-hedges, and Terry Heaps. Back matter. Marine Resource Economics, 10(3):??, Fall 1995. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Casey:1995:EIV

[CDTW95] Keith E. Casey, Christopher M. Dewees, Bruce R. Turris, and James E. Wilen. The effects of individual vessel quotas in the British Columbia halibut fishery. Marine Resource Economics, 10(3):211–230, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chavez:2021:EMR

[CER21] Carlos Chávez, Håkan Eggert, and Matthew Reimer. Economics of marine resources in the Global South — meeting the challenge of Agenda 2030. Marine Resource Economics, 36(4):307–318, October 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Crutchfield:1985:IEF

[CG85] Stephen R. Crutchfield and John M. Gates. The impact of extended fishery jurisdiction on the New England otter trawl fleet. Marine Resource Economics, 2(2):153–173, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chand:2003:MGF

[CGP03a] Satish Chand, R. Quentin Grafton, and Elizabeth Petersen. Multilateral governance of fisheries: Management and cooperation in the Western and Central Pacific tuna fisheries. Marine Resource Economics, 18(4):329–344, ???? 2003. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chand:2003:BM

[CGP^{+}03b] Satish Chand, R. Quentin Grafton, Elizabeth Petersen, Keith R. Criddle, Mark Herrmann, S. Todd Lee, Charles Hamel, Geir Grundvåg Ottesen, Kjell Grønhaug, Keith R. Criddle, Mark Herrmann, S. Todd Lee, Charles Hamel, Linda Nøstbakken, Trond Bjørndal, Alan Collins, Sean Pascoe, and David Whitmarsh. Back matter. Marine Resource Economics, 18(4):??, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

CastillaEspino:2005:CCU

[CGS05] David Castilla Espino, Juan José García Del Hoyo, and Basil M. H. Sharp. Capacity and capacity utilization of the “Voracera” fleet in the Strait of Gibraltar. Marine Resource Economics, 20 (4):367–384, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chen:2025:STB

[CH25] Y. Allen Chen and Alan C. Haynie. Size-targeting in the Bering Sea pollock catcher/processor fishery with heterogeneous incentives. Marine Resource Economics, 39(4):307–331, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chapman:1995:BRM

[Cha95] Duane Chapman. Book review: Mary Ann Tetreault, The Kuwait Petroleum Corporation and the Economics of the New World Order. Marine Resource Economics, 10(4):413–414, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Charles:2000:BRD

[Cha00] Anthony T. Charles. Book review: Daniel M. Kammen and David M. Hassenzahl, Should We Risk It? Exploring Environmental, Health and Technological Problem Solving. Marine Resource Economics, 15(1):73–75, Spring 2000. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Charles:2009:BRJ

[Cha09] Anthony Charles. Book review: Joseph Gough, Managing Canada’s Fisheries: from early days to the year 2000. Marine Resource Economics, 24(2):207–211, ???? 2009. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Criddle:2003:PDA

[CHLH03] Keith R. Criddle, Mark Herrmann, S. Todd Lee, and Charles Hamel. Participation decisions, angler welfare, and the regional economic impact of sportfishing. Marine Resource Economics, 18(4):291–312, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chidmi:2012:SBF

[CHN12] Benaissa Chidmi, Terry Hanson, and Giap Nguyen. Substitutions between fish and seafood products at the US national retail level. Marine Resource Economics, 27(4):359–370, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huang:2007:EVB

[cHPZ07] Ju chin Huang, P. Joan Poor, and Min Qiang Zhao. Economic valuation of beach erosion control. Marine Resource Economics, 22(3):221–238, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Christy:1996:DRO

[Chr96] Francis T. Christy. The death rattle of open access and the advent of property rights regimes in fisheries. Marine Resource Economics, 11(4):287–304, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Campbell:1993:BMM

[CHS93] H. F. Campbell, A. J. Hand, and A. D. M. Smith. A bioeconomic model for management of orange roughy stocks. Marine Resource Economics, 8(2):155–172, Summer 1993. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chu:1986:GPE

[Chu86] D. K. Y. Chu. Government policies, economic development, and possible environmental effects at the land–water interfaces of Guangdong Province, China. Marine Resource Economics, 3(1):29–44, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Carson:2009:NLM

[CHW09a] Richard T. Carson, W. Michael Hanemann, and Thomas C. Wegge. A nested logit model of recreational fishing demand in Alaska. Marine Resource Economics, 24(2):101–129, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Carson:2009:BM

[CHW^{+}09b] Richard T. Carson, W. Michael Hanemann, Thomas C. Wegge, Olli Tahvonen, Daniel R. Petrolia, Tae goun Kim, Olli Tahvonen, Richard T. Carson, W. Michael Hanemann, Thomas C. Wegge, O. A. Adeogun, P. O. Abohweyere, H. K. Ogunbadejo, Alhaji Tanko, and L. Jim-saiki. Back matter. Marine Resource Economics, 24(2):??, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Conklin:1994:CH

[CK94] James E. Conklin and William C. Kolberg. Chaos for the halibut? Marine Resource Economics, 9(2):159–182, Summer 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chhun:2015:AMP

[CKMT15] Sophal Chhun, Viktoria Kahui, Henrik Moller, and Paul Thorsnes. Advancing marine policy toward ecosystem-based management by eliciting public preferences. Marine Resource Economics, 30(3):261–275, July 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Campbell:1989:NOE

[CL89] H. F. Campbell and R. K. Lindner. A note on optimal effort in the Maldivian tuna fishery. Marine Resource Economics, 6 (2):173–176, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See comments [ST89].

Cato:1998:EUS

[CL98] James C. Cato and Carlos A. Lima Dos Santos. European Union 1997 seafood-safety ban: The economic impact on Bangladesh shrimp processing. Marine Resource Economics, 13(3):215–227, Fall 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Carter:2009:SVE

[CL09] David W. Carter and David Letson. Structural vector error correction modeling of integrated sportfishery data. Marine Resource Economics, 24(1):19–41, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Carter:2010:HVS

[CL10] David W. Carter and Christopher Liese. Hedonic valuation of sportfishing harvest. Marine Resource Economics, 25(4):391– 407, December 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clark:1985:EFQ

[Cla85] Colin W. Clark. The effect of fishermen’s quotas on expected catch rates. Marine Resource Economics, 1(4):419–427, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Carter:2022:OPR

[CLL22] David W. Carter, Christopher Liese, and Sabrina J. Lovell. The option price of recreational bag limits and the value of harvest. Marine Resource Economics, 37(1):35–52, January 2022. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chukwuone:2022:BDF

[CLSA22] Nnaemeka Andegbe Chukwuone, Kira Lancker, Jörn O. Schmidt, and Ebele Chinelo Amaechina. Biomass development and fishers’ vulnerability in Nigeria: Evidence from a survey dataset. Marine Resource Economics, 37(4):387–407, October 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Charles:1985:IIO

[CM85] Anthony T. Charles and Gordon R. Munro. Irreversible investment and optimal fisheries management: a stochastic analysis. Marine Resource Economics, 1(3):247–264, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Curtis:2004:IIE

[CM04] Rita E. Curtis and Kenneth E. McConnell. Incorporating information and expectations in fishermen’s spatial decisions. Marine Resource Economics, 19(1):131–143, ???? 2004. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clark:2017:CTE

[CM17] Colin W. Clark and Gordon R. Munro. Capital theory and the economics of fisheries: Implications for policy. Marine Resource Economics, 32(2):123–142, April 2017. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Charles:1999:EIF

[CMC99a] Anthony T. Charles, R. Leigh Mazany, and Melvin L. Cross. The economics of illegal fishing: a behavioral model. Marine Resource Economics, 14(2):95–110, Summer 1999. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Charles:1999:BM

[CMC^{+}99b] Anthony T. Charles, R. Leigh Mazany, Melvin L. Cross, William E. Schrank, Walter R. Keithly, Jr., Christopher Green, William E. Schrank, Walter R. Keithly, Jr., Anthony T. Charles, R. Leigh Mazany, Melvin L. Cross, and Gorazd Ruseski. Back matter. Marine Resource Economics, 14(2):??, Summer 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clarke:2007:SER

[CMGB07a] Shelley Clarke, E. J. Milner-Gulland, and Trond Bjørndal. Social, economic, and regulatory drivers of the shark fin trade. Marine Resource Economics, 22(3):305–327, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clarke:2007:BM

[CMGB^{+}07b] Shelley Clarke, E. J. Milner-Gulland, Trond Bjørndal, Frank Asche, Kristin H. Roll, Ragnar Tveteras, Ju chin Huang, P. Joan Poor, Min Qiang Zhao, Shelley Clarke, E. J. Milner-Gulland, Trond Bjørndal, Cathy A. Roheim, Lacey Gardiner, Frank Asche, Frank Asche, Kristin H. Roll, and Ragnar Tveteras. Back matter. Marine Resource Economics, 22(3):??, ???? 2007. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Carter:2016:VSH

[CMLC16] David W. Carter, Akbar Marvasti, Christopher Liese, and Scott Crosson. Valuing sportfishing harvest with the demand for boat fuel. Marine Resource Economics, 31(3):323–338, July 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clark:1988:DIN

[CMM88] Ian N. Clark, Philip J. Major, and Nina Mollett. Development and implementation of New Zealand’s ITQ management system. Marine Resource Economics, 5(4):325–349, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Charles:1985:BM

[CMS^{+}85] Anthony T. Charles, Gordon R. Munro, Dale Squires, Joseph Mueller, Robert Reidman, James Kirkley, S. B. Saila, E. Lorda, H. A. Walker, Anthony T. Charles, Gordon R. Munro, Daniel F. Spulber, S. B. Saila, E. Lorda, and H. A. Walker. Back matter. Marine Resource Economics, 1(3):??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Campbell:1995:AYT

[CN95] H. F. Campbell and R. B. Nicholl. Allocating yellowfin tuna between the multispecies purse seine and longline fleets. Marine Resource Economics, 10(1):35–58, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chakravorty:2000:MEA

[CN00] Ujjayant Chakravorty and Keiichi Nemoto. Modeling the effects of area closure and tax policies: a spatial-temporal model of the Hawaii longline fishery. Marine Resource Economics, 15(3):179– 204, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Crothers:2006:HSF

[CN06] G. T. (Stan) Crothers and Lindie Nelson. High seas fisheries governance: a framework for the future? Marine Resource Economics, 21(4):341–353, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cunningham:2009:WBF

[CNAB09] Stephen Cunningham, Arthur E. Neiland, Michael Arbuckle, and Tim Bostock. Wealth-based fisheries management: Using fisheries wealth to orchestrate sound fisheries policy in practice. Marine Resource Economics, 24(3):271–287, ???? 2009.

CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Comitini:1988:BRC

[Com88] Salvatore Comitini. Book review: C. Bailey, A. Dwiponggo, and F. Marahudin, Indonesian Marine Capture Fisheries. Marine Resource Economics, 5(3):279–283, ???? 1988. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Comitini:1989:BRH

[Com89] Salvatore Comitini. Book review: H. Campbell, K. Menz, and G. Waugh, Economics of Fishery Management in the Pacific Islands Region, ACIAR Proceedings No. 26. Marine Resource Economics, 6(4):375–380, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Comitini:1993:BRR

[Com93] Salvatore Comitini. Book review: Richard B. Pollnac, Conner Bailey, and Alie Poernomo, Contributions to Fishery Development Policy in Indonesia. Marine Resource Economics, 8(2): 177–181, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Conrad:1985:RMD

[Con85] Jon M. Conrad. Residuals management: Disposal of sewage sludge in the New York Bight. Marine Resource Economics, 1 (4):321–345, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cook:1988:DEC

[Coo88] B. A. Cook. Discount effects and Canada’s Pacific halibut fishery. Marine Resource Economics, 5(1):71–77, ???? 1988. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cox:2007:CAF

[Cox07] Anthony Cox. Capacity analysis and fisheries management: Is the tail wagging the dog? Marine Resource Economics, 22(1): 95–98, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Crowley:1992:RBF

[CP92] R. W. Crowley and H. Palsson. Rights based fisheries management in Canada. Marine Resource Economics, 7(2):1–21, Summer 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Coppola:1998:SPM

[CP98] Gianluigi Coppola and Sean Pascoe. A surplus production model with a nonlinear catch-effort relationship. Marine Resource Economics, 13(1):37–50, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chan:2016:SEE

[CP16] Hing Ling Chan and Minling Pan. Spillover effects of environmental regulation for sea turtle protection in the Hawaii longline swordfish fishery. Marine Resource Economics, 31(3):259–279, July 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Crowley:1992:BM

[CPB^{+}92] R. W. Crowley, H. Palsson, M. P. Burton, Antonio J. Díaz de León, Juan Carlos Seijo, Antonio J. Díaz de León, Juan Carlos Seijo, M. P. Burton, and J. E. Easley, Jr. Back matter. Marine Resource Economics, 7(2):??, Summer 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Coppola:1998:BM

[CPG^{+}98] Gianluigi Coppola, Sean Pascoe, Eyjólfur Gudmundsson, Stein Ivar Steinshamn, E. M. Thunberg, T. E. Helser, R. K. Mayo, Gianluigi Coppola, Sean Pascoe, and Peter W. Schuhmann. Back matter. Marine Resource Economics, 13(1):??, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Collins:2003:PEF

[CPW03] Alan Collins, Sean Pascoe, and David Whitmarsh. Pollution externalities and fisheries: Insights from a spatially explicit bioeconomic model. Marine Resource Economics, 18(4):313–328, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Conrad:2015:BMD

[CR15] Jon M. Conrad and Daniel Rondeau. Bioeconomics of a marine disease. Marine Resource Economics, 30(4):393–416, October 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Crosson:2016:ACA

[Cro16] Scott Crosson. The affordable care act and opportunities for change in North Carolina’s commercial fisheries. Marine Resource Economics, 31(2):121–129, April 2016. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chambers:1986:EPR

[CS86] Robert G. Chambers and Ivar E. Strand, Jr. Estimating parameters of a renewable resource model without population data. Marine Resource Economics, 2(3):263–274, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chavez:2013:WSP

[CS13] Carlos A. Chávez and John K. Stranlund. Who should pay the administrative costs of an ITQ fishery? Marine Resource Economics, 28(3):243–261, September 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Crosson:2023:EAC

[CSW23] Scott Crosson, Cameron Speir, and Christina Wiegand. The effects of the Affordable Care Act’s insurance subsidies on a commercial fishing fleet. Marine Resource Economics, 39(1):61– 71, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Chu:2014:LGO

[CT14] Jingjie Chu and Lacey Tudur. Looking to grow outside the United States. Marine Resource Economics, 29(4):323–337, December 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cumberland:1987:GES

[Cum87a] John H. Cumberland. Guest editorial special issue on the economics of Chesapeake Bay management. Marine Resource Economics, 4(2):79–80, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cumberland:1987:NED

[Cum87b] John H. Cumberland. Need economic development be hazardous to the health of the Chesapeake Bay? Marine Resource Economics, 4(2):81–93, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cunningham:1994:FIF

[Cun94] Stephen Cunningham. Fishermen’s incomes and fisheries management. Marine Resource Economics, 9(3):241–252, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Curlin:1985:BRG

[Cur85] James W. Curlin. Book review: Gerard J. Mangone, The Future of Gas and Oil from the Sea. Marine Resource Economics, 1 (4):429–434, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cancino:2007:TIC

[CUW07a] José P. Cancino, Hirotsugu Uchida, and James E. Wilen. TURFs and ITQs: Collective vs. individual decision making. Marine Resource Economics, 22(4):391–406, ???? 2007. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cancino:2007:BM

[CUW^{+}07b] José P. Cancino, Hirotsugu Uchida, James E. Wilen, Christopher Costello, Robert Deacon, Gary D. Libecap, José P. Cancino, Hirotsugu Uchida, James E. Wilen, Christopher Costello, Robert Deacon, and Ragnar Arnason. Back matter. Marine Resource Economics, 22(4):??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Christensen:1993:BAG

[CV93] Steen Christensen and Niels Vestergaard. A bioeconomic analysis of the Greenland shrimp fishery in the Davis Strait. Marine Resource Economics, 8(4):345–365, Winter 1993. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Clarke:1992:BAN

[CYP92] Raymond P. Clarke, Stacey S. Yoshimoto, and Samuel G. Pooley. A bioeconomic analysis of the Northwestern Hawaiian Islands lobster fishery. Marine Resource Economics, 7(3):115–140, Fall 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Danielsen:2020:PTP

[DAAY^{+}20] Rannvá Danielsen, Sveinn Agnarsson, Isaac Ankamah-Yeboah, Frank Asche, Julia Bronnmann, Max Nielsen, Rasmus Nielsen, Stefan B. Gunnlaugsson, Hördur Saevaldsson, Dadi M. Kristofersson, Sveinn Agnarsson, Rannvá Danielsen, Sveinn Agnarsson, Isaac Ankamah-Yeboah, Frank Asche, Julia Bronnmann, Max Nielsen, Rasmus Nielsen, Stefan B. Gunnlaugsson, Hördur Saevaldsson, Dadi M. Kristofersson, Sveinn Agnarsson, Rannvá Danielsen, and Sveinn Agnarsson. In pursuit of the three pillars of sustainability in fisheries: a Faroese case study. Marine Resource Economics, 35(2):177–193, April 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Danielsson:2000:EIP

[Dan00a] Asgeir Danielsson. Efficiency of ITQs in the presence of production externalities. Marine Resource Economics, 15(1):37–43, Spring 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Danielsson:2000:REI

[Dan00b] Asgeir Danielsson. Reply: Efficiency of ITQs in the presence of production externalities. Marine Resource Economics, 15 (3):245–251, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dey:2011:MBA

[DAP11] Madan Mohan Dey, Md. Ferdous Alam, and Ferdinand J. Paraguas. A multistage budgeting approach to the analysis of demand for fish: an application to inland areas of Bangladesh. Marine Resource Economics, 26(1):35–58, March 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Duarte:2000:MNA

[DBP00] Clara Costa Duarte, Ana Brasão, and Pedro Pintassilgo. Management of the Northern Atlantic bluefin tuna: an application of C-games. Marine Resource Economics, 15(1):21–36, Spring 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Davis:2019:EPV

[DBR^{+}19] Katrina J. Davis, Michael Burton, Abbie Rogers, Alaya Spencer-Cotton, and Ram Pandit. Eliciting public values for management of complex marine systems: an integrated choice experiment. Marine Resource Economics, 34(1):1–21, January 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dresdner:2015:CAF

[DCB15] Jorge Dresdner, Carlos Chávez, and Omar Barriga. Compliance in artisanal fisheries: Do morality, legitimacy, and peer effects matter? Marine Resource Economics, 30(4):349–370, October 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Duy:2016:PEF

[DF16] Nguyen Ngoc Duy and Ola Flaaten. Profitability effects and fishery subsidies: Average treatment effects based on propensity scores. Marine Resource Economics, 31(4):373–402, October 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

deFranca:2024:PFS

[dFdOdA24] Vinícius Fellype Cavalcanti de França, Luan Diego de Oliveira, and Humber Agrelli de Andrade. Price forecasting of a small pelagic species in a South American supply center: a machine learning approach. Marine Resource Economics, 39(2):145–162, ???? 2024. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dirlam:1994:RAE

[DG94] Joel Dirlam and Daniel Georgianna. Recent adjustments in the England fresh groundfish processing. Marine Resource Economics, 9(4):375–384, Winter 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dupont:2000:MSI

[DG00] Diane P. Dupont and R. Quentin Grafton. Multi-species individual transferable quotas: The Scotia–Fundy mobile gear groundfishery. Marine Resource Economics, 15(3):205–220, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Diaby:1996:EIA

[Dia96] Souleymane Diaby. Economic impact analysis of the ivorian sardinella fishery. Marine Resource Economics, 11(1):31–42, Spring 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Diop:1996:TMM

[DK96] Hamady Diop and Richard F. Kazmierczak, Jr. Technology and management in Mauritanian cephalopod fisheries. Marine Resource Economics, 11(2):71–84, Summer 1996. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Diagne:2004:EEC

[DKK04] Assane Diagne, Walter R. Keithly, Jr., and Richard F. Kazmierczak, Jr. The effect of environmental conditions and regulatory costs on oyster relaying in Louisiana. Marine Resource Economics, 19(2):211–224, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dedah:2011:AUO

[DKK11] Cheikhna Dedah, Walter R. Keithly, Jr., and Richard F. Kazmierczak, Jr. An analysis of US oyster demand and the influence of labeling requirements. Marine Resource Economics, 26(1):17–33, March 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

DePiper:2017:VES

[DLL17] Geret S. DePiper, Douglas W. Lipton, and Romuald N. Lipcius. Valuing ecosystem services: Oysters, denitrification, and nutrient trading programs. Marine Resource Economics, 32(1):1–20, January 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Diaz-de-Leon:1992:MCN

[dLS92] Antonio J. Díaz de León and Juan Carlos Seijo. A multi-criteria non-linear optimization model for the control and management of a tropical fishery. Marine Resource Economics, 7(2):23–40, Summer 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Duffield:2012:MBM

[DNA^{+}12] John Duffield, Chris Neher, Stewart Allen, David Patterson, and Brad Gentner. Modeling the behavior of marlin anglers in the Western Pacific. Marine Resource Economics, 27(4):343–357, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dahl:2014:FPV

[DO14] Roy Endre Dahl and Atle Oglend. Fish price volatility. Marine Resource Economics, 29(4):305–322, December 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Doll:1988:TEM

[Dol88] John P. Doll. Traditional economic models of fishing vessels: a review with discussion. Marine Resource Economics, 5(2):99– 123, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dow:1993:DRF

[Dow93] James P. Dow, Jr. Dynamic regulation of fisheries: the case of the bowhead whale. Marine Resource Economics, 8(2):145– 154, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dahl:2021:SSM

[DOY21] Roy Endré Dahl, Atle Oglend, and Muhammad Yahya. Salmon stock market prices revealing salmon price information. Marine Resource Economics, 36(2):173–190, April 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dresdner:2022:BRE

[DPP^{+}22] Jorge Dresdner, Lisa Pfeiffer, Tess Petesch, Thamanna Vasan, Ruth B. M. Pincinato, Frank Asche, Andreea L. Cojocaru, Yaqin Liu, Kristin H. Roll, David W. Carter, Christopher Liese, and Sabrina J. Lovell. Book review: The Economics of Fishing. By Rögnvaldur Hannesson. Marine Resource Economics, 37(1): 109–110, January 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dalton:2004:CRC

[DR04] Michael G. Dalton and Stephen Ralston. The California rockfish conservation area and groundfish trawlers at Moss Landing Harbor. Marine Resource Economics, 19(1):67–83, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

DeVoretz:1985:HCF

[DS85a] Don DeVoretz and Richard Schwindt. Harvesting Canadian fish and rents: a partial review of the report of the Commission on Canadian Pacific Fisheries Policy. Marine Resource Economics, 1(4):347–367, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Drynan:1985:IEO

[DS85b] Ross G. Drynan and Frances Sandiford. Incorporating economic objectives in goal programs for fishery management. Marine Resource Economics, 2(2):175–195, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dower:1987:CNR

[DS87] Roger C. Dower and Paul F. Scodari. Compensation for natural resource injury: an emerging Federal framework. Marine Resource Economics, 4(3):155–174, ???? 1987. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dunn:1988:ANS

[DS88] James W. Dunn and James S. Shortle. Agricultural nonpoint source pollution control in theory and practice. Marine Resource Economics, 5(3):259–270, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Danielsson:1997:UIC

[DSBT97] Asgeir Danielsson, Gunnar Stefansson, Fridrik M. Baldursson, and Kristjan Thorarinsson. Utilization of the Icelandic cod stock in a multispecies context. Marine Resource Economics, 12(4): 329–344, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Depalle:2020:AFH

[DTS20] Maxime Dépalle, Olivier Thébaud, and James N. Sanchirico. Accounting for fleet heterogeneity in estimating the impacts of large-scale fishery closures. Marine Resource Economics, 35(4): 361–378, October 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dupont:1993:PUE

[Dup93] Diane P. Dupont. Price uncertainty, expectations formation and fishers’ location choices. Marine Resource Economics, 8(3):219– 247, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Dundas:2018:RCE

[DvHM18] Steven J. Dundas, Roger H. von Haefen, and Carol Mansfield. Recreation costs of endangered species protection: Evidence from Cape Hatteras National Seashore. Marine Resource Economics, 33(1):1–25, January 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Emerson:1989:SAM

[EA89] William Emerson and James Anderson. A spatial allocation model for the New England fisheries. Marine Resource Economics, 6(2):123–144, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Elofsson:2025:APP

[EAÖH25] Katarina Elofsson, Tenaw G. Abate, Örjan Östman, and Magnus Huss. Accounting for predator–prey fisheries in the cost-effective management of eutrophicated coastal waters. Marine Resource Economics, 40(2):81–105, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Easley:1992:SIM

[Eas92] J. E. Easley, Jr. Selected issues in modeling allocation of fishery harvests. Marine Resource Economics, 7(2):41–56, Summer 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Edwards:1989:ECI

[EC89] Steven F. Edwards and Cynthia Carlson. On estimating compensation for injury to publicly owned marine resources. Marine Resource Economics, 6(1):27–42, ???? 1989. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Edwards:1992:ESC

[Edw92] Steven F. Edwards. Evidence of structural change in preferences for seafood. Marine Resource Economics, 7(3):141–151, Fall 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Edwards:1994:ORO

[Edw94] Steven F. Edwards. Ownership of renewable ocean resources. Marine Resource Economics, 9(3):253–273, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Edwards:2001:RSP

[Edw01] Steven E. Edwards. Rent-seeking and property rights formation in the U.S. Atlantic sea scallop fishery. Marine Resource Economics, 16(4):263–275, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Edwards:2005:ARU

[Edw05] Steve Edwards. Accounting for rents in the U.S. Atlantic sea scallop fishery. Marine Resource Economics, 20(1):61–76, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ekerhovd:2013:CSE

[EG13] Nils-Arne Ekerhovd and Daniel V. Gordon. Catch, stock elasticity, and an implicit index of fishing effort. Marine Resource Economics, 28(4):379–395, December 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eggert:2006:FEY

[Egg06] Håkan Eggert. Fisheries economics and 20 years with Marine Resource Economics: a citation analysis. Marine Resource Economics, 21(3):269–284, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eggert:2012:CJW

[Egg12] Håkan Eggert. The centenary of Jens Warming’s optimal landing tax in fisheries. Marine Resource Economics, 27(1):107–110, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Engle:1995:PCA

[EK95] Carole R. Engle and Pierre-Justin Kouka. Potential consumer acceptance of canned bighead carp: a structural model analysis. Marine Resource Economics, 10(2):101–116, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eggert:2007:SSF

[EL07] Håkan Eggert and Razack B. Lokina. Small-scale fishermen and risk preferences. Marine Resource Economics, 22(1):49–67, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eggert:2007:BM

[ELP^{+}07] Håkan Eggert, Razack B. Lokina, Sean Pascoe, Mitesh Kataria, Håkan Eggert, Razack B. Lokina, Rögnvaldur Hannesson, and Sean Pascoe. Back matter. Marine Resource Economics, 22(1): ??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ewald:2017:AFP

[EOA^{+}17] Christian-Oliver Ewald, Ruolan Ouyang, Claire W. Armstrong, Viktoria Kahui, Godwin K. Vondolia, Margrethe Aanesen, Mikołaj Czajkowski, Maryam Tabarestani, Walter R. Keithly, Jr., Hassan Marzoughi-Ardakani, Isaac Ankamah-Yeboah, Lisa Ståhl, Max Nielsen, Isaac Ankamah-Yeboah, Lisa Ståhl, Max Nielsen, Claire W. Armstrong, Viktoria Kahui, Godwin K. Vondolia, Margrethe Aanesen, Mikołaj Czajkowski, Carole Ropars-Collet, Mélody Leplat, and Philippe Le Goffe. An analysis of the fish pool market in the context of seasonality and stochastic convenience yield. Marine Resource Economics, 32(4):431–449, October 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Emami:1994:TRT

[EQJ94] Ali Emami, Lewis E. Queirolo, and Richard S. Johnston. Trade restrictions and trade reversal: Lessons from the U.S.–Canada herring dispute. Marine Resource Economics, 9(1):31–45, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ekerhovd:2016:EBM

[ES16] Nils-Arne Ekerhovd and Stein Ivar Steinshamn. Economic benefits of multi-species management: The pelagic fisheries in the Northeast Atlantic. Marine Resource Economics, 31(2):193–210, April 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Estrada:2018:ECR

[ESC18] Gloria Angélica Chávez Estrada, Miguel Ángel Quiroga Suazo, and Jorge David Dresdner Cid. The effect of collective rightsbased management on technical efficiency: The case of Chile’s common sardine and anchovy fishery. Marine Resource Economics, 33(1):87–112, January 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eide:2003:BM

[ESO^{+}03] Arne Eide, Frode Skjold, Frank Olsen, Ola Flaaten, Cathy A. Roheim, Kathleen P. Bell, Daniel Huppert, Rebecca L. Johnson, Arne Eide, Frode Skjold, Frank Olsen, Ola Flaaten, Kathleen P. Bell, Daniel Huppert, Rebecca L. Johnson, and Cathy A. Roheim. Back matter. Marine Resource Economics, 18(1):??, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eide:2003:HFN

[ESOF03] Arne Eide, Frode Skjold, Frank Olsen, and Ola Flaaten. Harvest functions: The Norwegian bottom trawl cod fisheries. Marine Resource Economics, 18(1):81–93, ???? 2003. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Espinoza:2015:SCC

[Esp15] César Salazar Espinoza. Share contract choices and economic performance: Empirical evidence from the artisanal fisheries sector in Chile. Marine Resource Economics, 30(1):71–95, January 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eggert:1999:BAS

[EU99] Håkan Eggert and Mats Ulmestrand. A bioeconomic analysis of the Swedish fishery for Norway lobster (Nephrops norvegicus). Marine Resource Economics, 14(3):225–244, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eero:2005:QAE

[EVH05] Margit Eero, Markus Vetemaa, and Rögnvaldur Hannesson. The quota auctions in Estonia and their effect on the trawler fleet. Marine Resource Economics, 20(1):101–112, ???? 2005. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eales:1986:EFL

[EW86] James Eales and James E. Wilen. An examination of fishing location choice in the pink shrimp fishery. Marine Resource Economics, 2(4):331–351, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Eales:1986:BM

[EWR^{+}86] James Eales, James E. Wilen, Edward J. Richardson, John M. Gates, John M. Gates, James Eales, James E. Wilen, Edward J. Richardson, and John M. Gates. Back matter. Marine Resource Economics, 2(4):??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Felthoven:2002:BM

[FAH^{+}02] Ronald G. Felthoven, Frank Asche, Rögnvaldur Hannesson, David Whitmarsh, Carl James, Helen Pickering, Carlo Pipitone, Fabio Badalamenti, Giovanni D’Anna, Ronald G. Felthoven, Frank Asche, Rögnvaldur Hannesson, and Lee G. Anderson. Back matter. Marine Resource Economics, 17(3):??, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Felixson:1994:BRW

[Fel94] Tryggvi Felixson. Book review: Will. C. van den Hoonaard, Reluctant Pioneers, Constraints and Opportunities in an Icelandic Fishing Community. Marine Resource Economics, 9(1):95–97, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Felthoven:2002:EAF

[Fel02] Ronald G. Felthoven. Effects of the American fisheries act on capacity, utilization and technical efficiency. Marine Resource Economics, 17(3):181–205, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fell:2009:EVP

[Fel09] Harrison Fell. Ex-vessel pricing and IFQs: a strategic approach. Marine Resource Economics, 24(4):311–328, ???? 2009. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Freeman:1995:BM

[FFH^{+}95] A. Myrick Freeman III, Ola Flaaten, Knut Heen, Kjell G. Salvanes, Ronald N. Johnson, A., Henry W. Kinnucan, Robert G. Nelson, Hui Xiao, and Ola Flaaten. Back matter. Marine Resource Economics, 10(4):??, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Funk:1998:MIS

[FGM^{+}98] Robert D. Funk, Wade L. Griffin, James W. Mjelde, Teofilo Ozuna, Jr., and John M. Ward. A method of imputing and simulating costs and returns in fisheries. Marine Resource Economics, 13(3):171–183, Fall 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Funk:2003:SML

[FGMW03] Robert D. Funk, Wade L. Griffin, Sr., James W. Mjelde, and John M. Ward. A simulation model of license buyback in the Texas bay shrimp fishery. Marine Resource Economics, 18(1): 33–53, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fletcher:1987:MMH

[FHJ87] Jerald J. Fletcher, Richard E. Howitt, and Warren E. Johnston. Management of multipurpose heterogeneous fishing fleets under uncertainty. Marine Resource Economics, 4(4):249–270, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Flaaten:2017:PRR

[FHM17] Ola Flaaten, Knut Heen, and Thórólfur Matthíasson. Profit and resource rent in fisheries. Marine Resource Economics, 32 (3):311–328, July 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Flaaten:1995:IRR

[FHS95] Ola Flaaten, Knut Heen, and Kjell G. Salvanes. The invisible resource rent in limited entry and quota managed fisheries: The case of Norwegian purse seine fisheries. Marine Resource Economics, 10(4):341–356, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Frost:2003:CWO

[FJ03] Hans Frost and Frank Jensen. Could welfare optimality be achieved with extensive data information? Marine Resource Economics, 18(3):275–286, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fofana:2008:MOP

[FJ08] Abdulai Fofana and Shabbar Jaffry. Measuring oligopsony power of UK salmon retailers. Marine Resource Economics, 23(4):485– 506, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Farolfi:2023:UPP

[FJ23] Giulio Farolfi and Robert J. Johnston. Understanding public preferences for molluscan shellfish aquaculture: The role of production technology and environmental impacts. Marine Resource Economics, 39(1):39–60, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Foley:2010:ELB

[FKAvR10] Naomi S. Foley, Viktoria Kahui, Claire W. Armstrong, and Tom M. van Rensburg. Estimating linkages between redfish and cold water coral on the Norwegian Coast. Marine Resource Economics, 25(1):105–120, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Felthoven:2014:CFP

[FLS14] Ronald G. Felthoven, Jean Lee, and Kurt E. Schnier. Cooperative formation and peer effects in fisheries. Marine Resource Economics, 29(2):133–156, June 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ferrara:1996:TRR

[FM96] Ida Ferrara and Paul C. Missios. Transboundary renewable resource management: a dynamic game with differing noncooperative payoffs. Marine Resource Economics, 11(4):239–245, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fousekis:2002:DVR

[Fou02] Panos Fousekis. Distance vs. ray functions: an application to the inshore fishery of Greece. Marine Resource Economics, 17 (4):251–267, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fielding:2025:PVP

[FP25] Samuel Fielding and George Parsons. Private versus public flood insurance: Differences in premiums and uptake observed in two coastal housing markets using survey data. Marine Resource Economics, 40(1):1–22, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Falk-Petersen:2013:OCE

[FPA13] Jannike Falk-Petersen and Claire W. Armstrong. To have one’s cake and eat it too managing the alien invasive red king crab. Marine Resource Economics, 28(1):65–81, January 2013. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ferrer-Perez:2020:APV

[FPdR20] Hugo Ferrer-Pérez and Pilar Gracia de Rentería. Asymmetric price volatility transmission in the Spanish fresh wild fish supply chain. Marine Resource Economics, 35(1):65–81, January 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Farrow:1992:PIE

[FR92] Scott Farrow and Marshall Rose. Public information externalities: Estimates for offshore energy exploration. Marine Resource Economics, 7(2):67–82, Summer 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fousekis:2004:RFD

[FR04] Panos Fousekis and Brian J. Revell. Retail fish demand in Great Britain and its fisheries management implications. Marine Resource Economics, 19(4):495–510, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Freeman:1995:BWQ

[Fre95] A. Myrick Freeman III. The benefits of water quality improvements for marine recreation: a review of the empirical evidence. Marine Resource Economics, 10(4):385–406, Winter 1995. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fousekis:2004:BM

[FRK^{+}04] Panos Fousekis, Brian J. Revell, Lone Groenbaek Kronbak, Shabbar Jaffry, Shabbar Jaffry, Panos Fousekis, Brian J. Revell, John Ward, Pamela Mace, and Eric Thunberg. Back matter. Marine Resource Economics, 19(4):??, ???? 2004. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Frost:1984:FMU

[Fro84] Hans Frost. Fisheries management and uncertainty within the EEC. Marine Resource Economics, 1(1):97–103, ???? 1984. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Fonner:2015:WPM

[FS15] Robert Fonner and Gil Sylvia. Willingness to pay for multiple seafood labels in a niche market. Marine Resource Economics, 30(1):51–70, January 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Freeman:2010:SLB

[FW10] Matthew A. Freeman and Richard T. Woodward. Sequential license buyback auctions: an experimental analysis. Marine Resource Economics, 25(4):355–372, December 2010. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gu:1995:DSS

[GA95] Guang Gu and James L. Anderson. Deseasonalized state-space time series forecasting with application to the US salmon market. Marine Resource Economics, 10(2):171–185, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grigalunas:1986:ECO

[GAB^{+}86a] Thomas A. Grigalunas, Robert C. Anderson, Gardner M. Brown, Jr., Richard Congar, Norman F. Meade, and Philip E. Sorensen. Estimating the cost of oil spills: Lessons from the Amoco Cadiz incident. Marine Resource Economics, 2(3):239– 262, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grigalunas:1986:BM

[GAB^{+}86b] Thomas A. Grigalunas, Robert C. Anderson, Gardner M. Brown, Jr., Richard Congar, Norman F. Meade, Philip E. Sorensen, Daniel L. Georgianna, William V. Hogan, George B. Assaf, Brent G. Kroetch, Brent C. Kroetch, Subodh C. Mathur, Thomas A. Grigalunas, Robert C. Anderson, Gardner M. Brown, Jr., Richard Congar, Norman F. Meade, Philip E. Sorensen, George B. Assaf, Brent G. Kroetch, Brent C. Kroetch, and Subodh C. Mathur. Back matter. Marine Resource Economics, 2(3):??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gardner:1988:EAS

[Gar88] Michael Gardner. Enterprise allocation system in the offshore groundfish sector in Atlantic Canada. Marine Resource Economics, 5(4):389–414, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Garrod:1993:PSE

[Gar93] Brian Garrod. Pavel salz, the European Atlantic fisheries: Structure, economic performance and policy. Marine Resource Economics, 8(1):82–84, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gates:1984:PTU

[Gat84] J. M. Gates. Principal types of uncertainty in fishing operations. Marine Resource Economics, 1(1):31–49, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gates:1986:BRP

[Gat86] John M. Gates. Book review: P. Geoffrey Allen, Louis W. Botsford, Anthonie M. Schuur, and Warren E. Johnston, Bioeconomics of Aquaculture Developments in Aquaculture and Fisheries Science, Vol. 13. Marine Resource Economics, 2(4):383– 385, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gates:1996:BRJ

[Gat96] John M. Gates. Book review: Jorgen W. Weibull, Evolutionary Game Theory. Marine Resource Economics, 11(4):313–317, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guldin:2018:CSS

[GAW^{+}18] Marie Guldin, Christopher M. Anderson, James E. Wilen, Bård Misund, Rune Nygård, Joshua K. Abbott, James N. Sanchirico, Martin D. Smith, Bård Misund, Rune Nygård, James E. Wilen, Marie Guldin, and Christopher M. Anderson. Catch shares and shoreside processors: a costs and earnings exploration into the downstream sector. Marine Resource Economics, 33(3):289–307, July 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gordon:2008:IFS

[GBDT08] Daniel V. Gordon, Trond Bjørndal, Madan M. Dey, and Rezaul Karim Talukder. An intra-farm study of production factors and productivity for shrimp farms in Bangladesh: an index approach. Marine Resource Economics, 23(4):411–424, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gokhale:2025:RNF

[GBL^{+}25] Sanmitra Gokhale, Johan Blomquist, Martin Lindegren, Andries Richter, and Staffan Waldo. The role of non-fishing and partner incomes in managing fishers’ economic risk. Marine Resource Economics, 39(4):285–305, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grainger:2016:DET

[GC16] Corbett A. Grainger and Christopher Costello. Distributional effects of the transition to property rights for a common-pool resource. Marine Resource Economics, 31(1):1–26, January 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gillig:1998:SEV

[GCG98] Dhazn Gillig, Oral Capps, Jr., and Wade L. Griffin. Shrimp exvessel prices landed from the Gulf of Mexico. Marine Resource Economics, 13(2):89–102, Summer 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guyader:2005:CSI

[GD05] Olivier Guyader and Fabienne Daurès. Capacity and scale inefficiency: Application of data envelopment analysis in the case of the French seaweed fleet. Marine Resource Economics, 20 (4):347–365, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guyader:2004:BAI

[GDF04] Olivier Guyader, Fabienne Daures, and Spyros Fifas. A bioeconomic analysis of the impact of decommissioning programs: Application to a limited-entry French scallop fishery. Marine Resource Economics, 19(2):225–242, ???? 2004. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Georgianna:1987:BRP

[Geo87] Daniel L. Georgianna. Book review: Peter B. Doeringer, Philip I. Moss, and David G. Terkla, The New England Fishing Economy: Jobs, Income, and Kinship. Marine Resource Economics, 4(2):139–143, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Georgianna:1996:BRP

[Geo96] Daniel L. Georgianna. Book review: Peter B. Doeringer and David G. Terkla, Troubled Waters: Economic Structure, Regulatory Reform, and Fisheries Trade. Marine Resource Economics, 11(3):219–220, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Georgianna:2006:BRB

[Geo06] Daniel Georgianna. Book review: Bonnie J. McCay, Oyster Wars and the Public Trust: Property, Law, and Ecology in New Jersey History. Marine Resource Economics, 21(1):125– 127, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guignet:2017:IPA

[GGKW17] Dennis Guignet, Charles Griffiths, Heather Klemick, and Patrick J. Walsh. The implicit price of aquatic grasses. Marine Resource Economics, 32(1):21–41, January 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Georgianna:1986:PCA

[GH86] Daniel L. Georgianna and William V. Hogan. Production costs in Atlantic fresh fish processing. Marine Resource Economics, 2(3):275–292, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gordon:1996:PFF

[GH96] Daniel V. Gordon and Rögnvaldur Hannesson. On prices of fresh and frozen cod fish in European and U.S. markets. Marine Resource Economics, 11(4):223–238, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gordon:1993:TOS

[GHB93] D. V. Gordon, R. Hannesson, and S. Bibb. Testing for output substitution possibilities in cod fish processing in Norway.

Marine Resource Economics, 8(1):17–30, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Greaker:2015:BRE

[GHH^{+}15] Mads Greaker, Pei Huang, Cecilia Hammarlund, Daniel K. Lew, Amber Himes-Cornell, Jean Lee, Cecilia Hammarlund, Jordi Guillen, Claire Macher, Mathieu Merzéréaud, Jean Boncoeur, Olivier Guyader, and Pei Huang. Book review: Ecofundamentalism: a Critique of Extreme Environmentalism. By Rögnvaldur Hannesson. Marine Resource Economics, 30(2):231–234, April 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Greenberg:1995:ISD

[GHM95] Joshua A. Greenberg, Mark Herrmann, and John McCracken. An international supply and demand model for Alaska snow crab. Marine Resource Economics, 10(3):231–246, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Greenville:2006:TEI

[GHM06] Jared Greenville, Jessica Hartmann, and T. Gordon Macaulay. Technical efficiency in input-controlled fisheries: The NSW ocean prawn trawl fishery. Marine Resource Economics, 21 (2):159–179, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gilbert:1988:USA

[Gil88] David John Gilbert. Use of a simple age-structured bioeconomic model to estimate optimal long-run surpluses. Marine Resource Economics, 5(1):23–42, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gallastegui:2002:BFR

[GIP02] M. Carmen Gallastegui, Elena Iñarra, and Raul Prellezo. Bankruptcy of fishing resources: The Northern European anglerfish fishery. Marine Resource Economics, 17(4):291–307, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gianessi:1988:WIR

[GKK^{+}88] Leonard P. Gianessi, Raymond J. Kopp, Raymond J. Koop, Peter Kuch, Cynthia Puffer, and Robert Torla. Welfare implications of restricted triazine herbicide use in the Chesapeake Bay region. Marine Resource Economics, 5(3):243–258, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Groeneveld:2010:FTI

[GLL^{+}10] Rolf A. Groeneveld, Daniel K. Lew, David F. Layton, Robert D. Rowe, Edward C. Waters, Chang K. Seung, Claire Macher, Jean Boncoeur, Daniel K. Lew, David F. Layton, Robert D. Rowe, Edward C. Waters, Chang K. Seung, Claire Macher, and Jean Boncoeur. Framing and training to induce preference learning in choice experiments. Marine Resource Economics, 25(2):233– 245, June 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gronbaek:2025:CCG

[GLM^{+}25] Lone Grønbæk, Marko Lindroos, Gordon Munro, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, FRANK ASCHE, Jay Abolofia, Frank Asche, James E. Wilen, Robert Fonner, and Gil Sylvia. Cooperation and club goods: Fisheries management in the spirit of Elinor Ostrom. Marine Resource Economics, 40(1):65–79, ???? 2025. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guillotreau:2023:PTB

[GLN^{+}23] Patrice Guillotreau, Frédéric Lantz, Lesya Nadzon, Jonathan Rault, and Olivier Maury. Price transmission between energy and fish markets: Are oil rates good predictors of tuna prices? Marine Resource Economics, 38(1):29–46, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Greenville:2007:UBP

[GM07] Jared Greenville and T. Gordon Macaulay. Untangling the benefits of protected areas in fisheries. Marine Resource Economics, 22(3):267–285, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guillen:2015:ESR

[GMM^{+}15] Jordi Guillen, Claire Macher, Mathieu Merzéréaud, Jean Boncoeur, and Olivier Guyader. Effects of the share remuneration system on fisheries management targets and rent distribution. Marine Resource Economics, 30(2):123–138, April 2015. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Greene:1997:DRF

[GMS97a] Gretchen Greene, Charles B. Moss, and Thomas H. Spreen. Demand for recreational fishing in Tampa Bay, Florida: a random utility approach. Marine Resource Economics, 12(4):293–305, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Greene:1997:BM

[GMS^{+}97b] Gretchen Greene, Charles B. Moss, Thomas H. Spreen, Julio Peña-Torres, Matthew Berman, Sharman Haley, Hongjin Kim, Julio Peña-Torres, Gretchen Greene, Charles B. Moss, Thomas H. Spreen, Asgeir Danielsson, Gunnar Stefansson, Fridrik M. Baldursson, and Kristjan Thorarinsson. Back matter. Marine Resource Economics, 12(4):??, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guttormsen:2011:ISC

[GMT11] Atle G. Guttormsen, Øystein Myrland, and Ragnar Tveteras. Innovations and structural change in seafood markets and production: Special issue introduction. Marine Resource Economics, 26(4):247–253, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gupta:2019:MPE

[GMTW19] Rangan Gupta, Zinnia Mukherjee, Mike G. Tsionas, and Peter Wanke. Measuring the productive efficiency of the Connecticut Long Island Lobster Sound fishery using a novel finite mixture model. Marine Resource Economics, 34(3):267–285, July 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Geen:1988:ITQ

[GN88] Gerry Geen and Mark Nayar. Individual transferable quotas in the Southern bluefin tuna fishery: an economic appraisal. Marine Resource Economics, 5(4):365–387, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Geen:1988:BM

[GNC^{+}88] Gerry Geen, Mark Nayar, Ian N. Clark, Philip J. Major, Nina Mollett, James E. Wilen, Kenneth Ruddle, Anthony Scott, and James E. Wilen. Back matter. Marine Resource Economics, 5 (4):??, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grigalunas:1988:MDM

[GOFR88] Thomas A. Grigalunas, James J. Opaluch, Deborah French, and Mark Reed. Measuring damages to marine natural resources from pollution incidents under CERCLA: Applications of an integrated ocean systems/economic model. Marine Resource Economics, 5(1):1–21, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gillig:2000:VGM

[GOG00] Dhazn Gillig, Teofilo Ozuna, Jr., and Wade L. Griffin. The value of the Gulf of Mexico recreational red snapper fishery. Marine Resource Economics, 15(2):127–139, Summer 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gomersall:1992:OAF

[Gom92] C. Nicholas Gomersall. Open-access fishery models: Relaxing a constraint and removing an econometric obstacle. Marine Resource Economics, 7(1):59–74, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gonzalez:1996:TUR

[Gon96] Exequiel González. Territorial use rights in Chilean fisheries. Marine Resource Economics, 11(3):211–218, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gordon:2020:SRA

[Gor20] Daniel V. Gordon. A short-run ARDL-bounds model for forecasting and simulating the price of lobster. Marine Resource Economics, 35(1):43–63, January 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gordon:2013:BRP

[GPC^{+}13] Daniel V. Gordon, George R. Parsons, Zhe Chen, Michael K. Hidrue, Naomi Standing, Jonathan Lilley, Véronique Le Bihan, Sophie Pardo, Patrice Guillotreau, James E. Wilen, George R.

Parsons, Zhe Chen, Michael K. Hidrue, Naomi Standing, Jonathan Lilley, Véronique Le Bihan, Sophie Pardo, Patrice Guillotreau, and James E. Wilen. Book review: Preference Data for Environmental Valuation: Combining revealed and stated approaches. Edited by John Whitehead, Tim Haab, and Ju-Chin Huang. Marine Resource Economics, 28(3):307–309, September 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guttormsen:2011:TEH

[GR11] Atle G. Guttormsen and Kristin H. Roll. Technical efficiency in a heterogeneous fishery: The case of Norwegian groundfish fisheries. Marine Resource Economics, 26(4):293–307, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grafton:1996:ITQ

[Gra96] R. Quentin Grafton. Implications of taxing quota value in an individual transferable quota fishery: Comment. Marine Resource Economics, 11(2):125–127, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Greaker:2025:ISF

[GRA^{+}25] Mads Greaker, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, FRANK ASCHE, Jay Abolofia, Frank Asche, James E. Wilen, Robert Fonner, and Gil Sylvia. Innovation in salmon farming: Enhancing profitability or reducing environmental hazards? Marine Resource Economics, 39(4):333–347, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Green:1999:BRR

[Gre99] Christopher Green. Book review: Robert Mendelsohn and James E. Neumann, The Impact of Climate Change on the United States Economy. Marine Resource Economics, 14(2): 175–177, Summer 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Griffin:1993:BRF

[Gri93] Wade L. Griffin. Book review: F. Harold, Peter Hoar, and Melissa Upton, The Gulf of Mexico Shrimp Fishery: Profile of a Valuable Natural Resource. Marine Resource Economics, 8 (4):407–408, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grigalunas:2000:BRI

[Gri00] Thomas A. Grigalunas. Book review: Ing-Marie Gren, R. Kerry Turner, and Fredrik Wulff, Managing a Sea: The Ecological Economics of the Baltic. Marine Resource Economics, 15(3):255– 259, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gudmundsson:1998:CMB

[GS98] Eyjólfur Gudmundsson and Jon G. Sutinen. Challenges of marine biodiversity. Marine Resource Economics, 13(1):75–83, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gordon:1993:FFF

[GSA93] Daniel V. Gordon, Kjell G. Salvanes, and Frank Atkins. A fish is a fish is a fish? Testing for market linkages on the Paris fish market. Marine Resource Economics, 8(4):331–343, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grafton:1997:HMN

[GSE97] R. Quentin Grafton and J. Silva-Echenique. How to manage nature? Strategies, predator–prey models, and chaos. Marine Resource Economics, 12(2):127–143, Summer 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Grafton:1997:BM

[GSEM^{+}97] R. Quentin Grafton, J. Silva-Echenique, Thorolfur Matthiasson, Ussif Rashid Sumaila, R. Quentin Grafton, J. Silva-Echenique, Thorolfur Matthiasson, and Ussif Rashid Sumaila. Back matter. Marine Resource Economics, 12(2):??, Summer 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gren:2013:CES

[GSJ13] Ing-Marie Gren, Oleg. P. Savchuk, and Torbjörn Jansson. Costeffective spatial and dynamic management of a eutrophied Baltic Sea. Marine Resource Economics, 28(3):263–284, September 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gunnlaugsson:2020:RRD

[GSKA20] Stefan B. Gunnlaugsson, Hördur Saevaldsson, Dadi M. Kristofersson, and Sveinn Agnarsson. Resource rent and its distribution in Iceland’s fisheries. Marine Resource Economics, 35(2):113–135, April 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gilliland:2022:BLG

[GST22] Ted E. Gilliland, James N. Sanchirico, and J. Edward Taylor. A bioeconomic local general equilibrium assessment of distributional consequences of small-scale fisheries reform in developing countries. Marine Resource Economics, 37(2):111–134, April 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Georgianna:2011:BR

[GTP^{+}11] Daniel Georgianna, Sigbjorn Tveteras, Carlos E. Paredes, Julio Peña-Torres, John C. Whitehead, Christopher F. Dumas, Craig E. Landry, Jim Herstine, Andrew Muhammad, Keithly G. Jones, Sigbjorn Tveteras, Carlos E. Paredes, Julio Peña-Torres, Zinnia Mukherjee, Kathleen Segerson, Andrew Muhammad, and Keithly G. Jones. Books reviewed. Marine Resource Economics, 26(3):243–245, September 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gudmundsson:1998:MKC

[Gud98] Eyjólfur Gudmundsson. Mark kurlansky, cod: a biography of the fish that changed the world. Marine Resource Economics, 13(1):85–87, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gulland:1989:CGP

[Gul89] J. A. Gulland. Comments on Giulio Pontecorvo’s “The State of Worldwide Fishery Statistics: a Modest Proposal”. Marine Resource Economics, 6(1):85–86, ???? 1989. CODEN JMREDD.

ISSN 0738-1360 (print), 2334-5985 (electronic). See [Pon88] and reply [Pon89].

Guttormsen:2002:IFS

[Gut02] Atle G. Guttormsen. Input factor substitutability in salmon aquaculture. Marine Resource Economics, 17(2):91–102, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Guttormsen:2008:FSO

[Gut08] Atle G. Guttormsen. Faustmann in the sea: Optimal rotation in aquaculture. Marine Resource Economics, 23(4):401–410, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Goodlad:1992:FMS

[GW92] John H. Goodlad and James Wilson. Fisheries management by sectoral quotes — an assessment of the UK system. Marine Resource Economics, 7(2):83–91, Summer 1992. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gudmundsson:2000:ESS

[GW00] Eyjólfur Gudmundsson and Cathy R. Wessells. Ecolabeling seafood for sustainable production: Implications for fisheries management. Marine Resource Economics, 15(2):97–113, Summer 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gauvin:1994:DEW

[GWB94a] John R. Gauvin, John M. Ward, and Edward E. Burgess. Description and evaluation of the wreckfish (Polyprion americanus) fishery under individual transferable quotas. Marine Resource Economics, 9(2):99–118, Summer 1994. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gauvin:1994:BM

[GWB^{+}94b] John R. Gauvin, John M. Ward, Edward E. Burgess, Di Jin, Yoshihiro Kuronuma, Clement A. Tisdell, John R. Gauvin, John M. Ward, Edward E. Burgess, Lee G. Anderson, James E. Conklin, and William C. Kolberg. Back matter. Marine Resource Economics, 9(2):??, Summer 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Gudmundsson:2000:BM

[GWR^{+}00] Eyjólfur Gudmundsson, Cathy R. Wessells, Julie A. Reynolds, James E. Wilen, Dhazn Gillig, Teofilo Ozuna, Jr., Wade L. Griffin, Eyjólfur Gudmundsson, Cathy R. Wessells, Julie A. Reynolds, James E. Wilen, Dhazn Gillig, Teofilo Ozuna, Jr., and Wade L. Griffin. Back matter. Marine Resource Economics, 15(2):??, Summer 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hallwood:2004:PAO

[Hal04] Paul Hallwood. Protected areas, optimal policing and optimal rent dissipation. Marine Resource Economics, 19(4):481–493, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hallwood:2016:IPL

[Hal16] Paul Hallwood. International public law and the failure to efficiently manage ocean living resources. Marine Resource Economics, 31(2):131–139, April 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hamilton:1998:SCS

[Ham98] Marcia Hamilton. A system for classifying small boat fishermen in Hawaii. Marine Resource Economics, 13(4):289–291, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hammarlund:2015:BBA

[Ham15] Cecilia Hammarlund. The big, the bad, and the average: Hedonic prices and inverse demand for Baltic cod. Marine Resource Economics, 30(2):157–177, April 2015. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1984:FMU

[Han84] Rögnvaldur Hannesson. Fisheries management and uncertainty. Marine Resource Economics, 1(1):89–96, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1985:EFM

[Han85a] Rögnvaldur Hannesson. The effects of a fishermen’s monopoly in the market for unprocessed fish. Marine Resource Economics, 2(1):75–85, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1985:ITG

[Han85b] Rögnvaldur Hannesson. Inefficiency through government regulations: The case of Norway’s fishery policy. Marine Resource Economics, 2(2):115–141, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1986:EDR

[Han86] Rögnvaldur Hannesson. The effect of the discount rate on the optimal exploitation of renewable resources. Marine Resource Economics, 3(4):319–329, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1988:FVC

[Han88] Rögnvaldur Hannesson. Fixed or variable catch quotas? The importance of population dynamics and stock dependent costs. Marine Resource Economics, 5(4):415–432, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1993:FCH

[Han93] Rögnvaldur Hannesson. Fishing capacity and harvest rules. Marine Resource Economics, 8(2):133–143, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1998:MRW

[Han98] Rögnvaldur Hannesson. Marine reserves: What would they accomplish? Marine Resource Economics, 13(3):159–170, Fall 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2007:NSE

[Han07a] Rögnvaldur Hannesson. A note on the “Stock Effect”. Marine Resource Economics, 22(1):69–75, ???? 2007. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2007:TII

[Han07b] Rögnvaldur Hannesson. Taxes, ITQs, investments, and revenue sharing. Marine Resource Economics, 22(4):363–371, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2010:RDG

[Han10] Rögnvaldur Hannesson. The “Rent Drain”: a good measure of the gains from better resource management? Marine Resource Economics, 25(1):3–10, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2013:SMF

[Han13] Rögnvaldur Hannesson. Sharing a migrating fish stock. Marine Resource Economics, 28(1):1–17, January 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2015:WFC

[Han15] Rögnvaldur Hannesson. World fisheries in crisis? Marine Resource Economics, 30(3):251–260, July 2015. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2025:SSS

[Han25] Rögnvaldur Hannesson. Stock size and stock density. Marine Resource Economics, 39(3):229–242, ???? 2025. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hatcher:2005:MQM

[Hat05] Aaron Hatcher. On the microeconomics of quota management in fisheries. Marine Resource Economics, 20(1):77–99, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:1996:MRF

[HB96] Daniel S. Holland and Richard J. Brazee. Marine reserves for fisheries management. Marine Resource Economics, 11(3):157– 171, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2016:DPG

[HBAY^{+}16] Daniel S. Holland, Julia Bronnmann, Isaac Ankamah-Yeboah, Max Nielsen, Daniel S. Holland, Nguyen Ngoc Duy, Ola Flaaten, Julia Bronnmann, Isaac Ankamah-Yeboah, Max Nielsen, Sturla Furunes Kvamsdal, and Daniel S. Holland. Development of the Pacific groundfish trawl IFQ market. Marine Resource Economics, 31(4):453–464, October 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:1996:BM

[HBB^{+}96] Daniel S. Holland, Richard J. Brazee, Shekar Bose, Alistair McIlgorm, Exequiel González, Daniel S. Holland, Richard J. Brazee, Shekar Bose, Alistair McIlgorm, and Exequiel González. Back matter. Marine Resource Economics, 11(3):??, Fall 1996. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hunt:2007:ACS

[HBB07a] Len M. Hunt, Peter C. Boxall, and Barry Boots. Accommodating complex substitution patterns in a random utility model of recreational fishing. Marine Resource Economics, 22(2):155– 172, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hunt:2007:BMD

[HBB^{+}07b] Len M. Hunt, Peter C. Boxall, Barry Boots, Martin D. Smith, Junjie Zhang, Felicia C. Coleman, Lone Grønbæk Kronbak, Marko Lindroos, Max Nielsen, Jari Setälä, Jukka Laitinen, Kaija Saarni, Jarno Virtanen, Asmo Honkanen, Lone Grønbæk Kronbak, Marko Lindroos, Len M. Hunt, Peter C. Boxall, and Barry Boots. Back matter: Dr. S.-Y. Hong Award for outstanding article in Marine Resoure Economics. Marine Resource Economics, 22(2):??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hu:2009:EVD

[HBCP09] Wuyang Hu, Katya Boehle, Linda Cox, and Minling Pan. Economic values of dolphin excursions in Hawaii: a stated choice analysis. Marine Resource Economics, 24(1):61–76, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hammarlund:2022:WWB

[HBW22] Cecilia Hammarlund, Johan Blomquist, and Staffan Waldo. The way the wind blows: Tracing out the demand for Norwegian lobster using instrumental variables. Marine Resource Economics, 37(3):263–282, July 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Herrmann:2006:EMM

[HC06] Mark Herrmann and Keith Criddle. An econometric market model for the Pacific halibut fishery. Marine Resource Economics, 21(2):129–158, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hsu:2023:MID

[HCC23] Chun-Fu Hsu, Yun-Ju Chen, and Kuo-I Chang. Modeling import demand for fishery products in Japan: a dynamic AIDS approach. Marine Resource Economics, 38(2):135–151, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1998:BM

[HCL^{+}98] Rögnvaldur Hannesson, James C. Cato, Carlos A. Lima Dos Santos, Carsten Lynge Jensen, Rögnvaldur Hannesson, James C. Cato, Carlos A. Lima Dos Santos, and Carsten Lynge Jensen. Back matter. Marine Resource Economics, 13(3):??, Fall 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hamilton:1996:HLV

[HCT96] Marcia S. Hamilton, Rita E. Curtis, and Michael D. Travis. Hawaii longline vessel economics. Marine Resource Economics, 11(2):137–140, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hashida:2025:BCM

[HD25] Yukiko Hashida and Steven J. Dundas. Barriers to coastal managed retreat: Evidence from New Jersey’s Blue Acres Program. Marine Resource Economics, 39(3):179–205, ???? 2025. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hanna:2011:ESF

[HDK^{+}11] Susan Hanna, Cheikhna Dedah, Walter R. Keithly, Jr., Richard F. Kazmierczak, Jr., Madan Mohan Dey, Md. Ferdous Alam, Ferdinand J. Paraguas, Anthony Scott, Madan Mohan Dey, Md. Ferdous Alam, Ferdinand J. Paraguas, Kwamena K. Quagrainie, Ai Xing, Kevin G. Hughes, Cheikhna Dedah, Walter R. Keithly, Jr., and Richard F. Kazmierczak, Jr. Economics in the service of fisheries policy and practice. Marine Resource Economics, 26(1):87–94, March 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hanson:1996:MCA

[HDR96] Gregory D. Hanson, James W. Dunn, and Ganesh P. Rauniyar. Marketing characteristics associated with seafood counters in grocery stores. Marine Resource Economics, 11(1):11–22, Spring 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1985:BM

[HDS^{+}85] Rögnvaldur Hannesson, Ross G. Drynan, Frances Sandiford, Rögnvaldur Hannesson, Stephen R. Crutchfield, and John M. Gates. Back matter. Marine Resource Economics, 2(2):??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2015:ESC

[HdSK15] Daniel S. Holland, Patricia Pinto da Silva, and Andrew W. Kitts. Evolution of social capital and economic performance in New England harvest cooperatives. Marine Resource Economics, 30 (4):371–392, October 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Heaps:1993:OFF

[Hea93] Terry Heaps. The optimal feeding of farmed fish. Marine Resource Economics, 8(2):89–99, Summer 1993. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Heaps:1995:DDG

[Hea95] Terry Heaps. Density dependent growth and the culling of farmed fish. Marine Resource Economics, 10(3):285–298, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Heen:1989:IAM

[Hee89] Knut Heen. Impact analysis of multispecies marine resource management. Marine Resource Economics, 6(4):331–348, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Helgesen:2022:IMP

[Hel22] Irmelin Slettemoen Helgesen. ITQs, market power, and efficiency loss. Marine Resource Economics, 37(4):409–435, October 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Herrick:1988:BRD

[Her88] Samuel F. Herrick, Jr. Book review: David Doulman, The Development of the Tuna Industry in the Pacific Islands Region: an Analysis of Options. Marine Resource Economics, 5(2):165– 169, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Herrmann:1993:UIE

[Her93] Mark Herrmann. Using an international econometric model to forecast Alaska salmon revenues. Marine Resource Economics, 8(3):249–271, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See erratum [Ano93b].

Herrick:1995:BRH

[Her95] Samuel F. Herrick, Jr. Book review: H. F. Campbell and A. D. Owen, The Economics of Papua New Guinea’s Tuna Fisheries. Marine Resource Economics, 10(4):409–412, Winter 1995. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Herrera:2006:BSR

[Her06] Guillermo E. Herrera. Benefits of spatial regulation in a multispecies system. Marine Resource Economics, 21(1):63–79, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Heykoop:2001:GPS

[HF01] Jerry Heykoop and Darren L. Frechette. Gatornomics: Profitable and sustainable use of alligators in the Southeastern United States. Marine Resource Economics, 16(2):127–142, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hoff:2007:OVQ

[HF07] Ayoe Hoff and Hans Frost. Optimal vessel quotas and capacity of a Danish trawler fleet segment: a dual approach. Marine Resource Economics, 22(1):1–14, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hogan:1989:UFP

[HG89] William Hogan and Daniel Georgianna. U.S. fish processing capacity and imports of whole groundfish from Canada. Marine Resource Economics, 6(3):213–225, ???? 1989. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Haab:1999:SII

[HGH99] Timothy Haab, Amy Gautam, and Robert Hicks. Special issue introduction. Marine Resource Economics, 14(4):??, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hoshino:2015:ELR

[HGJH15] Eriko Hoshino, Caleb Gardner, Sarah Jennings, and Klaas Hartmann. Examining the long-run relationship between the prices of imported abalone in Japan. Marine Resource Economics, 30 (2):179–192, April 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hicks:1999:INM

[HGV^{+}99] Robert L. Hicks, Amy B. Gautam, David Van Voorhees, Maury Osborn, and Brad Gentner. An introduction to the NMFS marine recreational fisheries statistics survey with an emphasis on economic valuation. Marine Resource Economics, 14(4):375– 385, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hagan:1986:PED

[HH86] Philip Hagan and Gary Henry. Potential effects of differing management programs on the Southern bluefin tuna fishery. Marine Resource Economics, 3(4):353–389, ???? 1986. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Haab:1999:CSC

[HH99] Timothy C. Haab and Robert L. Hicks. Choice set considerations in models of recreation demand: History and current state of the art. Marine Resource Economics, 14(4):271–281, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2009:UMF

[HH09] Daniel S. Holland and Guillermo (Ta) E. Herrera. Uncertainty in the management of fisheries: Contradictory implications and a new approach. Marine Resource Economics, 24(3):289–299, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huppert:1989:BM

[HHN^{+}89] Daniel D. Huppert, Jack E. Houston, Amelia E. Nieto, James E. Epperson, Ho shui Li, George W. Lewis, Lee G. Anderson, Jack E. Houston, Amelia E. Nieto, James E. Epperson, Ho shui Li, George W. Lewis, Lee G. Anderson, R. A. Androkovich, and K. R. Stollery. Back matter. Marine Resource Economics, 6 (2):??, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2012:BR

[HHN^{+}12] Rögnvaldur Hannesson, Ling Huang, Lauren A. B. Nichols, J. Kevin Craig, Martin D. Smith, Max Nielsen, Ola Flaaten, Staffan Waldo, Josh Nowlis, Arthur A. Van Benthem, Max Nielsen, Ola Flaaten, Staffan Waldo, Ling Huang, Lauren A. B. Nichols, J. Kevin Craig, Martin D. Smith, Jose M. Maroto, Manuel Moran, Leif K. Sandal, and Stein I. Steinshamn. Books reviewed. Marine Resource Economics, 27(1):111–113, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hoshino:2010:EOM

[HHP10] Eriko Hoshino, Richard M. Hillary, and John Pearce. Economically optimal management strategies for the South Georgia Patagonian toothfish fishery. Marine Resource Economics, 25 (3):265–280, September 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Haab:2012:AHS

[HHSW12] Timothy Haab, Robert Hicks, Kurt Schnier, and John C. Whitehead. Angler heterogeneity and the species-specific demand for marine recreational fishing. Marine Resource Economics, 27(3): 229–251, September 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hoagland:1997:MBI

[HJ97] Porter Hoagland and Di Jin. A model of bycatch involving a passive use stock. Marine Resource Economics, 12(1):11–28, Spring 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hoagland:2003:OAO

[HJKP03] Porter Hoagland, Di Jin, and Hauke Kite-Powell. The optimal allocation of ocean space: Aquaculture and wild-harvest fisheries. Marine Resource Economics, 18(2):129–147, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hackett:2005:DFI

[HKBH05] Steven C. Hackett, Matthew J. Krachey, Sharon Brown, and David Hankin. Derby fisheries, individual quotas, and transition in the fish processing industry. Marine Resource Economics, 20 (1):47–60, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2013:SCS

[HKPW13] Daniel S. Holland, Andrew W. Kitts, Patricia Pinto Da Silva, and Joshua Wiersma. Social capital and the success of harvest cooperatives in the New England groundfish fishery. Marine Resource Economics, 28(2):133–153, June 2013. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hicks:2004:SRW

[HKS04] Robert L. Hicks, James Kirkley, and Ivar E. Strand, Jr. Shortrun welfare losses from essential fish habitat designations for the surfclam and ocean quahog fisheries. Marine Resource Economics, 19(1):113–129, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2011:BR

[HLA^{+}11] Rögnvaldur Hannesson, Thomas A. Larsen, Frank Asche, Linda Nøstbakken, Olivier Thébaud, Lars-Christian Sørensen, William L. Huth, O. Ashton Morgan, Thomas A. Larsen, Frank Asche, Linda Nøstbakken, Olivier Thébaud, Lars-Christian Sørensen, Le Kim Long, and Ola Flaaten. Books reviewed. Marine Resource Economics, 26(2):171–172, June 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hara:2014:ASA

[HLH^{+}14] Mafaniso M. Hara, Min-Yang Lee, Barbara Hutniczak, Rasmus Nielsen, Jesper Levring Andersen, Peter Bogetoft, Min-Yang Lee, Rasmus Nielsen, Jesper Levring Andersen, Peter Bogetoft, and Barbara Hutniczak. Analysis of South African commercial traditional linefish snoek value chain. Marine Resource Economics, 29(3):279–299, September 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hindsley:2020:ICA

[HLM20] Paul Hindsley, Craig E. Landry, and O. Ashton Morgan. Incorporating certainty and attribute non-attendance in choice experiments: an application to valuation of coastal habitat. Marine Resource Economics, 35(3):241–262, July 2020. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lin:1987:AED

[hLRT87a] Biing hwan Lin, Hugh S. Richards, and Joseph M. Terry. An analysis of the exvessel demand for Pacific halibut. Marine Resource Economics, 4(4):305–314, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lin:1987:BM

[hLRT^{+}87b] Biing hwan Lin, Hugh S. Richards, Joseph M. Terry, Geoffrey Waugh, Theodore Panayotou, Jerald J. Fletcher, Richard E. Howitt, Warren E. Johnston, Andrew Staniford, Biing hwan Lin, Hugh S. Richards, and Joseph M. Terry. Back matter. Marine Resource Economics, 4(4):??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holm:1995:CON

[HM95] Petter Holm and Leigh Mazany. Changes in the organization of the Norwegian fishing industry. Marine Resource Economics, 10(3):299–312, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huth:2011:MWP

[HM11] William L. Huth and O. Ashton Morgan. Measuring the willingness to pay for cave diving. Marine Resource Economics, 26 (2):151–166, June 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Herrmann:1992:AAT

[HMMhL92] Mark Herrmann, Ron C. Mittelhammer, Ron C. Mittlehammer, and Biing hwan Lin. Applying Almon-type polynomials in modelling seasonality of the Japanese demand for salmon. Marine Resource Economics, 7(1):3–13, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:1986:BM

[HMS^{+}86] Rögnvaldur Hannesson, Gordon R. Munro, Kenneth R. Stollery, Gordon R. Munro, Daniel D. Huppert, Dale Squires, and Kenneth R. Stollery. Back matter. Marine Resource Economics, 3 (4):??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hossain:2021:MIB

[HNAY^{+}21] Afjal Hossain, Max Nielsen, Isaac Ankamah-Yeboah, Badiuzzaman, and Fakir Azmal Huda. Market integration between cultured and captured species in developing countries: Lessons from inland areas in Bangladesh. Marine Resource Economics, 36(2):155–172, April 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huang:2012:MWL

[HNCS12] Ling Huang, Lauren A. B. Nichols, J. Kevin Craig, and Martin D. Smith. Measuring welfare losses from hypoxia: The case of North Carolina brown shrimp. Marine Resource Economics, 27(1):3–23, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Houston:1989:FAL

[HNE^{+}89] Jack E. Houston, Amelia E. Nieto, James E. Epperson, Ho shui Li, and George W. Lewis. Factors affecting local prices of shrimp landings. Marine Resource Economics, 6(2):163–172, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hansen:2011:WDH

[HO11] Håvard Hansen and Yuko Onozaka. When diseases hit aquaculture: an experimental study of spillover effects from negative publicity. Marine Resource Economics, 26(4):281–291, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hoff:2004:LAC

[Hof04] Ayoe Hoff. The linear approximation of the CES function with n input variables. Marine Resource Economics, 19(3):295–306, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:1999:DIM

[Hol99] Daniel S. Holland. On direct and indirect management of fishing capacity. Marine Resource Economics, 14(3):263–267, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2000:FFC

[Hol00] Daniel S. Holland. Fencing the fisheries commons: Regulatory barbed wire in the Alaskan groundfish fisheries. Marine Resource Economics, 15(2):141–149, Summer 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2004:SFR

[Hol04] Daniel S. Holland. Spatial fishery rights and marine zoning: a discussion with reference to management of marine resources in New England. Marine Resource Economics, 19(1):21–40, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2013:MCB

[Hol13] Daniel S. Holland. Making cents out of barter data from the British Columbia groundfish ITQ market. Marine Resource Economics, 28(4):311–330, December 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Homans:1999:EMS

[HR99] Frances R. Homans and Jane A. Ruliffson. The effects of minimum size limits on recreational fishing. Marine Resource Economics, 14(1):1–14, Spring 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hannesson:2017:FQP

[HRA^{+}17] Rögnvaldur Hannesson, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Steve J. Miller, Robert T. Deacon, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, and Rögnvaldur Hannesson. Fish quota prices in Norway. Marine Resource Economics, 32(1):109–117, January 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hanna:2023:BRS

[HRA^{+}23] Susan Hanna, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, and Rasmus Nielsen. [Book review:] The Salt Roads: How Fish Made a Culture by John Goodlad. Marine Resource Economics, 39(1):73–75, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hamalainen:1986:MSE

[HRK86] Raimo P. Hämäläinen, Jukka Ruusunen, and Veijo Kaitala. Myopic Stackelberg equilibria and social coordination in a share contract fishery. Marine Resource Economics, 3(3):209–235, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huang:2018:ICR

[HRSW18] Ling Huang, Subhash Ray, Kathleen Segerson, and John Walden. Impact of collective rights-based fisheries management: Evidence from the New England groundfish fishery. Marine Resource Economics, 33(2):177–201, April 2018. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huppert:1986:PEB

[HS86] Daniel D. Huppert and Dale Squires. Potential economic benefits and optimum fleet size in the Pacific Coast trawl fleet. Marine Resource Economics, 3(4):297–318, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sun:1998:ONF

[hS98] Chin hwa Sun. Optimal number of fishing vessels for Taiwan’s offshore fisheries: a comparison of different fleet size reduction policies. Marine Resource Economics, 13(4):275–288, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holland:2004:ISM

[HSCH04] Daniel S. Holland, James N. Sanchirico, Rita E. Curtis, and Robert L. Hicks. An introduction to spatial modeling in fisheries economics. Marine Resource Economics, 19(1):1–6, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Holloway:2006:BRS

[HT06] Garth Holloway and David Tomberlin. Bayesian ranking and selection of fishing boat efficiencies. Marine Resource Economics, 21(4):415–432, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hayes:1992:EBW

[HTTA92] Karen M. Hayes, Timothy J. Tyrrell, Timothy J. Tyrell, and Glen Anderson. Estimating the benefits of water quality improvements in the Upper Narragansett Bay. Marine Resource Economics, 7(1):75–85, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huang:2015:IDS

[Hua15] Pei Huang. An inverse demand system for the differentiated blue crab market in Chesapeake Bay. Marine Resource Economics, 30(2):139–156, April 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Huppert:1989:MVF

[Hup89] Daniel D. Huppert. Measuring the value of fish to anglers: Application to Central California anadromous species. Marine Resource Economics, 6(2):89–107, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hutniczak:2014:IPU

[Hut14] Barbara Hutniczak. Increasing pressure on unregulated species due to changes in individual vessel quotas: an empirical application to trawler fishing in the Baltic Sea. Marine Resource Economics, 29(3):201–217, September 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Helstad:2005:PLB

[HVTY05] Kjetil Helstad, Terje Vassdal, Torbjørn Trondsen, and James A. Young. Price links between auction and direct sales of fresh and frozen fish in North Norway (1997–2003). Marine Resource Economics, 20(3):305–322, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hung:2012:FSI

[HW12] Chih-Ming Hung and Yungho Weng. Fishery subsidies, illegal fuel trading, and conservation. Marine Resource Economics, 27 (3):253–265, September 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Irz:2018:FCF

[ILRS18] Xavier Irz, Pascal Leroy, Vincent Réquillart, and Louis-Georges Soler. Fish in climate-friendly and healthy diets. Marine Resource Economics, 33(4):309–330, October 2018. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ikiara:1999:FRE

[IO99] Moses Muriira Ikiara and Joop G. Odink. Fishermen resistance to exit fisheries. Marine Resource Economics, 14(3):199–213, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ikiara:1999:BM

[IOE^{+}99] Moses Muriira Ikiara, Joop G. Odink, Håkan Eggert, Mats Ulmestrand, Daniel S. Holland, Moses Muriira Ikiara, Joop G. Odink, Håkan Eggert, Mats Ulmestrand, Sherry L. Larkin, and Gilbert Sylvia. Back matter. Marine Resource Economics, 14 (3):??, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Johnson:1989:MVF

[JA89] Neal S. Johnson and Richard M. Adams. On the marginal value of a fish: Some evidence from a steelhead fishery. Marine Resource Economics, 6(1):43–55, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jaffry:2004:APT

[Jaf04] Shabbar Jaffry. Asymmetric price transmission: a case study of the French hake value chain. Marine Resource Economics, 19 (4):511–523, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jardine:2024:E

[Jar24] Sunny L. Jardine. Editorial. Marine Resource Economics, 39(2): 77–78, ???? 2024. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jardine:2024:FPR

[JBH24] Sunny L. Jardine, Katy Bland, and Adam L. Hayes. Fishing portfolio response to a climate shock. Marine Resource Economics, 39(2):79–100, ???? 2024. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jensen:1998:IBT

[Jen98] Carsten Lynge Jensen. Investment behavior and tax policy. Marine Resource Economics, 13(3):185–196, Fall 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jensen:2002:ADT

[Jen02] Carsten Lynge Jensen. Applications of dual theory in fisheries: a survey. Marine Resource Economics, 17(4):309–334, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jin:2005:SBM

[JH05] Di Jin and Guillermo E. Herrera. A stochastic bioeconomic model with research. Marine Resource Economics, 20(3):249– 261, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jin:2006:ARB

[JHT06] Di Jin, Porter Hoagland, and Eric Thunberg. An analysis of the relationship between fish harvesting and processing sectors in New England. Marine Resource Economics, 21(1):47–62, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jin:1994:MWD

[Jin94] Di Jin. Multimedia waste disposal optimization under uncertainty with an ocean option. Marine Resource Economics, 9(2): 119–139, Summer 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jin:1996:AIF

[Jin96] Di Jin. ADB/INFOFISH, fishery export industry profile: China. Marine Resource Economics, 11(1):57–58, Spring 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jinji:2012:FSM

[Jin12] Naoto Jinji. Fisheries subsidies and management in open economies. Marine Resource Economics, 27(1):25–41, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jin:1995:ELM

[JKP95] Di Jin and Hauke L. Kite-Powell. Environmental liability, marine insurance and an optimal risk sharing strategy for marine oil transport. Marine Resource Economics, 10(1):1–19, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jin:2012:BAT

[JKPHS12] Di Jin, Hauke Kite-Powell, Porter Hoagland, and Andrew Solow. A bioeconomic analysis of traditional fisheries in the Red Sea. Marine Resource Economics, 27(2):137–148, June 2012. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jones:1999:EMS

[JL99] Carol A. Jones and Frank Lupi. The effect of modeling substitute activities on recreational benefit estimates. Marine Resource Economics, 14(4):357–374, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Junqueira-Lopes:1996:BMR

[JLMR96] Rui Junqueira-Lopes, Philippe Michel, and Gilles Rotillon. Bioeconomic management of red swamp crayfish (Procambarus clarkii) in the presence of environmental externalities. Marine Resource Economics, 11(1):1–9, Spring 1996. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jin:2019:EAI

[JLT19] Di Jin, Min-Yang Lee, and Eric Thunberg. An empirical analysis of individual fishing quota market trading. Marine Resource Economics, 34(1):39–57, January 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jensen:2002:BM

[JMF^{+}02] Carsten Lynge Jensen, Andrew D. Murray, Abdulai Fofana, M. Carmen Gallastegui, Elena Iñarra, Raul Prellezo, M. Carmen Gallastegui, Elena Iñarra, Raul Prellezo, Panos Fousekis, and Carsten Lynge Jensen. Back matter. Marine Resource Economics, 17(4):??, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Johnston:1985:ICM

[Joh85] Richard S. Johnston. Ian chaston, marketing in fisheries and aquaculture. Marine Resource Economics, 2(2):197–201, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Johnson:1995:ITQ

[Joh95] Ronald N. Johnson. Implications of taxing quota value in an individual transferable quota fishery. Marine Resource Economics, 10(4):327–340, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Johnson:1996:ITQ

[Joh96a] Ronald N. Johnson. Implications of taxing quota value in an individual transferable quota fishery: Reply. Marine Resource Economics, 11(2):129–130, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Johnson:1996:CFR

[Joh96b] Terry Johnson. Changing fisheries of the Russian Far East. Marine Resource Economics, 11(2):131–135, Summer 1996. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Johnston:2006:BM

[JRB^{+}06] Robert J. Johnston, Matthew H. Ranson, Elena Y. Besedin, Erik C. Helm, Chang K. Seung, Edward C. Waters, Harald Bergland, Pål Andreas Pedersen, Robert J. Johnston, Matthew H. Ranson, Elena Y. Besedin, Erik C. Helm, Chang K. Seung, Edward C. Waters, and Guillermo E. Herrera. Back matter. Marine Resource Economics, 21(1):??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Johnston:2006:WDW

[JRBH06] Robert J. Johnston, Matthew H. Ranson, Elena Y. Besedin, and Erik C. Helm. What determines willingness to pay per fish? A meta-analysis of recreational fishing values. Marine Resource Economics, 21(1):1–32, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Jensen:2001:MFE

[JV01] Frank Jensen and Niels Vestergaard. Management of fisheries in the EU: a principal-agent analysis. Marine Resource Economics, 16(4):277–291, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knapp:1997:BM

[KAC^{+}97] Gunnar Knapp, Claire W. Armstrong, Derek J. Clark, Sam Pooley, Claire W. Armstrong, Derek J. Clark, Gunnar Knapp, and Frank Asche. Back matter. Marine Resource Economics, 12 (3):??, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kahn:1987:MED

[Kah87] James R. Kahn. Measuring the economic damages associated with terrestrial pollution of marine ecosystems. Marine Resource Economics, 4(3):193–209, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kataria:2007:CBA

[Kat07] Mitesh Kataria. A cost-benefit analysis of introducing a nonnative species: The case of signal crayfish in Sweden. Marine Resource Economics, 22(1):15–28, ???? 2007. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knowler:2005:MBS

[KB05] Duncan Knowler and Edward B. Barbier. Managing the Black Sea anchovy fishery with nutrient enrichment and a biological invader. Marine Resource Economics, 20(3):263–285, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

KhanhNgoc:2023:IML

[KBBN23] Quach Thi Khanh Ngoc, Bui Bich Xuan, Tobias Börger, and Nguyen Ngoc Duy. The impact of marine litter on production risk and technical efficiency in the trawl fisheries of Vietnam. Marine Resource Economics, 38(1):65–84, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kidane:2021:GSD

[KBK^{+}21] Dejene Gizaw Kidane, Eivind Hestvik Brækkan, Dejene Gizaw Kidane, Eivind Hestvik Brækkan, Tongzhe Li, Ahsanuzzaman, Kent D. Messer, César Salazar, Marcela Jaime, Miguel Quiroga, Anders Skonhoft, Melina Kourantidou, Dejene Gizaw Kidane, and Eivind Hestvik Brækkan. Global seafood demand growth differences across regions, income levels, and time. Marine Resource Economics, 36(3):289–305, July 2021. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knowler:2001:OAM

[KBS01] Duncan Knowler, Edward B. Barbier, and Ivar Strand. An openaccess model of fisheries and nutrient enrichment in the Black Sea. Marine Resource Economics, 16(3):195–217, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See errata [KBS02].

Knowler:2002:EOA

[KBS02] Duncan Knowler, Edward B. Barbier, and Ivar Strand. Errata: An Open-Access Model of Fisheries and Nutrient Enrichment in the Black Sea. Marine Resource Economics, 17(4):347–349, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [KBS01].

Kazmierczak:1995:MAE

[KC95] Richard F. Kazmierczak, Jr. and Rex H. Caffey. Management ability and the economics of recirculating aquaculture production systems. Marine Resource Economics, 10(2):187–209, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kennedy:1987:CGT

[Ken87] John O. S. Kennedy. A computable game theoretic approach to modelling competitive fishing. Marine Resource Economics, 4(1):1–14, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kennedy:1992:OAC

[Ken92] John O. S. Kennedy. Optimal annual changes in harvests from multicohort fish stocks: The case of Western mackerel. Marine Resource Economics, 7(3):95–114, Fall 1992. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kennedy:2003:SEM

[Ken03] John Kennedy. Scope for efficient multinational exploitation of North-East Atlantic mackerel. Marine Resource Economics, 18(1):55–80, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kennedy:2006:CTE

[Ken06] John Kennedy. Clem tisdell, economics and ecology in agriculture and marine production: Bioeconomics and resource use. Marine Resource Economics, 21(3):333–335, ???? 2006. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kahui:2015:BRE

[KFS^{+}15] Viktoria Kahui, Robert Fonner, Gil Sylvia, Joshua K. Abbott, César Salazar Espinoza, Robert Fonner, Gil Sylvia, Joshua K. Abbott, and César Salazar Espinoza. Book review: Economics of the Oceans. Rights, Rents and Resources. By Paul Hallwood. Marine Resource Economics, 30(1):121–122, January 2015. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kaufmann:1997:CRF

[KG97] Barry Kaufmann and Gerry Geen. Cost-recovery as a fisheries management tool. Marine Resource Economics, 12(1):57–66, Spring 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kaufmann:1998:QAL

[KG98] Barry Kaufmann and Gerry Geen. Quota allocation and litigation: an economic perspective. Marine Resource Economics, 13 (2):143–157, Summer 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kung:2022:CPW

[KGW22] Megan Kung, Dennis Guignet, and Patrick Walsh. Comparing pollution where you live and play: a hedonic analysis of enterococcus in the Long Island Sound. Marine Resource Economics, 37(1):65–89, January 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kennedy:2006:WSR

[KH06] John Kennedy and Rögnvaldur Hannesson. Within-season rents: Maximised or dissipated in an open-access fishery? Marine Resource Economics, 21(3):251–267, ???? 2006. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kourantidou:2025:NNA

[KHB^{+}25] Melina Kourantidou, Porter Hoagland, Megan Bailey, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, FRANK ASCHE, Jay Abolofia, Frank Asche, James E. Wilen, Robert Fonner, and Gil Sylvia. Navigating Nunatsiavut’s Arctic charr: a simultaneous commercial and subsistence fishery with many unknowns. Marine Resource Economics, 39(3):243–261, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knapp:1996:BM

[KHD^{+}96] Gunnar Knapp, Gregory D. Hanson, James W. Dunn, Ganesh P. Rauniyar, J. J. Poggie, Jr., R. B. Pollnac, C. Van Dusen, J. J. Poggie, Jr., R. B. Pollnac, C. Van Dusen, Gregory D. Hanson, James W. Dunn, Ganesh P. Rauniyar, Rui Junqueira-Lopes, Philippe Michel, and Gilles Rotillon. Back matter. Marine Resource Economics, 11(1):??, Spring 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kinnucan:1993:BRR

[Kin93] Henry Kinnucan. Book review: R. B. Pollnac and P. Weeks, Coastal Aquaculture in Developing Countries: Problems and Perspectives. Marine Resource Economics, 8(1):85–88, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kinnucan:2003:FTF

[Kin03] Henry W. Kinnucan. Futility of targeted fish tariffs and an alternative. Marine Resource Economics, 18(3):211–224, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kaitala:2003:HSF

[KJE03] Veijo Kaitala, Niclas Jonzén, and Katja Enberg. Harvesting strategies in a fish stock dominated by low-frequency variability: The Norwegian spring-spawning herring (Clupea harengus). Marine Resource Economics, 18(3):263–274, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kronbak:2007:SRS

[KL07] Lone Grønbæk Kronbak and Marko Lindroos. Sharing rules and stability in coalition games with externalities. Marine Resource Economics, 22(2):137–154, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kling:1987:SAC

[Kli87] Catherine L. Kling. A simulation approach to comparing multiple site recreation demand models using Chesapeake Bay survey data. Marine Resource Economics, 4(2):95–109, ???? 1987. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kaitala:1993:MHS

[KM93] Veijo Kaitala and Gordon R. Munro. The management of high seas fisheries. Marine Resource Economics, 8(4):313–329, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kinnucan:2002:OSA

[KM02] Henry W. Kinnucan and Øystein Myrland. Optimal seasonal allocation of generic advertising expenditures with product substitution: Salmon in France. Marine Resource Economics, 17 (2):103–120, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knauss:1995:BRM

[Kna95] John Knauss. Book review: E. M. Borgese, N. Ginsburg, and J. R. Morgen, Ocean Yearbook 11. Marine Resource Economics, 10 (4):412–413, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knapp:1996:AHC

[Kna96] Gunnar Knapp. Alaska halibut captains’ attitudes towards IFQs. Marine Resource Economics, 11(1):43–55, Spring 1996.

CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knapp:1997:IEA

[Kna97] Gunnar Knapp. Initial effects of the Alaska halibut IFQ program: Survey comments of Alaska fishermen. Marine Resource Economics, 12(3):239–248, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knapp:2012:EFP

[Kna12] Gunnar Knapp. Economics and fisheries policy: a survey of NAAFE members. Marine Resource Economics, 27(4):389–395, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kinnucan:1993:UPF

[KNH93] Henry W. Kinnucan, Robert G. Nelson, and Johanis Hiariey. U.S. preferences for fish and seafood: an evoked set analysis. Marine Resource Economics, 8(3):273–291, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kinnucan:1995:CAR

[KNX95] Henry W. Kinnucan, Robert G. Nelson, and Hui Xiao. Cooperative advertising rent dissipation. Marine Resource Economics, 10(4):373–384, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Koemle:2023:STL

[KNYA23] Dieter Koemle, Thang Dao Nguyen, Xiaohua Yu, and Robert Arlinghaus. Subsidies, temporary laying-up, and efficiency in a coastal commercial fishery. Marine Resource Economics, 38 (2):153–179, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Koenig:1984:FRU

[Koe84] Evan F. Koenig. Fisheries regulation under uncertainty: a dynamic analysis. Marine Resource Economics, 1(2):193–208, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kouka:1995:EMP

[Kou95] Pierre-Justin Kouka. An empirical model of pricing in the catfish industry. Marine Resource Economics, 10(2):161–169, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Knoppe:2025:SSR

[KOZ25] Clemens Knoppe, Mariia Okuneva, and Mikaella Zitti. Salmon stock returns around market news. Marine Resource Economics, 40(2):107–140, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kinnucan:2001:UEG

[KP01] Henry W. Kinnucan and Laxmi Paudel. Upstream effects of generic advertising: The case of catfish. Marine Resource Economics, 16(2):83–107, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Keithly:2008:SUS

[KP08] Walter R. Keithly, Jr. and Pawan Poudel. The Southeast U.S.A. shrimp industry: Issues related to trade and antidumping duties. Marine Resource Economics, 23(4):459–483, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kirkley:2004:DSC

[KPS04] James Kirkley, Catherine J. Morrison Paul, and Dale Squires. Deterministic and stochastic capacity estimation for fishery capacity reduction. Marine Resource Economics, 19(3):271–294, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kristofersson:2009:BRF

[Kri09] Dadi Mar Kristofersson. Book review: Frank Asche, Primary Industries Facing Global Markets: The Supply Chains and Markets for Norwegian Food. Marine Resource Economics, 24(1): 97–98, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kristofersson:2019:BRO

[Kri19] Dadi Kristofersson. Book review: The Omega Principle: Seafood and the Quest for a Long Life and a Healthier Planet. By Paul Greenberg. Marine Resource Economics, 34(3):287–288, July 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kronbak:2004:DOA

[Kro04] Lone Groenbaek Kronbak. The dynamics of an open-access fishery: Baltic Sea cod. Marine Resource Economics, 19(4):459–479, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kroetz:2017:ECJ

[KSPT^{+}17] Kailin Kroetz, James N. Sanchirico, Julio Peña-Torres, David Corderi Novoa, Min-Yang Lee, Scott Steinback, Kristy Wallmo, Colin W. Clark, Gordon R. Munro, Matthew N. Reimer, Joshua K. Abbott, James E. Wilen, Min-Yang Lee, Scott Steinback, Kristy Wallmo, Matthew N. Reimer, Joshua K. Abbott, James E. Wilen, Matthew N. Reimer, Joshua K. Abbott, and Alan C. Haynie. Evaluation of the Chilean jack mackerel ITQ system. Marine Resource Economics, 32(2):217–241, April 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kerstens:2005:MRS

[KSV05] Kristiaan Kerstens, Dale Squires, and Niels Vestergaard. Methodological reflections on the Short-Run Johansen Industry Model in relation to capacity management. Marine Resource Economics, 20(4):425–443, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kuronuma:1994:EAM

[KT94] Yoshihiro Kuronuma and Clement A. Tisdell. Economics of Antarctic minke whale catches: Sustainability and welfare considerations. Marine Resource Economics, 9(2):141–158, Summer 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kitts:2000:WPB

[KTR00] Andrew Kitts, Eric Thunberg, and John Robertson. Willingness to participate and bids in a fishing vessel buyout program: a case study of New England groundfish. Marine Resource Economics, 15(3):221–232, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kumbhakar:2002:RPT

[Kum02] Subal C. Kumbhakar. Risk preferences and technology: a joint analysis. Marine Resource Economics, 17(2):77–89, ???? 2002.

CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kvamsdal:2016:TCS

[Kva16] Sturla Furunes Kvamsdal. Technical change as a stochastic trend in a fisheries model. Marine Resource Economics, 31(4):403– 419, October 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kvamsdal:2018:BRA

[Kva18] Sturla F. Kvamsdal. Book review: All the Boats on the Ocean. How Government Subsidies Led to Global Overfishing. By Carmel Finley. Marine Resource Economics, 33(1):113–117, January 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kennedy:1986:TDQ

[KW86] John O. S. Kennedy and James W. Watkins. Time-dependent quotas for the Southern bluefin tuna fishery. Marine Resource Economics, 2(4):293–313, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Kasperski:2009:WIO

[KW09] Stephen Kasperski and Robert Wieland. When is it optimal to delay harvesting? The role of ecological services in the Northern Chesapeake Bay oyster fishery. Marine Resource Economics, 24 (4):361–385, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lewis:2006:EQH

[LA06] Lynne Y. Lewis and Gayatri Acharya. Environmental quality and housing markets: Does lot size matter? Marine Resource Economics, 21(3):317–330, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Larsen:2011:CSA

[LA11] Thomas A. Larsen and Frank Asche. Contracts in the salmon aquaculture industry: an analysis of Norwegian salmon exports. Marine Resource Economics, 26(2):141–150, June 2011. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lipton:2004:BM

[LAG^{+}04] Douglas Lipton, Frank Asche, Daniel V. Gordon, Rögnvaldur Hannesson, Olivier Guyader, Fabienne Daures, Spyros Fifas, Frank Asche, Daniel V. Gordon, Rögnvaldur Hannesson, Douglas Lipton, and Shekar Bose. Back matter. Marine Resource Economics, 19(2):??, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Li:2021:TPU

[LAM21] Tongzhe Li, Ahsanuzzaman, and Kent D. Messer. Is there a potential US market for seaweed-based products? A framed field experiment on consumer acceptance. Marine Resource Economics, 36(3):255–268, July 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lindner:1992:RGD

[LCB92] R. K. Lindner, H. F. Campbell, and G. F. Bevin. Rent generation during the transition to a managed fishery: The case of the New Zealand ITQ system. Marine Resource Economics, 7 (4):229–248, Winter 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lee:2014:HPA

[Lee14] Min-Yang Lee. Hedonic pricing of Atlantic cod: Effects of size, freshness, and gear. Marine Resource Economics, 29(3):259–277, September 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lew:2019:PRC

[Lew19] Daniel K. Lew. Place of residence and cost attribute nonattendance in a stated preference choice experiment involving a marine endangered species. Marine Resource Economics, 34 (3):225–245, July 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Long:2011:SAP

[LF11] Le Kim Long and Ola Flaaten. A Stackelberg analysis of the potential for cooperation in straddling stock fisheries. Marine Resource Economics, 26(2):119–139, June 2011. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Letelier:1998:IDM

[LG98] Eduardo Letelier and Exequiel González. Impact of the drawback mechanisms in the Chilean salmon industry. Marine Resource Economics, 13(4):293–309, Winter 1998. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lee:2007:VPU

[LG07] Jungsam Lee and John M. Gates. Virtual population units: a new institutional approach to fisheries management. Marine Resource Economics, 22(1):29–47, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lucas:2019:SCL

[LGL19] Sterenn Lucas, Stéphane Gouin, and Marie Lesueur. Seaweed consumption and label preferences in France. Marine Resource Economics, 34(2):143–162, April 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Loomis:1988:BM

[LGO^{+}88] John B. Loomis, Thomas A. Grigalunas, James J. Opaluch, Deborah French, Mark Reed, Stanley D. H. Wang, Christopher B. Kellogg, Thomas A. Grigalunas, James J. Opaluch, Deborah French, Mark Reed, Stanley D. H. Wang, Christopher B. Kellogg, and John B. Loomis. Back matter. Marine Resource Economics, 5(1):??, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Luna:2021:HDI

[LGOS21] Manuel Luna, Myriam García-Olalla, and José L. Fernández Sánchez. How do investors value firms’ decisions on obtaining an eco-label? Evidence from the fishing industry. Marine Resource Economics, 36(3):211–228, July 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Leung:2012:SII

[LH12] PingSun Leung and Daniel Huppert. Special issue introduction. Marine Resource Economics, 27(4):287–288, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lew:2015:WIM

[LHCL15] Daniel K. Lew, Amber Himes-Cornell, and Jean Lee. Weighting and imputation for missing data in a cost and earnings fishery survey. Marine Resource Economics, 30(2):219–230, April 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Larson:1996:TEB

[LHT96] Douglas M. Larson, Brett W. House, and Joseph M. Terry. Toward efficient bycatch management in multispecies fisheries: a nonparametric approach. Marine Resource Economics, 11(3): 181–201, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Li:1998:CHS

[Li98a] Eric Li. Cooperative high-seas straddling stock agreement as a characteristic function game. Marine Resource Economics, 13 (4):247–258, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Li:1998:OVH

[Li98b] Eric Li. Option value of harvesting: Theory and evidence. Marine Resource Economics, 13(2):135–142, Summer 1998. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Li:2009:ROV

[Li09] Eric A. L. Li. A reply to “Option Value of Harvesting” revisited (the comment). Marine Resource Economics, 24(1):83–87, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [Sar09].

Libecap:2007:APR

[Lib07] Gary D. Libecap. Assigning property rights in the common pool: Implications of the prevalence of first-possession rules for ITQs in fisheries. Marine Resource Economics, 22(4):407–423, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lindebo:2005:MNI

[Lin05a] Erik Lindebo. Multi-national industry capacity in the North Sea flatfish fishery. Marine Resource Economics, 20(4):385–406, ????

2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lindebo:2005:RSE

[Lin05b] Erik Lindebo. Role of subsidies in EU fleet capacity management. Marine Resource Economics, 20(4):445–466, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lindroos:2013:BRC

[Lin13] Marko Lindroos. Book review: Conservation on the High Seas — Harmonizing International Regimes for the Sustainable Use of Living Resources. By Simone Borg. Marine Resource Economics, 28(2):199–200, June 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lipton:2004:VIW

[Lip04] Douglas Lipton. The value of improved water quality to Chesapeake Bay boaters. Marine Resource Economics, 19(2):265–270, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lindroos:2000:NEC

[LK00] Marko Lindroos and Veijo Kaitala. Nash equilibria in a coalition game of the Norwegian spring-spawning herring fishery. Marine Resource Economics, 15(4):321–339, ???? 2000. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Landry:2003:EEB

[LKK03a] Craig E. Landry, Andrew G. Keeler, and Warren Kriesel. An economic evaluation of beach erosion management alternatives. Marine Resource Economics, 18(2):105–127, ???? 2003. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Landry:2003:BM

[LKK^{+}03b] Craig E. Landry, Andrew G. Keeler, Warren Kriesel, Porter Hoagland, Di Jin, Hauke Kite-Powell, Frank Asche, Trond Bjørndal, Elin H. Sissener, Craig E. Landry, Andrew G. Keeler, Warren Kriesel, Porter Hoagland, Di Jin, Hauke Kite-Powell, Scott C. Matulich, and Michael L. Clark. Back matter. Marine Resource Economics, 18(2):??, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Loomis:1994:TEV

[LL94] John B. Loomis and Douglas M. Larson. Total economic values of increasing gray whale populations: Results from a contingent valuation survey of visitors and households. Marine Resource Economics, 9(3):275–286, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Larson:2013:HDH

[LL13] Douglas M. Larson and Daniel K. Lew. How do harvest rates affect angler trip patterns? Marine Resource Economics, 28 (2):155–173, June 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lee:2000:BAM

[LLA00] Donna J. Lee, Sherry L. Larkin, and Charles M. Adams. A bioeconomic analysis of management alternatives for the U.S. North Atlantic swordfish fishery. Marine Resource Economics, 15(2):77–96, Summer 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Landry:2016:AOS

[LLLV16] Craig E. Landry, Alyson R. Lewis, Haiyong Liu, and Hans Vogelsong. Addressing onsite sampling in analysis of recreation demand: Economic value and impact of visitation to Cape Hatteras National Seashore. Marine Resource Economics, 31(3): 301–322, July 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lew:2010:VEE

[LLR10] Daniel K. Lew, David F. Layton, and Robert D. Rowe. Valuing enhancements to endangered species protection under alternative baseline futures: The case of the Steller sea lion. Marine Resource Economics, 25(2):133–154, June 2010. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Liu:2025:IRD

[LLS25] Yaqin Liu, Qingran Li, and Andrew R. Solow. Impacts of reef degradation on commercial fisheries. Marine Resource Economics, 40(1):23–44, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lazkano:2016:QEC

[LN16] Itziar Lazkano and Linda Nøstbakken. Quota enforcement and capital investment in natural resource industries. Marine Resource Economics, 31(3):339–354, July 2016. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Loomis:1988:BET

[Loo88] John B. Loomis. The bioeconomic effects of timber harvesting on recreational and commercial salmon and steelhead fishing: a case study of the siuslaw national forest. Marine Resource Economics, 5(1):43–60, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Loomis:2001:BRF

[Loo01] John Loomis. Book review: Frank Ward and Diana Beal, Valuing Nature with Travel Cost Models: a Manual. Marine Resource Economics, 16(4):315–317, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Leung:2001:REI

[LP01] Pingsun Leung and Sam Pooley. Regional economic impacts of reductions in fisheries production: a supply-driven approach. Marine Resource Economics, 16(4):251–262, ???? 2001. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

LeBihan:2013:RPR

[LPG13] Véronique Le Bihan, Sophie Pardo, and Patrice Guillotreau. Risk perception and risk management strategies of oyster farmers. Marine Resource Economics, 28(3):285–304, September 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Leung:2001:BM

[LPV^{+}01] Pingsun Leung, Sam Pooley, K. Kuperan Viswanathan, Ishak Haji Omar, Yongil Jeon, James Kirkley, Dale Squires, Indah Susilowati, Steven E. Edwards, Pingsun Leung, Sam Pooley, K. Kuperan Viswanathan, Ishak Haji Omar, Yongil Jeon, James Kirkley, Dale Squires, Indah Susilowati, and Steven E. Edwards. Back matter. Marine Resource Economics, 16(4): ??, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Leibbrandt:2021:RRO

[LPV21] Andreas Leibbrandt, Sergio Puerto, and Maria Alejandra Vélez. Regrouping to reduce overfishing: Evidence from a series of labin-the-field experiments in Mexico. Marine Resource Economics, 36(4):319–340, October 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Leplat:2025:RCF

[LRCL25] Mélody Leplat, Carole Ropars-Collet, and Philippe Le Goffe. Regulating congestion with fee-based versus nonfee management measures: Application of a repeated random utility model to outdoor recreation. Marine Resource Economics, 39(3):207–228, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Larkin:1999:FLH

[LS99] Sherry L. Larkin and Gilbert Sylvia. Firm-level hedonic analysis of U.S. produced surimi: Implications for processors and resource managers. Marine Resource Economics, 14(3):179–197, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lew:2014:SSR

[LS14] Daniel K. Lew and Chang K. Seung. On the statistical significance of regional economic impacts from recreational fishing harvest limits in Southern Alaska. Marine Resource Economics, 29(3):241–257, September 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Li:2021:FCR

[LS21] Qingran Li and Martin D. Smith. Fishery collapse revisited. Marine Resource Economics, 36(1):1–22, January 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Liu:2021:DDS

[LSBH21] Xiaozi Liu, Daigee Shaw, Trond Bjørndal, and Mikko Heino. The day-to-day supply responses of a limited-entry mixed fishery. Marine Resource Economics, 36(1):71–90, January 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Leggett:2018:AEB

[LSH^{+}18] Christopher G. Leggett, Nora Scherer, Timothy C. Haab, Ryan Bailey, Jason P. Landrum, and Adam Domanski. Assessing the economic benefits of reductions in marine debris at Southern California beaches: a random utility travel cost model. Marine Resource Economics, 33(2):133–153, April 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Larkin:2006:ORF

[LSHQ06] Sherry L. Larkin, Gil Sylvia, Michael Harte, and Kathryn Quigley. Optimal rebuilding of fish stocks in different nations: Bioeconomic lessons for regulators. Marine Resource Economics, 21(4):395–413, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lewis-Smith:2024:DHI

[LSJN^{+}24] Connor Lewis-Smith, Sunny Jardine, Karma Norman, Daniel S. Holland, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, and Rasmus Nielsen. Drivers of health insurance status in the US West Coast commercial fishing fleet. Marine Resource Economics, 39(2):163–177, ???? 2024. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lian:2009:FRR

[LSW09] Carl Lian, Rajesh Singh, and Quinn Weninger. Fleet restructuring, rent generation, and the design of individual fishing quota programs: Empirical evidence from the Pacific Coast groundfish fishery. Marine Resource Economics, 24(4):329–359, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lee:2017:ABM

[LSW17] Min-Yang Lee, Scott Steinback, and Kristy Wallmo. Applying a bioeconomic model to recreational fisheries management: Groundfish in the Northeast United States. Marine Resource Economics, 32(2):191–216, April 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lindebo:2005:BM

[LVV^{+}05] Erik Lindebo, Niels Vestergaard, Luc Van Hoof, Jan Willem De Wilde, Kristiaan Kerstens, Dale Squires, Niels Vestergaard, Olivier Guyader, Fabienne Daurès, Luc Van Hoof, and Jan Willem De Wilde. Back matter. Marine Resource Economics, 20(4):??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lew:2020:ANAa

[LW20] Daniel K. Lew and John C. Whitehead. Attribute nonattendance in choice experiments of marine ecosystem goods and services: Special issue introduction. Marine Resource Economics, 35(3):195–200, July 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lew:2020:ANAb

[LWL^{+}20] Daniel K. Lew, John C. Whitehead, Daniel K. Lew, John C. Whitehead, Julide Ceren Ahi, Gorm Kipperberg, Daniel K. Lew, John C. Whitehead, Xumin Zhang, Yingkai Fang, Zhifeng Gao, Paul Hindsley, Craig E. Landry, and O. Ashton Morgan. Attribute non-attendance as an information processing strategy in stated preference choice experiments: Origins, current practices, and future directions. Marine Resource Economics, 35(3):285– 317, July 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Lichtenberg:1987:RMC

[LZ87] Erik Lichtenberg and David Zilberman. Regulation of marine contamination under environmental uncertainty: Shellfish contamination in California. Marine Resource Economics, 4(3):211– 225, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mainardi:2010:MSY

[Mai10] Stefano Mainardi. Maximum sustainable yields and non-linear paths to steady-state equilibrium. Marine Resource Economics, 25(2):185–212, June 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mainardi:2019:AFE

[Mai19] Stefano Mainardi. Access fees and efficiency frontiers with selectivity and latent classes: Falkland Islands fisheries. Marine Resource Economics, 34(2):163–195, April 2019. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Marr:1986:BRG

[Mar86a] John Marr. Book review: George Kent and Mark Valencia, Jr., Marine Policy in Southeast Asia. Marine Resource Economics, 3(3):265–270, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Marsh:1986:GEI

[Mar86b] James Barney Marsh. Guest Editor’s introduction. Marine Resource Economics, 3(1):1–2, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Matthiasson:1996:WFF

[Mat96] Thorolfur Matthiasson. Why fishing fleets tend to be “Too Big”. Marine Resource Economics, 11(3):173–179, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Matthiasson:1997:CLG

[Mat97] Thorolfur Matthiasson. Consequences of local government involvement in the Icelandic ITQ market. Marine Resource Economics, 12(2):107–126, Summer 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Matulich:2009:VIP

[Mat09] Scott C. Matulich. The value of individual processing quota in the Alaska red king crab fishery: a preliminary analysis. Marine Resource Economics, 24(2):187–193, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Matulich:2010:SCF

[Mat10] Scott C. Matulich. Stranded capital in fisheries: The Pacific Coast groundfish/whiting case. Marine Resource Economics, 25 (1):121–128, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See reply [WCA^{+}10].

Maza:2004:CPU

[Maz04] Arantza Murillas Maza. Common property under management flexibility: Valuation, optimal exploitation, and regulation. Marine Resource Economics, 19(2):173–194, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Macher:2010:OSE

[MB10] Claire Macher and Jean Boncoeur. Optimal selectivity and effort cost a simple bioeconomic model with an application to the Bay of Biscay Nephrops fishery. Marine Resource Economics, 25 (2):213–232, June 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Munro:2000:BM

[MBD^{+}00] Gordon R. Munro, Ana Brasão, Clara Costa Duarte, Clara Costa-Duarte, Maria A. Cunha e Sá, Ragnar Arnason, Gylfi Magnusson, Sveinn Agnarsson, Trond Bjørndal, Daniel V. Gordon, Gordon R. Munro, Pedro Pintassilgo, and Clara Costa Duarte. Back matter. Marine Resource Economics, 15(4):??, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Memon:1992:FPT

[MC92] Pyar Ali Memon and Ross Cullen. Fishery policies and their impact on the New Zealand Maori. Marine Resource Economics, 7(3):153–167, Fall 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Macnair:1999:HNR

[MC99] Douglas J. Macnair and Sean D. Cox. A heteroskedastic nested RUM of freshwater fishing. Marine Resource Economics, 14(4): 333–341, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Matulich:2003:NPH

[MC03] Scott C. Matulich and Michael L. Clark. North Pacific halibut and sablefish IFQ policy design: Quantifying the impacts on processors. Marine Resource Economics, 18(2):149–166, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Memon:1992:BM

[MCC^{+}92] Pyar Ali Memon, Ross Cullen, Raymond P. Clarke, Stacey S. Yoshimoto, Samuel G. Pooley, Steven F. Edwards, Raymond P. Clarke, Stacey S. Yoshimoto, Samuel G. Pooley, Pyar Ali Memon, Ross Cullen, and Steven F. Edwards. Back matter. Marine Resource Economics, 7(3):??, Fall 1992. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

McCay:2002:BRR

[McC02] Bonnie J. McCay. Book review: Rosemary E. Ommer, The Resilient Outport: Ecology, Economy, and Society in Rural Newfoundland. Marine Resource Economics, 17(4):341–343, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Martinez-Cordero:2009:MEO

[MCFH09] Francisco Javier Martínez-Cordero, Quentin Sai Wing Fong, and Maria C. Haws. Marketing extension and outreach in Sinaloa, Mexico: a preliminary analysis of preferences for oysters. Marine Resource Economics, 24(1):89–95, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

McConnell:1999:SPE

[McT99] Kenneth E. McConnell and Wei chun Tseng. Some preliminary evidence on sampling of alternatives with the random parameters logit. Marine Resource Economics, 14(4):317–332, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Miller:2017:PME

[MD17] Steve J. Miller and Robert T. Deacon. Protecting marine ecosystems: Regulation versus market incentives. Marine Resource Economics, 32(1):83–107, January 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Munro:2015:TAD

[MDC^{+}15] Gordon R. Munro, Jorge Dresdner, Carlos Chávez, Omar Barriga, Jon M. Conrad, Daniel Rondeau, Daniel S. Holland, Patricia Pinto da Silva, Andrew W. Kitts, Daniel S. Holland, Patricia Pinto da Silva, Andrew W. Kitts, Jon M. Conrad, Daniel Rondeau, Andrew K. G. Tan, Steven T. Yen, and Abdul Rahman Hasan. Tribute to Anthony Dalton Scott. Marine Resource Economics, 30(4):439–441, October 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Milon:1988:BM

[MDS^{+}88] J. Walter Milon, James W. Dunn, James S. Shortle, John J. Poggie, Jr., Richard B. Pollnac, Miguel Fierro, J. Walter Milon, James W. Dunn, and James S. Shortle. Back matter. Marine Resource Economics, 5(3):??, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Madsen:2023:UFB

[MEA^{+}23] Jens Koed Madsen, Rani Ekawaty, Aarthi Ananthanarayanan, Richard Bailey, Ernesto Carrella, Chris Dorsett, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Min-Yang Lee, Scott Steinback, and Kristy Wallmo. Understanding fisher behavior: The case of snapper fishers in Indonesia. Marine Resource Economics, 38(1):85–100, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Meredith:2023:DAR

[Mer23] Jennifer Meredith. Drivers of access right sales: The role of resource volatility and individual shocks in the Alaska salmon fishery. Marine Resource Economics, 38(4):413–434, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Meuriot:1995:BRM

[Meu95] Eric Meuriot. Book review: M. J. Holden, The Common Fisheries Policy — Origin, Evaluation and Future. Marine Resource Economics, 10(3):321–322, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Murray:2002:CNU

[MF02] Andrew D. Murray and Abdulai Fofana. The changing nature of UK fish retailing. Marine Resource Economics, 17(4):335– 339, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Moore:2022:EIO

[MF22] Chris Moore and Jasmine Fuller. Economic impacts of ocean acidification: a meta-analysis. Marine Resource Economics, 37 (2):201–219, April 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Moeltner:2023:HAB

[MFF^{+}23] Klaus Moeltner, Tracy Fanara, Hosein Foroutan, Regina Hanlon, Vince Lovko, Shane Ross, and David Schmale III. Harmful algal blooms and toxic air: The economic value of improved forecasts. Marine Resource Economics, 38(1):1–28, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mishra:2016:HPN

[MG16] Ashok K. Mishra and Jeffrey M. Gillespie. Hiring preferences for nonimmigrant labor: The case of the seafood processing industry. Marine Resource Economics, 31(1):83–99, January 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

McCay:1989:LLP

[MGC89] Bonnie J. McCay, John B. Gatewood, and Carolyn F. Creed. Labor and the labor process in a limited entry fishery. Marine Resource Economics, 6(4):311–330, ???? 1989. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Hung:2006:IUC

[mHS06] Chih ming Hung and Daigee Shaw. The impact of upstream catch and global warming on the grey mullet fishery in Taiwan: a non-cooperative game analysis. Marine Resource Economics, 21(3):285–300, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mikkelsen:2007:AFI

[Mik07] Eirik Mikkelsen. Aquaculture-fisheries interactions. Marine Resource Economics, 22(3):287–303, ???? 2007. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Milon:1988:NDS

[Mil88] J. Walter Milon. A nested demand shares model of artificial marine habitat choice by sport anglers. Marine Resource Economics, 5(3):191–213, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Muhammad:2011:SBP

[MJ11] Andrew Muhammad and Keithly G. Jones. Source-based preferences and U.S. salmon imports. Marine Resource Economics, 26 (3):191–209, September 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mao:2020:MIT

[MJ20] Junwei Mao and Sunny L. Jardine. Market impacts of a toxic algae event: The case of California Dungeness crab. Marine Resource Economics, 35(1):1–20, January 2020. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mediodia:2023:SST

[MKN23] Hanny John Mediodia, Viktoria Kahui, and Ilan Noy. Sea surface temperature and tuna catch in the Eastern Pacific Ocean under climate change. Marine Resource Economics, 38(4):329– 351, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Miller:2004:CCN

[MM04] Kathleen A. Miller and Gordon R. Munro. Climate and cooperation: a new perspective on the management of shared fish stocks. Marine Resource Economics, 19(3):367–393, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Madani:2013:CVC

[MMCM13] Shima Madani, Adán L. Martínez-Cruz, and Kenneth E. Mc-Connell. Conservation value of coral reefs around Kish Island, Iran. Marine Resource Economics, 28(4):331–343, December 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Morgan:2009:DDL

[MMH09a] O. Ashton Morgan, D. Matthew Massey, and William L. Huth. Diving demand for large ship artificial reefs. Marine Resource Economics, 24(1):43–59, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Morgan:2009:BM

[MMH^{+}09b] O. Ashton Morgan, D. Matthew Massey, William L. Huth, Wuyang Hu, Katya Boehle, Linda Cox, Minling Pan, James E. Wilen, Wuyang Hu, Katya Boehle, Linda Cox, Minling Pan, O. Ashton Morgan, D. Matthew Massey, William L. Huth, and James E. Wilen. Back matter. Marine Resource Economics, 24 (1):??, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Miller:2004:BM

[MMP^{+}04] Kathleen A. Miller, Gordon R. Munro, Hoanjae Park, Walter N. Thurman, J. E. Easley, Jr., Ayoe Hoff, Kathleen A. Miller, Gordon R. Munro, James Kirkley, Catherine J. Morrison Paul, Dale Squires, and Ayoe Hoff. Back matter. Marine Resource Economics, 19(3):??, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Maroto:2012:PCF

[MMSS12] Jose M. Maroto, Manuel Moran, Leif K. Sandal, and Stein I. Steinshamn. Potential collapse in fisheries with increasing returns and stock-dependent costs. Marine Resource Economics, 27(1):43–63, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Misund:2018:BFV

[MN18] Bård Misund and Rune Nygård. Big fish: Valuation of the world’s largest salmon farming companies. Marine Resource Economics, 33(3):245–261, July 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Managi:2004:FES

[MOJG04] Shunsuke Managi, James J. Opaluch, Di Jin, and Thomas A. Grigalunas. Forecasting energy supply and pollution from the offshore oil and gas industry. Marine Resource Economics, 19 (3):307–332, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Misund:2023:PES

[MOOT23] Bård Misund, Marit Schei Olsen, Tonje C. Osmundsen, and Ragnar Tveterås. The political economy of salmon aquaculture: Value sharing and societal support for aquaculture in Norway. Marine Resource Economics, 38(4):365–390, ???? 2023. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mangel:1985:RMI

[MP85] Marc Mangel and Richard E. Plant. Regulatory mechanisms and information processing in uncertain fisheries. Marine Resource Economics, 1(4):389–418, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Missios:1997:TRR

[MP97] Paul C. Missios and Charles Plourde. Transboundary renewable resource management and conservation motives. Marine Resource Economics, 12(1):29–36, Spring 1997. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mardle:1999:RAM

[MP99] Simon Mardle and Sean Pascoe. A review of applications of multiple-criteria decision-making techniques to fisheries. Marine Resource Economics, 14(1):41–63, Spring 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Myers:2010:MRU

[MPE10] Kelley H. Myers, George R. Parsons, and Peter E. T. Edwards. Measuring the recreational use value of migratory shorebirds on the Delaware Bay. Marine Resource Economics, 25(3):247–264, September 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mardle:1999:BM

[MPH^{+}99] Simon Mardle, Sean Pascoe, Frances R. Homans, Jane A. Ruliffson, Nerissa D. Salayo, Thomas J. P. Voon, Saroja Selvanathan, Simon Mardle, Sean Pascoe, Frances R. Homans, Jane A. Ruliffson, Nerissa D. Salayo, Thomas J. P. Voon, and Saroja Selvanathan. Back matter. Marine Resource Economics, 14(1): ??, Spring 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

MacNair:2022:TEE

[MPTB22] Doug MacNair, George Parsons, Theodore Tomasi, and Heath Byrd. Trip equivalency for economic valuation in recreation demand models: Implications for compensatory restoration and benefits transfer. Marine Resource Economics, 37(1):91–107, January 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mistiaen:1998:OFH

[MS98] Johan A. Mistiaen and Ivar Strand. Optimal feeding and harvest time for fish with weight-dependent prices. Marine Resource Economics, 13(4):231–246, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Martinez:2001:ACS

[MS01] J. Antonio Martinez and J. Carlos Seijo. Alternative cycling strategies for shrimp farming in arid zones of Mexico: Dealing with risk and uncertainty. Marine Resource Economics, 16(1): 51–63, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mukherjee:2011:TED

[MS11] Zinnia Mukherjee and Kathleen Segerson. Turtle excluder device regulation and shrimp harvest: The role of behavioral and market responses. Marine Resource Economics, 26(3):173–189, September 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

McConnell:1995:RUM

[MSBh95] Kenneth E. McConnell, Ivar E. Strand, and Lynne Blake-hedges. Random utility models of recreational fishing: catching fish using a Poisson process. Marine Resource Economics, 10(3):247– 261, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Matulich:2001:FCA

[MSI01] Scott C. Matulich, Murat Sever, and Fred Inaba. Fishery cooperatives as an alternative to ITQs: Implications of the American fisheries act. Marine Resource Economics, 16(1):1–16, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Manning:2016:TRB

[MU16] Dale T. Manning and Hirotsugu Uchida. Are two rents better than none? When monopolies correct ill-defined property rights. Marine Resource Economics, 31(2):141–164, April 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Montanio:2023:ICC

[MUK23] Kyle Montanio, Emi Uchida, and Katrina Kosec. Impacts of conditional cash transfers on fish consumption: Evidence from Tanzania. Marine Resource Economics, 38(4):391–411, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Munro:1986:MSF

[Mun86] Gordon R. Munro. The management of shared fishery resources under extended jurisdiction. Marine Resource Economics, 3 (4):271–296, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Munro:2000:UNF

[Mun00] Gordon R. Munro. The United Nations Fish Stocks Agreement of 1995: History and problems of implementation. Marine Resource Economics, 15(4):265–280, ???? 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Munro:2007:ISF

[Mun07] Gordon R. Munro. Internationally shared fish stocks, the high seas, and property rights in fisheries. Marine Resource Economics, 22(4):425–443, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mazzocco:2023:RBC

[MUWB23] Vivianne Mazzocco, Hirotsugu Uchida, Michael J. Weir, and David Bidwell. Risky business: Can oyster farmers defend themselves against foodborne illness-related demand shocks? Marine Resource Economics, 39(1):1–20, ???? 2023. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Marco:2021:IUQ

[MVRRP21] Jorge Marco, Diego Valderrama, Mario Rueda, and Maykol Rodríguez-Prieto. Improving utilization of the queen conch (Aliger gigas) resource in Colombia. Marine Resource Economics, 36(4):411–437, October 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Mariojouls:2002:CQS

[MW02] Catherine Mariojouls and Cathy Roheim Wessells. Certification and quality signals in the aquaculture sector in France. Marine Resource Economics, 17(2):175–180, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Morgan:2018:MIT

[MWH^{+}18] O. Ashton Morgan, John C. Whitehead, William L. Huth, Xavier Irz, Pascal Leroy, Vincent Réquillart, Louis-Georges Soler, Kentaka Aruga, Hiroki Wakamatsu, Ly Nguyen, Henry W.

Kinnucan, O. Ashton Morgan, John C. Whitehead, William L. Huth, Ly Nguyen, Henry W. Kinnucan, Kentaka Aruga, and Hiroki Wakamatsu. Measuring the impact of traceability information on oyster consumer behavior following a contamination event. Marine Resource Economics, 33(4):387–400, October 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Neher:1988:I

[NA88] Philip A. Neher and Ragnar Arnason. Introduction. Marine Resource Economics, 5(4):285–287, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nielsen:2014:DRM

[NAB14] Rasmus Nielsen, Jesper Levring Andersen, and Peter Bogetoft. Dynamic reallocation of marketable nitrogen emission permits in Danish freshwater aquaculture. Marine Resource Economics, 29(3):219–239, September 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nostbakken:2003:SFN

[NB03] Linda Nøstbakken and Trond Bjørndal. Supply functions for North Sea herring. Marine Resource Economics, 18(4):345–361, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Neugebauer:2007:BSH

[NC07] Tibor Neugebauer and Begoña Casino. Bioeconomics of sustainable harvest of competing species: a comment. Marine Resource Economics, 22(1):109–113, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nepf:2022:EVC

[NDW22] Mark Nepf, Anthony Dvarskas, and Patrick J. Walsh. Economic valuation for coastal water infrastructure planning: Analysis of the housing market and nutrient pollution in Suffolk County, NY. Marine Resource Economics, 37(4):369–386, October 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Neher:1993:BRT

[Neh93] Philip A. Neher. Book review: Tadashi Yamamoto and Kevin Short, International Perspectives on Fisheries Management.

With Special Emphasis on Community-Based Management Systems Developed in Japen. Marine Resource Economics, 8(4): 403–405, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ngoc:2010:PAC

[NF10] Quach Thi Khanh Ngoc and Ola Flaaten. Protected areas for conflict resolution and management of recreational and commercial fisheries. Marine Resource Economics, 25(4):409–426, December 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nielsen:2012:MER

[NFW12] Max Nielsen, Ola Flaaten, and Staffan Waldo. Management of and economic returns from selected fisheries in the Nordic countries. Marine Resource Economics, 27(1):65–88, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nauman:1995:BM

[NGB^{+}95] Fayyaz A. Nauman, Conrado M. Gempesaw, J. Richard Bacon, Alberto Manalo, Cathy Roheim Wessells, Christopher J. Miller, Priscilla M. Brooks, Richard F. Kazmierczak, Jr., Rex H. Caffey, Fayyaz A. Nauman, Conrado M. Gempesaw, J. Richard Bacon, Alberto Manalo, Cathy Roheim Wessells, Christopher J. Miller, Priscilla M. Brooks, Richard F. Kazmierczak, Jr., and Rex H. Caffey. Back matter. Marine Resource Economics, 10(2):??, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nauman:1995:CCF

[NGBM95] Fayyaz A. Nauman, Conrado M. Gempesaw, J. Richard Bacon, and Alberto Manalo. Consumer choice for fresh fish: Factors affecting purchase decisions. Marine Resource Economics, 10 (2):117–142, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nielsen:2005:PFM

[Nie05] Max Nielsen. Price formation and market integration on the European first-hand market for whitefish. Marine Resource Economics, 20(2):185–202, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nguyen:2018:WPT

[NK18] Ly Nguyen and Henry W. Kinnucan. World price transmission for differentiated products: The case of shrimp in the US market. Marine Resource Economics, 33(4):351–372, October 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Norman-Lopez:2009:CBD

[NL09] Ana Norman-López. Competition between different farmed and wild species: The US tilapia market. Marine Resource Economics, 24(3):237–251, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nieminen:2012:OBM

[NLH12] Emmi Nieminen, Marko Lindroos, and Outi Heikinheimo. Optimal bioeconomic multispecies fisheries management: a Baltic Sea case study. Marine Resource Economics, 27(2):115–136, June 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nguyen:2022:MLA

[NNNJ22] Kim Anh Thi Nguyen, Tram Anh Thi Nguyen, Brice M. Nguelifack, and Curtis M. Jolly. Machine learning approaches for predicting willingness to pay for shrimp insurance in Vietnam. Marine Resource Economics, 37(2):155–182, April 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Naito:1997:AHM

[NP97] Toyokazu Naito and Stephen Polasky. Analysis of a highly migratory fish stocks fishery: a game theoretic approach. Marine Resource Economics, 12(3):179–201, Fall 1997. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Neira:2006:ARB

[NQ06] Ivano Neira and Kwamena Quagrainie. Analysis of risk behavior in the U.S. farm-raised catfish market. Marine Resource Economics, 21(4):433–443, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nielsen:2012:PEC

[NSG12] Max Nielsen, Jos Smit, and Jordi Guillen. Price effects of changing quantities supplied at the integrated European fish market.

Marine Resource Economics, 27(2):165–180, June 2012. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nielsen:2007:MIF

[NSL^{+}07] Max Nielsen, Jari Setälä, Jukka Laitinen, Kaija Saarni, Jarno Virtanen, and Asmo Honkanen. Market integration of farmed trout in Germany. Marine Resource Economics, 22(2):195–213, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nguyen:2022:VMC

[NSV^{+}22] Thanh Viet Nguyen, Michel Simioni, Hung Trung Vo, Anna M. Birkenbach, Andreea L. Cojocaru, Martin D. Smith, Frank Asche, Kristin H. Roll, Rune Nygaard, Benjamin Fissel, James Hilger, Daniel R. Petrolia, William C. Walton, and Just Cebrian. Valuing mangrove conservation attributes in Red River Delta, Vietnam: a choice experiment approach. Marine Resource Economics, 37(3):349–368, July 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nostbakken:2011:IBC

[NTS11] Linda Nøstbakken, Olivier Thébaud, and Lars-Christian Sørensen. Investment behaviour and capacity adjustment in fisheries: a survey of the literature. Marine Resource Economics, 26(2):95–117, June 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Nowlis:2012:DPR

[NV12] Josh Nowlis and Arthur A. Van Benthem. Do property rights lead to sustainable catch increases? Marine Resource Economics, 27(1):89–105, March 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ogundari:2010:MTE

[OA10] Kolawole Ogundari and Olajide O. Akinbogun. Modeling technical efficiency with production risk: a study of fish farms in Nigeria. Marine Resource Economics, 25(3):295–308, September 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Omezzine:2004:MSF

[OAA04] Abdallah Omezzine, Hamed Al Oufi, and Younis Al Akhzami. The marketing system of fresh fish products on Masirah Island in the Sultanate of Oman: a comprehensive reassessment. Marine Resource Economics, 19(3):407–416, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Opaluch:1984:BMF

[OB84] James J. Opaluch and Nancy E. Bockstael. Behavioral modeling and fisheries management. Marine Resource Economics, 1 (1):105–115, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Opaluch:1984:BM

[OBB^{+}84] James J. Opaluch, Nancy E. Bockstael, Nancy E. Bockstael, M. P. Sissenwine, James J. Opaluch, Nancy E. Bockstael, Fred J. Prochaska, and J. M. Gates. Back matter. Marine Resource Economics, 1(1):??, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ottesen:2003:PUS

[OG03] Geir Grundvåg Ottesen and Kjell Grønhaug. Primary uncertainty in the seafood industry: an exploratory study of how processing firms cope. Marine Resource Economics, 18(4):363– 371, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Oinonen:2016:IFM

[OGL^{+}16] Soile Oinonen, Lone Grønbæk, Marita Laukkanen, Polina Levontin, Marko Lindroos, Emmi Nieminen, Katja Parkkila, Pedro Pintassilgo, Henni Pulkkinen, and Atso Romakkaniemi. International fisheries management and recreational benefits: The case of Baltic salmon. Marine Resource Economics, 31(4):433– 451, October 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

OHara:2020:FMS

[O’H20] Jeffrey K. O’Hara. Farmers markets and seafood: Where is it feasible? Marine Resource Economics, 35(4):411–426, October 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Onozaka:2014:CPP

[OHS14] Yuko Onozaka, Håvard Hansen, and Arne Sørvig. Consumer product perceptions and salmon consumption frequency: The role of heterogeneity based on food lifestyle segments. Marine Resource Economics, 29(4):351–374, December 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Olaussen:2007:PCS

[Ola07] Jon Olaf Olaussen. Playing chicken with salmon. Marine Resource Economics, 22(2):173–193, ???? 2007. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Olaussen:2009:BSA

[Ola09] Jon Olaf Olaussen. Bandwagon or snob anglers? Evidence from Atlantic salmon recreational fishing. Marine Resource Economics, 24(4):387–403, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ovando:2021:CAA

[OLM^{+}21] Daniel Ovando, Gary D. Libecap, Katherine D. Millage, Lennon Thomas, Daniel Ovando, Gary D. Libecap, Katherine D. Millage, Lennon Thomas, Qiujie Zheng, H. Holly Wang, Jason F. Shogren, Qingran Li, Martin D. Smith, Md Takibur Rahman, Rasmus Nielsen, Md Akhtaruzzaman Khan, Dewan Ahsan, Qiujie Zheng, H. Holly Wang, Jason F. Shogren, Daniel Ovando, Gary D. Libecap, Katherine D. Millage, and Lennon Thomas. Coasean approaches to address overfishing: Bigeye tuna conservation in the Western and Central Pacific Ocean. Marine Resource Economics, 36(1):91–109, January 2021. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Opaluch:1987:MPE

[Opa87] James J. Opaluch. Marine pollution and environmental damage assessment: Introduction. Marine Resource Economics, 4 (3):151–154, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Oglend:2008:BSP

[OS08] Atle Oglend and Marius Sikveland. The behaviour of salmon price volatility. Marine Resource Economics, 23(4):507–526, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ostbye:1999:TNI

[Øst99] Stein Østbye. A technical note on input price proxies used in salmon farming industry studies. Marine Resource Economics, 14(3):215–223, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Omezzine:1996:MSF

[OZAo96] Abdallah Omezzine, Lokman Zaibet, and Hamad Al-oufi. The marketing system of fresh fish products on the Masirah Island in the Sultanate of Oman. Marine Resource Economics, 11(3): 203–210, Fall 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parsons:2009:PNL

[PA09] George R. Parsons and James L. Anderson. Preface: a nested logit model of recreational fishing demand in Alaska. Marine Resource Economics, 24(2):99–100, ???? 2009. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pettersen:2020:HPA

[PA20] Ingrid Kristine Pettersen and Frank Asche. Hedonic price analysis of ex-vessel cod markets in Norway. Marine Resource Economics, 35(4):343–359, October 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pincinato:2022:ITF

[PAC^{+}22] Ruth B. M. Pincinato, Frank Asche, Andreea L. Cojocaru, Yaqin Liu, and Kristin H. Roll. The impact of transferable fishing quotas on cost, price, and season length. Marine Resource Economics, 37(1):53–63, January 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Panayotou:1987:BRR

[Pan87] Theodore Panayotou. Book review: Rowena Lawson, Economics of Fisheries Development. Marine Resource Economics, 4(4): 323–328, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parsons:1987:OCR

[Par87] George R. Parsons. The opportunity costs of residential displacement due to coastal land use restrictions: a conceptual framework. Marine Resource Economics, 4(2):111–122, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pascoe:1993:T

[Pas93] Sean Pascoe. Thalassorama. Marine Resource Economics, 8 (4):395–401, Winter 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pascoe:2007:CAF

[Pas07] Sean Pascoe. Capacity analysis and fisheries policy: Theory versus practice. Marine Resource Economics, 22(1):83–87, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pascoe:2012:BR

[Pas12] Sean Pascoe. Books reviewed. Marine Resource Economics, 27 (2):193–194, June 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Plourde:1984:UFE

[PB84] Charles Plourde and Richard Bodell. Uncertainty in fisheries economics: The role of the discount rate. Marine Resource Economics, 1(2):155–170, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parks:1994:NUR

[PB94] Peter J. Parks and Manuel Bonifaz. Nonsustainable use of renewable resources: Mangrove deforestation and mariculture in Ecuador. Marine Resource Economics, 9(1):1–18, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parks:1994:BM

[PBA^{+}94] Peter J. Parks, Manuel Bonifaz, Robert A. Androkovich, Kenneth R. Stollery, Ali Emami, Lewis E. Queirolo, Richard S. Johnston, Peter J. Parks, Manuel Bonifaz, Robert A. Androkovich, and Kenneth R. Stollery. Back matter. Marine Resource Economics, 9(1):??, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Padilla:1994:BAM

[PC94] Jose E. Padilla and Parzival Copes. Bioeconomic analysis of management options for tropical fisheries using a bicriteria programming model. Marine Resource Economics, 9(1):47–66, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Perez-Cobb:2014:ARF

[PCAIGO^{+}14] Addiel U. Perez-Cobb, A. Minerva Arce-Ibarra, Martha García-Ortega, Martha Valdéz-Moreno, and James Oscar Azueta. Artisanal recreational fisheries: Using a combined approach to fishery assessment aimed at providing insights for fishery managers. Marine Resource Economics, 29(2):89–109, June 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parsons:2013:VBW

[PCH^{+}13] George R. Parsons, Zhe Chen, Michael K. Hidrue, Naomi Standing, and Jonathan Lilley. Valuing beach width for recreational use: Combining revealed and stated preference data. Marine Resource Economics, 28(3):221–241, September 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pintassilgo:2000:NMP

[PD00] Pedro Pintassilgo and Clara Costa Duarte. The new-member problem in the cooperative management of high seas fisheries. Marine Resource Economics, 15(4):361–378, ???? 2000. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pintassilgo:2002:OMN

[PD02] Pedro Pintassilgo and Clara Costa Duarte. Optimal management of the Northern Atlantic bluefin tuna. Marine Resource Economics, 17(1):47–67, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Petesch:2021:IDS

[PDS^{+}21] Tess Petesch, Bradford Dubik, Martin D. Smith, Roy Endré Dahl, Atle Oglend, Muhammad Yahya, Anna Andersson, Johan Blomquist, Staffan Waldo, Tess Petesch, Bradford Dubik, Martin D. Smith, Afjal Hossain, Max Nielsen, Isaac Ankamah-Yeboah, Badiuzzaman, Fakir Azmal Huda, Roy Endré Dahl, Atle Oglend, Muhammad Yahya, Max T. Stoeven, Florian K. Diekert, and Martin F. Quaas. Implications of disease in shrimp aquaculture for wild-caught shrimp. Marine Resource Economics, 36(2):191–209, April 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Petrolia:2025:VOR

[PECY25] Daniel R. Petrolia, Freedom Enyetornye, Zhenshan Chen, and Seong Yun. The value of oyster reef restoration. Marine Resource Economics, 40(1):45–64, ???? 2025. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pham:2013:RSE

[PFN13] Thi Thanh Thuy Pham, Ola Flaaten, and Thi Kim Anh Nguyen. Remuneration systems and economic performance: Theory and Vietnamese small-scale purse seine fisheries. Marine Resource Economics, 28(1):19–41, January 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pham:2021:SHH

[PFNV21] Thuy Thi Thanh Pham, Ola Flaaten, Luong Trong Nguyen, and Nghiep Ke Vu. Subsidies — help or hurt? A study from Vietnamese fisheries. Marine Resource Economics, 36(4):369– 387, October 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Peridy:2000:IPS

[PGB00] Nicolas Péridy, Patrice Guillotreau, and Pascal Bernard. The impact of prices on seafood trade: a panel data analysis of the French seafood market. Marine Resource Economics, 15(1):45– 66, Spring 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Petrolia:2009:WBI

[PgK09] Daniel R. Petrolia and Tae goun Kim. What are barrier islands worth? Estimates of willingness to pay for restoration. Marine Resource Economics, 24(2):131–146, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Phaneuf:1999:CSD

[PH99] Daniel J. Phaneuf and Joseph A. Herriges. Choice set definition issues in a Kuhn–Tucker model of recreation demand. Marine Resource Economics, 14(4):343–355, Winter 1999. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Petrolia:2020:AAN

[PH20] Daniel R. Petrolia and Joonghyun Hwang. Accounting for attribute non-attendance in three previously published choice studies of coastal resources. Marine Resource Economics, 35 (3):219–240, July 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Penn:2016:VRB

[PHCK16] Jerrod Penn, Wuyang Hu, Linda Cox, and Lara Kozloff. Values for recreational beach quality in Oahu, Hawaii. Marine Resource Economics, 31(1):47–62, January 2016. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Petrolia:2014:AWN

[PIH14] Daniel R. Petrolia, Matthew G. Interis, and Joonghyun Hwang. America’s wetland? A national survey of willingness to pay for restoration of Louisiana’s coastal wetlands. Marine Resource Economics, 29(1):17–37, March 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parsons:2009:VBC

[PKLB09] George R. Parsons, Ami K. Kang, Christopher G. Leggett, and Kevin J. Boyle. Valuing beach closures on the Padre Island National Seashore. Marine Resource Economics, 24(3):213–235, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parsons:1999:FFS

[PMT99a] George R. Parsons, D. Matthew Massey, and Ted Tomasi. Familiar and favorite sites in a random utility model of beach recreation. Marine Resource Economics, 14(4):299–315, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Parsons:1999:BM

[PMT^{+}99b] George R. Parsons, D. Matthew Massey, Ted Tomasi, John C. Whitehead, Timothy C. Haab, Robert L. Hicks, Amy B. Gautam, David Van Voorhees, Maury Osborn, Brad Gentner, George R. Parsons, D. Matthew Massey, Ted Tomasi, Carol A. Jones, Frank Lupi, Kenneth E. McConnell, and Wei chun Tseng. Back matter. Marine Resource Economics, 14(4):??, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pollnac:1986:BRW

[Pol86] Richard B. Pollnac. Book review: W. H. L. Allsopp, Fishery Development Experiences. Marine Resource Economics, 3(4):391– 392, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pontecorvo:1988:SWF

[Pon88] Giulio Pontecorvo. The state of worldwide fishery statistics: a modest proposal. Marine Resource Economics, 5(2):79–81, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See comments [RC89, Gul89] and reply [Pon89].

Pontecorvo:1989:RCF

[Pon89] Giulio Pontecorvo. Reply to comments by Francis T. Christy, Jr. and Michael Robinson and John Gulland on “The State of Worldwide Fishery Statistics: a Modest Proposal”. Marine Resource Economics, 6(1):87–88, ???? 1989. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [Pon88, RC89, Gul89].

Pooley:1994:MLF

[Poo94] Sam Pooley. Managing longline fishing in Hawaii — practical aspects of regulatory economics. Marine Resource Economics, 9(1):77–86, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pooley:1997:BRA

[Poo97] Sam Pooley. Book review: Alan Christopher Finlayson, Fishing for Truth: a Sociological Analysis of Northern Cod Stock Assessments from 1977–1990. Marine Resource Economics, 12(3): 249–251, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Powers:2004:SIU

[Pow04] Joseph E. Powers. Strategic interaction in United States fishery management councils. Marine Resource Economics, 19(4):417– 438, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Poggie:1988:FIS

[PPF88] John J. Poggie, Jr., Richard B. Pollnac, and Miguel Fierro. Factors influencing the success of fishermen’s cooperatives in Ecuador. Marine Resource Economics, 5(3):231–242, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Petrolia:2019:KTB

[PPQ^{+}19] Daniel R. Petrolia, Jerrod Penn, Ruth Quainoo, Rex H. Caffey, and J. Matthew Fannin. Know thy beach: Values of beach condition information. Marine Resource Economics, 34(4):331– 359, October 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Poggie:1996:IVC

[PPV96] J. J. Poggie, Jr., R. B. Pollnac, and C. Van Dusen. Intracultural variability in the cognition of danger among Southern New England fishers. Marine Resource Economics, 11(1):23–30, Spring 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pfeiffer:2022:SCR

[PPV22] Lisa Pfeiffer, Tess Petesch, and Thamanna Vasan. A safer catch? The role of fisheries management in fishing safety. Marine Resource Economics, 37(1):1–33, January 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Park:1999:NPP

[PR99] Seong K. Park and Jeong G. Ryu. New policy paradigms for Korean fisheries’ transition to responsible practices. Marine Resource Economics, 14(1):79–93, Spring 1999. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Prochaska:1984:PTU

[Pro84] Fred J. Prochaska. Principal types of uncertainty in seafood processing and marketing. Marine Resource Economics, 1(1): 51–66, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Placenti:1992:BEM

[PRS92] V. Placenti, G. Rizzo, and M. Spagnolo. A bio-economic model for the optimization of a multi-species, multi-gear fishery: The Italian case. Marine Resource Economics, 7(4):275–295, Winter 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Plourde:1989:CSF

[PS89] Charles Plourde and J. Barry Smith. Crop sharing in the fishery and industry equilibrium. Marine Resource Economics, 6 (3):179–193, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Plourde:1994:CSF

[PS94] Charles Plourde and J. Barry Smith. Crop sharing in the fishery and industry equilibrium: Reply. Marine Resource Economics, 9(1):73–76, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Phaneuf:2003:AKT

[PS03] Daniel J. Phaneuf and Chris Siderelis. An application of the Kuhn–Tucker model to the demand for water trail trips in North Carolina. Marine Resource Economics, 18(1):1–14, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Poudel:2015:SIC

[PSK15] Diwakar Poudel, Leif K. Sandal, and Sturla F. Kvamsdal. Stochastically induced critical depensation and risk of stock collapse. Marine Resource Economics, 30(3):297–313, July 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Poudel:2013:FMU

[PSKS13] Diwakar Poudel, Leif K. Sandal, Sturla F. Kvamsdal, and Stein I. Steinshamn. Fisheries management under irreversible investment: Does stochasticity matter? Marine Resource Economics, 28(1):83–103, January 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pradhan:2003:ATE

[PSL03] Naresh C. Pradhan, Khem R. Sharma, and Pingsun Leung. Analyzing technical and economic interrelationships in Hawaii’s longline fishery. Marine Resource Economics, 18(2):167–193, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pena-Torres:1997:PEF

[PT97] Julio Peña-Torres. The political economy of fishing regulation: The case of Chile. Marine Resource Economics, 12(4):253–280, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pena-Torres:2019:CSQ

[PTDQ^{+}19] Julio Peña-Torres, Jorge Dresdner, Felipe Quezada, Iván Luzardo, Julio Peña-Torres, Jorge Dresdner, Felipe Quezada, Iván Luzardo, Peggy Schrobback, Sean Pascoe, Rui Zhang, Peggy Schrobback, Sean Pascoe, Rui Zhang, Martin F. Quaas, Olli Tahvonen, Julio Peña-Torres, Jorge Dresdner, Felipe Quezada, and Iván Luzardo. Collective share quotas and the role of fishermen’s organizations in ex-vessel price determination. Marine Resource Economics, 34(4):361–385, October 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pena-Torres:2017:NNF

[PTDV17] Julio Peña-Torres, Jorge Dresdner, and Felipe Vasquez. El Niño and fishing location decisions: The Chilean straddling jack mackerel fishery. Marine Resource Economics, 32(3):249–275, July 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Park:2004:MID

[PTE04] Hoanjae Park, Walter N. Thurman, and J. E. Easley, Jr. Modeling inverse demands for fish: Empirical welfare measurement in Gulf and South Atlantic fisheries. Marine Resource Economics, 19(3):333–351, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pena-Torres:2022:EDC

[PTMQ22] Julio Peña-Torres, Roberto Muñoz, and Felipe Quezada. Entry deterrence and collusion at repeated multiunit auctions of ITQs. Marine Resource Economics, 37(4):437–465, October 2022. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Pena-Torres:1999:WCF

[PTSB99] Julio Peña-Torres, Rodolfo Serra, and Michael Basch. Who can fish what and where: Chile’s tradeoffs in high seas fishing of straddling stocks. Marine Resource Economics, 14(3):245–262, Fall 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Petrolia:2022:OES

[PWC22] Daniel R. Petrolia, William C. Walton, and Just Cebrian. Oyster economics: Simulated costs, market returns, and nonmarket ecosystem benefits of harvested and nonharvested reefs, offbottom aquaculture, and living shorelines. Marine Resource Economics, 37(3):325–347, July 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Perruso:2005:POT

[PWL05a] Larry Perruso, Richard N. Weldon, and Sherry L. Larkin. Predicting optimal targeting strategies in multispecies fisheries: a portfolio approach. Marine Resource Economics, 20(1):25–45, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Perruso:2005:BM

[PWL^{+}05b] Larry Perruso, Richard N. Weldon, Sherry L. Larkin, Steven C. Hackett, Matthew J. Krachey, Sharon Brown, David Hankin, Aaron Hatcher, Jarno Virtanen, Jari Setälä, Kaija Saarni, Asmo Honkanen, Larry Perruso, Richard N. Weldon, Sherry L. Larkin, and Steve Edwards. Back matter. Marine Resource Economics, 20(1):??, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Plourde:1989:HTR

[PY89] Charles Plourde and David Yeung. Harvesting of a transboundary replenishable fish stock: a noncooperative game solution. Marine Resource Economics, 6(1):57–70, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Quaas:2019:SHA

[QT19] Martin F. Quaas and Olli Tahvonen. Strategic harvesting of age-structured populations. Marine Resource Economics, 34(4): 291–309, October 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Quagrainie:2006:AUC

[Qua06] Kwamena K. Quagrainie. Analysis of U.S. catfish fillet market share using a flexible logistic model. Marine Resource Economics, 21(1):33–45, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Quagrainie:2011:FIP

[QXH11] Kwamena K. Quagrainie, Ai Xing, and Kevin G. Hughes. Factors influencing the purchase of live seafood in the North Central region of the United States. Marine Resource Economics, 26(1): 59–74, March 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rickertsen:2017:FCA

[RAC^{+}17] Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, and Jason F. Shogren. French consumers’ attitudes and preferences toward wild and farmed fish. Marine Resource Economics, 32(1):59–81, January 2017. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Reimer:2017:EMF

[RAH17] Matthew N. Reimer, Joshua K. Abbott, and Alan C. Haynie. Empirical models of fisheries production: Conflating technology with incentives? Marine Resource Economics, 32(2):169–190, April 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Reimer:2017:FPM

[RAW17] Matthew N. Reimer, Joshua K. Abbott, and James E. Wilen. Fisheries production: Management institutions, spatial choice, and the quest for policy invariance. Marine Resource Economics, 32(2):143–168, April 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Robinson:1989:CPP

[RC89] Michael A. Robinson and Francis T. Christy, Jr. Comment on Professor Pontecorvo’s “The State of Worldwide Fishery Statistics: a Modest Proposal”. Marine Resource Economics, 6(1): 83–84, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [Pon88] and reply [Pon89].

Ropars-Collet:2017:CFA

[RCLL17] Carole Ropars-Collet, Mélody Leplat, and Philippe Le Goffe. Commercial fisheries as an asset for recreational demand on the coast: Evidence from a choice experiment. Marine Resource Economics, 32(4):391–409, October 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roheim:2009:BM

[RCN^{+}09] Cathy A. Roheim, Stephen Cunningham, Arthur E. Neiland, Michael Arbuckle, Tim Bostock, George R. Parsons, Ami K. Kang, Christopher G. Leggett, Kevin J. Boyle, Cathy A. Roheim, Stephen Cunningham, Arthur E. Neiland, Michael Arbuckle, Tim Bostock, and Ana Norman-López. Back matter. Marine Resource Economics, 24(3):??, ???? 2009. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Resosudarmo:1995:CBM

[Res95] Budy P. Resosudarmo. The construction of a bioeconomic model of the Indonesian flying fish fishery. Marine Resource Economics, 10(4):357–372, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rettig:1988:BRE

[Ret88] R. Bruce Rettig. Book review: Elmer A. Keen, Ownership and Productivity of Marine Fishery Resources. Marine Resource Economics, 5(3):275–278, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rettig:1994:WSP

[Ret94] R. Bruce Rettig. Who should pay to preserve the marine environment? Marine Resource Economics, 9(1):87–94, Spring 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Richardson:1986:EBA

[RG86] Edward J. Richardson and John M. Gates. Economic benefits of American lobster fishery management regulations. Marine Resource Economics, 2(4):353–382, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roheim:2007:VBO

[RGA07] Cathy A. Roheim, Lacey Gardiner, and Frank Asche. Value of brands and other attributes: Hedonic analysis of retail frozen fish in the UK. Marine Resource Economics, 22(3):239–253, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ryan:2010:BEB

[RHH10] Richard W. Ryan, Daniel S. Holland, and Guillermo E. Herrera. Bioeconomic equilibrium in a bait-constrained fishery. Marine Resource Economics, 25(3):281–293, September 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ryan:2014:EEF

[RHH14] Richard W. Ryan, Daniel S. Holland, and Guillermo E. Herrera. Ecosystem externalities in fisheries. Marine Resource Economics, 29(1):39–53, March 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ruddle:1992:MRM

[RHJ92a] K. Ruddle, E. Hviding, and R. E. Johannes. Marine resources management in the context of customary tenure. Marine Resource Economics, 7(4):249–273, Winter 1992. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ruddle:1992:BM

[RHJ^{+}92b] K. Ruddle, E. Hviding, R. E. Johannes, Cathy Roheim Wessells, James L. Anderson, R. K. Lindner, H. F. Campbell, G. F. Bevin, K. Ruddle, E. Hviding, R. E. Johannes, R. K. Lindner, H. F. Campbell, G. F. Bevin, Cathy Roheim Wessells, and James L. Anderson. Back matter. Marine Resource Economics, 7(4):??, Winter 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rust:2016:ECC

[RJY16] Steven Rust, Sarah Jennings, and Satoshi Yamazaki. Excess capacity and capital malleability in a fishery with myopic expectations. Marine Resource Economics, 31(1):63–81, January 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ropicki:2014:SNA

[RL14] Andrew J. Ropicki and Sherry L. Larkin. Social network analysis of price dispersion in fishing quota lease markets. Marine Resource Economics, 29(2):157–176, June 2014. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ropicki:2010:SSM

[RLA10] Andrew J. Ropicki, Sherry L. Larkin, and Charles M. Adams. Seafood substitution and mislabeling: WTP for a locally caught grouper labeling program in Florida. Marine Resource Economics, 25(1):77–93, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ray:2022:HPF

[RLK22] Kevin D. Ray, Daniel K. Lew, and Rosemary Kosaka. Hedonic price functions and market structure: an analysis of supplymotivated submarkets for salmon in California. Marine Resource Economics, 37(2):135–154, April 2022. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ray:2023:RCMb

[RLK23] Kevin D. Ray, Daniel K. Lew, and Rosemary Kosaka. A random coefficients model of seafood demand: Implications for consumer preferences and substitution patterns. Marine Resource Economics, 38(2):113–133, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roll:2022:UWS

[RNFH22] Kristin H. Roll, Rune Nygaard, Benjamin Fissel, and James Hilger. Are US wild salmon products affected by farmed salmon? A cointegration analysis. Marine Resource Economics, 37(3): 283–303, July 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rahman:2021:PRR

[RNKA21] Md Takibur Rahman, Rasmus Nielsen, Md Akhtaruzzaman Khan, and Dewan Ahsan. Perceived risk and risk management strategies in pond aquaculture. Marine Resource Economics, 36 (1):43–69, January 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Robinson:1993:BRJ

[Rob93] Michael A. Robinson. Book review: James R. McGoodwin, Crisis in the World’s Fisheries. Marine Resource Economics, 8(2): 173–175, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roheim:2003:BRB

[Roh03a] Cathy A. Roheim. Book review: Bruce Phillips, Trevor Ward, and Chet Chaffee, Eco-labelling in Fisheries: What is it all about? Marine Resource Economics, 18(4):375–378, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roheim:2003:EIM

[Roh03b] Cathy A. Roheim. Early indications of market impacts from the Marine Stewardship Council’s ecolabeling of seafood. Marine Resource Economics, 18(1):95–104, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roheim:2007:DCA

[Roh07] Cathy A. Roheim. Does capacity analysis help us meet fishery policy and management objectives? Marine Resource Economics, 22(1):??, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roheim:2009:ESS

[Roh09] Cathy A. Roheim. An evaluation of sustainable seafood guides: Implications for environmental groups and the seafood industry. Marine Resource Economics, 24(3):301–310, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roheim:2011:RJM

[Roh11] Cathy A. Roheim. Remembering John M. Gates. Marine Resource Economics, 26(2):169–170, June 2011. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rowse:2004:UWD

[Row04] John Rowse. Using the wrong discount rate to allocate a marine resource. Marine Resource Economics, 19(2):243–264, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Roy:1998:FBL

[Roy98] Noel Roy. Fishing behavior and the length of the fishing season. Marine Resource Economics, 13(3):197–213, Fall 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rosenthal:1988:EVO

[RRS88] Donald H. Rosenthal, Marshall B. Rose, and Lawrence J. Slaski. Economic value of the oil and gas resources on the Outer Continential Shelf. Marine Resource Economics, 5(3):171–189, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Rizzo:1996:MOM

[RS96] Gianfranco Rizzo and Massimo Spagnolo. A model for the optimal management of sea bass Dicentrarchus labrax aquaculture. Marine Resource Economics, 11(4):267–286, Winter 1996. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ruddle:1988:SPU

[Rud88] Kenneth Ruddle. Social principles underlying traditional inshore fishery management systems in the Pacific Basin. Marine Resource Economics, 5(4):351–363, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ruseski:1999:MPM

[Rus99] Gorazd Ruseski. Market power, management regimes, and strategic conservation of fisheries. Marine Resource Economics, 14(2):111–127, Summer 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ruseski:2000:BRR

[Rus00] Gorazd Ruseski. Book review: Ragnar Arnason, The Icelandic Fisheries: Evolution and Management of a Fishing Industry. Marine Resource Economics, 15(3):253–254, ???? 2000. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Reynolds:2000:SUF

[RW00] Julie A. Reynolds and James E. Wilen. The sea urchin fishery: Harvesting, processing and the market. Marine Resource Economics, 15(2):115–126, Summer 2000. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Straume:2020:DGD

[SAA^{+}20] Hans-Martin Straume, James L. Anderson, Frank Asche, Ivar Gaasland, Junwei Mao, Sunny L. Jardine, Hans-Martin Straume, James L. Anderson, Frank Asche, Ivar Gaasland, Hans-Martin Straume, James L. Anderson, Frank Asche, Ivar Gaasland, Daniel V. Gordon, Junwei Mao, and Sunny L. Jardine. Delivering the goods: The determinants of Norwegian seafood exports. Marine Resource Economics, 35(1):83–96, January 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Stage:2012:BR

[SAB^{+}12] Jesper Stage, Frank Asche, Lori S. Bennear, Atle Oglend, Martin D. Smith, Nguyen Tien Thong, Di Jin, Hauke Kite-Powell, Porter Hoagland, Andrew Solow, Frank Asche, Lori S. Bennear, Atle Oglend, Martin D. Smith, Max Nielsen, Jos Smit, Jordi Guillen, Di Jin, Hauke Kite-Powell, Porter Hoagland, and Andrew Solow. Books reviewed. Marine Resource Economics, 27 (2):195–196, June 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2014:SDH

[SABO14] Martin D. Smith, Frank Asche, Lori S. Bennear, and Atle Oglend. Spatial-dynamics of hypoxia and fisheries: The case of Gulf of Mexico brown shrimp. Marine Resource Economics, 29(2):111–131, June 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Saila:1988:BRD

[Sai88] Saul B. Saila. Book review: D. Pauly and I. Tsukayama, The Peruvian Anchoaveta and Its Upwelling Ecosystem: Three Decades of Change. Marine Resource Economics, 5(3):273–274, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Saila:1997:BRD

[Sai97] Saul B. Saila. Book review: D. Pauly and P. Martosubroto, Baseline Studies of Biodiversity: The Fish Resources of Western Indonesia. Marine Resource Economics, 12(3):251–252, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Salvanes:1989:SNF

[Sal89] Kjell Gunnar Salvanes. The structure of the Norwegian fish farming industry: an empirical analysis of economies of scale and substitution possibilities. Marine Resource Economics, 6 (4):349–373, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Salvanes:1993:PRP

[Sal93] Kjell G. Salvanes. Public regulation and production factor misallocation. A restricted cost function for the Norwegian aquaculture industry. Marine Resource Economics, 8(1):50–64, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sampson:1992:FTF

[Sam92] David B. Sampson. Fishing technology and fleet dynamics: Predictions from a bioeconomic model. Marine Resource Economics, 7(1):37–58, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sanchirico:2004:DCE

[San04] James N. Sanchirico. Designing a cost-effective marine reserve network: a bioeconomic metapopulation analysis. Marine Resource Economics, 19(1):41–65, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sandberg:2005:HRW

[San05] Per Sandberg. Harvest rules when price depends on quantity: The case of Norwegian spring spawning herring (Clupea harengus L.). Marine Resource Economics, 20(3):287–303, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Steine:2005:ECC

[SAR05] Gro Steine, Frode Alfnes, and Mia Bencze Rørå. The effect of color on consumer WTP for farmed salmon. Marine Resource Economics, 20(2):211–219, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sarkar:2009:OVH

[Sar09] Sudipto Sarkar. “Option Value of Harvesting” revisited. Marine Resource Economics, 24(1):77–81, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See reply [Li09].

Swallow:1994:BM

[SAS^{+}94] Stephen K. Swallow, C. W. Armstrong, Gilbert Sylvia, Roberto R. Enríquez, C. W. Armstrong, Stephen K. Swallow, Gilbert Sylvia, and Roberto R. Enríquez. Back matter. Marine Resource Economics, 9(4):??, Winter 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Scheld:2012:EEC

[SAU12] Andrew M. Scheld, Christopher M. Anderson, and Hirotsugu Uchida. The economic effects of catch share management: The Rhode Island fluke sector pilot program. Marine Resource Economics, 27(3):203–228, September 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Samples:1985:EVV

[SB85] Karl C. Samples and Richard C. Bishop. Estimating the value of variations in anglers’ success rates: an application of the multiple-site travel cost method. Marine Resource Economics, 2(1):55–74, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Shabman:1987:MDC

[SB87] Leonard A. Shabman and Sandra S. Batie. Mitigating damages from coastal wetlands development: Policy, economics and financing. Marine Resource Economics, 4(3):227–248, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Shabman:1987:BM

[SBK^{+}87] Leonard A. Shabman, Sandra S. Batie, James R. Kahn, Kathleen Segerson, James R. Kahn, James J. Opaluch, Leonard A. Shabman, and Sandra S. Batie. Back matter. Marine Resource Economics, 4(3):??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sylvia:1995:GPI

[SC95] Gilbert Sylvia and Deqin Cai. Generating policy information for fisheries management: a comparison of alternative approaches. Marine Resource Economics, 10(1):77–91, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schrank:1995:BRT

[Sch95] William Schrank. Book review: Tim D. Smith, Scaling Fisheries: The Science of Measuring the Effects of Fishing, 1855– 1955. Marine Resource Economics, 10(3):322–324, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schulz:1997:TSE

[Sch97] Carl-Erik Schulz. Trade sanctions and effects on long-run stocks of marine mammals. Marine Resource Economics, 12(3):159– 178, Fall 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schuhmann:1998:DSS

[Sch98] Peter W. Schuhmann. Deriving species-specific benefits measures for expected catch improvements in a random utility framework. Marine Resource Economics, 13(1):1–21, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schott:2001:OSH

[Sch01] Stephan Schott. The optimal selection and harvest of multiple cohorts. Marine Resource Economics, 16(2):109–126, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schrank:2007:BRR

[Sch07] William E. Schrank. Book review: Rögnvaldur Hannesson, Manuel Barange, and Samuel F. Herrick, Jr., Climate Change and the Economics of the World’s Fisheries: Examples of Small Pelagic Stocks. Marine Resource Economics, 22(2):215–220, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Seung:2015:UEI

[SCK^{+}15] Chang K. Seung, Sophal Chhun, Viktoria Kahui, Henrik Moller, Paul Thorsnes, Eric Sjöberg, Lee G. Anderson, Eric Sjöberg, Chang K. Seung, Johan Blomquist, Cecilia Hammarlund, and Staffan Waldo. Untangling economic impacts for Alaska fisheries: a structural path analysis. Marine Resource Economics, 30(3):331–347, July 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Scott:1988:DPF

[Sco88] Anthony Scott. Development of property in the fishery. Marine Resource Economics, 5(4):289–311, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Scott:1993:OFS

[Sco93] Anthony Scott. Obstacles to fishery self-government. Marine Resource Economics, 8(3):187–199, Fall 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Scott:2011:PFE

[Sco11] Anthony Scott. The pedigree of fishery economics. Marine Resource Economics, 26(1):75–85, March 2011. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sumaila:2006:TPF

[SCS06] Ussif Rashid Sumaila, Anthony Charles, and Gil Sylvia. Topical problems in fishery economics: an introduction. Marine Resource Economics, 21(4):337–340, ???? 2006. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Salvanes:1997:HDF

[SD97] Kjell G. Salvanes and Don J. Devoretz. Household demand for fish and meat products: Separability and demographic effects. Marine Resource Economics, 12(1):37–55, Spring 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Stoeven:2021:SFQ

[SDQ21] Max T. Stoeven, Florian K. Diekert, and Martin F. Quaas. Should fishing quotas be measured in terms of numbers? Marine Resource Economics, 36(2):133–153, April 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Singh:2012:ADS

[SDS12] Kehar Singh, Madan M. Dey, and Prasanna Surathkal. Analysis of a demand system for unbreaded frozen seafood in the United States using store-level scanner data. Marine Resource Economics, 27(4):371–387, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sylvia:1994:MBA

[SE94] Gilbert Sylvia and Roberto R. Enríquez. Multiobjective bioeconomic analysis: an application to the Pacific whiting fishery. Marine Resource Economics, 9(4):311–328, Winter 1994. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Segerson:1987:RSL

[Seg87] Kathleen Segerson. Risk-sharing and liability in the control of stochastic externalities. Marine Resource Economics, 4(3):175– 192, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Seijo:1993:ITG

[Sei93] Juan Carlos Seijo. Individual transferable grounds in a community managed artisanal fishery. Marine Resource Economics, 8 (1):78–81, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Seung:2024:DFM

[Seu24] Chang K. Seung. Does factor mobility matter? A general equilibrium analysis of a fishery rationalization. Marine Resource Economics, 39(2):123–144, ???? 2024. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Springborn:2019:MFS

[SF19] Michael R. Springborn and Amanda Faig. Moving forward: a simulation-based approach for solving dynamic resource management problems. Marine Resource Economics, 34(3):199–224, July 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2016:BRA

[SGC^{+}16] Martin D. Smith, Corbett A. Grainger, Christopher Costello, Hiroki Wakamatsu, Tsutomu Miyata, Jerrod Penn, Wuyang Hu, Linda Cox, Lara Kozloff, Corbett A. Grainger, Christopher Costello, Jerrod Penn, Wuyang Hu, Linda Cox, Lara Kozloff, Hiroki Wakamatsu, and Tsutomu Miyata. Book review: American Catch: The Fight for Our Local Seafood. By Paul Greenberg. Marine Resource Economics, 31(1):117–119, January 2016. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sanchirico:2004:NUF

[SH04] James N. Sanchirico and Susan S. Hanna. Navigating U.S. fishery management into the 21st century. Marine Resource Economics, 19(3):395–406, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Shamshak:2011:EEC

[Sha11] Gina Louise Shamshak. Economic evaluation of capture-based bluefin tuna aquaculture on the US East Coast. Marine Resource Economics, 26(4):309–328, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2013:BRD

[SHK^{+}13] Martin D. Smith, Daniel S. Holland, Andrew W. Kitts, Patricia Pinto Da Silva, Joshua Wiersma, Leif E. Anderson, S. Todd Lee, Douglas M. Larson, Daniel K. Lew, Leif E. Anderson, S. Todd Lee, Daniel S. Holland, Andrew W. Kitts, Patricia Pinto Da Silva, Joshua Wiersma, Douglas M. Larson, and Daniel K. Lew. Book review: Demon Fish: Travels through the Hidden World of Sharks. By Juliet Eilperin. Marine Resource Economics, 28 (2):201–202, June 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Simnitt:2020:UMC

[SHL^{+}20a] Skyler Simnitt, Lisa House, Sherry L. Larkin, Jennifer Sweeney Tookes, and Tracy Yandle. Using markets to control invasive species: Lionfish in the US Virgin Islands. Marine Resource Economics, 35(4):319–341, October 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Simnitt:2020:AA

[SHL^{+}20b] Skyler Simnitt, Lisa House, Sherry L. Larkin, Jennifer Sweeney Tookes, Tracy Yandle, Ingrid Kristine Pettersen, Frank Asche, Maxime Dépalle, Olivier Thébaud, James N. Sanchirico, Ingrid Kristine Pettersen, Frank Asche, Jeffrey K. O’Hara, Skyler Simnitt, Lisa House, Sherry L. Larkin, Jennifer Sweeney Tookes, and Tracy Yandle. Award announcements. Marine Resource Economics, 35(4):427–430, October 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sissenwine:1984:UEF

[Sis84] M. P. Sissenwine. The uncertain environment of fishery scientists and managers. Marine Resource Economics, 1(1):1–30, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sjoberg:2015:PFM

[Sjö15] Eric Sjöberg. Pricing on the fish market — does size matter? Marine Resource Economics, 30(3):277–296, July 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Salazar:2021:TPF

[SJQ21] César Salazar, Marcela Jaime, and Miguel Quiroga. Transition patterns of fishermen and land farmers into small-scale seaweed aquaculture: The role of risk and time preferences. Marine Resource Economics, 36(3):269–288, July 2021. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schrank:1999:CS

[SK99] William E. Schrank and Walter R. Keithly, Jr. The concept of subsidies. Marine Resource Economics, 14(2):151–164, Summer 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Skonhoft:2021:MNA

[SK21] Anders Skonhoft and Melina Kourantidou. Managing a natural asset that is both a value and a nuisance: Competition versus cooperation for the Barents Sea red king crab. Marine Resource Economics, 36(3):229–254, July 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sharma:1998:TEL

[SL98] Khem R. Sharma and Pingsun Leung. Technical efficiency of the longline fishery in Hawaii: an application of a stochastic production frontier. Marine Resource Economics, 13(4):259– 274, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Slaibi:2007:BRK

[Sla07] Ahmad Slaibi. Book review: Kenneth S. Deffeyes, Hubbert’s Peak: The Impending World Oil Shortage. Marine Resource Economics, 22(1):115–117, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Seijo:2014:NFS

[SLA14] Juan Carlos Seijo, Sherry L. Larkin, and Charles M. Adams. 2013 NAAFE forum: Special issue introduction. Marine Resource Economics, 29(2):85–87, June 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sharma:1998:BM

[SLM^{+}98] Khem R. Sharma, Pingsun Leung, Johan A. Mistiaen, Ivar Strand, Chin hwa Sun, Khem R. Sharma, Pingsun Leung, Johan A. Mistiaen, Ivar Strand, and Chin hwa Sun. Back matter. Marine Resource Economics, 13(4):??, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Saila:1985:APE

[SLW85] S. B. Saila, E. Lorda, and H. A. Walker. The analysis of parameter error propagation in simple fishery models. Marine Resource Economics, 1(3):235–246, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:1985:DMR

[Smi85] J. Barry Smith. A discrete model of replenishable resource management under uncertainty. Marine Resource Economics, 1(3):283–308, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:1986:SSS

[Smi86] J. Barry Smith. Stochastic steady-state replenishable resource management policies. Marine Resource Economics, 3(2):155– 168, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:1989:TSP

[Smi89] V. Kerry Smith. Taking stock of progress with travel cost recreation demand methods: Theory and implementation. Marine Resource Economics, 6(4):279–310, ???? 1989. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smit:1996:EAM

[Smi96] Wil Smit. An economic approach to measuring fishing effort: Application to a Dutch cutter fleet. Marine Resource Economics, 11(4):305–311, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smit:1997:CFP

[Smi97] Wil Smit. Common fishery policy and national fisheries management. Marine Resource Economics, 12(4):355–359, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2012:RMR

[Smi12] Martin D. Smith. Reflections on Marine Resource Economics: Editor’s introduction. Marine Resource Economics, 27(3):197– 201, September 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2017:EBO

[Smi17] Martin D. Smith. Editorial: The breadth of ocean and coastal economics. Marine Resource Economics, 32(2):119–121, April 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2023:EAF

[Smi23] Martin D. Smith. Economics of aquatic foods: Combining bioeconomics and market analysis to inform regulations that deliver value. Marine Resource Economics, 38(4):305–327, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Squires:1985:BRM

[SMRK85] Dale Squires, Joseph Mueller, Robert Reidman, and James Kirkley. Book review: Margaret E. Dewar, Industry in Trouble: The Federal Government and the New England Fisheries. Marine Resource Economics, 1(3):309–312, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:1985:BRD

[SP85] Ian R. Smith and Daniel Pauly. Book review: D. Stevenson, R. Pollnac, and P. Logan, A Guide for the Small-Scale Fishery Administrator: Information from the Harvest Sector. Marine Resource Economics, 1(3):313–319, ???? 1985. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schrobback:2018:QEI

[SPC18] Peggy Schrobback, Sean Pascoe, and Louisa Coglan. Quantifying the economic impact of climate change and market dynamics: The case of Australia’s Sydney rock oyster industry. Marine Resource Economics, 33(2):155–175, April 2018. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Spulber:1985:MFU

[Spu85] Daniel F. Spulber. The multicohort fishery under uncertainty. Marine Resource Economics, 1(3):265–282, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schrobback:2019:MID

[SPZ19] Peggy Schrobback, Sean Pascoe, and Rui Zhang. Market integration and demand for prawns in Australia. Marine Resource Economics, 34(4):311–329, October 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sutinen:2023:MCA

[SRA^{+}23] Jon G. Sutinen, Kyrre Rickertsen, Frode Alfnes, Pierre Combris, Géraldine Enderli, Sylvie Issanchou, Jason F. Shogren, Geret S. DePiper, Douglas W. Lipton, Romuald N. Lipcius, Isaac Ankamah-Yeboah, Jette Bredahl Jacobsen, Søren Bøye Olsen, Max Nielsen, Rasmus Nielsen, Anna M. Birkenbach, David J. Kaczan, Martin D. Smith, Greg Ardini, Daniel S. Holland, Min-Yang Lee, Doug Lipton, and Michael D. Travis. Musings on a career in applied fisheries economics research. Marine Resource Economics, 38(3):285–303, ???? 2023. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schrank:2013:NFE

[SRS^{+}13] William E. Schrank, Noel Roy, William E. Schrank, Noel Roy, Nils-Arne Ekerhovd, Daniel V. Gordon, Ragnar Arnason, William E. Schrank, Noel Roy, Nils-Arne Ekerhovd, Daniel V. Gordon, and Daniel S. Holland. The Newfoundland fishery and economy twenty years after the Northern Cod Moratorium. Marine Resource Economics, 28(4):397–413, December 2013. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Schrank:1986:EPC

[SRTS86] William E. Schrank, Noel Roy, Eugene Tsoa, and Wiliam E. Shrank. Employment prospects in a commercially viable Newfoundland fishery: an application of ‘an econometric model of the Newfoundland groundfishery’. Marine Resource Economics, 3(3):237–263, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sandal:1997:OSS

[SS97a] Leif K. Sandal and Stein I. Steinshamn. Optimal steady states and the effects of discounting. Marine Resource Economics, 12 (2):95–105, Summer 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sandal:1997:GE

[SS97b] Leif K. Sandal and Stein Ivar Steinshamn. Guest editorial. Marine Resource Economics, 12(2):??, Summer 1997. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Skonhoft:2001:ELC

[SS01] Anders Skonhoft and Jan Tore Solstad. Exploiting a local common: Egoistic vs. altruistic behavior. Marine Resource Economics, 16(1):39–49, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Stefanski:2016:VMB

[SS16] Stephanie F. Stefanski and Jay P. Shimshack. Valuing marine biodiversity in the Gulf of Mexico: Evidence from the proposed boundary expansion of the Flower Garden Banks National Marine Sanctuary. Marine Resource Economics, 31(2):211–232, April 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Salvanes:1989:BM

[SSH^{+}89] Kjell Gunnar Salvanes, V. Kerry Smith, Knut Heen, Kjell Gunnar Salvanes, V. Kerry Smith, and Knut Heen. Back matter. Marine Resource Economics, 6(4):??, ???? 1989. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sandal:2007:IIR

[SSH07] Leif K. Sandal, Stein I. Steinshamn, and Ayoe Hoff. Irreversible investments revisited. Marine Resource Economics, 22(3):255– 266, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sathiendrakumar:1987:OEF

[ST87] R. Sathiendrakumar and C. A. Tisdell. Optimal economic fishery effort in the Maldivian tuna fishery: an appropriate model. Marine Resource Economics, 4(1):15–44, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sathiendrakumar:1989:CNO

[ST89] R. Sathiendrakumar and C. A. Tisdell. Comments on a note on optimal effort in the Maldivian tuna industry. Marine Resource Economics, 6(2):177–178, ???? 1989. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [CL89].

Staniford:1987:EPR

[Sta87] Andrew Staniford. The effects of the pot reduction in the South Australian Southern Zone rock lobster fishery. Marine Resource Economics, 4(4):271–288, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Stage:2015:VSE

[Sta15] Jesper Stage. The value of the Swedish eel fishery. Marine Resource Economics, 30(1):21–34, January 2015. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Steinshamn:1998:IHS

[Ste98] Stein Ivar Steinshamn. Implications of harvesting strategies on population and profitability in fisheries. Marine Resource Economics, 13(1):23–36, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sakai:2023:WVF

[STNY23] Yutaro Sakai, Tomoki Tada, Tsubasa Nomura, and Nobuyuki Yagi. The welfare value of freshness: a hedonic price analysis in the retail seafood market in Japan. Marine Resource Economics, 39(1):21–38, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Stollery:1986:SRM

[Sto86a] Kenneth Stollery. A short-run model of capital stuffing in the Pacific halibut fishery. Marine Resource Economics, 3(2):137– 153, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Stollery:1986:MPO

[Sto86b] Kenneth R. Stollery. Monopsony processing in an open-access fishery. Marine Resource Economics, 3(4):331–351, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Stollery:1987:CAR

[Sto87] Kenneth R. Stollery. Cooperatives as an alternative to regulation in commercial fisheries. Marine Resource Economics, 4(4):289– 304, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Strand:2004:SVR

[Str04] Ivar E. Strand, Jr. Spatial variation in risk preferences among Atlantic and Gulf of Mexico pelagic longline fishermen. Marine Resource Economics, 19(1):145–160, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Straume:2014:CIN

[Str14] Hans-Martin Straume. Currency invoicing in Norwegian salmon export. Marine Resource Economics, 29(4):391–409, December 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Shamshak:2020:EEV

[STY20] Gina Louise Shamshak, Jennifer Sweeney Tookes, and Tracy Yandle. Evaluating the economic viability of inland seafood markets in Georgia: a two-pronged approach. Marine Resource Economics, 35(1):21–42, January 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sumaila:1995:ICI

[Sum95] Ussif Rashid Sumaila. Irreversible capital investment in a twostage bimatrix fishery game model. Marine Resource Economics, 10(3):263–283, Fall 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sumaila:1997:SDI

[Sum97] Ussif Rashid Sumaila. Strategic dynamic interaction: The case of Barents Sea fisheries. Marine Resource Economics, 12(2):77– 93, Summer 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sutinen:1984:E

[Sut84] Jon G. Sutinen. Editorial. Marine Resource Economics, 1(1): ??, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sutinen:1992:E

[Sut92] Jon G. Sutinen. Editorial. Marine Resource Economics, 7(1): ??, Spring 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sutinen:1995:E

[Sut95] Jon G. Sutinen. Editorial. Marine Resource Economics, 10(4): ??, Winter 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sutinen:2011:JMG

[Sut11] Jon G. Sutinen. John M. Gates and Marine Resource Economics. Marine Resource Economics, 26(2):167–168, June 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Squires:2016:PEM

[SV16] Dale Squires and Niels Vestergaard. Putting economics into maximum economic yield. Marine Resource Economics, 31(1): 101–116, January 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Salayo:1999:IPP

[SVS99] Nerissa D. Salayo, Thomas J. P. Voon, and Saroja Selvanathan. Implicit prices of prawn and shrimp attributes in the Philippine domestic market. Marine Resource Economics, 14(1):65– 78, Spring 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2004:MRE

[SW04] Martin D. Smith and James E. Wilen. Marine reserves with endogenous ports: Empirical bioeconomics of the California sea urchin fishery. Marine Resource Economics, 19(1):85–112, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Seung:2006:RRE

[SW06] Chang K. Seung and Edward C. Waters. A review of regional economic models for fisheries management in the U.S. Marine Resource Economics, 21(1):101–124, ???? 2006. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Singh:2015:CTB

[SW15] Rajesh Singh and Quinn Weninger. Cap-and-trade bycatch management with costly avoidance and stock uncertainty. Marine Resource Economics, 30(1):97–119, January 2015. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Scheld:2018:AFS

[SW18] Andrew M. Scheld and John Walden. An analysis of fishing selectivity for Northeast US multispecies bottom trawlers. Marine Resource Economics, 33(4):331–350, October 2018. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Swallow:1994:RNR

[Swa94] Stephen K. Swallow. Renewable and nonrenewable resource theory applied to coastal agriculture, forest, wetland, and fishery linkages. Marine Resource Economics, 9(4):291–310, Winter 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Swierzbinski:1985:SMA

[Swi85] Joseph Swierzbinski. Statistical methods applicable to selected problems in fisheries biology and economics. Marine Resource Economics, 1(3):209–233, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Sin:2019:EDA

[SYN^{+}19] Moe Shwe Sin, Tai Shzee Yew, Kusairi Mohd Noh, Frank Asche, Bård Misund, Atle Oglend, Katrina J. Davis, Michael Burton, Abbie Rogers, Alaya Spencer-Cotton, Ram Pandit, Ping Wang, Nhuong Tran, Norbert L. W. Wilson, Chin Yee Chan, Danh Dao, Frank Asche, Bård Misund, Atle Oglend, Ping Wang, Nhuong Tran, Norbert L. W. Wilson, Chin Yee Chan, Danh Dao, Katrina J. Davis, Michael Burton, Abbie Rogers, Alaya Spencer-Cotton, and Ram Pandit. Effort dynamics and alternative management policies: The case of the West Coast Zone B trawl fishery in Peninsular Malaysia. Marine Resource Economics, 34(1):77–103, January 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Smith:2007:SMM

[SZC07] Martin D. Smith, Junjie Zhang, and Felicia C. Coleman. Structural modeling of marine reserves with Bayesian estimation. Marine Resource Economics, 22(2):121–136, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tahvonen:2009:OHA

[Tah09] Olli Tahvonen. Optimal harvesting of age-structured fish populations. Marine Resource Economics, 24(2):147–169, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tveteraas:2002:BM

[TAW^{+}02] Sigbjørn Tveterås, James L. Anderson, Cathy Roheim Wessells, James L. Anderson, Sigbjørn Tveterås, and Cathy Roheim Wessells. Back matter. Marine Resource Economics, 17(2):??, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Teisl:1997:NHC

[TB97] Mario F. Teisl and Kevin J. Boyle. Needles in a haystack: Costeffective sampling of marine sport anglers. Marine Resource Economics, 12(1):1–10, Spring 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Thorpe:2001:GSW

[TB01] Andy Thorpe and Elizabeth Bennett. Globalisation and the sustainability of world fisheries: a view from Latin America. Marine Resource Economics, 16(2):143–164, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Thunberg:1995:EAT

[TBA95] Eric M. Thunberg, Edward W. Bresnyan, and Charles M. Adams. Economic analysis of technical interdependencies and the value of effort in a multi-species fishery. Marine Resource Economics, 10(1):59–76, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tisdell:1986:BM

[TBC^{+}86] Clem Tisdell, J. M. Broadus, Allen L. Clark, Jennifer Cook Clark, J. M. Broadus, Allen L. Clark, and Jennifer Cook Clark. Back matter. Marine Resource Economics, 3(1):??, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Thorpe:2001:BM

[TBH^{+}01] Andy Thorpe, Elizabeth Bennett, Jerry Heykoop, Darren L. Frechette, Claire W. Armstrong, Siv Reithe, Jerry Heykoop, Darren L. Frechette, Henry W. Kinnucan, Laxmi Paudel, Andy Thorpe, and Elizabeth Bennett. Back matter. Marine Resource Economics, 16(2):??, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cheng:1993:PIS

[tCT93] Hsiang tai Cheng and Ralph E. Townsend. Potential impact of seasonal closures in the U.S. lobster fishery. Marine Resource Economics, 8(2):101–117, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Cheng:1993:BM

[tCTH^{+}93] Hsiang tai Cheng, Ralph E. Townsend, Terry Heaps, John C. Whitehead, John C. Whitehead, Terry Heaps, H. F. Campbell, A. J. Hand, and A. D. M. Smith. Back matter. Marine Resource Economics, 8(2):??, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Terry:2007:AUC

[Ter07] Joseph M. Terry. An assessment of the use of capacity analysis in U.S. federal fishery management. Marine Resource Economics, 22(1):99–103, ???? 2007. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Toufique:2018:MFF

[TFB18] Kazi Ali Toufique, Sami Naim Farook, and Ben Belton. Managing fisheries for food security: Implications from demand analysis. Marine Resource Economics, 33(1):61–85, January 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tyrrell:1999:MTR

[TgKtC99] Timothy Tyrrell, Sung gwi Kim, and Young tae Chang. Marine tourism resource development in Korea. Marine Resource Economics, 14(2):165–174, Summer 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Theophille:2012:EAT

[TH12] Brent Theophille and Sharon D. Hutchinson. Economic analysis of the targeting strategy of artisanal fishers in Dominica. Marine Resource Economics, 27(4):323–342, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Thunberg:1998:BAA

[THM98] E. M. Thunberg, T. E. Helser, and R. K. Mayo. Bioeconomic analysis of alternative selection patterns in the United States Atlantic silver hake fishery. Marine Resource Economics, 13(1): 51–74, Spring 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Thong:2012:IAI

[Tho12] Nguyen Tien Thong. An inverse almost ideal demand system for mussels in Europe. Marine Resource Economics, 27(2):149– 164, June 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tisdell:1986:CAL

[Tis86] Clem Tisdell. Conflicts about living marine resources in Southeast Asian and Australian waters: Turtles and dugong as cases. Marine Resource Economics, 3(1):89–109, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tabarestani:2017:AUS

[TKMA17] Maryam Tabarestani, Walter R. Keithly, Jr., and Hassan Marzoughi-Ardakani. An analysis of the US shrimp market: a mixed demand approach. Marine Resource Economics, 32(4): 411–429, October 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Thiao:2012:SIS

[TL12] Djiga Thiao and Francis Laloë. A system of indicators for sustainability: an example from the Senegalese fisheries. Marine Resource Economics, 27(3):267–282, September 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tokunaga:2018:EED

[Tok18] Kanae Tokunaga. Estimating elasticity of demand for Pacific bluefin tuna in tsukiji fish market. Marine Resource Economics, 33(1):27–60, January 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tomberlin:2010:ESH

[Tom10] David Tomberlin. Endangered seabird habitat management as a partially observable Markov decision process. Marine Resource Economics, 25(1):93–104, March 2010. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Townsend:2003:BHU

[Tow03] Ralph Townsend. Book review: Bjørn Hersoug, Unfinished Business: New Zealand’s Experience with Rights-based Fisheries Management, Delft, Netherlands: Eburon, 2002, 256 pp. ISBN 90-5166-882-1. Marine Resource Economics, 18(4):373–375, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Townsend:2006:BRL

[Tow06] Ralph Townsend. Book review: Laura Jones and Miriam Bixby, Managing Fish: Ten Case Studies from Canada’s Pacific Coast. Marine Resource Economics, 21(3):331–332, ???? 2006. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Townsend:2009:BRG

[Tow09] Ralph Townsend. Book review: G. Bruce Knecht, Hooked: Pirates, Poaching, and the Perfect Fish. Marine Resource Economics, 24(2):205–206, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Taylor:1985:FPF

[TP85] Timothy G. Taylor and Fred J. Prochaska. Fishing power functions in aggregate bioeconomic models. Marine Resource Economics, 2(1):87–107, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tingley:2005:EEC

[TP05] Diana Tingley and Sean Pascoe. Eliminating excess capacity: Implications for the Scottish fishing industry. Marine Resource Economics, 20(4):407–424, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Townsend:2003:EPB

[TPC03] Ralph E. Townsend, Samuel G. Pooley, and Raymond Clarke. Evidence on producer bargaining in the Northwestern Hawaiian Islands lobster fishery. Marine Resource Economics, 18(2):195– 203, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tveteras:2011:IVQ

[TPPT11] Sigbjorn Tveteras, Carlos E. Paredes, and Julio Peña-Torres. Individual vessel quotas in Peru: Stopping the race for anchovies. Marine Resource Economics, 26(3):225–232, September 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Trondsen:2001:FMM

[Tro01] Torbjørn Trondsen. Fisheries management and market-oriented value adding (MOVA). Marine Resource Economics, 16(1):17– 37, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Thomas:2002:CVR

[TS02] Michael Thomas and Nicholas Stratis. Compensating variation for recreational policy: a random utility approach to boating in Florida. Marine Resource Economics, 17(1):23–33, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tveteras:2002:BM

[TTS^{+}02] Ragnar Tveteras, Michael Thomas, Nicholas Stratis, Gerald R. Visgilio, Ragnar Tveteras, Michael Thomas, Nicholas Stratis, Pedro Pintassilgo, and Clara Costa Duarte. Back matter. Marine Resource Economics, 17(1):??, ???? 2002. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Turner:1996:VBI

[Tur96] Matthew A. Turner. Value-based ITQ’s. Marine Resource Economics, 11(2):59–69, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tveteras:2002:IAP

[Tve02a] Ragnar Tveteras. Industrial agglomeration and production costs in Norwegian salmon aquaculture. Marine Resource Economics, 17(1):1–22, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tveteraas:2002:NSA

[Tve02b] Sigbjørn Tveterås. Norwegian salmon aquaculture and sustainability: The relationship between environmental quality and industry growth. Marine Resource Economics, 17(2):121–132, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Townsend:2005:ELA

[TY05] Ralph E. Townsend and Michael D. Young. Evergreen leasing of aquaculture sites. Marine Resource Economics, 20(2):203– 210, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tan:2015:HCF

[TYH15] Andrew K. G. Tan, Steven T. Yen, and Abdul Rahman Hasan. At-home consumption of fish products in Malaysia: an analysis of household expenditure survey data. Marine Resource Economics, 30(4):417–433, October 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Tyrrell:2001:BRC

[Tyr01] Timothy Tyrrell. Book review: Clem Tisdell, The Economics of Tourism (Volumes I and II). Marine Resource Economics, 16 (3):249–250, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Uchida:2010:DSM

[UUL^{+}10] Hirotsugu Uchida, Emi Uchida, Jung-Sam Lee, Jeong-Gon Ryu, and Dae-Young Kim. Does self management in fisheries enhance profitability? Examination of Korea’s coastal fisheries. Marine Resource Economics, 25(1):37–59, March 2010. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

VanDyke:1996:BRD

[Van96] Jon M. Van Dyke. Book review: Dennis W. Nixon, Marine and Coastal Law: Cases and Materials. Marine Resource Economics, 11(2):??, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

vanBinh:2010:ICM

[vBDSD10] Tu van Binh, Marijke D’Haese, Stijn Speelman, and Luc D’Haese. The influence of changes in the market environment on economic production characteristics of Pangasius farming in the Mekong Delta (Vietnam). Marine Resource Economics, 25(4): 373–390, December 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

VanHoof:2005:CAD

[VD05] Luc Van Hoof and Jan Willem De Wilde. Capacity assessment of the Dutch beam-trawler fleet using Data Envelopment Analysis (DEA). Marine Resource Economics, 20(4):327–345, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Vestergaard:1996:DBH

[Ves96] Niels Vestergaard. Discard behavior, highgrading and regulation: The case of the Greenland shrimp fishery. Marine Resource Economics, 11(4):247–266, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Vestergaard:2005:FCE

[Ves05] Niels Vestergaard. Fishing capacity in Europe: Special issue introduction. Marine Resource Economics, 20(4):323–326, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Vestergaard:1996:BM

[VGH^{+}96] Niels Vestergaard, Daniel V. Gordon, Rögnvaldur Hannesson, Francis T. Christy, Niels Vestergaard, Daniel V. Gordon, Rögnvaldur Hannesson, Gianfranco Rizzo, and Massimo Spagnolo. Back matter. Marine Resource Economics, 11(4):??, Winter 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Vassdal:2011:TPR

[VH11] Terje Vassdal and Helen Marita Sørensen Holst. Technical progress and regress in Norwegian salmon farming: a Malmquist Index approach. Marine Resource Economics, 26(4):329–341, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Visgilio:2001:BRC

[Vis01] Gerald R. Visgilio. Book review: Christopher H. Foreman, Jr., The Promise and Peril of Environmental Justice. Marine Resource Economics, 16(1):79–81, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Visgilio:2002:BRD

[Vis02] Gerald R. Visgilio. Book review: David D. Friedman, Law’s Order — What Economics Has To Do With Law and Why It Matters. Marine Resource Economics, 17(1):69–71, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Valencia:1986:SAM

[VM86] Mark J. Valencia and James Barney Marsh. Southeast Asia: Marine resources, extended maritime jurisdiction, and development. Marine Resource Economics, 3(1):3–27, ???? 1986. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Viswanathan:2001:FSD

[VOJ^{+}01] K. Kuperan Viswanathan, Ishak Haji Omar, Yongil Jeon, James Kirkley, Dale Squires, and Indah Susilowati. Fishing skill in developing country fisheries: The Kedah, Malaysia trawl fishery. Marine Resource Economics, 16(4):293–314, ???? 2001. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Virtanen:2005:FST

[VSSH05] Jarno Virtanen, Jari Setälä, Kaija Saarni, and Asmo Honkanen. Finnish salmon trout — discriminated in the European market. Marine Resource Economics, 20(1):113–119, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Valcu:2013:MPR

[VW13] Adriana Valcu and Quinn Weninger. Markov-perfect rent dissipation in rights-based fisheries. Marine Resource Economics, 28(2):111–131, June 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wessells:1992:IPS

[WA92] Cathy Roheim Wessells and James L. Anderson. Innovations and progress in seafood demand and market analysis. Marine Resource Economics, 7(4):209–228, Winter 1992. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Walden:2023:MFC

[WA23] John B. Walden and Joseph Atwood. Measuring fishing capacity using quantile data envelopment analysis. Marine Resource Economics, 38(3):229–247, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wakamatsu:2014:IMC

[Wak14] Hiroki Wakamatsu. The impact of MSC certification on a Japanese certified fishery. Marine Resource Economics, 29(1): 55–67, March 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Walden:2006:EVE

[Wal06] John B. Walden. Estimating vessel efficiency using a bootstrapped data envelopment analysis model. Marine Resource Economics, 21(2):181–192, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wang:1992:NRE

[Wan92] Stanley D. H. Wang. Natural resource and environmental economics database. Marine Resource Economics, 7(4):297–304, Winter 1992. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wang:1995:SCI

[Wan95] Stanley Wang. The surf clam ITQ management: an evaluation. Marine Resource Economics, 10(1):93–98, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ward:1994:BIB

[War94] John M. Ward. The bioeconomic implications of a bycatch reduction device as a stock conservation management measure. Marine Resource Economics, 9(3):227–240, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Waters:2001:QMC

[Wat01] James R. Waters. Quota management in the commercial red snapper fishery. Marine Resource Economics, 16(1):65–78, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Waugh:1987:BRL

[Wau87] Geoffrey Waugh. Book review: Lee G. Anderson, The Economics of Fisheries Management. Marine Resource Economics, 4(4):319–321, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Waugh:1988:BRD

[Wau88] Geoffrey Waugh. Book review: David J. Doulman, Tuna Issues and Perspective in the Pacific Islands Region. Marine Resource Economics, 5(2):161–163, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Waugh:1994:RHB

[Wau94] Geoffrey Waugh. Rognvaldur hannesson, bioeconomic analysis of fisheries. Marine Resource Economics, 9(3):287–290, Fall 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wakamatsu:2017:PES

[WAUR17] Hiroki Wakamatsu, Christopher M. Anderson, Hirotsugu Uchida, and Cathy A. Roheim. Pricing ecolabeled seafood products with heterogeneous preferences: an auction experiment in Japan. Marine Resource Economics, 32(3):277–294, July 2017. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wallace:1986:OFS

[WB86] Stein W. Wallace and Karl Brekke. Optimal fleet size when national quotas can be traded. Marine Resource Economics, 2 (4):315–329, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Watson:2009:SCL

[WB09] Philip Watson and Nick Beleiciks. Small community level social accounting matrices and their application to determining marine resource dependency. Marine Resource Economics, 24(3):253– 270, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Whitehead:2013:BRC

[WBC^{+}13] John C. Whitehead, Okmyung Bin, Jeffrey Czajkowski, Jannike Falk-Petersen, Claire W. Armstrong, Göran Bostedt, Thi Thanh Thuy Pham, Ola Flaaten, Thi Kim Anh Nguyen, Okmyung Bin, Jeffrey Czajkowski, and Rögnvaldur Hannesson. Book review: Contingent Valuation: a Comprehensive Bibliography and History By Richard T. Carson. Marine Resource Economics, 28(1):107–109, January 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:2010:RSC

[WCA^{+}10] James E. Wilen, Jingjie Chu, James L. Anderson, Frank Asche, Lacey Tudur, Andrew J. Ropicki, Sherry L. Larkin, Charles M. Adams, Peder Andersen, Jesper Levring Andersen, Hans Frost, Jingjie Chu, James L. Anderson, Frank Asche, Lacey Tudur, Peder Andersen, Jesper Levring Andersen, Hans Frost, Andrew J. Ropicki, Sherry L. Larkin, and Charles M. Adams. A reply to “Stranded Capital in Fisheries: The Pacific Coast Groundfish/Whiting Case” (the comment). Marine Resource Economics, 25(1):129–132, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [Mat10].

Wielgus:2003:ECR

[WCFZS03] Jeffrey Wielgus, Nanette E. Chadwick-Furman, Naomi Zeitouni, and Mordechai Shechter. Effects of coral reef attribute damage on recreational welfare. Marine Resource Economics, 18(3):225– 237, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Whitehead:2001:WPS

[WCH01] John C. Whitehead, William B. Clifford, and Thomas J. Hoban. Willingness to pay for a saltwater recreational fishing license: a comparison of angler groups. Marine Resource Economics, 16 (3):177–194, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Werner:2020:ECF

[WDJK20] Samantha Werner, Geret DePiper, Di Jin, and Andrew Kitts. Estimation of commercial fishing trip costs using sea sampling data. Marine Resource Economics, 35(4):379–410, October 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Whitehead:2011:VBL

[WDLH11] John C. Whitehead, Christopher F. Dumas, Craig E. Landry, and Jim Herstine. Valuing bag limits in the North Carolina charter boat fishery with combined revealed and stated preference data. Marine Resource Economics, 26(3):233–241, September 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:1985:BM

[WDS^{+}85] James E. Wilen, Don DeVoretz, Richard Schwindt, Marc Mangel, Richard E. Plant, James E. Wilen, Marc Mangel, and Richard E. Plant. Back matter. Marine Resource Economics, 1 (4):??, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wessells:1998:E

[Wes98] Cathy R. Wessells. Editorial. Marine Resource Economics, 13(4):??, Winter 1998. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wessells:2002:EIM

[Wes02] Cathy Roheim Wessells. The economics of information: Markets for seafood attributes. Marine Resource Economics, 17(2):153– 162, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Woodward:2003:SBL

[WG03] Richard T. Woodward and Wade L. Griffin. Size and bag limits in recreational fisheries: Theoretical and empirical analysis. Marine Resource Economics, 18(3):239–262, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wang:1986:EAS

[WGM86] Der Hsiung Wang, Louis J. Goodreau, and Joseph J. Mueller. Economics of Atlantic sea scallop management. Marine Resource Economics, 3(2):111–135, ???? 1986. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Woodward:2003:BM

[WGW^{+}03] Richard T. Woodward, Wade L. Griffin, Jeffrey Wielgus, Nanette E. Chadwick-Furman, Naomi Zeitouni, Mordechai Shechter, Henry W. Kinnucan, Richard T. Woodward, Wade L. Griffin, Jeffrey Wielgus, Nanette E. Chadwick-Furman, Naomi Zeitouni, Mordechai Shechter, and Henry W. Kinnucan. Back matter. Marine Resource Economics, 18(3):??, ???? 2003. CO-DEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Whitehead:1999:SMR

[WH99] John C. Whitehead and Timothy C. Haab. Southeast marine recreational fishery statistical survey: Distance and catch based choice sets. Marine Resource Economics, 14(4):283–298, Winter 1999. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Whitehead:1993:TEV

[Whi93] John C. Whitehead. Total economic values for coastal and marine wildlife: Specification, validity, and valuation issues. Marine Resource Economics, 8(2):119–132, Summer 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Whitehead:2018:ELR

[WHL^{+}18] John C. Whitehead, Tim Haab, Sherry L. Larkin, John B. Loomis, Sergio Alvarez, and Andrew Ropicki. Estimating lost recreational use values of visitors to Northwest Florida due to the Deepwater Horizon oil spill using cancelled trip data. Marine Resource Economics, 33(2):119–132, April 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:1985:TTR

[Wil85] James E. Wilen. Towards a theory of the regulated fishery. Marine Resource Economics, 1(4):369–388, ???? 1985. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:1987:CWA

[Wil87] James E. Wilen. Carl walters, adaptive management of renewable resources. Marine Resource Economics, 4(2):133–137, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:1988:LEL

[Wil88] James E. Wilen. Limited entry licensing: a retrospective assessment. Marine Resource Economics, 5(4):313–324, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:2004:SMF

[Wil04] James E. Wilen. Spatial management of fisheries. Marine Resource Economics, 19(1):7–19, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:2007:TCA

[Wil07] James E. Wilen. Thoughts on capacity analysis: Is capacity analysis giving policy makers information they need? Marine Resource Economics, 22(1):79–82, ???? 2007. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:2009:SCF

[Wil09] James E. Wilen. Stranded capital in fisheries: The Pacific Coast groundfish/whiting case. Marine Resource Economics, 24(1):1– 18, ???? 2009. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:2013:CPP

[Wil13] James E. Wilen. The challenges of Pro-poor fisheries reform. Marine Resource Economics, 28(3):203–220, September 2013. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:2018:AIC

[Wil18a] James E. Wilen. Author’s introductory comment on “Common Property Resources and the Dynamics of Overexploitation: The Case of the North Pacific Fur Seal”. Marine Resource Economics, 33(3):213–215, July 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic). See [ASS18].

Wilen:2018:CPR

[Wil18b] James E. Wilen. Common property resources and the dynamics of overexploitation: The case of the North Pacific fur seal. Marine Resource Economics, 33(3):217–243, July 2018. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Watson:2012:FFP

[WJ12] Philip Watson and Timothy Johnson. Federal fishery policy and the geographic distribution of commercial U.S. West Coast fish landings: Insights from the 2003 Federal groundfish permit buyback. Marine Resource Economics, 27(4):289–301, December 2012. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Waldo:2016:RME

[WJN^{+}16] Staffan Waldo, Frank Jensen, Max Nielsen, Hans Ellefsen, Jónas Hallgrimsson, Cecilia Hammarlund, Øystein Hermansen, John Isaksen, Scott Crosson, Stephanie F. Stefanski, Jay P. Shimshack, Staffan Waldo, Frank Jensen, Max Nielsen, Hans Ellefsen, Jónas Hallgrimsson, Cecilia Hammarlund, Øystein Hermansen, John Isaksen, Rong Zhou, Kathleen Segerson, Paul Hallwood, Staffan Waldo, Frank Jensen, Max Nielsen, Hans Ellefsen, Jónas Hallgrimsson, Cecilia Hammarlund, Øystein Hermansen, and John Isaksen. Regulating multiple externalities: The case of Nordic fisheries. Marine Resource Economics, 31(2):233–257, April 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Whitmarsh:2002:EEF

[WJP^{+}02] David Whitmarsh, Carl James, Helen Pickering, Carlo Pipitone, Fabio Badalamenti, and Giovanni D’anna. Economic effects of fisheries exclusion zones: a Sicilian case study. Marine Resource Economics, 17(3):239–250, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wang:1988:EMA

[WK88] Stanley D. H. Wang and Christopher B. Kellogg. An econometric model for American lobster. Marine Resource Economics, 5(1):61–70, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:1987:BM

[WKGK87] James E. Wilen, Catherine L. Kling, Daniel L. Georgianna, and Catherine L. Kling. Back matter. Marine Resource Economics, 4(2):??, ???? 1987. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilson:1994:EPE

[WL94] James R. Wilson and Rebecca Lent. Economic perspective and the evolution of fisheries management: Towards subjectivist methodology. Marine Resource Economics, 9(4):353–373, Winter 1994. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wakamatsu:2016:DRS

[WM16] Hiroki Wakamatsu and Tsutomu Miyata. Do radioactive spills from the Fukushima disaster have any influence on the Japanese seafood market? Marine Resource Economics, 31(1):27–45, January 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wessells:1995:TAC

[WMB95] Cathy Roheim Wessells, Christopher J. Miller, and Priscilla M. Brooks. Toxic algae contamination and demand for shellfish: a case study of demand for mussels in Montreal. Marine Resource Economics, 10(2):143–159, Summer 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Waters:2001:BM

[WMS^{+}01] James R. Waters, Scott C. Matulich, Murat Sever, Fred Inaba, Torbjørn Trondsen, James R. Waters, J. Antonio Martinez, J. Carlos Seijo, and Torbjørn Trondsen. Back matter. Marine Resource Economics, 16(1):??, ???? 2001. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ward:2004:RFH

[WMT04] John Ward, Pamela Mace, and Eric Thunberg. The relationship of fish harvesting capacity to excess capacity and overcapacity. Marine Resource Economics, 19(4):525–529, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wang:1997:UNE

[WR97] Stanley D. H. Wang and Andrew A. Rosenberg. U.S. New England groundfish management under the Magnuson–Stevens Fishery Conservation and Management Act. Marine Resource Economics, 12(4):361–366, Winter 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Waters:2010:IRS

[WS10] Edward C. Waters and Chang K. Seung. Impacts of recent shocks to Alaska fisheries: a computable general equilibrium (CGE) model analysis. Marine Resource Economics, 25(2):155– 183, June 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wilen:2004:BM

[WSW^{+}04] James E. Wilen, Martin D. Smith, James E. Wilen, Daniel S. Holland, James N. Sanchirico, Rita E. Curtis, Robert L. Hicks, James E. Wilen, Rita E. Curtis, Kenneth E. McConnell, and Daniel S. Holland. Back matter. Marine Resource Economics, 19 (1):??, ???? 2004. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wu:2005:RNR

[WT05] Junjie Wu and Katsuya Tanaka. Reducing nitrogen runoff from the Upper Mississippi River Basin to control hypoxia in the Gulf of Mexico: Easements or taxes? Marine Resource Economics, 20(2):121–144, ???? 2005. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Walden:2010:EFV

[WT10] John B. Walden and David Tomberlin. Estimating fishing vessel capacity: a comparison of nonparametric frontier approaches. Marine Resource Economics, 25(1):23–36, March 2010. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wang:2019:AST

[WTW^{+}19] Ping Wang, Nhuong Tran, Norbert L. W. Wilson, Chin Yee Chan, and Danh Dao. An analysis of seafood trade duration: The case of ASEAN. Marine Resource Economics, 34(1):59– 76, January 2019. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Wessells:1993:IDJ

[WW93] Cathy Roheim Wessells and James E. Wilen. Inventory dissipation in the Japanese wholesale salmon market. Marine Resource Economics, 8(1):1–16, Spring 1993. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Xie:2008:EER

[XKM08] Jinghua Xie, Henry W. Kinnucan, and Øystein Myrland. The effects of exchange rates on export prices of farmed salmon. Marine Resource Economics, 23(4):439–457, ???? 2008. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Xie:2011:CAF

[XM11] Jinghua Xie and Øystein Myrland. Consistent aggregation in fish demand: a study of French salmon demand. Marine Resource Economics, 26(4):267–280, December 2011. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Xie:2014:ICS

[XZ14] Jinghua Xie and Dengjun Zhang. Imperfect competition and structural changes in the US salmon import market. Marine Resource Economics, 29(4):375–389, December 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yang:2020:DCS

[YAA20] Bixuan Yang, James L. Anderson, and Frank Asche. Determinants of China’s seafood trade patterns. Marine Resource Economics, 35(2):97–112, April 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yahaya:1988:FMR

[Yah88] Jahara Yahaya. Fishery management and regulation in Peninsular Malaysia: Issues and constraints. Marine Resource Economics, 5(2):83–98, ???? 1988. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yamamoto:1995:DCB

[Yam95] Tadashi Yamamoto. Development of a community-based fishery management system in Japan. Marine Resource Economics, 10 (1):21–34, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ye:1996:BIB

[YB96] Yimin Ye and John R. Beddington. Bioeconomic interactions between the capture fishery and aquaculture. Marine Resource Economics, 11(2):105–123, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Ye:1996:BM

[YBG^{+}96] Yimin Ye, John R. Beddington, R. Quentin Grafton, Hamady Diop, Richard F. Kazmierczak, Jr., Hamady Diop, Richard F. Kazmierczak, Jr., Yimin Ye, John R. Beddington, Tai Shzee Yew, and Terry Heaps. Back matter. Marine Resource Economics, 11(2):??, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yandle:2015:WHW

[YC15] Tracy Yandle and Scott Crosson. Whatever happened to the wreckfish fishery? An evaluation of the oldest finfish ITQ program in the United States. Marine Resource Economics, 30 (2):193–217, April 2015. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yamamoto:1995:BM

[YCN^{+}95] Tadashi Yamamoto, H. F. Campbell, R. B. Nicholl, Stanley Wang, Tadashi Yamamoto, Stanley Wang, H. F. Campbell, and R. B. Nicholl. Back matter. Marine Resource Economics, 10(1): ??, Spring 1995. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yeto:1997:PGC

[YGB97] Susana Cabrera Yeto, Rosario Gómez Garcia, and Gumersindo Ruíz Bravo. Potential gains from cooperation for vessels and countries. Marine Resource Economics, 12(2):145–158, Summer 1997. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yew:1996:EDA

[YH96] Tai Shzee Yew and Terry Heaps. Effort dynamics and alternative management policies for the small pelagic fisheries of Northwest Peninsular Malaysia. Marine Resource Economics, 11(2):85– 103, Summer 1996. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yu:2006:EOP

[YL06a] Run Yu and Pingsun Leung. Errata: Optimal partial harvesting schedule for aquaculture operations. Marine Resource Economics, 21(4):??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yu:2006:OPH

[YL06b] Run Yu and Pingsun Leung. Optimal partial harvesting schedule for aquaculture operations. Marine Resource Economics, 21 (3):301–315, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yu:2006:BM

[YLB^{+}06] Run Yu, Pingsun Leung, Kathryn D. Bisack, Jon G. Sutinen, Chih ming Hung, Daigee Shaw, Run Yu, Pingsun Leung, Chih ming Hung, Daigee Shaw, Kathryn D. Bisack, and Jon G. Sutinen. Back matter. Marine Resource Economics, 21(3):??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Young:2002:TBF

[YM02] James A. Young and James F. Muir. Tilapia: Both fish and fowl? Marine Resource Economics, 17(2):163–173, ???? 2002. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yohe:1984:RUU

[Yoh84] Gary W. Yohe. Regulation under uncertainty: an intuitive survey and application to fisheries. Marine Resource Economics, 1 (2):171–192, ???? 1984. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Yohe:2003:BRN

[Yoh03] Gary W. Yohe. Book review: Norbert P. Psuty and Douglas D. Ofiara, Coastal Hazard Management: Lessons and Future Directions from New Jersey. Marine Resource Economics, 18 (3):287–289, ???? 2003. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zhuang:2024:SAA

[ZAS^{+}24] Haijing (Jennifer) Zhuang, Jeffery Adkins, Michael D. Smith, Monica Grasso, Christopher Lauer, Kate Quigley, Lauren Knapp, Charles Colgan, and William Nicolls. The satellite account approach for measuring the US marine economy. Marine Resource Economics, 39(2):101–122, ???? 2024. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zeller:2006:FCG

[ZBP06a] Dirk Zeller, Shawn Booth, and Daniel Pauly. Fisheries contributions to the gross domestic product: Underestimating smallscale fisheries in the Pacific. Marine Resource Economics, 21 (4):355–374, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zeller:2006:BM

[ZBP^{+}06b] Dirk Zeller, Shawn Booth, Daniel Pauly, Matthew Berman, G. T. (Stan) Crothers, Lindie Nelson, Dirk Zeller, Shawn Booth, Daniel Pauly, G. T. (Stan) Crothers, Lindie Nelson, and Matthew Berman. Back matter. Marine Resource Economics, 21 (4):??, ???? 2006. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zhang:2020:AAN

[ZFG20] Xumin Zhang, Yingkai Fang, and Zhifeng Gao. Accounting for attribute non-attendance (ANA) in Chinese consumers’ awayfrom-home sustainable salmon consumption. Marine Resource Economics, 35(3):263–284, July 2020. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zhang:2014:ERV

[ZKD^{+}14] Dengjun Zhang, Henry W. Kinnucan, Roy Endre Dahl, Atle Oglend, Eivind Hestvik Brækkan, Sverre Braathen Thyholdt, Yuko Onozaka, Håvard Hansen, Arne Sørvig, Roy Endre Dahl, Atle Oglend, Yuko Onozaka, Håvard Hansen, Arne Sørvig, and Hans-Martin Straume. Exchange rate volatility and US import demand for salmon. Marine Resource Economics, 29(4): 411–430, December 2014. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zhai:2009:IBT

[ZS09] Guofang Zhai and Takeshi Suzuki. International benefit transfer related to coastal zones: Evidence from Northeast Asia. Marine Resource Economics, 24(2):171–186, ???? 2009. CODEN JM-REDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zhou:2016:IVC

[ZS16] Rong Zhou and Kathleen Segerson. Individual vs. collective approaches to fisheries management. Marine Resource Economics, 31(2):165–192, April 2016. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zhang:2023:RPB

[ZSG23] Dengjun Zhang and Geir Sogn-Grundvåg. Reserve prices and bidder participation in an English auction: The case of Atlantic cod in Norway. Marine Resource Economics, 38(3):249–261, ???? 2023. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zhang:2022:IPV

[ZT22] Dengjun Zhang and Ragnar Tveterås. Influence of price variability and financial ratios on business failure in the Atlantic salmon industry. Marine Resource Economics, 37(2):183–200, April 2022. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).

Zheng:2021:FAC

[ZWS21] Qiujie Zheng, H. Holly Wang, and Jason F. Shogren. Fishing or aquaculture? Chinese consumers’ stated preference for the growing environment of salmon through a choice experiment and the consequentiality effect. Marine Resource Economics, 36(1):23– 42, January 2021. CODEN JMREDD. ISSN 0738-1360 (print), 2334-5985 (electronic).